A Complete Bibliography of Publications in History of Science

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

15 April 2025 Version 1.31

Title word cross-reference

e [Slo94]. F = MA [Ver96].

-96 [WGvA22].

1 [Bat98a, Cla97a, Fis82a, Kol06a, Mag99a, Old79a]. 14th [Has79]. 1535- [WGvA22]. 15th [Kel66a]. 1730s [Ili93]. 1760-1840 [Kus16]. 1790s [Gol16, Kno99]. 17th [Hei81, Ver14]. 1870s [SS15, SM24]. 18th [ES15, Hei81, Mil81]. 18th-century [ES15]. 1910s [Pag16]. 1920s [Are19]. 1929 [Cro01]. 1930S [AA87, SM24]. 1939-1967 [Flo22]. 1940s [Por92]. 195 [Kus16]. 1950s [Fyf22, Pag16]. 1955 [Pal96]. 1960s [Lee23]. 19th [Haj15, Urb23]. 19th-century [Haj15].

2 [Bat98b, Cla97b, Des85a, Fis82b, Fra73a, Kol06b, Lon78a, Mag99b, Old79b]. 20th [She15]. 20th-century [She15].

50s [Are19].

88 [Slo94].

Abandoned [Val12]. Abhandlungen [Whi64b]. Abiding [Ort05]. Ability [Dal97f, Goo04]. aboard [Bil22]. above [Pum91]. Abraham [Eyl64]. Abroad [McE13]. absolute [Oak18]. Absolutism [Kno79]. Absorption [Jam83]. Abuses [Jar00b]. Academic [Car68, SL86]. academies [Car16]. Academy [Out83, Rab13, Wei88, Car16, Gol16]. Acclimatization [Liv87, Sam23]. Accumulation [Kin14, Rob14a, Web14]. Accurate [Sib20]. Achievement [Lon81]. Achievements [Ano62d]. Achromaticity [Sor01]. Acoustic [Val12]. Acquired [Hol92]. across [Chi15b, Gao15, Mea24]. Actes [Slo94]. Action [Rav62a, SJ21]. Active [Hen86]. Activity [GG04]. Actualism [Ham93]. Adam [Buc09, Cle89]. Adamachi [Chi23]. Adaptation [DT10]. Admirable [AN13]. Admiralty [Bar17]. Advancement [Nev86, Por82]. Advent [Eam84]. advertisement [Mer20]. Aesthetic [Sen00a]. Aether [Nav21, Wil74]. Affair [Mil08a]. Affective [Goo23]. Afflictions [Dal94]. Africa [Wor94]. African [Gou25, Osb16]. Afro [Pal93]. Afro-Centrism [Pal93]. After [Whi62a, App96, Byn75, Hom85, PPT17]. Aftermath [Mas02]. Afterword [Som22, Bru17, Lav22]. Afterwords [Pic11b]. against [Can76, Pag16]. Agassiz [BdMF01]. Age [Buc09, Hes62, Kno79, Por79, Sho94, Mun21, You67, All17, Hei78, Woo91]. agency [Bil22]. Agendas [Sil12]. Ages [Eas86, Gra11, Hes62, Mol81, Sch63b, Sch65, Sio88, Was65, You67]. Ago [Wil02]. Agricultural [Pal94, Fag21]. agriculture [Jon16, ND22]. Agronomes [Rap69]. Agronomie [Rap69]. AHA [Gru81]. Aided [DT10]. Aims [Smi82]. Air [Boa13, Mar98, Woo88, Sin22]. Air-Engine [Mar98]. Air-Pump [Woo88, Sin22]. Airy [Ash98, Gil20]. Akademgorodok [Tan98]. Al [Kin75, Has79, Joh98, Kin75, Par88, Kin75]. Al-Din [Has79]. Al-Handasiya [Kin75]. Al-Hiyal [Kin75]. Al-Jazari [Kin75]. Al-Khwarizmi [Par88]. Al-Razzaz [Kin75]. Al-Shatir [Has79]. Alamos [Sla24]. Albert [Mil75]. Album [Eas86]. Alchemical [Kas11, Fig73b]. alchemischen [Fig73b]. Alchemist [Ano77b]. Alchemy [Ano77b, Deb67, Loe70, ND22]. Alcmaeon [Nut74]. Alexander [Nic87, Wil65b]. Alexandrian [Lon81]. Alfonso [Mel96]. Alfred [Cro98]. Algebra [Par88]. Alienists [Gua91]. Alistair [Nor96]. Allegro [Not17]. Almada [Car16]. Almanacs [Mor16]. Almeida [Sil12]. Alternative [Nav21]. Aluminum [Osb16]. Ambivalence [Dal97i]. America [Ang12, Byn73, Hew16, Hol25, Liv91, Pet10, Ras97]. American [Ste65, Ang12, Are19, Byn73, Del05, Doe03, Gou74, HPtH24, Mun21, Nic88, Oli24, Pal96]. Americana [Pos15]. Americans [Tam16]. Ammal [Dam13]. among [Has87, Roq22, Ver14, Was64]. Amusing [Lac09]. Anachronism [Jar00b]. Analyses [Pic11b]. Analysis [Ham11, Jam83, Old86, Por91].

Analytic [Aic09]. Analytical [Bro67, Pic94]. Analytical/Comparative [Pic94]. Anatomical [Mae07, Rup90, Wil65a]. Anatomie [Wil65a]. Anatomist [Wea83]. Anatomy [Pag63, Ber14, Sto18]. Ancestors [Mol76]. Ancestral [FN71]. Ancien [Can01]. Ancient [Big69, Cro75, Kei77, Llo68, Swe93, Tur75, Fle00, Gol19, Nut74, Poy63]. and/as [RRH23b]. André [Car16]. Andreae [Pag75b]. Andreas [Cla65]. Andrei [Lin06]. Andrews [Mor16]. Andreyevich [Lin06]. Anemics [Jar04]. Anfängen [Pag74a]. Angel [Hut12]. Anglican [Whi05]. Animal [Far95a, Gib12, Ish06, Jor78, Pal97, Wil02, Hol25, Uys24, WM22]. Animals [Bar22, Cha10, Geo80]. Annotated [Ano11a, Hos03a]. Answers [Sch74]. antagonism [GH20]. anthropologie [Roq22]. Anthropology [Hes62, Web74, SS15]. Anthropometric [SS15]. anthropometry [Smi20]. Anti [Cal98, Edg05, Ols78]. Anti-Darwin [Cal98]. Anti-Historian [Edg05]. Anti-Science [Ols78]. Anti-Spencer [Cal98]. Antiquarianism [Rou79]. Antiquities [Sch10b]. Antiquity [Eas86, Fle00, Mat63, Sio88, Gol19]. Antisepsis [Fox88]. Antoine [Liv05]. Antonio [NG11, Rap91]. Aplatisseur [Ili93]. Apparatus [Tur70]. Appear [Wes88]. Application [Swe93]. Applied [Bud13, Luc99]. Appraisal [Byn75]. Appraising [Wal02]. Approaches [Ang12, McG73, Smi82]. Appropriation [Sab87, OEMB19]. Arab [Has79, Has87]. Arabia [Con95]. Arabic [Ano76f, Has79, Kin77, Ras89, Ste65]. Arabs [Has87]. Arcana [Eam84]. Archaeology [Coo82, GCM05, GCM06, Por12, Sch10b, Rig17]. archeological [Van17]. archéologie [Poy64]. Archeology [Car17, Sav17, Bru17, CV17]. Archibald [Fri03]. Archimedean [Mol81, Van06]. Archimedes [CDG02, Mol81]. architects [Hol25]. Architecture [For94b, WB14]. Archiv [Pag74a]. Archive [Bal89, Hei68]. Archives [Ano77a, MF73, Bar20b]. area [Fig73b]. Areas [Gue79]. Aretini [Pag75b]. Argument [Ili98, JJ76]. Aristarchus [CDG02]. Aristoteles [Mau64]. Aristotelian [Duc05, Gra87, Sio88, Van06]. Aristotelianism [Gra78, Gra87, Sch73, Sch63b]. Aristotle [Mau64, Dal97a, Hal91a, Hal91b, Tem63, Wil65a]. Arithmetica [Eva78]. Armagh [Tur91]. army [Sto22]. Arnold [Whi05]. Aron [Kat13]. array [Sha22]. Arrive [Old84]. Arruda [FTC16]. Arsen’ev [SP07]. Art [Par88, Ste91, vdR04, Jac17]. Article [Sch95]. artificial [Ame24]. Artisan [HP07, Sec94]. artisanal [Osb16]. Artisans [Gau06]. artisanship [Lee19]. Artistry [Ili98]. Arts [Cro80, Ili98, Ksi15]. Asepsis [Fox88]. Ash [Has79]. Ash-Shatir [Has79]. Ashmole [Deb67, Tur86]. Ashmolean [Tur86]. Asia [Fin00]. Asimov [Jen81]. Aspect [Buc70]. Aspects [SS80]. assemblage [Bul18]. Assessment [Coh99]. assessments [You67]. Assistant [Hos02]. Associated [Ste05]. Association [Fen15, Nev86, Por82]. Associations [Nea99]. Astrology [Dal97d, LN17, Ver14]. Astronomer [Has79, Hos02, Wes80]. astronomers [Ver14]. Astronomical [Ano11a, Cha06, Hos02, Kei77, Mad63, Bel22]. Astronomie [Hos66]. astronomique [Hel64]. Astronomy [Alp91, App96, Cle89, Hua16, Kum23, McC85, Rav63a, Smi81, Tur91, Whi83, Wil83, Spe15]. Astrophysics [Smi81]. Athena [Pal93, Mok07]. Athens [Lon80, Ols78]. Atlantic [Can17, Del05, Gou25]. Atlantis [Tan98]. atmosphere [Fra24]. Atom [Fis73]. Atomic [Ber66]. Atomism [Han70, Hen82, Lan09]. Atoms [Kim92b]. atque [Pag75b]. attained [Fle18]. Attempt [AT12, Bro67]. Attempts [Cor14]. Attenborough [Gou11b]. attentiveness [Yil23]. Attitudes [Cro75]. Atttractive [Far95a]. Audience [Sha74b]. August [Jac06]. Augustan [Rou79]. Augustine [Sch63b, Sch65]. Augustus [Ric96]. aurora [Ame24]. Ausgabe [Pag73]. Australian [Hom95]. Author [Bon14, Ano84a]. Authorities [Goo08]. Authority [AA87, Dac01, Mye92, Pal07, Pic88, Cha18, Haj15, Hel20, LN17]. Authors [Tur78]. authorship [Bil22]. Autobiographies [Win03, Cha08]. automobility [Tet23]. Autonomy [Mae07]. Autrecourt [Sti92]. Average [Don14]. Avian [Hol25]. Aviation [Bra23]. Avogadro [Fis82a, Fis82b, Bro81b]. Avoiding [Nea99]. awareness [Chi23]. Aztec [Gue69].

Baartman [Qur04]. Babbage [Dol98]. Babylon [Her62b, Stu77]. Back [BZ09, Duc05]. Background [Fra73b, Fra73a, Geo80, May02, Nic88, Ols90, Smi73, Whi66b, VFD21]. Bacon [Phi15, Rav66b, Ste05]. Baconianism [Yeo85, Yeo07]. Bacteriology [Ham11]. Baedeker [Sen92]. balance [Bia19]. Balloon [Pet10]. Balls [Sec85]. Banal [Liv05]. Band [Whi62b, Whi64b]. Banks [Mil81]. Barcelona [Bar22]. Barometer [Chi65]. Baroque [Kno79]. Barrett [Noa04]. Basel [Hos03a]. Basic [Sta86]. Basis [Sch63b]. Bassi [Row98]. Bath [Hos03c]. Battery [Kat13]. battles [Ara19]. bauxite [Osb16]. BBC [Gou11a]. Beach [Sch09a]. bearbeitete [Pag73]. Beasts [Cha10, Mil08c]. Beautiful [Sen00a]. Beauty [Sen00a]. Became [Hun95, Sch95, Chi15b]. Becoming [Cro01]. Bedeutung [Mau64]. Bedlam [And03]. Been [Coh99]. Beer [Ver90]. Bees [Bel09]. before [Buc09, Kas84, Kel10, Key63, Por76, SP07]. Began [Tei03]. Beginning [McG85]. Beginnings [Nee73]. Begriffe [Pag73]. Behaviour [Bro71, Wil02]. Behavioural [You66]. Behemoth [Woo88]. beiden [Mau64]. Beighton [Cos02]. Being [Byn75, Gib12, Jac03, Mil00, SLT17, Urb23]. Belgium [Don14]. Belief [Sho87]. Believe [McG85]. Bell [Ber14]. Belong [Yox79]. Beloved [Lew62]. benefactor [Leo24]. benefits [Bia19]. Berlin [Ter90, tH04]. Bermuda [Ton19]. Bernal [Tei90]. Bernard [Jor78]. best [Hel20]. Between [HV19, Yeo07, Ano20, Chu03, Gru81, Gua91, Han70, Hop97, Lev70, Maa10, Noa04, Pau94, Rab13, Sch65, Sha81, Sta21, Sur19, Wer23, Whi96, vdR04]. Bewegungslehre [Mau64]. Beyond [Dal97g, Kra24, Noa05, SLT17]. Bible [Kno96]. Bibliographical [Ege77, Lau68]. Bibliographie [Hos66]. Bibliography [Fig79, Out76, Por81, Whi73]. Bibliotheca [Tur94]. Bibliothèque [Kel66a]. Bicentenaire [Slo94]. Bichat [Pic81]. Biddell [Gil20]. Big [AA06, Gro13]. bigotry [Nav21]. bilans [You67]. Bild [D’A90].

Bild-Conception [D’A90]. Bilingualism [Sil12]. Bio [Fig76, Rey08]. Bio-Medical [Fig76]. Bio-Political [Rey08]. Biochemical [Bro92a]. Biochemistry [Bro92a, Hol75, Nee73, Tei80, vH03]. Biographical [Jen81]. Biographies [Col89]. Biography [Han79, Har83, McG66, Ost93, Out76, SP90]. Biological [Ack67, Gou74, Har96b, Men64, Ras97, Ton19]. Biology [AA85, AA87, AA06, Bel09, Dal97a, Hop97, Koh02b, Pal96, Pal97, Ree06, Sch74, SdC11, Yox79, Kon19, Mun21, Car88]. Biomedical [Ili98, San17]. Biometrics [Mag99a, Mag99b, Tuc20]. Biophysics [Ras97]. biosocial [CP21]. Biosynthesis [Tei80]. Biotechnics [VT24]. Biotheoretical [AA87]. Birds [Bel09, Bro98b, Hal91a, Hal91b, Bru17]. Birth [Dal98a, Gol02, Pet23]. Birthday [Nor66]. bis [Pag74a]. Bitter [Web14]. Black [SLT17, VFD21, Pal93]. blaming [Hyu21]. Bletchley [Smi14]. Blindness [Mus00]. Blood [Mas03, Pel01, dB13]. Blowing [Mar98]. blunder [Nav21]. Bodies [Han70, Jac95c, Kim92a, Löw11, San17, Sho86, Fle18, Gia15]. Body [Fle00, Mor99, Pic81, Sen00a, Sho86, Tom84, Sch18, WM22]. Boethius [Whi64b, Eva78]. Bohr [Rot87]. Boltzmann [D’A90]. Bombe [Smi14]. bones [Tam16]. Book [Ait75, All76, Arm62, Bar64, Bee75, Ber66, Bro62, Bro81c, Bro98b, Bry77, Byn73, Cai65, Chi65, Cla64, Cla65, Coc63, Con95, Cro64, Cro98, Cuo97, Dal95, Dal97a, Dal97b, Dal97h, Dal97c, Dal97d, Dal97e, Dal97f, Dal97g, Dal98a, Dal98b, Dal99, Dav62, Deb63, Duk63, Fig73a, Fle00, Gin00, Goo76, Gow73, Har62, Har63c, Har63a, Har63b, Hel64, Her62a, Her62b, Hes62, Hom03, Hos62, Hos63a, Hos66, Hos73, Hos03a, Jac95b, Jac95c, Jen81, Jon74, Kel62a, Kel62b, Kel66a, Kel66b, Kel76, Key63, Kin75, Kin77, Kni95, Las73, Lew62, Luc95, Mat63, Mau64, Nee73, Nev76, Nor66, Old98, Pag75a, Pal97, Por76, Por81, Por01, Poy64, Rav63a, Rav63b, Rav66a, Rav66b, RS62, Row98, Row99, Rud63, Rya66, Sch77, Sen00a, Shr66, Sme64a, Sme64b, Smi81, Spe62, Sta63, Ste65, Ste05]. Book [Sut75, Tal65, Tan98, Tur73, Tur75, War65, Whi66a, Whi64a, Whi64b, Whi63b, Wig63, Wil65b, Win03, Zub64]. Bookmaking [Ste65]. Books [APT20, Ano62a, Ano62d, Ano62e, Ano63a, Ano63c, Ano64a, Ano64c, Ano65a, Ano66a, Ano66c, Ano67a, Ano67c, Ano68b, Ano69b, Ano70b, Ano71, Ano73f, Ano73g, Ano73h, Ano74g, Ano74h, Ano74i, Ano75e, Ano75f, Ano75g, Ano75h, Ano76d, Ano76e, Ano77g, Ano77h, Ano77i, Ano78e, Ano79e, Ano79f, Ano79g, Ano80d, Ano80e, Ano80f, Ano80g, Ano81e, Ano81f, Ano81g, Ano82e, Ano82f, Ano82g, Ano83e, Ano83f, Ano83g, Ano83h, Ano84g, Ano84h, Ano84i, Ano84b, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano88e, Ano89f, Ano89g, Ano91e, Ano91f, Ano91g, Ano92e, Ano92f, Ano95e, Ano95f, Ano98e, Ano98f, Ano99e, Ano99f, Ano99g, Ano99h, Ano00e, Ano00f, Ano01e, Ano01f, Ano01g, Ano02e, Ano02f, Ano02g, Ano03e]. Books [Ano03f, Ano03g, Ano04e, Ano04f, Ano04g, Ano05e, Ano05f, Ano06e, Ano07e, Ano07f, Ano08e, Ano08f, Ano08g, Ano08h, Ano09f, Ano09g, Ano09h, Ano10d, Ano10e, Ano11e, Ano11f, Ano11g, Bro81c, Eam84, Kas11, Key63, Por76, Rav62b, Ste02, Die22a]. borderline [JS24]. Borelli [Hel64, Mel96].

Boscovich [Rav63b]. Boston [Nor66]. Botanical [Bro84, Por76, Bec22, ES15, WGvA22]. Botanist [Bar79a, Mil81]. Botanists [Bro84, Sec94]. Botany [Bil22, Bro84, Liv05, Ogb13, Die22a, Die24, Lab21]. bottom [All17]. Boubée [dMF23]. bound [Chi15b]. Boundaries [Str95]. Boundary [Löw92, GGN16]. boundary-crossing [GGN16]. Bounding [Sha92]. box [ES15]. Boyle [Hal70, BC00, Hal70, Hun95, Ili97, Jac95a, Ost93, Por01, Win97, Woo88]. Brahe [Chr98, Tho73]. Brahmin [Kum23]. Brain [Kas84, Smi88]. Brazil [BdMF01]. Breadfruit [Siv01]. Brewster [Hom85]. Bridge [Noa04]. Brief [Bra77, Mad63, Pic11a, Tur69, Not17]. Briefwechsel [Dou88]. Britain [WS89b, WS90, Bal12, Ber14, Ber66, Biv20, Bro77a, Can85, Dal94, For94b, Fre71, GG97, Gib12, Gou11a, Hom85, Hua16, Kee14, Leo24, Lev70, Liv91, Mil12, Mor94, Mor18, O’C07, Pag71, Pag16, Pal94, Por80a, Por91, Por92, RS62, RC13, Sec86, Smi20, SC98, SS80, Tur78, Ver90, Whi63b, Wil80a, Wil02, WS89a, Yeo85, dB13]. Britannica [Fal21]. British [Aga13, Alb96, Ano73i, Ash08, Ash14, Biv99, Bro81c, Bro92c, Cap23, Coh99, Edg05, Jan00, Jub85, Kol06a, Kol06b, Lab21, Luc95, MF73, May02, Nev86, Pag16, Pal96, Por76, Por82, ST74, Smi78, Sta86, Woo13]. broader [RSM20]. Broadsides [Wal99]. Brokering [Hyu21]. Brothers [GG97]. Brought [Sch03]. Brunelleschi [Ste91]. Bruno [Clu08]. Brussels [Cla65]. Bubble [Ras97]. Buffon [Slo94]. Building [Mou21, Rob21]. bullet [Uys24]. Burden [Cha10]. Bureaucracy [Pic81]. Bureaucratic [Ham93, Fel18]. Burmese [Ren23]. burn [Rab19]. Bursting [Buc09]. Business [Alb96, Hos05, MP14]. Butterfield [Jar03, MA09].

C [Cro98, vH03, Alg23, Cro07, Dal97f, GC10, Jon16, Lab21, SC98, Tet23, Uys24, Woo13, tH04]. C. [Bra88, Edg05]. Cabanis [Lev84]. Cabinet [vdR04]. Caesalpini [Pag75b]. Cain [Mer00]. California [Mor93]. Calling [RH99]. Calouste [Rap21]. Cambridge [Buc62c, DV01, Gas85, Kno96, Nav06, Phi06, Woo91]. Cameron [Nor96]. Camillo [Rob05]. campaign [Sto22]. Can [Bia19, Tam16]. canaries [Far23]. cannot [WB14]. Capitalism [Par84]. Captain [Leo24]. Care [Gol02]. Career [Mag99a, Mag99b, Mor16]. Careers [Ben88, Dea03, Gay08]. Cargo [Sch06]. Carl [Web14, Drö15]. Carnap [Nor66]. Caroline [Cha08, Hos02, Win03, Cha08]. Cartelli [Kel76]. Casaubon [Ste02]. Case [Bro81b, Cro08, Fox88, Gas85, GG90, HM10, Hom95, Löw92, Mil00, Out76, Rap91, SP90, Ste95, Whi71, Wil00, Par20, Web14]. Case-Studies [Bro81b]. Case-Study [Bro81b]. Cases [GC10]. Caspar [Web14]. Cassini [Ili93]. cast [Rie15]. caste [Kum23]. Catadioptrics [Ben06]. cataloging [Fig73b]. Catalogue [Key63, Tur86]. Cathedrals [SP87]. Catholic [Mar82]. Causal [Sti92]. Causality [Hen80]. Causes [Duc05]. CE [dP18]. Celestial [Ait81, Rob05]. Celestino [OEMB19]. Cell [Rey08]. Censoring [Tar14]. Census [Hig85, Hos03a, May12]. Centenary [Ano74a]. center [Mun21]. Central [Wor94, Tet23, LJG^{+}24]. centralization [Rob14b]. Centre [Ano77a, Rob14b]. Centrism [Pal93]. Centuries [Hei81, Geo80, Liv88, Rob66, Suz95]. Century [Ang12, Cos02, Gau79, GG04, Hal68, Har98, Hua16, Mae07, Mil81, Mil12, Mor16, O’C07, Pet10, Por80a, Por91, Ras97, Rup90, Sec94, Sha74b, Sha91b, Sta13, SS80, Vla99, Wil80a, WS89a, WS89b, WS90, Yeo85, vH94, Ame24, Ano20, Bar20a, BB21, Bri20, Car16, Cha18, Cha25, Coo18, Die22a, ES15, Eve23, Fag21, Fle21, Gil24, Glu24, Gor15, Haj15, JS24, LP23, ND22, Por20, Rie15, Rob21, She15, Sto18, Sur19, Tam15, Urb23, Ver14, Whi17, cW18, Ben06, Bos05, Bri09, Bro13, Byn73, Can82, Can85, Car62, Car63, Chr87, Clu08, Coo82, Cor14, Csi10, Cur12, Dac01, Dal98b, Dam13, Dob77, Dol84, Eam84, Ebr00, Eva78, Fan00, Far95a, Fig75, Fig76, For84a, For94b, Frä74, GG97, Gib12, Gol87, Hal63, Hal66, Has79, HP07]. Century [HSS10, Jac04, JJ76, Jac83, Jac95c, Kee14, Kel66a, Kel66b, Lac09, Lev70, McE13, McG85, Men64, Mil89, Mil08c, Nev76, Ogb13, Out78, Per87, Pic94, Rap69, RH99, Row98, Sch83, Sch68, Sch74, SP87, Sim12b, Smi75, Smi78, Smi88, Sou92, Ste96, Ste02, Tar14, Tom91a, Tur75, Wal97, Wal99, Web74, Wes67, Wes80, Wil62, Wil83, Wil74, Yeo07, dB13, vdR04, Lew62]. Certainty [Cos02, Cro80]. Cesalpino [Pag75b]. Châtelet [Ter95]. Chadwick [Coc63]. Chain [Byn75]. Chair [Dac01]. Challenge [Jan00]. Challenges [Kel76, De 17]. Chameleons [Rab13]. Change [Dol88, GC10, Gei81, Hop05, Lev84, Mat71, Per87, RH88, Sou63, Wil80b]. Changes [Mar82]. Changing [Arm62, Hes64, Høy96, Por92, GS16]. chapter [Gui17]. Characters [Eyl64]. Charcot [Gil83]. Charles [Ano74a, Bel09, Ber14, BdMF01, Byn88, Col89, Dol98, Rud63, Rus74, Sta63]. Chart [Smi78]. Cheerfully [Jac03]. Chemical [Bro92b, Deb74, Far23, Fis73, Kim92a, Kim92b, McE97, Per87, Jac17, Lee23]. chemicals [GGN16]. Chemie [Pag74a, Kra24, Fig73a]. Chemist [Kik12, Pag16]. Chemistry [Bro67, Bro92b, Cha11, CG82, Chr87, Deb63, Fig73a, Fis82a, Fis82b, Gol87, Hes62, Hom03, Jam86, Jor78, Kim92b, Kni93, Kni95, Le 16, McE97, Mil04, Pag66, Rob16a, Sme64a, Sme64b, Ste65, SM24, Tha66, Ber16, Jon16, Kra24, Osb16]. Chemists [Fis82a, Fis82b]. chemopolitics [Muk16]. Childbirth [Dal97c]. Children [Dal95, Zil14]. Chile [Wal22]. Chimie [VB87]. China [Bra23, Cha25, Chi15a, Chu03, Cul93, Dik91, Fen15, Fra24, Gao15, Mou21, Sch06, She15, Swe93, pY15, dP18]. Chinese [Ano11a, Biv99, Fen15, JS24, Kei77, Loe70, SV18, pY15]. Choice [Fox88]. Cholera [Pic84, Sch24]. Chorographic [Jan00]. choses [Rob14b]. Choson [Lee19]. Christian [Tur75]. chromosomes [San17]. Church [Tur91]. Chymia [Cro64]. Chymist [Woo88]. chymistry [ND22]. circa [Fel18]. Circle [Bat98a, Bat98b, Mel96, APT20]. Circles [Ait81, LC98]. Circulating [BZ11, San17]. Circulation [Bro13, McE13, dP18, MB23, Ren23]. circumnavigation [Wer23]. Citing [Win97]. Citizen [Jac95b]. Citizen-Patient [Jac95b]. citizenship [Tuc20]. City [Smi75, Tan98]. Civil [Maa10, Smi75]. Civilization [Ort05, Pye93]. Civilizing [Siv01]. Claim [Kol06a, Kol06b, Pag75b]. Clarifications [Pep76].

Clarissimi [Pag75b]. Class [MJ01]. Classical [Cas84, Har83, Sza90, Yol71]. Classics [Kus16]. Classification [Hal91a, Hal91b, Löw11, Tur89]. Claude [Jor78]. Clavius [Cla09]. clay [Gil24]. Clean [Sim17]. clear [Are19]. Clergy [Goo83]. Clergyman [Bar79a]. Clergymen [Mor80b]. climate [Fra24]. Clinic [Pel01]. Clinical [Duk63, GH20, Mar00, Sta86]. clinique [Poy64]. Clocks [BZ11, Glu24, Kin77]. Closed [LC98, Spe15]. Closet [Tur86]. Closing [Bat98a, Bat98b]. Closure [Sil99]. Clouds [Sch77]. Clubs [Liv91]. cluster [Eve23]. Cnidus [Lon78b, Lon78a]. Co [Kol06a, Kol06b]. Co-Prediction [Kol06a, Kol06b]. Coerced [Wal22]. Coffee [Hom03]. Cold [Mar98, Car17, GR22, Pal96]. Coleridge [Sha74a]. Collected [Coc63, Por96]. Collecting [Bro81c, Del11, Koh07, Siv01, Win97]. Collection [Tur70, Tur94]. Collections [Tur86, tH04]. Collective [Dac01, Gin08, Rig17]. Collectors [Buc09, tH04]. College [Mag99a, Mag99b, Rob66]. Collegio [Rap14]. Colloque [Slo94]. Colonial [Cha10, Del05, Muk16, SP87, Ano21, GM18, Kum23, Sin22]. colonization [Can17]. Colour [Mus00]. Combinatorial [Kno79]. Comes [Ang12]. Comet [SZDC13]. Comment [Bru07, Stu77]. Commerce [Ben06, Fyf22]. Commercial [Mil00]. Commitments [Cro95]. Committed [Lav70, Woo85]. Common [MJ01]. commons [GÁ25]. Communicating [LC98]. Communication [Bal12, Edg79, Tam15, Are19, ES15]. Communications [Doo95]. communist [Pet23]. Community [Bar03, Chr74, ST74, GÁ25, VFD21]. Companion [Law90, Law95]. Companions [Law95]. Company [GH20]. comparability [cW18]. Comparative [Lon81, Pic94, Pye02a, SdC11, Wor94, LJG^{+}24]. Compared [Luc95]. Comparison [Sch65]. Competition [Bec22]. compiling [De 17]. Completed [Ano67b, Ano68a, Ano69a, Ano70a, Ano73i]. complexities [Glu24]. complexity [Vig21]. Composite [Mor92a]. Compositionism [Cha11]. compositions [Bil22]. Computer [DT10, Kas84, Kum23]. Computer-Aided [DT10]. computers [Pet23]. computing [Pet23]. Concept [Ait71, Han70, Har63c, Hes64, LeB01, Nic88, Pei65, Web67]. Conception [D’A90]. Conceptions [Sch82a]. Concepts [Har63b, Kal83, Löw92, Per87, Smi88, Swe93, Wil00]. Conceptual [Bud13, vH94, GS16]. Conceptualization [Lev84]. Concerns [Eme88]. Concise [All76]. Condenser [Mil04]. Conditions [Rup90]. conference [Som21]. configuration [Bun19]. configurationibus [Mol69]. Conflict [Gua91, Han70, Hos03c]. Conflicts [Boa13]. Confrontations [Mas02]. Confronting [Dal94]. Confusion [Lev67]. Congress [Cro01, May02]. Coniques [Whi64a]. Connecting [Sav17]. Connection [Gin08]. connoisseur [Byc22]. connoisseurship [BWM22, WM22]. Conquerors [Fin00]. Consequences [Smi94]. Conservation [Han70]. Considered [Law90]. Consolidation [AA85]. Constitution [Kim92a]. Constructed [For84a]. Constructing [Bec04, Rob14b, Gil20]. Construction [Gin08, Gru81, Kin77, Pal05, Sti64]. Constructions [Sen00b]. Contact [Lev67]. Contemporaine [Car63]. Contemporaries [Bat98a, Bat98b].

Contemporary [Ano77a, Bro75, Bro77b, Har63a, CP21]. contest [Muk16]. Contested [Liv87, Tet23]. Context [Hoc91, Kin95, Lev84, McG90, Tei03, Yat73]. Contextual [Liv88]. Continental [Hei79, Rup96]. Contingency [Col81, Mar82]. Continuity [Fag21, vH94]. Continuum [Sio88]. Contribution [Car88, Kin79, Vla99]. Contributions [Rob20a]. Contributors [Ano73b, Ano73c, Ano73d, Ano73e, Ano74c, Ano74d, Ano74e, Ano74f, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano77c, Ano77d, Ano77e, Ano77f, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80a, Ano80b, Ano80c, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84c, Ano84d, Ano84e, Ano84f, Ano85b, Ano85c, Ano85d, Ano85e, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89b, Ano89c, Ano89d, Ano89e, Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a, Ano95b]. Contributors [Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14]. Control [Ash01, Sch07]. Controlling [Fin93]. controversial [Sav17]. Controversies [Llo68, GP18, Gil24]. Controversy [BdMF01, Dol84, Gib12, Hei79, Kno96, Pie86, Ree09, Rus74, Wil80a, Mor18, Mil12]. Convention [Mil08b]. Convergence [Gua91]. Conversation [Wal97]. cooking [Osb16]. coopers [Eve23]. coordination [Bec22]. Copernic [Hel64]. Copernicus [Hos73, Hos03a, Rag07]. Core [Col81, Jam83]. Core-Set [Col81, Jam83]. Cornish [Gil24]. Correcting [Dal95]. Correction [Ano89a]. Correlation [Mag99a, Mag99b]. Correspondence [Dou88, Hal68, HT73, HH73, Hei79, Mil81, Nev86, Whi62a]. Corrigendum [Ano85a, Ano20, Ano21]. Corruption [Loh67]. Cos [Lon78b, Lon78a]. Cosmographia [Wat74]. Cosmography [Wat74]. Cosmological [Mas02, Whi74]. cosmologies [Oli24]. Cosmology [Hut12, Joh98, Sti64, Oli24]. Costa [RH00]. costs [Bia19]. Cottonopolis [Pic84]. counseling [Por20]. Counter [Net02]. Counting [Hig85]. countries [You67]. Course [Ewe63]. Court [Fin93, Coo18, LN17, RH00, Val12]. Cows [Wad01]. Crafting [Gor15]. Crane [RH00]. Crania [Pos15]. Creating [Ame24, Sen00a]. Creation [Jac06, Jam85, Sio88, Web67, Wal22]. Creationists [O’C07, PC18].

Creative [Dal98a, Rot87]. Creativity [Hol81]. Credibility [Cro08, Mal18]. credit [Cha18]. Creed [PV11]. Crisis [Ort05, Gol16, SR20]. Criteria [Car79]. Critical [Edg79, Hak95, Out76, Rus74]. Criticism [Bro13]. Critics [Gru64]. Critique [Cal98, For85]. Critiques [Ols90]. Crombie [Nor96]. Cross [Kik12, Sim12a]. Cross-National [Kik12, Sim12a]. crossing [GGN16]. Crow [Mar11]. Cultivating [Siv01]. Cultural [Del05, Ham11, Jac87, Pic99, Pye82, Rup90, Sen00a, Biv20, GS16]. Culture [Ash08, CP94, Dal97g, Fre95, GG97, Har96b, Har98, Hen80, HP07, Mor80b, Net02, Ost93, Rud80, Sho94, Ter90, Whi05, Win94, Chi15b, Hyu21]. culture-bound [Chi15b]. Cultures [Por05]. Cumulative [Whi73]. Curatorship [Pum91]. Cure [Sen00a]. Curing [Kin95]. Curiosities [vdR04]. Curious [Pug04]. Current [FSB^{+}23, McG73, Whi67, Tur73]. customs [SM24]. Cuvier [Out76, Out78, Pic88]. Cycle [Nic88]. cycles [Spe15].

D [Fal21]. D.176 [Bar17]. daily [Zil14]. Dalton [Mus00]. Danger [Por80a]. dans [Slo94, Wil67, You67]. d’Apollonius [Whi64a]. Dark [Ram10, Whi71, Mou21]. Darkness [Mil89, Sho94]. Darwin [Bar79a, Bel09, Bow09, Byn88, Cal98, Col89, Col00, Dea16, FTC16, Geo85, GM89, Gru64, Has87, Kjæ10, Loe65, Mar82, McG85, Old84, Old97, Por87a, Ree09, Ros24, Rus74, Sta63, vW09]. Darwinian [Por12, Rus74, Rus09, Sho87]. Darwinism [Bow09, Cal98, Flo22, Geo85, Gru64, Has87, Old97, Sho87]. Data [Por87b, Rhe11]. date [She15]. Daughter [Pye90]. David [Fri03]. Davis [GH20, Dav75]. Dawson [SP90]. Dead [Ano62d, Feu93, Rav62b]. Deaf [Aic09]. Deaf-Mutes [Aic09]. Death [Pag81, Pic88, Rup90]. Debate [Dav75, Jam83, May05, Mor94, Por12, Sha92, cW18, Kaa14]. debates [FSB^{+}23, dP18]. Debt [Wil65a]. Decade [Nic88]. Decay [Pag71]. December [Hos03b]. Declaration [Sil99]. Decline [GG04, Hen08, Wea75]. Deconstructing [AA85, Slo90]. Dedicated [Nor66]. Deep [Dea16]. Defence [GM89, Han79]. Defending [Ano21, GM18]. Defining [Mor94, SP07, LJG^{+}24]. Deformities [Dal95]. Deliver [Cor14]. Demilitarized [Hyu21]. democracies [FNG22, Som22]. Demonic [McG73]. Demonstrations [Bri09]. Dephlogisticating [Kno96]. depots [Kin14]. Depth [Can64]. Derek [Stu77]. Descartes [Eri22, Gau06, Kel10, Lev67, Tom84, Yol71]. Descent [Cal88]. descriptions [LS17]. Design [Liv91]. Designed [Cro88]. desselben [Tur73]. Determinism [Gou74, Wal02]. deutschen [Fig73b]. Development [Car79, Chr74, Dav75, Eam84, FN71, Hes62, Kat13, Mat71, RC13, SP87, Smi78, Smi14, Tha66, Was65, Woo13, You75, Col24, GR22, Kon19]. Developments [Bow09]. Devices [Kin75, Kin75]. Devotees [Pal94]. Dew [Tay94]. d’Histoire [Lav70]. diabetes [Tuc20]. Diagnosis [DT10]. Diagrams [Loh67]. Dialectical [Cla89]. Dialogue [Clu08, Mye92]. Dials [Tur89]. Diamond [Mor94]. Diary [Gay03, Gay08]. Diatribae [Haw11]. dictatorships [FNG22, Som22]. Dictionaries [Jen81, Por80b, BP19]. Dictionary [Dal97a, Fox83, Pag75a]. Did [Gin01, Old84]. Diego [APT20]. Different [Mor92b, Tur78]. differentiation [Gao15]. Diffident [Ili97]. Difformity [Mol69]. digital [Kan19]. Dilworth [Hew16]. Dimension [Pag73]. Dimensions [Ano11a, Geo85]. Din [Has79]. Dining [Liv91]. dinosaur [Tam16]. Direction [Coc63, Gin00]. Directions [CG82, SP90]. Disassembling [CV17]. disaster [Sch24]. Discharge [Ram10]. disciple [FTC16]. discipleship [FTC16]. Disciplinary [AA87, Hop05, Löw92, Rap14, Sha22]. Discipline [Bro98b, Gun78, Sha92, Sil99, Rob21]. Disciplines [AA85, Por92, Str95, Gao15]. Disciplining [Fen15]. Disclosed [Eam84]. Discontents [Gin08, Ham11]. Discourse [Dik91, Fan00, Hoc90a, Hoy01]. Discovered [KS03]. Discoveries [Dal99, Geo80, Mas02]. Discovering [Bro98b]. Discovery [Hos63a, Mil08b, Whi74, Cap23]. Discovery/Invention [Mil08b]. Discursive [Ara19]. Discussing [Hom03]. discussion [Pre22]. discussions [RSM20]. Disease [Dal97b, Dal98b, Feu93, Pic84, Pic88, Por12, Wil00]. Disease-Concepts [Wil00]. Disease-Systems [Feu93]. Diseases [Löw11]. Display [For94b, Fle21, SJ21]. Displaying [Qur04]. Dispute [dB13]. disputes [Haj15]. Dissected [Nut74]. Dissemination [Rei07]. Dissertation [Fal21]. Dissociation [Ebr02, LeB01]. Distance [Gin00, LC98, Rav62a]. Distillation [Tom11]. Disunity [Elw03]. Divers [Del11, Kik12]. Diverse [Kel63]. divide [Mea24]. Divided [For94a]. Divine [Rud62, Oak18]. Divinity [Sho87]. Do [Gin01, Hol92, May71, Smi20, Tam16]. Dobzhansky [Kon19]. Doctoral [Ano62b, Ano63b, Ano64b, Ano65b, Ano66b, Ano67b, Ano68a, Ano69a, Ano70a, Ano73i]. Doctors [Dal97h]. Doctrine [BC98]. Documentary [Bro92a, Gia15]. Does [GG90, Yox79]. Dollond [Sor01]. Domain [Geo85]. Domains [Nav06]. dome [Rap21]. domestication [Hol25]. domesticity [Ton19]. Dominant [Cha11]. Dons [Woo91]. Don’t [Bel09]. d’orient [You67]. Dormitive [Cla93, Hut91]. Double [Smi88, Tei75]. doubting [Sim17]. down [dMF23, Sch03]. Dr [Bro94, Don01, Rou79]. Draft [Mil08b]. Drebbel [Kel10]. Drift [Rup96, Wil74]. Drinking [Por91]. Drosophila} [Koh91]. Duae [Haw11]. Dualism [Suz95, Tom84]. Duchenne [Dal97b]. Duels [Kel76]. Dugald [Por96, Por96]. Duhem [Mar82, Mar87]. Duhemian [Mar87]. d’un [Kel66a]. during [BD23, Csi10, GG04, LC98, Rob21, SP87]. Durkheim [Jor78]. Dutch [Maa10, Ver14, vdR04]. Dutrochet [Goo76, Goo76]. Dynamics [Car79, Dal98a, Whi66b, WB14]. Dystrophy [Dal97b].

E. [Dam13, vH03]. Early [Bia06, Cro75, FN71, Hal91a, Hal91b, Har02, Hei81, Hua16, Hun88, Jac74, Jac07, Joh98, Kas82, Keb05, Mad63, Nev86, Osl98, Pic88, Pie86, Por82, Rup98, Sha91b, Spa08, Spe15, Ste65, SdC11, Tei80, Tur78, Tur86, Car16, Car17, Cha18, Cha25, Die24, Fag21, Hal15, Hel20, Leo24, Oak18, Pre22, Sch18, She15, SH18, Tam15, Uys24, Whi17, cW18, Biv99, Can64, Cos02, Dol98, Fig76, Fra73b, Fra73a, Har98, HW86, Ili92b, Kas11, Mor80b, Mor94, O’C07, Rab13, Sec94, Sha91a, Sio88, SS80, Tre10, Ung77, Vla99, Wal99, Win94, vdR04]. Early-Modern [Keb05, Oak18, Kas11]. early-twentieth-century [cW18]. earth [All17, Bar20b, Ksi15, Not17, AT12, Bon08, dMF23, Fle18, Gia15, Ili93, O’C07, Old98, Pag71, Por79, Sch03]. earthquake [Esp24]. Earthquakes [Con95]. easily [Are19]. East [Bro77b, Con95, LJG^{+}24, WGvA22, Wor94]. East-Central [LJG^{+}24]. Eastern [Sch65, You67]. Easy [Dal97a]. Eclipse [Bow09]. Eclipses [Wal99]. Eco [McG90]. Eco-Prophet [McG90]. Ecological [Nic88, Hyu21]. Ecology [Ank07, Ege77, Feu93, Nic88, Pag81, Ree06, Rab19]. Economics [Dal97h]. Economy [Koh91, Nic88, Pic88, Ste07, WS89a, WS89b, WS90, SM24]. Ed [Has79]. Ed-Din [Has79]. Edge [Buc62a, Mar82]. Edinburgh [Edd01, Sha74b]. Editing [Ano76f, Hal02]. Edition [Jor78]. Editions [Pag75b, Bec22]. Editor [Rob16b, Rob22]. Editorial [Ano62c, Ano73a, Ano74b, Mor16, Mor09b, Mor14, Rob15]. Editors [Ano09a]. Education [Ang12, Cur12, Dal97f, Sim12a, Sut75, dMF23, Mal18, VFD21]. Educational [Mar11]. Educationalist [Bar79a]. Educationism [May02]. Edward [Kik12]. Edwards [Ste20]. Effects [Duc05]. efficiency [Mor18]. effort [Rig17]. Egypt [Car17, Con95, Gol19]. Eighteenth [Cos02, Fig75, Mae07, Mor16, Pet10, vdR04, Bri20, Die22a, Fag21, Glu24, Whi17, Ben06, Can82, Can85, Car62, Chr87, Coo82, Dac01, Far95a, Frä74, Gib12, HP07, Kel66b, Lev70, McE13, Mil89, Ogb13, Per87, Rap69, RH99, Row98, Sch83, Sch74, Sha74b, Ste96, Suz95, Tom91a, Tur75, Wal97, Wal99, Wil80a, Wil83]. Eighteenth-Century [Cos02, Mae07, Pet10, Bri20, Die22a, Glu24, Ben06, Can85, Far95a, Gib12, Lev70, McE13, Ogb13, Rap69, Row98, Wal97, Wal99]. Einstein [Gin08, Mil75, Nav21, Pye90, Rot87]. Eisler [Pag73]. Elasticity [Ish06]. Elder [Hos63b]. Electric [Ram10, Esp24]. electrical [Gil24]. Electricity [Hei81, Ile24, Urb23]. Electromagnetic [Ste95]. Electron [RC13]. Elements [PV11, Ung77]. Eleven [Bec22]. Elias [Deb67, Tur86]. Eliot [McG85]. elite [Flo22]. Elizabeth [Far02]. Ellery [Oli24]. Éloges [Out78, Wei88]. Elusive [Wal73, Whi74]. emancipation [Bri20]. Emanuel [RH00]. Embodied [Boa13, Fle21]. embryologist [Buk20]. Embryology [Hop05]. Emergence [Bro98b, Gun78, Kus16, Pal97, Pic88, Rud76, Sch10b]. Emerging [Gei81, FSB^{+}23]. Emetics [Jar04]. Emics [Jar04]. Emil [Jac17]. Émilie [Ter95]. emperor [WGvA22]. Empire [Cro06, Del05, Mou21, Pag16]. Empires [Sur19, Ano20]. Empirical [Sti92]. Encounter [All17]. Encountering [Hal15]. Encyclopaedia [Car91, Fal21]. Encyclopaedias [Pag75a]. Encyclopaedic [Rei07]. Encyclopedia [Jen81, Law95]. End [Jac73, Pye02b, Rup98, Sil99, Ren23]. Endeavour [Bil22]. Ends [Jac73]. Energy [Ber66, Smi78, Tei80]. Engine [Mar98]. Engineering [Has79, Mar13, Mil12, Smi75]. Engineers [Bri09]. England [Cos02, Cro01, Dal97f, Far95a, Fra73b, Fra73a, Hei78, Hes62, Kas11, Lus68, Mor09a, PD04, Rou79, Sha91a, Sha91b, Sim12b, Sou92, Wal97, Wer00, Wil85, Win94, Yeo07].

English [Dal97h, Fri03, Hei78, ND22, Pag66, Wal99]. Enlightened [Jac03]. Enlightenment [Bow74, Gol03, Liv05, Sch03, Woo91, Boa13, Bro98a, Byc22, BWM22, Cor14, Eme88, Gol86, Gol11, Goo83, Hei78, Ish06, Mok07, Mor80a, Tei03, Vla99, Wal73, Woo90, Woo91]. enough [Sib20]. entanglements [Ara19]. Enterprise [Kjæ10, Luc95]. Entertainment [Pet10]. Entrepreneur [Mil81]. environment [Sam23]. environmental [Can17, Dav19, HV19, Kra24, Le 16, MB23]. Ephemeral [Wal99]. ephemeralities [Fle21]. Epigenesis [Wil67]. epigenetics [CP21]. Episode [Tay94]. Episodes [Nic88]. Epistemic [Kel10]. Epistemology [Jor78, GH20, MH23]. era [Gao15, Pet23, San17, Str95]. Ernest [Dea03]. Ernst [Joh05, Rey08, Sta21]. errors [LP23]. Especially [Ili98]. Essay [Ack67, Ait71, Ait73, Alp91, Ano62d, Ano77b, Ano11a, Bar79a, Ben88, Boy67, Bro67, Bro77a, Bro92a, Bro92b, Bro84, Buc62a, Buc62b, Buc70, Buc09, Byn88, Can82, Can85, Car63, Car77, Car79, Car91, Cha08, Coo82, Cur85, Dal94, Dal97i, Dav62, Dav75, Deb67, Del05, Dob77, Dol88, Dou88, Eas86, End09, Eyl64, Fig73b, Fig79, Fis73, For80, For84b, Fox83, Gau79, Geo85, Gil83, Gol86, Gol90, Gou74, Gru64, Hal68, Hal70, HT73, HH73, Hal73a, Hal73b, Han70, Har83, Har98, Has79, Has87, Hei78, Hei79, Hei81, Hen83, Hen93, Hes63, Hes64, Hoc90b, Hoc90a, Hoc91, Hol75, Hom85, Hos63b, Ili92a, Ili97, Ili98, Jac73, Jam86, Jor78, Kel63, Kin75, Kin79, Kni66, Kni93, Lau66, Lav70, Law90, Law95, Llo68, Loe70, Mar82, Mat71, May71, McC85]. Essay [McG85, McG90, McG66, Mil81, Mol69, Mol81, Moo92, Mor77, Mor80b, Mor92a, Mor93, Mor94, Mor09c, Nev86, Nic88, Nut74, Old97, Pag71, Pag63, Pag66, Pag73, Pag74a, Pag75b, Pag81, Pau94, Pei65, Pic88, Por73, Por79, Por80b, Por80a, Por82, Por87a, Por87b, Por91, Por96, Poy63, Pug04, Rap69, Rav62a, Rav62b, Rob66, Rou79, Rud62, Rud69, Rud80, Rus74, Rus94, Sch82a, Sch82b, Sch84a, Sch78, Sch63b, Sch65, Sha81, Sho83, Sho86, Sho87, Sio88, Slo90, Slo94, Smi75, Smi81, Smi82, Smi88, Smi94, Sou63, Spa08, Ste91, Ste02, Ste05, Sti64, Sut86, Tei75, Tei80, Tem63, Tha66, Tom91a, Tur70, Tur73, Tur78, Tur86, Tur91, Tur94, Ung77, Wal73, Was64, Was65, Wat74, Wea75, Web67, Wes67, Wes88, Whi62a, Whi62b, Whi63a, Whi66b]. Essay [Whi75, Whi71, Whi73, Whi74, Whi83, Wil65a, Wil67, Wil80b, Wil85, Wil74, Woo88, Woo91, Yat73, Yol71, You70, You75]. Essays [Bro92b, Car91, McG90, Sho87, Tur86, Yat73]. Essential [Wil80b]. Essentially [Alv13]. Establish [Bro67]. Establishing [Bal12, Ksi15]. Estado [Rap21]. Ether [Sch82a]. Ethereal [Wil74]. Ethers [Noa05]. Ethics [McG90]. ethnobotany [Bri20, Die24]. Ethnographic [Hoy01]. Ethnomathematics [AA86]. Etics [Jar04]. étude [Jor78]. Études [Lav70]. Euclid [Bri09]. Euclidean [Ung77]. Eudiometry [Boa13]. Eugenic [Por12, San17]. Eugenics [Mag99b, GR22, Wal22, Mag99a, RH88]. Eurasianism [GP18]. Eurocentric [Rup96]. Europe [Bia06, Chu03, Hoy01, LJG^{+}24, Mat71, Osl98, Rap91, Spa08]. European [Bro75, Bro77b, BWM22, Cro95, GPP^{+}08, GP18, Ili98, Pye93]. Evaluating [Can01]. Evaluation [Rus74]. Events [Wal99]. Everything [Mae07].

Evidence [Cul93, Jac78]. Evolution [Eng78, Gli99, McG90, Phi06, RS62, Sho87, Slo90, Sta63, Ung77, vW09, Pre20]. Evolutionary [Bru87, Geo85, Koh91, McG85, Old97, Som10, Kon19]. Evolutionism [Bow74]. Exact [Pye82]. exactitude [Cha25]. exakten [Whi64b]. Examining [Bau19, Jac06, SH18]. Examples [Ras89]. Excess [RH00]. Exchange [Gay08]. Exchanges [Chu03]. Excreta [Dac01]. executioner [Wal22]. Exemplary [SdC11]. Exhibiting [Kni95]. exhibitions [CG21]. Exit [Ver96]. Expanding [KS03, Whi62c]. Expansion [Pye82]. Expectations [Biv99]. Expeditions [Bel14]. experience [Hel20]. Experiences [Gru81]. Experiment [Bos05, Dob77, Fin93, Gol90, Hen93, Ili97, Sin22]. Experimental [BC00, Boa13, Eam84, Fin93, Hen86, Hol92, Löw92, Mor92b, Pic88, Sin22, Ste07, Str95, Wer00, Wil02, Win97, Woo88, Gia15, ND22, OEMB19]. experimentale [Jor78]. Experimentality [Rhe11]. Experimentation [Pic11b, Jac17]. Experiments [Dol98, Pum91, Wil74, Coo18]. expert [Are19]. expert-layman [Are19]. Expertise [Ash01, Biv99, Woo13, Can17, LN17, Ren23]. Experts [Aga13, Goo08, Woo13]. Explaining [Ree06]. Explanation [Ili98]. Explanations [Sou92]. Exploration [Cap23, Wer23, JS24]. Explorations [Coo82]. Explorer [Mil81]. Exploring [Rob16a, SM24]. Exponential [Ber73]. Exposed [Cro98]. Expression [Hoc91]. Extension [Pye02b]. External [Eyl64]. Externalism [Sha92, SL86]. Extractions [Don01]. Extreme [Wal02]. Eye [Ash98, Can01, Sen00a]. Ezekiel [Mor10].

F. [Ort05]. F.R.S. [Hal70, Zui07]. Face [Whi67]. Face-Lift [Whi67]. Facets [Bro94]. Facsimile [Hos73, Whi67]. Fact [Buc70, Har83, SH18]. Factory [Jac07]. Facts [BC00, Mye92, Rhe11, Win97]. Failed [Pel01, Vig21]. Failure [Bro81b, Mar98]. Fall [Boa13, Pal05, WB14, Zil14]. falling [Gia15]. Familiar [Kee14, Bri09]. Family [Dal98a, Goo23]. Fantasy [Kei77]. Faraday [McG66, Mer00, Mor92b, Ste95]. Farewell [Ili92a]. farmer [Are19]. farmers [Lee23]. Fashioning [Gou11b]. Fatal [May02]. Fate [Bro81b, Mil08b]. Father [GM89]. Fe [Vig21]. fear [Haj15]. Federal [For74]. Federative [Löw92]. Felix [Lew62]. Female [Fle00, Row98, Str20]. Feminism [Ols90]. Feminist [Dal97c, Ols90]. Ferguson [Alp91]. Ferment [Bro92a, Can82]. Ferrari [Kel76, Kel76]. Ferriar [Pic84]. fertilizer [Lee23]. fertilizers [ND22, Pag16]. Festival [NG22]. Fever [Pic84]. Feyerabendian [Ili92a]. fi [Kin75]. Fibre [Ish06]. Fiction [Har83, McG85]. Fictions [Mye92, Sin22]. Field [Koh02b, Liv87, Rab13, Rab19, WGvA22]. Fields [Bon08, Pic99, Rav62a]. Fifteenth [Tur75]. fight [Flo22, Pag16]. Figurations [Alg23]. Film [Gou11a]. Film-Making [Gou11a]. Fils [Liv05]. Fin [Hoy01, Buk20, Sen92]. Fin-de-Siècle [Hoy01, Buk20]. Financial [Cro07, dP18]. Finders [Koh07]. Finding [Gin00, McE13]. Findings [SL86]. Fine [Hol81]. Fire [GÁ25]. First [Bro67, Pag75b, Sch68, Tom84, Wal02, Wil62, Eri22, Hal15, LS17, Ver14, Far23].

Fischer [Jac17]. Fit [Far95b, Pie86]. Flames [Noa04]. Flat [All17, Not17]. flavor [Ber16]. Fletcher [Noa04]. flickering [Bar20b]. Flies [Mar11]. Flight [Pet10]. Floor [Jac07]. floras [MP14]. Florence [Mae07]. flow [Sch18, dP18]. Flower [Bel09]. Fluids [Pel01]. fluoridation [Zel19]. focus [Vig21]. fois [Whi64a]. Folkes [RH99]. Fontana [Kni93]. Forbes [Fal21]. Force [Ait71]. Forces [Rav62a]. forestry [GÁ25]. Foreword [Ano73a, Ano74b]. Form [Ewe63, Mye92, Pal97]. Forman [Rad83]. Formation [Haw11, Ste02]. Formative [Ost93]. former [Fle18]. forming [Chi23]. Forms [Mag99a, Mag99b]. Fortunes [Mol81]. Forty [Byn75, Doe03]. Forty-Year [Doe03]. Fossils [Por73, Rup83]. Foucault [For80, Sho86]. Foundation [Tur86]. Foundations [Ano77b, Fre71, Mor93]. Founder [Sta63, Tur86]. Founding [Nic88]. Four [Mor09c, PC18]. Fourcroy [Lev84]. Fourier [Car77]. Fourteenth [Eva78]. Fragile [Cap23]. Fragment [Ili97]. Fragmentation [Hen08]. fragmented [Cha25]. Framed [Bud13]. framing [Fra24]. francais [Whi64a]. Francaise [Slo94, Wil67]. France [Bri20, Cro06, Cro07, Dal98b, For84b, Fre95, Hal75, Jac87, Jac95c, Le 16, Pic81, Pic88, Pic99, Rap69, Rob14b, Sim12b, Ste20]. Francesco [Fin93]. Francis [Wal02, Phi15, Rav66b, SS15, Tur78]. Francisco [FTC16]. Franco [Flo22]. Francois [Maa20]. Francois-Marc-Louis [Maa20]. Frank [Cha18]. Frans [Whi64b]. Franz [Drö15]. Fraternal [GG97]. fraud [GH20, Par20, Rob20b, RSM20]. Frédéric [Ste20]. Frederick [Ter90]. Free [Tei80, Sch18]. Freetown [Eve23]. French [Ang12, Bau19, Bri09, Cai65, Haj15, Jac78, Kel66a, Kin14, Lac09, Lev84, Mar82, Tre10, Wei88, Woo91, You67]. Freud [Cap76]. Friedrich [Cal98]. Friends [Ili97, tH04]. friendship [Goo23]. frontiers [Sim17]. FRS [Alp91]. fruits [Web14]. Function [Sch65, Tei90]. Functioning [Kas84]. fundamental [GH20]. fürchtet [Row98]. Furies [McG73]. Furtado [FTC16]. Further [Eyl66, Gue79]. Fust [SS80]. Future [Col89, Tho73, CG21, RRH23a, Tam15].

Gabry [Zui07]. Gain [Zui07]. Galaxy [Whi74]. Galen [Goo04, Nut74, Wil65a]. Galenic [Lan09]. Galenism [Wea75]. Galilean [BZ09, Bos05]. Galileo [Bia90, Clu08, MH23, Mar87, Mas02, May12, Mil08a, Pal05, Sch63b, Sch65, Smi15, Val12, Wes88]. Gallery [Far95b]. Gallstones [Roo08]. Galton [FN71, Goo04, SS15, Wal02]. Galvanic [Str95]. Galway [Kik12]. Game [Bat98a, Bat98b]. games [Lig15]. garden [Lab21]. gardens [Bri20, Hal15]. Gaslight [Tom11]. Gathering [AA87]. Gebrauch [Tur73]. gegenwärtiger [Tur73]. Gems [Byc22]. Gender [Dam13, GC10, Sen92, Ton19]. Gendering [Gou25, Ter95]. genealogy [VT24]. General [Ege77, Lev84, McC85, Ste02]. Générale [Car63, Goo76, Hos66, DV01]. Generalization [vW04]. Generall [Ste02]. generation [Eri22]. generic [BP19]. Genesis [Mil75, Ber07]. Genetic [Dal97b, Pie86]. Geneticists [Pal96]. Genetics [FN71, Gou74, Koh91]. Geneva [Gui17]. Genius [Alb96]. Genombrott [Jon74]. Genomic [Por12].

Genres [AA85, Mea24]. Gentleman [Sar04, Sha91a]. Gentlemen [Nev86, Por82]. Geocentric [Hut12]. Geocentrism [CDG02]. geographies [Goo23]. Geography [Büt79, Del05, Liv84, Liv87, Liv88, Liv91, Wit99]. Geohistory [Buc09]. Geological [Eyl64, Kus16, Rud76, Rup98, Sec86, Gol19]. Geologists [Old98, Por81]. Geology [Dol98, Edd01, Eyl66, Gun78, Ham93, Luc99, O’C07, Old79a, Old79b, Por79, Por81, Rap64, Rud03, Sho94]. Geometry [Bri09, Mol69, Ung77]. Geomorphology [Dav75, Pag71]. Georg [Web14]. George [Far95b, Ash98, Gil20, McG85, Oli24, PV11, Ros24]. Georges [Out76, Out78, Pic88]. Georgian [Kno96, Can78]. German [Mau64, Pag75a, Cla97a, Cla97b, Fig73a, Fig73b, Har98, Luc95, Pal97, Pau94, Pre20, Pye82, Spe62, Tam16, Tur73, WGvA22, Wei86, dB13]. German-language [Fig73b]. Germanies [Woo85]. Germans [Phi15]. Germany [Büt79, For74, Gau79, Har98, Jac04, Mer20]. Gesammelte [Whi62b]. Geschichte [Pag74a, Tur73, Whi64b, Fig73a]. Gestalt [Har98, Mil75]. Ghastly [Gue16]. Ghost [Ash08]. Ghosts [Kan19]. Giacomo [Pal07]. Gianantonio [Gia15]. Gift [BC98]. Gifts [Mok07]. Gilbert [Bon08]. Giovanni [Mel96]. Giulio [Rob05]. Giuseppe [Esp24]. Give [AT12]. gives [Cha18]. Glacial [Rud69]. Glaciation [BdMF01]. Glass [Bel14, Win97]. Global [Kjæ10, Ren23, Ber16, Del19, GGN16, Rob14a, Rob16a]. Goa [BZ09]. God [SHW03]. Goda [May71]. Gods [Sch95]. Goethezeit [Cla97a, Cla97b]. Golden [Sho94, Tei90]. Good [End09, Law95, Pye89]. Goodness [Pie86]. Goodness-of-Fit [Pie86]. Got [Cun91, SLT17]. Gotha [Coo18]. Gothic [WB14]. Gottlob [Eyl64]. governance [GÁ25, GGN16, Le 16, Rob16a]. Governing [McC13]. Government [Rap69, Rei94, Woo13]. Gramsci [NG11]. Granada [MP14, OEMB19]. Grand [And03, Chu03, Mat71, PPT17]. Granting [Ano11a]. Gravitating [Van06]. Gravity [Hen93]. Great [Byn75, Dal97d, Lon80, Fre71, Pag16, Pal94, Sec86, Ter90]. greater [Mal18, Chi23]. Greatest [Dal99, Pag16]. greatness [Van17]. grec [Whi64a]. Greece [Fle00, Tam15, Vla99]. Greek [Fle00, Net02, Nut74, Ras89, Sab87, Ung77, Vla99, Was65]. Greeks [Was64]. Green [MP14, Kra24]. Greene [Ano77b]. Greenwich [Bel22, Tur78]. Gregor [Cal88, Drö15, Gli99, Pie86]. Gregory [Fri03]. Grid [Old86]. Grid/Group [Old86]. Ground [Bal12, HV19]. Group [Old86]. grouping [Por20]. Growth [Ber73, Cla64, Löw92, RH88, Sho83, Stu77]. grumettas [Eve23]. Guerre [For84b]. Guidance [Por20]. Guide [Ege77]. Guidobaldo [Van06]. Gulbenkian [Rap21]. Guys [Pye77].

H [McG90]. Habermasian [Bro98a]. Habsburg [Som21]. Haeckel [Rey08]. Haire [Cro01]. Hale [Oli24]. Half [Sch68, Wil62, Kum23, Ver14]. half-caste [Kum23]. Halley [SZDC13]. hand [Ber14, Byc22]. Handasiya [Kin75]. Handbook [Bec22]. Handlist [Bro81c]. hands [SJ21]. hands-on [SJ21].

Handschriften [Fig73b]. Hanover [Cha08]. Harbor [Pal96]. Hard [Han70, Sha22, Wal02]. hardness [Byc22]. Harriot [Hen82, Pep67, Pep76, Smi15, Tan67, Whi75]. Harrison [Hen09, Par84]. Hartman [Bec22]. Hartmann [Kel10]. Hartsoeker [AN13]. Hartt [BdMF01]. Harvard [Tur70]. Harvey [Ack67, Bat98a, Bat98b, Eri22, Pag63, Pag69, Pag70, Wea83, Wil65a]. Harveyan [Pag63]. Having [CDG02, Fle18]. Hawaii [Shi19]. Heads [Hig85]. Healers [Cul93]. Healing [Wer00]. Health [Dac01, Dol88, Pau94, Pic88, Wei86]. Heart [Nut74]. Heat [Car77, Kni66]. Heathorn [Whi96]. Heaven [Hut12]. Heavenly [Urb23]. Heavens [Alp91, Sti64, Bon08]. Hegemony [NG11]. height [Fle18]. Hélène [Chr87, Gol87, MA09, VB87]. Heliocentricity [Hut12]. Heliocentrism [CDG02]. Helix [Tei75]. Hellenic [Ols78]. Helmont [Deb74]. help [Fyf22]. Henri [Goo76, Goo76, PV11]. Henrietta [Whi96]. Henry [Eng78, Cos02, HH73, Har09, Hen93, Hun88, Mor16, Poy63]. Henslow [Bar79a]. Herbaria [Die24, APT20, Coo18]. herbarium [WGvA22]. Herbert [Elw03, Gay03, Gay08, Jar03, MA09]. Hereditarianism [Wal02]. Heredity [FN71]. Heretic [Wes88]. Hermann [Kat13]. Hermaphroditism [GC10]. Hermeneutist [BZ09]. Hermes [Par84, Vla99]. Hermeticism [Sch78]. Herschel [Cha08, Ash98, Cha08, Hes62, Hos63b, Hos02, Hos03c, Hos05, Hos11, Sti64, Win03]. Herschelian [Hos11]. Herschels [Cha08]. Herstmonceux [Tur78]. Hertz [D’A90]. Herwart [Bon14]. Hiatus [Kol06a, Kol06b]. Hides [Mae07]. hiding [Bar20b]. hierarchy [Ren23]. high [Sim17, Fin00]. Higher [Ang12]. Hill [RH99]. Himalaya [Fle18]. Himself [Ili97]. Hindi [Sin22]. Hippocrates [Fle00, Whi80]. Hippocratic [Lon89]. Hiroshima [Ras97]. Histoire [Car63, You67, Cai65]. Histology [Bra77]. Historian [Deb67, Edg05, Hol75, RP12, Ric96, Rud03, MA09]. Historians [Bel09, Buc62b, Fis82a, Fis82b, Lon78b, Lon78a, NG11, Old86, Old98, Pye93, Rei81]. Historical [AA87, Chu03, Dav78, Hoc90b, Key63, Lon81, Moo92, Old79a, Old79b, Ols90, Pag73, Pag75a, Rag07, Rud80, Sha74a, Sme64b, Tur70, Ung77, Van75, Whi83, Wil67, tH04, Rob14a]. Historically [Hol92]. Historicism [Old79a, Old79b]. Historicizing [CP21, Por05, Rob20b, SR20]. Histories [Lau93, Pic11b, Sch63a]. Historiographical [Car88, Fig76, GPP^{+}08, Mil81]. Historiographies [For85, RRH23a, Fle21]. Historiography [AA85, AA06, Bro75, Bro77b, Buc65, Can78, Can82, CG82, Chr87, Dea12, Goo09, Hak91, Hak95, Hei68, Høy96, Jam85, Jar03, Jor78, McE97, Mic89a, Mic89b, Sch84b, Smi82, Sta13, Sut86, Spe15]. Historisches [Pag73, Pag75a]. History [All76, Alv13, And03, Ank07, Ano09a, Ash14, Bec04, Bra77, Bro67, Bro77a, Bro92a, Bro92b, Bro81c, Bro84, Bro71, Buc62a, Buc62c, Buc70, Buc09, Bud13, Can64, Can76, Car68, Car91, Chi15a, Chi65, CP94, Cro62, Cro08, Dal97b, Dal97d, Dam13, Dan83, Dav75, Deb63, Doe03, Duk63, Edd01, Ege77, Ewe63, Eyl64, Eyl66, Fan00, Feu93, Fis73, For74, For94a, For80, Fox83, Goo09, Gou11a, GG90, GG04, Hal66, Hal73a, Ham93, Han70, Han79, Hei68, Hes63, Hoc90a, Hos63a, Ili92a, Ili98, Jac73, Jan00, Jar00b, Jar04, Joh05, Kni93, Kol06a, Kol06b, Lav70, Law90, Law95, Lev67, Liv84, May71, May02, McC85, McG66, McG73, Mil08c, Mor09a, Mor74, Mor93, Mor09c, Mye92, Nav21, Nee73, Net02, Nic88, Old97, Old98, Pag74a, Pal93, Pic95]. History [Por73, Por80b, Por81, Poy62, Pye77, Pye89, Pye02a, Ram10, Rap64, Rav62b, Rav63a, Ree09, Rey08, Rob66, Rou79, RH00, Rup83, Rup98, RTL00, Sch82a, SV18, Sch95, Sen00a, ST74, Sha82, Sha92, SP87, Sho86, Shr66, Siv01, Sme64a, Smi82, Smi94, Smi12, Som10, Sou63, Spa08, Sta13, SdC11, Sut75, Tal65, Tei80, Tei03, Tha66, Tom91b, Tur69, Tur73, Tur78, Tur89, Wea75, Wei86, Whi84, Whi71, Whi63b, Wil74, Wil00, Wit99, Woo90, Woo85, Wor94, Yeo07, You66, Yox79, vH94, Bru17, Dav19, Del19, Die22b, Dor23, FSB^{+}23, Fel18, dMF23, GH20, Gia15, GS16, GP24, HV19, Hew16, MB23, MSR20, Rob16a, RSM20, Sha22, SH18, SM24, Sys20, Uys24, You67, Cai65, Fig73a, Kus16]. Hitcham [Bar79a]. Hitler [Har98]. Hitting [Bru17]. Hiyal [Kin75]. HMS [Bil22]. Hobbes [Woo88, Sar04]. Hoff [Ham93]. Hofmann [Jac06]. Hohenburg [Bon14]. Holism [Har98]. Holistic [Har98]. Holland [Rup90]. Home [Whi96]. Homeopathy [Dea01]. Homo [For94a]. honesty [HPtH24]. Honor [Yat73]. Honourable [Hal70]. Hooke [AT12, Pug04, Pum91]. Horse [For85]. Horticultural [Por76]. horticulturalists [Lab21]. Hoskin [Ben22]. Hot [Mar98]. Hottentot [Qur04]. House [Hom03, HW86, Hal15]. Howard [Ros24]. Human [FN71, Goo09, Liv87, Chi15a, Ste20]. Humanism [Hal73b, Sha91b]. humanist [Kel66a]. Humaniste [Kel66a]. Humankind [Cro88]. Humboldt [Dau25, Nic87, Wil65b]. Humboldtian [Dau25, Hom95, Nic87]. Humoral [Fox88]. Humorous [Bro92c]. Hunter [Bro94, Jac83, McG90]. Hunting [Ano77b, Cha18]. Hurtado [APT20]. Hustlers [Mor93]. Hutchinsonianism [Wil80a]. Hutchison [Cla93]. Huxley [Eng78, Kaa14, McG90, Whi96, Whi05]. Hwang [Par20]. Hybrid [Fan00]. Hydrogen [Gor15]. Hydrographic [Bar17]. hypnosis [cW18]. Hypnotic [LeB01]. hypnotism [Haj15]. Hypothesis [Bro81b, Bru87, Hew16]. Hysteria [Jac95c, Mic89a, Mic89b]. hysterie [Gil83].

Iatrochemistry [Roo08]. Iberian [PPT17, Por17]. Ibn [Kin75, Has79]. Icon [Kjæ10]. Iconic [Ree09]. iconographie [Gil83]. Idea [Lev84, Sch65]. Idealization [Cla09]. Idealized [Pal05]. Ideals [Boa13]. Ideas [Ack67, Bro92b, Buc62a, Gru81, Nev76, Nic88, Pal96, Par88, Pum91, Smi94]. Identification [Tei80]. Identities [GC10, Str95, Rob16a]. Identity [Bar03, Dol88, Gou11b, Haw11, Jor03, Keb05, Sha91a, Ano21, GM18, GS16, Rap14]. Ideological [Jac95a, Mil00, Ash14]. Ideologies [Jac80]. Ideology [Hoc90a, RH88, Rup96]. Idol [Yeo85]. If [Tam16]. II [Hal91b, Mic89b, Whi64b, WS89b, Har98, Kim92b, Pep67, Ste05]. III [Far95b, WS90]. Illness [Col00, Hew16]. Illustrated [Eas86]. Illustrations [Lev67]. im [Fig73b]. Image [Hes62]. Images [Dal97c, GCM05, Jac83, San17, Sav17]. Imagination [Dea16, Gra11, Sou92].

Imagined [Kni95]. Imagining [Whi17]. Imitations [SHW03]. Immanent [dP18]. Immediate [Wil65a]. Immunology [Löw92]. Impact [Har63a, Has87]. impacts [GGN16]. Imperial [Cul93, Cap23, Mer20, Rap21, Sim17, Bel14, Jac95b, Nav06]. Imperialism [Pye82, pY15]. Impersonation [Bar20a]. Implements [Sch10b]. Implications [Liv84]. Importance [Kin95, Mil00, Mau64]. Imprimatur [Don01]. Imprimeur [Kel66a]. Impurity [Por91]. in-between [Ano20, Sur19]. Incipient [AA87]. Incommensurability [Ung77]. Incommensurable [Ung77]. Increasing [Loh67]. Incursions [Nav06]. Independence [Sch68]. India [All76, Cha10, Dam13, Gur21, Sim17, Sin22, Tet23]. Indian [All76, SR20]. Indicator [Mil12]. Indies [LS17]. Indigenous [Spa08]. Individuals [Rud81]. IndoGermanism [GP18]. Induction [Alb96]. Industrial [Ash08, Jac07, Mat71, Pic88, Tha70, Tom11, Ash14, Le 16, Mok07]. Industries [Luc95]. Industry [Alb96, Ili97, Jor78, Rus74, SJ21, vW09, CG21, Mor18]. infant [LJG^{+}24]. Inference [Sti92]. Inferential [Mar11]. infinite [Spe15]. Influence [FN71, Was65, Mor18]. Influenced [Ang12]. Information [Sch09a]. Infra [Rhe11]. Infra-Experimentality [Rhe11]. infrastructure [Col24, Wer23]. infrastructures [Bra23, Cap23, Tet23]. Ingenious [Kin75]. Inheritance [Pal94]. Initiative [Wil02]. Innovation [Gin00]. Inquiry [Gra11, Liv87]. Inquisition [LN17]. Insane [Gil83]. Insanity [Nev76]. Inscribing [Dea03]. inscription [BB21]. Inscriptions [Kei77, Fle18]. insect [JS24]. Insider [Høy96]. Instauratio [Ste05]. Institute [Mor93, Vig21]. Instituting [Gou11a]. Institution [Aic09, Tur78]. Institutional [Des85b, Des85a, Ebr00]. institutionalization [Vig21]. Institutions [Pal96]. Instruction [Ang12]. Instrument [Kas84]. instrumental [Kum23]. Instrumentation [Cha06]. Instruments [Bia06, Bry77, Dav78, Fle18, Kel10, Mad63, Sch06, Tur69, Tur70, Val12, Str20]. Intangible [She19]. Integrating [MSR20, RSM20]. integrity [HPtH24, MSR20, Rob20b, RSM20, Sib20]. Intellect [Pye93]. Intellectual [Ebr00, Goo04, McG66, Mil08b, Sha91b, Tur75, Bia19]. Intellectuals [Dal97g, NG11]. Intelligence [Cha25]. Intelligencer [Gro13]. intelligent [Por20]. Intensities [Mol69]. Interactions [Lev70, Liv84, Rud81]. Interest [Mor09a]. Interim [May12]. Internalism [Sha92, SL86]. International [Cur12, May02, Por81, Slo94, Chi23, Mou21, GR22]. Internationales [Bee75]. Interpret [Gra87]. Interpretation [Cla09, Gol87, Ols78, Ung77, Nav21, Ros24]. Interpreting [Stu77]. Intersection [Esh01]. interwar [Tet23, Tuc20, Cro01]. Intimate [Ben88]. Introduction [Ait73, BWM22, FNG22, GRW22, GT15, Pic11a, Por05, Rob19, Rud81, RC13, SV18, Tal65, Rob14a, RRH23b, Sys20, Whi64a, Jor78]. Introductions [Eva78]. introductory [Rob16b]. Intuitive [CDG02]. Inventing [Spa08]. Invention [Bun19, Gil83, Mil08b, Van75, Leo24]. Inventioni [Kel63]. inventions [Bau19, Kan19]. Inventor [Keb05].

Inventory [Pye95, Coo18]. Inverse [GCM05, GCM06]. investigating [Fle21]. Investment [tH04]. invisibility [Fag21]. Invisible [RC13, ABC^{+}23]. Invoking [Zel19]. involved [De 17]. Inward [Buc62a]. Iowa [Mar11]. Iran [Ebr02]. Iraq [Col24]. Ireland [Hom85, Tur91]. Irish [Gol16]. iron [Gou25]. Irrationality [Ung77]. Irréductions [For84b]. Isaac [You70, Ait73, Ait75, Boy67, Hal68, HT73, Hei79, Whi62a, Whi63a, You75]. Isis [Whi73]. Islam [Dol88, Sab87]. Islamic [Cha06, Dol88, Gin00, Kin79, Lan09, Rag07]. Islamicate [FSB^{+}23]. Islanders [Siv01]. islands [FTC16]. Isle [Bri20, Rob14b]. issue [Rob20a, RRH23b]. Issues [Pie86]. Istanbul [Ile24]. Istoriya [May71]. Italian [Bia89, Bri09, Doo95, Kel76, Spe62]. Italy [APT20, Bos05, Fag21, Mel96, Rap91, Tar14]. Iterative [Die22a]. IX [Whi62b]. Izmir [Sav17].

J [Mar13, Nav06]. J. [Fal21, Nav06]. Jamaica [Ogb13]. James [Coc63, Alp91, Mil00, Mil04]. Jamie [Mil00]. Janaki [Dam13]. Janet [LeB01]. Japan [Mea24, Wer23, cW18]. jar [Pre22, SH18]. Jardin [Slo94]. Jazari [Kin75]. Jerusalem [Mil08c]. Jesuit [Mal18]. Jew [RH00]. Jewish [Fre95]. Jinpingmei [Cul93]. Jobs [Hig85]. Johannes [Ait76]. John [Don01, Ash98, Bel14, Edd01, Har09, Jac83, Mil08c, Mor18, Mus00, Por92, Ram08, RH99]. Joint [Ano09a]. Jonas [Rus94, Rus94]. Jonston [Mil08c]. Jordan [Joh05]. José [OEMB19]. Joseph [Mil81, Car77, Fox88, Mou21, Rav63b]. Journal [Lew62, Kus16]. Journals [Ano09a, Fyf22]. Jubilee [Tei90]. Judge [Mor92a]. July [Ste20]. June [Pal96]. Jüngere [Whi64b]. Justify [Por01]. Justus [DV01].

K. [Ham93]. Kant [PO02, Yol71]. Karl [Joh05, Mag99a, Mag99b]. Katalogisierung [Fig73b]. Kay [Pic84]. Kearton [Gou11b]. Keepers [Koh07]. Keeping [Keb05]. Kelvin [Por79]. Kepler [Bee75, Jar00a, Ait76, Alg23, App96, Bee75, Bon08, Hel64, Whi62b]. Kepler-Symposium [Bee75]. Kepler/Kepler [Jar00a]. Keplerian [App96]. Kettle [Mil04]. Kezhen [Fra24]. Khwarizmi [Par88]. King [Far95b]. Kingdom [Coo82, Wad01]. Kinsey [Cro98]. Kirwan [Gol16]. Kitab [Kin75]. Kitchen [Gue16]. Knack [CDG02]. Knight [Som10]. Know [Bel09]. Knowing [Cha11, Löw11, Pic11a, Sim17, Del19, Shi19]. Knowledge [Ber07, Bru07, Can82, Car79, Cro01, HM10, Hoc90a, Hol92, Jac73, Kin75, Kno99, McE13, MJ01, Mok07, Mor94, Pye02b, Ren23, RH88, Spa08, Ste07, BD23, Bri20, Can17, Fel18, Gil24, Goo23, GP24, Gur21, MB23, Sav17, Whi17]. Knowledges [Pic95]. Known [Mol69]. Konstantin [SP07]. Korea [Hyu21, Lee19, PC18]. Korean [Lee23]. koro [Chi15b]. Koyré [Jar00a]. Kramers [Rad83]. Krusenstern [Wer23]. Kuhn [Hen83, Wil80b]. Künlün [Fin00].

Lab [Koh02a, Pel01]. Labelled [Löw11]. Labor [GR22, ABC^{+}23, Bel22, Dor23, Lee23, Ren23, RRH23a, Sam23, Yil23]. Laboratories [Sch07, SM24]. Laboratory [Cha10, Mag99a, Mag99b, Old97, Ton19, Ame24, Shi19, Mac21]. labors [Alg23]. Labour [For94a]. Labscapes [Koh02a]. Laby [Dea03]. Lamarck [Pic88]. Lancashire [Sec94]. land [Rab19]. Landforms [Dav75]. Landmarks [Bro84]. Lands [dB13]. Language [Alb96, Bar64, Bar03, Kim92a, Kim92b, Kus16, Nor66, Out78, Rud76, Sme64b, Sur19, Suz95, Ano20, Fig73b, Gor15, Lig15, Tam15]. languages [GT15, Gur21]. Last [Ram08, Sme63]. Late [Mat63, Noa05, Sio88, Ame24, Col24, ES15, Fag21, Gor15, Haj15, Ile24, Lee19, SS15, Smi20, Sto18, Bro13, Csi10, Cul93, Fig75, Gra11, Kno96, Mae07, MJ01, Mil12, Nav06, Pet10, SP87]. Late-Victorian [Noa05, Smi20, Nav06]. Later [Was65]. Latest [Cap76]. Latin [Wal22, GRW22, GR22, Lav22, Roq22]. Launching [Mil08c]. Laura [Row98]. Lausanne [Gui17]. Lavoisier [Lev84, Sme63]. Law [FN71, GCM05, GCM06, Gua91, Rud62, BB19, Kan19]. law/law [BB19]. Laws [Gli99]. Layers [Kim92a, Kim92b]. layman [Are19]. League [Cro01, Mil08b]. Learned [Zui07, Fyf22, Smi14]. Learning [Bal89, Sto18, Whi05, Ste02]. Leavis [Ort05]. Lectures [Edd01, Hua16, Pag63, Wil65a]. Leeds [Jac07]. Legacy [DV01, Gol03, HP07]. legal [Ara19]. Legend [SLT17]. Legitimation [AA85, AA87]. legitimization [Vig21]. Leibniz [Sha81, Smi12]. Leiden [Pre22, SH18]. lemons [Cha18]. lens [Rob16a]. Lenses [Lev67, MH23]. Leonardo [Pug04]. Leone [Eve23]. Leonhard [WGvA22]. Leopold [Buk20]. Lettere [Spe62]. Letters [May12, Nev86, dB13, Bon14, Spe62]. Leviathan [Del05, Woo88]. Lexical [BP19]. Liars [Goo08]. Libavius [Kel10]. Liberal [Whi05]. Liberalism [Pic81]. library [Kel66a, Key63]. Libri [Pag75b]. Libro [Wat74]. Lichtenberg [Dou88]. Licking [Sha81]. Liebigian [Jac06]. Liebigiana [Bro81a]. Life [Alp91, Bon08, Bos05, Car88, Cro98, Dav75, Hal70, Has79, Hol75, Ish06, Kni66, Nic88, Pug04, Ree06, Rup90, Sha91b, Woo88, Yox79, Cha18, Eri22, Mun21]. Lift [Whi67]. Light [Ait76, Hom85, Jam83, Joh96, Pug04, Sch77, Sme63, Wes67, Smi15]. Like [Mor09c]. Lilio [Not17]. Limits [Buc09, Cla09, Sho83]. lineage [Gou25]. Linguistics [GP18]. Linkage [Pie86]. Links [RTL00]. Linnaeus [Bro84, Cai65]. Linné [Cai65]. Lipsius [DV01]. Liquid [Bon08]. Lisbon [Rap21]. Lisianskii [Wer23]. Lister [Fox88, Roo08]. Lists [Coo18]. Literary [Rab13, Sha74a, Mea24]. Literature [Bar64, Dal97c, Hoc91, O’M65, Old84, Por76, Rab13, Tal65, Tom91a]. Literatures [Del05]. livebirth [LJG^{+}24]. Lives [Ben88, Dal98a, Mor92b]. Living [Bia06, Dac01, Dal97h, Kel10]. Local [GGN16, Mus00, Pye02b, Spa08, Bar22]. Locales [Cha06]. Localism [Kuu12]. Localizationists [Sta86]. Locke [Hes62, Tom84]. Lodestones [Roo08]. Lodovico [Kel76, Kel76]. L’oeuvre [You70]. Logic [Nor66]. London [Cro01, Ben06, Des85b, Des85a, Gay03, Gay08, Hal15, Jac06, Kik12, Kno99, Mag99a, Mag99b, Pel01, Pug04, Rob66, Slo90]. long [Van17, Rob21]. Longer [Nic88]. Longevity [Gra78]. Looking [Bel14, Ste95, Win97]. Loose [Löw92]. Lord [Coc63, Por79]. Loss [Wil02]. Lost [Ili97]. Lösungen [Mau64]. Louis [BdMF01, Maa20]. love [She15, Smi14]. Lovejoy [Sch10a]. low [Ren23]. low-end [Ren23]. LSD [Hew16]. Lu [Fen15]. Lunacy [Nev76]. Lunatick [Kno99]. lungs [Sam23]. Lusitanians [Roq22]. Lutheran [Büt79]. Lyell [Ano74a, Bar79b, Dol98, Rud63]. Lyon [Ano77b].

M [Mar13]. maarifat [Kin75]. Mach [Lau68, Sta21]. Machina [For94a]. Machine [Smi14]. Machinery [Smi94]. Machines [Gau06, Has79, Jac04, Kel10, Whi66a, She19]. Made [Dal97a, Wei88]. Madness [Dal94, Nev76, Pet10]. Madras [Kum23]. Madrid [NG22]. Magazines [Bro92c]. Magic [Cur85, Hen93, Hen08, Liv88, Drö15]. Magna [Ste05]. Magnet [Hom77]. Magnetic [Roo08]. Magnetism [Far95a]. Magnifying [Eri22]. Magnitudes [Ung77]. Main [Bel22]. maintenance [Wer23]. Major [Ano62d]. Make [McG85]. Make-Believe [McG85]. Maker [Kel62a]. Making [Ber16, Boy67, Bro62, Bun19, Cur85, Dal97h, Dal98b, Dea03, Des85b, Des85a, Gou11a, Hop97, Joh96, Jon16, Mar13, Mor09a, Par20, Por92, Sen00a, Shi19, Sim12a, Sim12b, SHW03, Som10, Bar22, Van17, FTC16, Lab21, Por20, ABC^{+}23]. Malthus [Por87a]. Mammography [DT10]. Man [Bro84, Car77, Don14, Kas84, Kel62a, Mor80a, Mor99, Por01, Pug04, Sho83, Woo90, Gol19]. Management [Har96a, SC98, Bel22, GÁ25, Rob14a]. managing [BD23, Rob14b]. Mandarin [Wei88]. Manhattan [Sla24]. Mankind [Old98, Leo24, Smi20]. Manners [dB13]. Manuscript [Hos73]. Manuscripts [Pep67, Qui62, Tan67, Die24, Fig73b, Pre22]. Many [Ano11a, Bro94]. Maoist [Gao15]. Map [Pic99, Kin14, Sim17]. Maps [Gin00]. Marc [Maa20]. March [Hos03b]. Marconi [Ara19]. marginalization [Ver14]. Maria [Fag21]. Marin [Gro13]. maritime [Sam23]. mark [LP23]. Market [Dal97h, Yeo85]. Market-Place [Yeo85]. Marketing [Cos02, Smi20]. Marketplace [Mar00]. Markov [Lin06]. marriage [NG22]. Martin [Kus16, Roo08, RH99]. Marvels [For94a]. Masculine [Gol02]. Master [And03]. Mastering [Mar24]. Maszyny [Pet23]. Matematica [Kel76]. Matematiki [May71]. Matematyczne [Pet23]. material [Ara19, Urb23]. Materialism [Can85, Gau79]. Matérialisme [Goo76]. Materials [Rob16a, Bul18, dMF23]. Mathematica [Ait73, Sch09a]. Mathematical [Ait75, Boy67, GCM05, Gro13, Ili98, Kel76, Mad63, You70, You75]. Mathematician [Boy67, Whi62b]. Mathematicians [Bia89, Gil20]. Mathematics [Ait75, Cos02, GCM06, Gin01, GG90, GG04, Har63c, Høy90, Høy96, Kel63, Kin79, May71, Nea99, Phi06, Ras89, Rus94, Bar20a, You67]. mathématique [You70]. mathématiques [You67]. Mathematische [Whi62b]. Mathematization [Lin06]. Matter [Can85, Gau79, Hen86, Hes64, Kim92a, Lau66, Sch82a, Oak18]. Matters [Rup90, Sim12b, Par20]. Maupertuis [Ili93]. Max [Mil75]. May [Tay94].

May-Dew [Tay94]. mayor [Bar22]. Mayr [Joh05]. McLeod [Gay03, Gay08]. Me [AT12, Wes88]. Mead [Jor03]. Meaning [Hes63, Por73, Pye02b]. Meanings [Jac95a, Sen00b, Ber16]. meant [Phi15]. Measure [Mor99, Smi20]. Measured [Pug04]. Measurement [Hes63, Ili93, Mar11, Por20, Sys20, Sib20]. Measures [Edg79]. Measuring [Dal97f, Mor18, Str20, Tuc20]. Mécaniste [Goo76]. Mecca [Gin00]. Mechanic [Ben86]. Mechanica [Tur94]. Mechanical [BZ11, Ben86, Has79, Hut83, Jac07, Kin75, Lau66]. Mechanics [Gil20, Tur94, Mor18, Smi15]. Mechanisation [Buc62b]. médecine [Jor78]. Mediaeval [Ste65]. mediation [Kin14]. mediations [Kan19]. Medical [Cla64, Cro01, Cur12, Dac01, Dal97h, Dav78, Fig76, HM10, Jac87, Jor03, Key63, Lew62, Lon81, Löw11, RS62, Wea75, Poy64]. Medicalization [Dik91]. Medici [Pag75b, Val12]. Medicine [Ano09a, Big69, Biv99, Byn73, Dal95, Dal97h, Deb74, Dol88, Ebr00, Ebr02, Fra73b, Fra73a, Har96a, Ish06, Law95, Liv87, Llo68, Lon89, Lus68, Nut74, Pic81, Pic88, Pic94, Por92, Poy62, Poy63, Shr66, Sim12a, Slo90, Smi88, SC98, Wei86, WM22, Yat73, LP23, Dal99, Wei88]. Medieval [Ano76f, Cro75, Eas86, Gra78, Gra87, Hut12, Kin75, Lus68, Mol69, Rya66, Sab87, Sou63, Tal65, Fre95]. Medina [Wat74]. Mediterranean [BZ11]. Medizin [Pag74a]. meets [Hua16]. Meiji [Mea24]. Memory [Chu03, Gin08, Yeo07, Gou25]. Men [Cro07, Har63b, Jen81, MF73, Woo13, Bar03]. Mendel [Cal88, Drö15, Gli99, Kal83, Olb79, Pie86, RB83]. Mendelian [Olb79, Pal94]. Mendes [RH00]. Mendoza [APT20]. Mental [Kas84, Smi94, Hew16]. Mentalities [Cur85]. Mention [Jar04]. Mercator [Pep76]. mercy [Uys24]. Meric [Ste02]. Mersenne [Gro13]. Merton [Hal63]. mesh [Die24]. Mesmerism [Win94]. Mesopotamia [Big69]. Mesopotamian [Høy96]. Messenger [BZ09]. metallurgists [Gou25]. Metaphor [Fig76, Rey08, War65, MH23]. Metaphors [Osl98, dP18]. Metaphysical [GCM05]. Metaphysics [Yol71]. Meteorology [And99]. Method [Can76, Fin93, Lau66, Lau68, Rus74, Wal02]. Methodist [Don01]. Methodological [Col81, De 17, Pre20]. Methodologies [Fle21]. Methodology [Buc70, Mar87, RB83]. Methods [Has79]. metropolitan [Ano21, Van17, GM18]. metropolitanism [Gol16]. Metzger [Chr87, Gol87, MA09, VB87]. Mexico [Cor14]. Meyer [Cha18]. Michael [Ben22, McG66, Mor92b]. Michel [For80, Sho86]. Michelson [Wil74]. Microbes [For84b]. Microbiology [Ver90]. Microcosm [Kel10]. Microscope [AN13, Tur73]. Microscopes [Sch09b]. Microscopy [RC13]. Mid [Ang12, Bar03, Ras97, RH99, Sou92, Bar20a, Coo18, Por20]. Mid-Century [Ras97]. Mid-Eighteenth [RH99]. Mid-Nineteenth [Ang12]. Mid-Seventeenth-Century [Sou92, Coo18]. mid-twentieth [Bar20a, Por20]. Mid-Victorian [Bar03]. Middle [Con95, Gra11, Sio88, Was65, Eas86, Hes62, Mol81, Sch63b, Sch65, WGvA22, Was65, You67]. Might [Pic99]. migration [Ren23]. migrations [GR22]. Mikroskop [Tur73]. Military [Hoc90a]. Milk [Ver90]. Milky [Whi74]. Millenarian [Mil08c]. Millenarianism [Dal97d]. Miller [Wil74]. Millikan [Mor93]. Mind [Cro88, Dal97f, Fig75, Har98, Smi88, Tom84]. Mind-Body [Tom84]. minds [Hol25, Feu93]. Mineral [Pag74a]. Mineralnamen [Pag74a]. Mineralogical [Eyl64]. Mineralogy [Edd01, Byc22]. Minerals [Eyl64]. Mining [Gil24, Fel18]. Ministers [Whi05]. ministry [Kin14]. Minnesota [Mar11]. Minutes [Hom03]. Miracle [Rud62]. misconduct [Buk20, SR20]. Mismeasure [Sho83]. Missing [Hak91, RTL00]. Mission [Fin00]. Mistrust [Mar00]. Misuse [Kei77]. Mixing [Osl98]. Model [Har63c, Kas82, Mae07, McC13, Sch95]. Modeling [Nic88]. Models [Kas84, Ksi15]. Modern [Bia06, Bro62, Cha11, Col81, Cro75, Cur85, Dal95, Dan83, Dik91, Fra73b, Fra73a, Fre71, Gru64, Hal91a, Hal91b, Har96a, Har02, Har09, Hei81, Hen09, Hoc90a, Høy90, Joh98, Keb05, Las73, Law90, Llo68, Lon81, Mad63, Ols90, Ort05, Osl98, Sen00b, Sha91a, Sha91b, Smi81, Smi82, Spa08, Tha66, Tur69, CV17, Chi15a, Die24, Fra24, Hel20, Mal18, Mar24, Mea24, Oak18, PPT17, Sav17, Sch18, Spe15, Kas11]. Modernist [PV11]. Modernity [Del05, For85, Sch07, Lee23, Wer23]. Modernization [Wei86]. Modes [Hoc91]. Modification [Cal88]. Molecular [AA85, AA87, AA06, SdC11, Yox79]. Molecules [Hol75, Kim92b]. Moment [Edg05]. Monarchy [Mor09a, Ste20]. monde [Ili93]. Monitoring [Maa20]. Montpellier [Lew62, Liv05]. monumental [Car17]. Moore [Rus94]. Moral [Jac87, Koh91, Maa20]. Morals [Nev76]. Morgan [Ric96]. Morley [Wil74]. Morley-Miller [Wil74]. Morphology [Pal97, Slo90]. Morris [Dav75]. Mort [Slo94]. mortality [LJG^{+}24]. mortification [Cap23]. Most [Dal94, Hel20, Shi19]. Mother [Pag81]. Motion [AT12, Dan83, Mau64, Smi15]. Motions [Mol69]. Motivations [Bel14]. motuum [Mol69]. Mount [Oli24]. mountaintop [Ame24]. Mouth [Lin06]. Move [AT12]. Movement [Ish06, Kni70]. moyen [You67]. Much [End09, LS17]. mule [Oli24]. Multi [HT73, HH73, Whi73]. Multi-Volume [HT73, HH73, Whi73]. Multiple [Dob77]. municipalities [Can17]. Munitions [Far23]. Muscular [Dal97b]. Musculis [Wil65a]. Museological [Pic94]. Museum [Bec04, SJ21, CG21, Far95b, Tur86]. Museums [For94b, SP87]. Music [Kas82, Sza90]. Musical [Jac04, Kas84]. Muslim [Kin79]. Musschenbroek [Pre22]. Musurgia [Kno79]. Mutes [Aic09]. Mutilations [May02]. Mutis [MP14, OEMB19]. Mystical [Hut12]. Mysticism [Dob77]. Myth [Jam85, Jam86, Sut86, War65, Rap21]. mythology [GP18]. Myths [Mil04].

Nafia [Col24]. Nägeli [Drö15]. Nahua [De 17]. Naissance [Poy64]. Name [Cun91, Cur12]. Naples [BZ09]. Napoleon [Fag21]. Napoleonic [Cro06]. Narrative [Chr87, LS17, McG85, Ree09, Pre20]. Narratives [Jac95c, Mil08b, Bar22, Cap23, PPT17]. Nasser [Car17]. Nation [Wer00]. National [Fra24, Kik12, Pau94, Pos15, Pye02b, Shi19, Sim12a, Smi12, Tei03, Bar22, Chi23, Mac21]. nationalism [pY15]. Nations [Mil08b]. Natural [AN13, Bec04, Bro81c, Clu08, Cun91, Edd01, Fan00, Gol90, Gou11a, Gra87, Gra11, Har62, Har63c, Hei73, Hei81, Hes63, Jon74, Kno96, Ksi15, Mil08c, Mor93, Mor94, Old97, Osl98, Out78, Par88, Pic11b, Pye90, Ram10, Rav66a, RH00, Rud62, Rud80, Rus74, Sch83, Sha74a, SP87, Siv01, Smi73, Smi78, Smi12, Spa08, Sta63, Wil80a, Win97, WS89a, WS89b, WS90, Woo90, Yeo07, Zui07, tH04, LS17, Shi19, Zel19]. Naturalis [Ait73]. Naturalist [Bro77a, Ili97, Dea16]. Naturalists [Bro77a, Hal91a, Hal91b, Ano21, GM18, Lig15]. Naturalization [Sab87]. Naturalizing [Koh02a]. Nature [Bon08, Buc62a, Cro88, Fis73, Jam83, Jan00, Nic88, Pag16, Pag81, Rup83, San79, Sch10b, tH04, vdR04, Yil23, Zel19, AN13, Bec04, Coo82, Nic88, Bal12]. Naturphilosophie [Can97]. naval [Cap23]. Navigation [McE13]. Navigator [Wat74]. Naville [Maa20]. Nazism [Pau94]. Near [Sil99]. Neat [vdR04]. Nebulae [Hos05, Sch77]. Nebular [Bru87]. Needham [Mou21]. Needs [Sch63a]. Neglect [Lon81, May05]. Neglect-of-Science [May05]. neglected [Pre22]. Nelson [Coc63]. Neoliberalism [Ash14]. Neolithic [For83]. Neoterics [Mel96]. Nephew [LeB01]. Neptune [Kol06a, Kol06b]. Nérée [dMF23]. Netherlanders [Kel62b]. Netherlands [Lev70, vH03]. Networks [Gro13, LC98, ES15, Sch18]. neu [Pag73]. Neugebauer [Pye95, RP12]. Neuroscience [Smi88]. Neuroscientific [Smi88]. newborns [LJG^{+}24]. News [Bro13]. newspaper [Zil14]. Newton [You70, Ait71, Ait73, Ait75, Ano77b, BC98, Boy67, Cas84, Cur85, DV01, Duc05, Fig79, For85, Gas85, Gui17, Hal68, HT73, Hal75, Hal02, Hei79, Hen93, Hes62, Hom85, Jac78, Jac95a, Loh67, McG73, Mor10, OEMB19, Sch84a, Sch09a, Sec85, Sha81, Ste96, Whi62a, Whi63a, Whi66b, Whi67, You70, You75]. Newtonian [Far02, Gue79, Hei73, Hen86, Her62a, Hom85, Sch82b, Sza90, Ver96, Whi62c, Whi66b]. Newtoniana [Fig79]. Newtonianism [Fri03, Gui17, Hom77, Jac74]. Newtonians [Hei78]. NHS [Biv20]. Niccoló [Kel76, Kel63, Kel76, Cuo97]. Nicholas [Hos73, Sti92]. Nicolas [AN13]. Nicole [Mol69]. niemandem [Row98]. Nietzsche [Cal98]. Night [Whi71, Har83]. Nineteenth [Csi10, Fig76, Gil24, Hua16, Mil12, O’C07, Sec94, SP87, Vla99, Ame24, Ano20, Eve23, Fag21, Gor15, JS24, Rie15, Rob21, Sur19, Tam15, Ang12, Byn73, Car63, Cor14, Cur12, Dal98b, Dol84, Ebr00, Fan00, For94b, Gau79, GG97, HSS10, Jac04, Jac83, Jac95c, Kee14, Lac09, McG85, Men64, Nev76, Out78, Pic94, Por80a, Por91, Sch68, Sim12b, Smi75, Smi78, Smi88, Web74, Wil62, Wil74, WS89a, WS89b, WS90, Yeo85, dB13]. Nineteenth-Century [Hua16, Mil12, O’C07, Sec94, Vla99, Gil24, Eve23, Rie15, Cor14, Cur12, Dal98b, Ebr00, For94b, GG97, Jac04, Jac95c, Kee14, Lac09, McG85, Pic94, Sim12b, Smi88, dB13]. Nitrogenous [Pag16]. Nitrous [Boa13]. No [Nic88, Olb79, Fle18]. nobleman [Car16]. Noblewomen [Lab21]. Nodes [BD23]. Non [CDG02, GCM06, Mag99a, Mag99b, Ile24]. Non-Correlation [Mag99a, Mag99b]. Non-Intuitive [CDG02]. Non-Use [GCM06, Ile24]. nonhuman [VT24]. Normal [Hop05]. Norman [Cro01]. Norms [Car79]. north [Sin22, pY15]. northern [Fag21, dP18].

Northumberland [Ash01]. Norway [Ank07]. Not-So-Recent [Dea12, Tar14]. Note [Ano62c, Rob22]. Notebook [Sch84a]. Notes [Ano73b, Ano73c, Ano73d, Ano73e, Ano74c, Ano74d, Ano74e, Ano74f, Ano75a, Ano75b, Ano75c, Ano75d, Ano76a, Ano76b, Ano76c, Ano77c, Ano77d, Ano77e, Ano77f, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80a, Ano80b, Ano80c, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83a, Ano83b, Ano83c, Ano83d, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano85b, Ano85c, Ano85d, Ano85e, Ano86a, Ano86b, Ano86c, Ano86d, Ano87a, Ano87b, Ano87c, Ano87d, Ano88a, Ano88b, Ano88c, Ano88d, Ano89b, Ano89c, Ano89d, Ano89e, Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano91d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93a, Ano93b, Ano93c, Ano93d, Ano94a, Ano94b, Ano94c, Ano94d, Ano95a]. Notes [Ano95b, Ano95c, Ano95d, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano99a, Ano99b, Ano99c, Ano99d, Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d, Ano09b, Ano09c, Ano09d, Ano09e, Ano10a, Ano10b, Ano10c, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano13a, Ano13b, Ano13c, Ano13d, Ano14, Cuo97, Whi64a]. nothingness [cW18]. Notices [Ano62e, Ano63c, Ano64c, Ano66c, Ano67c, Ano68b, Ano69b, Ano70b, Ano71, Ano73f, Ano73g, Ano73h, Ano74g, Ano74h, Ano74i, Ano75e, Ano75f, Ano75g, Ano75h, Ano76d, Ano76e, Ano77g, Ano77h, Ano77i, Ano78e, Ano79e, Ano79f, Ano79g, Ano80d, Ano80e, Ano80f, Ano80g, Ano81e, Ano81f, Ano81g, Ano82e, Ano82f, Ano82g, Ano83e, Ano83f, Ano83g, Ano83h, Ano84g, Ano84h, Ano84i, Ano85f, Ano85g, Ano85h, Ano86e, Ano86f, Ano86g, Ano86h, Ano87e, Ano87f, Ano88e, Ano89f, Ano89g, Ano91e, Ano91f, Ano91g, Ano92e, Ano92f, Ano95e, Ano95f, Ano98e, Ano98f, Ano99e, Ano99f, Ano99g, Ano99h, Ano00e, Ano00f, Ano01e, Ano01f, Ano01g, Ano02e, Ano02f, Ano02g, Ano03e, Ano03f, Ano03g, Ano04e, Ano04f, Ano04g, Ano05e, Ano05f, Ano06e, Ano07e]. Notices [Ano07f, Ano08e, Ano08f, Ano08g, Ano08h, Ano09f, Ano09g, Ano09h, Ano10d, Ano10e, Ano11e, Ano11f, Ano11g]. Nova [Cuo97]. Novelty [Dol98]. Novo [Rap21]. Novum [Ste05]. Nuclear [Aga13]. numbers [Bar17, Drö15]. Numeracy [Net02]. Numismatics [Sch10b]. Nuremberg [Hos03a]. Nursery [Sec85].

obesity [Biv20]. Objectivity [Buc62a, HPtH24, Har98]. Objects [BZ11, Bry77, For94b, HSS10, Sch09b, Str95, Van17, ES15]. Observation [Hos63a, Rab13]. Observations [AT12, Høy90]. observatories [Glu24]. Observatory [Bel22, Tur78, Tur78, Tur91]. Obstacle [Sch63b]. Occasion [Nor66]. Occult [Cur85, Hen86, Nea99, Tur75]. Occultism [Hen08]. Occupation [Hig85]. Ocean [SHW03]. Ockham [Col77]. Odyssey [Kik12].

OEconomy [AN13]. oeuvres [Whi64a]. Office [Bar17, Jub85]. Official [SS80]. oilfield [Ren23]. Oklahoma [Hal73a]. Olber [Whi71]. Old [Bro81a, Cro80]. Oldenburg [HH73, Hun88]. Oldest [Tur78]. Omnia [Gui17]. Omniologists [Por82]. once [Tam15]. One [Tur86, Bru17, Sen00a]. Online [vW09]. Only [Cur12]. Open [LC98]. Opera [Byn88, Gui17]. Operating [Sch07]. Opponent [Cal88]. Opportunities [Sch63a]. Opportunity [Gay03]. Optical [Ste91, Tur69]. Optics [Can78, Hom85, Ras89, Ste91, Wes67]. Optimizing [Zel19]. Oral [Doe03]. ordained [Oak18]. Ordeal [Out83]. Order [Cor14, Csi10, Dal98a, For80, Rud80, Sch09a, Tre10, Bel22]. Ordering [Chi15a]. Orders [AA85]. Oresme [Mol69]. Organ [Bal12]. Organic [Kim92b]. Organicism [McG85]. Organism [Elw03]. Organisms [Pic99, Tei80]. Organization [Dal95, Fig76, Mae07, GR22]. Organized [Bro98b]. Organizing [Cro01]. Organon [Gau06]. Organum [Ste05]. Origin [Ram08, Sza90]. origines [Cai65]. Origins [Chr74, Cla89, Dan83, Eme88, Fri03, Goo09, Har09, Hen09, Jac75, Kal83, LeB01, Nic87, Por81, Ree06, Smi88, Smi94, Was64, Was65, Web67, Yol71, Cai65, GH20, Mac21, She15]. Ornithologists [Bro98b]. Ornithology [Bro98b]. Orthopaedics [Dal95]. Osborn [Som10]. Osiris [Row99]. Other [Don14, Wil85]. Otto [RP12, Sen92]. Ottoman [Col24, Ile24, Sch24, Yil23]. Ottomans [Mar24]. Our [Chu03, Whi74]. Outlines [Ili92a]. Ovary [Sen00b]. Overseas [Pye82]. Overview [GG04]. Own [Sen00a, SHW03]. Oxford [Cro62, Kaa14, Tur86]. Oxygenovich [Gor15].

P [Bra88, Edg05]. Pacific [Chi15b, Old97]. pacifism [Som21]. Padua [Sto18]. Pagel [Yat73]. Pain [Uys24]. Paix [For84b]. Palace [Bec04]. palaeoanthropology [pY15]. Palaeontology [Por73]. paleontology [Rie15, Tam16]. Palter [Ber94]. Paper [Bud13, HM10, Coo18, MP14]. Paperbooks [Yeo07]. papermaking [Lee19]. Papers [Ait75, Boy67, Bra88, Coc63, Hal02, Pep67, Whi63a, You70, You75]. Paracelsian [Pag74b]. Paracelsians [Pag66]. Paracelsus [Pag74a, Cur85, Deb74]. Parachemistries [Muk16]. Paradox [For83, Whi71, Mau64, Whi71]. paradoxes [PPT17]. Paradoxie [Mau64]. Paris [Aic09, Van17, Jac95b, Out83, Pic81]. Parisian [Rab13]. Park [Shi19, Smi14]. Parke [GH20]. Parliament [Gil20]. Parson [McG90]. Parson-Hunter [McG90]. Part [Bat98a, Bat98b, Cla97a, Cla97b, Des85b, Des85a, Fis82a, Fis82b, Fra73b, Fra73a, Kol06a, Kol06b, Lon78b, Lon78a, Mag99a, Mag99b, Old79a, Old79b, Ste05]. Partners [Mor93]. Partnership [Cha08, Win03]. Past [Col89, Mic89a, Mic89b, Pic95, Rud81, Tho73, Car17, Rig17, RRH23a]. Pasteurization [For84b]. Patent [Kan19, She19, Bau19]. patenting [Bia19]. Patents [Bia06]. Path [Tei75]. Pathology [Duk63]. Patient [Jac95b]. Patients [Cul93, Dal97h]. Patron [Bon14]. Patronage [Bia90, Fin93, PD04, Rap69, Car16]. Patrons [Mor93]. Patterning [Rhe11].

Patterns [DT10, Lev84, Tur75]. Paul [LeB01]. Pavlovianism [Gao15]. pays [You67]. Peace [Dal95]. Pearson [Mag99a, Mag99b]. Peasants [Tet23]. Pedagogy [Jac04, Mil08c]. Pedantry [Joh98]. Pedro [Wat74]. Pensée [Slo94, Wil67]. People [Alp91, Jor03, Pal96, SP07, Sut75]. Perception [Fig75]. Père [Liv05]. perfect [Shi19]. Performance [Goo04]. Performing [Sla24]. Perge [Whi64a]. perils [Dav19]. Period [Can64, Mor80b]. periodical [Kaa14]. Periodicals [Rei07]. Peripatetic [Pag75b]. Peripheral [Old97]. Periphery [GPP^{+}08]. Peritissimi [Pag75b]. Perpateticarum [Pag75b]. Persian [Ebr00]. Person [Tom84]. Personal [Coh99]. personification [Bar20a]. Perspective [AA87, Fig76, Mil81, MA09, Win03, Rob14a]. Perspectives [Bro81a, Dob77, Gur21, Hal73a, Liv87, Moo92, Sho86, LJG^{+}24, RRH23a, You67]. Peru [GR22]. Pesticides [Dav19]. Peter [Hen09]. Petrus [Pre22]. Petty [McC13]. Pharmaceutical [Fig73a, GH20]. Pharmacology [Ste65]. pharmacopeia [De 17]. Pharmacy [Whi63b, Dor23]. pharmazeutischen [Fig73a]. Pharmazie [Pag74a]. Phases [Fal21, Rob21]. Phenomena [Mar87]. Phenomenon [Høy90, Sil99, SH18]. Philo [DV01]. Philology [Die24]. Philoponus [Sio88]. Philosophe [Slo94]. Philosopher [AN13, Mer00, Mil00, Mor92a, Rav66a]. Philosophers [Deb74, Sha81, Hel20]. Philosophi [Pag75b]. Philosophiae [Ait73]. Philosophical [Dal97a, Har63a, Hom03, Ste65, Ksi15]. Philosophie [Lav70, Pag73, Pag75a]. Philosophies [Hut91]. philosophischen [Pag73]. Philosophy [BC00, Ben86, Boa13, Buc62c, Buc70, Car79, Clu08, Cro62, Cun91, Ewe63, Gra87, Gra11, Har62, Hei73, Hei81, Hen86, Hoc90b, Hut83, Ili97, Jon74, Kno96, Lau66, Lon89, Mye92, Nor66, Osl98, Pag73, Pag75a, Rav66b, Rup83, Sch83, Sec85, Sha74a, Sio88, Smi73, Smi78, Wea75, Wer00, Wil80a, Win97, WS89a, WS89b, WS90, Woo85, Yol71, Zui07]. Photographic [Bel14, SS15]. photographique [Gil83]. Photography [Bel14, Gil83, Rig17]. Photometry [Joh96]. Phrenology [Coo76, vW04]. Physical [D’A90, Har63b, Hes62, Kni70, Mac21, Mat63, Nav06, Noa04, Tem63, Wil62]. physician [WGvA22]. Physicians [Byn73, Rob66, VFD21]. Physicist [Car77, Har83]. physicists [Mac21]. Physics [Ait71, Can97, Car77, Cla97a, Cla97b, Dea03, Gin01, Han70, Har63a, Hei68, Hei81, Jac04, Lac09, Mol76, Noa05, Por79, Ram10, Sim12b, Smi78, Sta13, Sza90, Tem63, Fal21, OEMB19, Rob21, Sta21]. Physikern [Row98]. Physiological [Smi73]. Physiologie [Goo76]. Physiology [Bro71, Fig75, Fre71, Goo76, Jac87, Kni66, Pic81, Pic88, Mun21, Sta21, Sch68]. Picture [AA06, Buc62b]. Pictures [Kin95]. Picturing [Rup98]. Pieces [Wal97]. Pierre [LeB01]. Pieter [Zui07]. pillar [Van17]. Pillow [Cro08]. Piltdown [Whi80]. Pioneers [Hes62]. Pirenne [PV11]. Piss [Haw11]. Pitcairne [Fri03]. Pitfalls [Hal02]. Pixels [DT10]. Place [Col81, Goo09, Jan00, Koh02b, Pic94, PO02, Yeo85, Tur78]. Places [CP94, Kin14]. Placing [Coo18]. Plague [Lon80, Zui07]. Planet [Hos63a]. Planetarium [Rap21]. Planets [BZ11]. Plant [Nic88, Bri20, Cha18, Mun21].

Plants [Coo82, Ogb13, Por76, APT20, Coo18, Fag21, Yil23]. Plaster [Rie15]. plasticity [CP21]. Plates [Hop05]. Plato [Lau68]. Platonic [DV01]. Platonism [Tur75]. Platonists [DV01]. Platter [Lew62]. Played [Bat98a, Bat98b]. Plea [Luc99]. Pleasure [Nea99]. Plebeians [Can17]. Pleistocene [BdMF01]. Pleurisy [Wil00]. Plots [McG85]. Plotting [McG85]. Poetry [Dav62]. Poincaré [Gin08]. points [Eri22]. Poland [Urb23]. Poli [Esp24]. policies [Bar22]. Polish [Pet23]. Politeness [Wal97]. Political [Fen15, Lev84, Mel96, Pic88, Rey08, WS89a, WS89b, WS90, Sha22]. Politics [Dal97e, Esh01, Gao15, Mus00, Noa05, Pau94, Pet10, Slo90, Woo91, Ano21, Buk20, GM18, Ile24, Leo24, SJ21, VT24, Pic99]. Polyphonic [Sza90]. Polytechnic [Ang12]. Popular [CP94, Mor16, VB87, Win94, dMF23]. Popularising [Ano62d, Rav62b]. Popularization [CP94, FNG22, Flo22, Som22]. Popularizer [Sil12]. Population [Ege77, Nic88]. Populations [McC13]. Porter [Hos03b, And03, Rud03]. Portrait [Mor92a]. Portraits [Jor03]. portraiture [SS15]. Portugal [Car16, Mal18]. Portuguese [Ano21, GM18, Rei07, Sil12, SZDC13, Zil14]. Position [Eyl64]. Positivism [McE97]. Positron [Pei65]. possibilities [RRH23a]. Post [Bos05, Hom85, LeB01, Pic81, San17, Pic88]. post-eugenic [San17]. Post-Galilean [Bos05]. Post-Hypnotic [LeB01]. Post-Newtonian [Hom85]. Post-Revolutionary [Pic81, Pic88]. Posthumous [Die22a]. Postmodernists [Jac95a]. Postwar [vH03, Hew16]. pots [Osb16]. pour [Slo94, Whi64a]. Power [Hoc90a, Kel66b, Mel96, Mil12, Out78, Sou92, FTC16, Flo22, Oak18]. Powers [Cla93]. pp [Kus16]. Practical [Rus94, Woo13, Hel20, Ksi15, Mor18]. Practice [Dal97h, DT10, Ebr00, Fan00, Gol90, Kat13, Kim92b, Koh91, Koh02b, Koh07, Mil12, Pic95, RS62, RTL00, Rap14, Tam15]. Practices [GCM05, Joh96, Lee23, Sto22]. Practitioner [Sha91a]. Practitioners [Cur12, Joh96]. Prairies [Nic88]. Pre [Dan83, Hen86, Høy90]. Pre-History [Dan83]. Pre-Modern [Høy90]. Pre-Newtonian [Hen86]. Preaching [BC98]. Precision [Ili93, Glu24]. Precursor [San79]. Predecessors [Rag07]. Prediction [Kol06a, Kol06b]. Preface [Ano84a]. Preformation [Wil67]. Prehistoric [Sch10b]. Prehistory [Hes62, GP18]. Prelectiones [Wil65a]. Preliminary [Loe70, PD04, Sab87, Wes80]. Prelude [Pag63]. première [Whi64a]. premodern [FSB^{+}23]. Preparation [Ano62b, Ano63b, Ano64b, Ano65b, Ano66b, Ano67b, Ano68a, Ano69a, Ano70a]. Preparatory [Lee23]. Prerogatives [Pye93]. Present [Mat71, Mic89a, Mic89b, Pic95, Rud81]. Presented [SZDC13]. Presentism [MA09]. Presocratic [Lon89]. press [Kaa14, Bro13, MJ01, SZDC13]. Prestige [Zui07]. Presuppositions [Buc62b]. Price [Stu77]. Priestley [Esh01, Sch84b]. Primate [Ree06]. Primitive [Hoy01]. princess [Som21, Spe62]. Principessa [Spe62]. Principia [Sch09a, Ait73, Ait75, Cun91, Whi62a, Whi66b]. Principle [Rud62]. Principles [Bro67, Hen86]. Print [Bia06, Csi10]. Printed [Key63]. printer [Kel66a]. Printing [Eam84]. Priority [Ili92b, Bun19]. Privacy [Ili92b]. Private [Cro98, Dac01, Dik91, Lab21]. privileges [Bun19]. Privy [Col24]. Problem [Can82, LeB01, Pep76, Sti92, Tal65, Van75, vH94, She15]. Problematic [Sha91a]. Problems [App96, Csi10, DT10, Fre71, Goo09, Jac73, McG73, Men64, Rap64, Ras89, Sch74, Sho86, Wes88, Wil62]. Procedures [Ano76f]. Proclus [Cla09]. Product [Kjæ10]. Production [Hün10, Mae07, Goo23, Osb16]. Professional [Haw11]. Professionalization [Bar03, Pal94, SS80]. professor [Car16, Hop97, Nor66, Row98]. professoriate [Buk20]. Profitability [Nea99]. Profits [Haw11]. Progress [Coo76, Cor14, Cro75, Dea01, HT73, HH73, RH88, Rus09, Sch10b, Sch63b, Whi73, Spe15]. Progressing [Sch24]. Project [Val12, GR22, Aga13, Fig73b, Sla24]. projectile [Smi15]. Projectionism [Cla09]. projects [ND22]. Projekt [Fig73b]. Proletariat [Hop97]. Prometheus [Mat71]. Promoting [Hun88, Sch18]. Promotion [May05, Pye93]. proofs [Gia15]. properties [BP19]. Property [Ili92b, Mil08b, Bau19, Bia19]. Prophecy [Kno99]. Prophet [McG90]. Prophets [And99, Jac78, Tre10]. Propriety [Col81]. Prosopography [Pye77, ST74]. prospects [You67]. Protean [San79]. Protection [Bal89, She19]. Prove [AT12]. Provincial [Gol16, Jan00, Liv05, Sch24]. Provocation [Coo76]. Prudence [Joh98]. Prussian [Sto22]. Pseudo [Bow09]. Pseudo-Darwinism [Bow09]. Psychiatric [Suz95]. Psychiatry [And03, Gua91, Chi15b]. Psychical [Noa04, Wil85]. Psychoanalysis [For80]. Psychological [Smi94, Sou92]. Psychologist [Mil75, SS80]. Psychology [Bro71, Bro75, Bro77b, Dal97f, Dan83, Har98, Sen00a, Smi73, Smi82, Smi94, SS80, Tom91a, Sta21]. Psychometry [Str20]. psychophysics [Sta21]. Pub [Sec94]. Public [Bro98a, Bud13, CG21, CP94, Cos02, Cro98, Dik91, HP07, Jub85, Lab21, Mor94, Out78, Pet10, Pic88, Sch83, Wit99, Ros24]. Public/Private [Cro98]. publication [Bec22, Gui17, Pos15]. Publicity [Buk20]. Publics [Sil12]. publishing [Die22a, Die22b, Fyf22, Rie15]. Puffing [Mil00]. Pulmonary [Mas03]. Pump [Woo88, Sin22]. Puno [GR22]. Puno-Tambopata [GR22]. Purifier [Mor92a]. Purifying [vH03]. Puritanism [Par84]. purity [Ber16]. Purse [Col24]. Pursuit [Sor01, Ksi15]. Pus [Ver90]. Put [LP23]. Putting [Wal02]. puzzling [NG22].

qualitatum [Mol69]. Qualities [Cla93, Hen86, Hut91, Mol69]. Quantification [Hes63, McC13, Biv20]. Quantifying [Dal97f]. Quantitative [Edg79, Pal96]. Quantum [Hei68, Hen80]. Quarrel [Sha81]. Queen [Ash01]. Queries [McC13, Sch74]. Quesiti [Kel63]. Quest [AA06, Har98, MF73]. questionnaire [Pre20]. Questions [Sch84b, Tal65, Pag75b]. Questionum [Pag75b]. Quetelet [Don14]. Quinque [Pag75b].

R [Ort05]. Rôle [Har63c]. Race [Dam13, Dik91, GRW22, GR22, Lav22, Pau94, Sen00a, Ton19]. races [Ste20]. Radical [BZ09, Kra24, Slo90]. Radiological [DT10]. railroad [Sch24]. Ranked [Dal99]. Rankine [Mar13]. Ranking [Mar13]. Rarities [Tur86].

Rassenhygiene [Pau94]. Rational [Ili98]. Rauwolf [WGvA22]. Ravished [Feu93]. Razzaz [Kin75]. Re [Can01, Jac06, SH18, Wal02]. Re-Appraising [Wal02]. Re-Creation [Jac06]. Re-Evaluating [Can01]. Re-Examining [Jac06, SH18]. Reaction [Mar82]. readers [Sin22]. Reading [Fle00, Kin95, Ste96]. Really [Mor09c]. Reanimation [Hoy01]. Reappraisal [Hen82]. Reappraisals [Sch78]. Reason [Dob77, Hei78, Hes62, Ili92a, Sto86]. reassembling [CV17]. Reassessment [Liv88, O’C07, Sch73]. Received [Ano62a, Ano63a, Ano64a, Ano65a, Ano66a, Ano67a]. Recently [Ano67b, Ano68a, Ano69a, Ano70a, Ano73i]. Reception [Mar11, Gui17, Pos15]. Recherches [You67]. Reconnaissance [Can97]. Reconsidered [Ili97]. Reconstruction [Buc09, Gia15]. Reconstructions [Sha82]. Record [Jub85]. Records [Ano11a, Jub85]. Recovery [GG04]. Red [Con95, Hop97]. Redi [Fin93]. Redivivus [Mol69]. Redondi [Wes88]. reduce [LJG^{+}24]. Reduction [Kni66]. Reenchanted [Har98]. Reference [Lev67, Rup90]. Referencing [Fox83]. refining [GP24]. Reflecting [Tei90]. reflection [Smi15]. Reflections [CP94, GPP^{+}08, Jac95a, Kni95, Mil89, Mok07, Mor74, Ols90, Por17, Rya66, San79, Tom91b, Bow09]. Reflectors [Hos11]. Reform [Alb96, Ano11a, Buc09, Cro01, Mas03, Per87, RH99, Slo90, Whi05, vW04]. Reformation [Büt79, Jac75, Las73]. Reforming [HW86]. Reforms [Ebr00]. Reframing [Can17]. regard [Poy64]. Régime [Can01]. regulation [Bul18]. Rehabilitating [Hew16]. Reimagining [Sch18]. Reinwardt [Web14]. Rejection [Par84, Sio88]. Rejoinder [Jam86]. Reklamegutachten [Mer20]. Relation [Gru81, Ste65, vdR04]. Relations [Jac87, Lev70, Mil12]. Relationship [Ebr02]. Relativity [Mil75]. Religion [Ebr02, Hei78, Hen93, JJ76, Las73, Liv88, Moo92, Noa05, Osl98, Sha91b, Sio88, Ste96, Ver96, Wil85, Wil77, Woo91]. Religious [Kno96, Mas03, Sho87, Wil80a]. Reluctant [May05]. Remarks [Rag07, Rey08]. Remembering [Chu03]. Renaissance [Cro80, Cur85, Gra87, Hal73b, Hen08, Kel76, Sch73, Sch78, Tal65, Tur75, Wat74, Whi75, Wig63, Wig64, Yat73, Zub64]. reorganization [Mou21]. Reorientation [Büt79]. repair [BD23, Eve23, MB23]. Reply [Cla93, Har09]. Report [May12]. Reports [Mar98]. Representing [Dol98]. Reprinted [Por96]. Reprints [Whi67]. Reproducing [Dal97c]. reptile [Hal15]. Republic [For74, Bon14, Bra23]. Republican [SZDC13]. Reputation [And99, ES15, Mil00]. Rereading [Esh01]. Research [Ait76, Cap76, Cha10, Cla89, Dav75, Eyl66, Gei81, Goo09, Jac06, Lon81, Mar00, Noa04, Sch63a, Sec86, Sha74a, ST74, Sme63, Smi14, Sta86, Tar14, Tho73, Wad01, Whi62c, Wil85, tH04, vH94, GS16, MSR20, Rob20b, RSM20, Sib20, Ton19, You67]. Researches [App96]. Resistance [SL86]. Response [Ano09a, Ber94, Hen09, Not17]. Responses [Biv99, Biv20]. Restoration [Gas85, Jac75, Jac80, Rus94, Jac87, Pic99, Ste20, Wer00, Woo91]. Restorations [Som10]. restraint [GÁ25]. Retorts [Mar98]. Return [SZDC13]. Rev. [Don01]. Revealed [Kas11]. Review [Ack67, Ait71, Ait73, Ait75, All76, Alp91, Ano62d, Ano77b, Ano11a, Arm62, Bar64, Bar79a, Bee75, Ben88, Ber66, Boy67, Bro62, Bro67, Bro77a, Bro92a, Bro92b, Bro81c, Bro84, Bro98b, Bry77, Buc62a, Buc62b, Buc70, Buc09, Bud13, Byn73, Byn88, Cai65, Can82, Can85, Car63, Car77, Car79, Car91, Cha08, Chi65, Cla64, Cla65, Coc63, Con95, Coo82, Cro64, Cro98, Cur85, Dal94, Dal95, Dal97a, Dal97b, Dal97h, Dal97c, Dal97d, Dal97e, Dal97f, Dal97g, Dal97i, Dal98a, Dal98b, Dal99, Dav62, Dav75, Deb63, Deb67, Del05, Dob77, Doe03, Dol88, Dou88, Duk63, Eas86, Edg79, End09, Eyl64, Fig73a, Fig73b, Fig79, Fis73, Fle00, For80, For84b, Fox83, Gau79, Geo85, Gil83, Gin00, Gol86, Gol90, Goo76, Gou74, Gow73, Gru64, Hak95, Hal68, Hal70, HT73]. Review [HH73, Hal73a, Hal73b, Han70, Har83, Har62, Har63c, Har63a, Har63b, Har98, Has79, Has87, Hei78, Hei79, Hei81, Hel64, Hen83, Hen93, Her62a, Her62b, Hes62, Hes63, Hes64, Hoc90b, Hoc90a, Hoc91, Hol75, Hom85, Hom03, Hos62, Hos63a, Hos63b, Hos66, Hos73, Hos03a, Ili92a, Ili97, Ili98, Jac73, Jac95b, Jac95c, Jam86, Jen81, Jon74, Jor78, Kel62a, Kel62b, Kel63, Kel66a, Kel66b, Kel76, Key63, Kin75, Kin77, Kin79, Kni66, Kni93, Kni95, Las73, Lau66, Lau68, Lau93, Lav70, Law90, Law95, Lew62, Llo68, Loe70, Luc95, Mar82, Mat71, Mat63, Mau64, May71, McC85, McG85, McG90, McG66, Mic89a, Mic89b, Mil81, Mol69, Mol81, Moo92, Mor77, Mor80b, Mor92a, Mor93, Mor94, Mor09c, Nee73, Nev76, Nev86, Nic88, Nor66, Nut74, O’M65, Old97, Old98, Pag71, Pag63]. Review [Pag66, Pag73, Pag74a, Pag75a, Pag75b, Pag81, Pal97, Pau94, Pei65, Pic88, Por73, Por76, Por79, Por80b, Por80a, Por81, Por82, Por87a, Por87b, Por91, Por96, Por01, Poy63, Poy64, Pug04, Rap69, Rav62a, Rav62b, Rav63a, Rav63b, Rav66a, Rav66b, RS62, Rob66, Rou79, Row98, Row99, Rud62, Rud63, Rud69, Rud80, Rus74, Rus94, Sch77, Sch82a, Sch82b, Sch84a, Sch78, Sch63b, Sch65, Sen00a, Sha74a, Sha81, Sho83, Sho86, Sho87, Shr66, Sio88, Slo90, Slo94, Sme64a, Sme64b, Smi75, Smi81, Smi82, Smi88, Smi94, Sou63, Spa08, Spe62, Sta63, Ste65, Ste91, Ste02, Ste05, Sti64, Sut75, Sut86, Tal65, Tan98, Tei75, Tei80, Tem63, Tha66, Tom91a, Tur70, Tur73, Tur75, Tur78, Tur86, Tur91, Tur94, Ung77, Wal73, War65, Was64, Was65]. Review [Wat74, Wea75, Web67, Wes67, Wes88, Whi66a, Whi62a, Whi62b, Whi63a, Whi64a, Whi64b, Whi66b, Whi75, Whi71, Whi73, Whi74, Whi83, Whi63b, Wig63, Wil65a, Wil67, Wil65b, Wil80b, Wil85, Wil74, Win03, Woo88, Woo91, Yat73, Yol71, You70, You75, Zub64]. Revised [Jor78]. Revisions [Cur85]. Revisited [Gol11, Hal63, Has79, Hom95, Mar13, NG11, Pag69, Pag70, Ras97, Sch82b, Tan98, Ter04]. Revisiting [Edg05, Kaa14, SdC11]. Revista [Chi23]. Revolution [Ash08, Buc65, Buc09, Coh99, Dal97g, Dob77, Doo95, For84a, Gas85, Hei73, Hei78, Jac07, Lev84, LC98, McE97, Pag81, Per87, Ras97, Sch82b, Sch78, Tei03, Tha70, Tom11, Ver96, Ash14, Hel64, Mok07, Woo91]. Revolutionary [Dal98a, Jac95b, Pic81, Pic88]. Revolutions [For94a, Hos73]. Revolving [BZ11]. Reynolds [Mil12]. Rhetoric [Chr87, Clu08, Fin93, Nea99, Win97, MH23]. Rhetorical [Ili92a]. Richard [Gol16, Har96b, Jor03]. Riddle [Bel09]. Ridicule [Hei78]. Right [Swe93].

Riotous [Bul18]. Rise [Boa13, Car77, Las73, Old79a, Old79b, Rup90, Wea75, GÁ25, Zil14]. Risky [GR22]. Ritual [Wer00]. Rival [HW86, Mag99a, Mag99b]. Rivalry [Lev70]. Road [Whi74, Tet23]. Roaming [Ash98]. Robert [Hal70, Por87a, Bel22, Ber94, Hal70, Ili97, Ost93, Phi06, Por01, Pug04, Win97]. Roberts [Tur94]. Roger [Rav63b]. Roi [Slo94]. Role [Wes80, Ano21, Chi23, GM18]. Roman [Was65]. Romania [Chi23]. Romano [Rap14]. Romans [Was65]. Romantic [Kni70, Rup83]. Romanticism [Str95]. Room [Mye92]. Rooms [Sch07]. Roots [Nic88]. Rosicrucian [Wal73]. Rosicrucians [Wal73]. Ross [Joh98]. Rossii [May71]. Rostow [Ash08]. Rotation [Ste95, Gia15]. Rough [Mor94]. Route [Pye95]. Roy [And03, Hos03b, Rud03]. Royal [HW86, Ili92b, Wer00, Bel22, Gol16, Hal66, Hun88, Jac75, Jac80, Mil89, Mor09a, Ram08, Rob66, RH99, Sar04, Tur78, Web67]. Royston [Mor16]. Rudolf [Nor66]. Rudolph [Pag73]. Rudwick [Kus16]. rule [Kum23]. Rural [Lee23]. Russia [May71, Rya66, Wer23]. Russian [Gor15, SP07]. Rusticall [ND22]. Rutherford [Coc63, Dea03]. Ryle [Por92].

S [Hen83]. Sacrificial [Aga13]. sail [BD23]. Saints [Sch95]. Salernitan [Tal65]. Salpêtrière [Gil83]. saltpeter [ND22]. Salubrity [McC13]. Samuel [Buk20, Sar04]. Sanfte [Kra24]. Santa [Vig21]. Sara [Qur04]. Sarton [Pye95, PV11]. Satire [Rou79]. Saturn [Hos63a]. Savants [Mor80b, Bau19]. Saverio [Esp24]. Saving [Aga13, LJG^{+}24, Mar87, Nic88]. Saxon [Fel18]. Says [Mor10]. Scale [Mat71, McC13, Eri22]. Scales [Dea16]. Scaling [dMF23]. Scandinavian [Chr98]. Sceptical [Woo88]. Schatten [Par20]. Schema [Dan83]. Schenk [Buk20]. Schlagintweit [Fin00]. Scholar [Bon14, Mor10, Sha91a, Ste02, Vla99]. Scholarly [Bro13, Alg23, Rap14]. Scholars [Bry77, Fig79]. Scholarship [Chr98, You66]. Scholastic [Cla93, Hut91, Kal83]. scholasticism [WB14]. Scholia [Cas84]. Scholium [Ste96, DV01]. School [Jac06, Mor93, Nic88, Pic81, Sec86]. Schools [Ano62d, Gei81, Rav62b]. Schooten [Whi64b]. Schriften [Whi62b]. Schrödinger [Yox79]. Science [AA87, All76, Alv13, And99, Ang12, Ank07, Ano62d, Ano09a, Ano20, Ash08, Ash14, Bal89, Bar64, Bar03, Bel14, Ben88, Ber73, Bro62, Bro92b, Bro98a, Bro92c, Buc62b, Buc62c, Buc65, Buc70, Bud13, BWM22, Byn73, Can01, Can82, Car62, Car63, Car79, Car16, Col81, CP94, Cro62, Cro80, Cro88, Cro06, Cro07, Cur85, Dal97c, Dal97e, Dal97i, Dam13, Dav62, Dav95, Dea01, Dea03, Dea12, Deb67, Doe03, Dol77, Don14, Doo95, Eam84, Eas86, Edg79, Edg05, Eme88, Esh01, Ewe63, Fen15, FNG22, For74, For83, For84a, Fox83, Frä74, Fra73b, Fra73a, Fre95, Gau79, GPP^{+}08, GG97, Gay03, Gay08, Gin00, Gol86, Gol02, Gol11, Goo83, Goo09, Gow73, GG90, Gue69, Hak91, Hal63, Hal73a, Hal73b, Han79, Har63b, Har02, Har09, Har98, Has87, Hei78]. Science [Hen83, Hen09, Her62a, Her62b, Hes63, Hig85, Hoc90b, Hoc90a, Hoc91, Hom95, Hom03, Hun88, Ili92a, Jac73, JJ76, Jac95a, Jac07, Jac87, Jan00, Jar03, Jen81, Kas82, Kee14, Kni95, Kus16, Lac09, Las73, Lav70, Law90, Lev70, Liv84, Liv87, Liv88, Lon81, MF73, Mar13, May02, May05, McG85, McG73, MJ01, Mea24, Mil12, Moo92, Mor09a, Mor16, Mor74, Mor77, Mor80b, Mor92a, Mor93, Mor94, Mor09c, Mye92, Nev86, Nic87, NG11, Nor66, Old86, Ols78, Ols90, Ort05, Osl98, Ost93, Out78, Pal94, Pal07, PC18, Par84, Pet10, Pic88, Pic94, Pic95, Pic99, Por80b, Por80a, Por82, Por91, Por01, PO02, PD04, Pye77, Pye89, Pye93, Pye02a, Pye02b, Rab13, Rap69, Rav62b, Ree09, Rei81, Rei94, Ric96, RRH23b, Rot87, Rou79, Row99]. Science [Rud76, Rud81, Rus94, RTL00, Rya66, Sab87, SJ21, SV18, Sch07, Sch78, Sch65, Sec94, Sen92, Sha74b, ST74, Sha82, Sha92, Sha91b, SP87, Sil12, Sim12a, Sio88, Som21, Som22, Ste91, Str95, SC98, Sur19, Sut75, Tan98, Tar14, Tei90, Ter90, Ter95, Tha70, Tom91b, Tre10, Tur70, Tur75, Tur86, Vla99, Wad01, Wal97, Was64, Was65, Web74, Wes67, Whi84, Whi96, Whi74, Wig64, Wil85, Wil77, Woo91, Yat73, pY15, Yol71, Zub64, vH03, vW04, All17, Are19, Bar20b, BB21, Bru17, CG21, Dau25, Dav19, Del19, Fag21, FSB^{+}23, Fel18, Fle21, GRW22, Gil24, GÁ25, Gor15, Gur21, Hel20, HV19, Hyu21, JS24, Jon16, Lab21, Lav22, LP23, Mal18, MSR20, Mou21, Mun21, Oak18, Oli24, Phi15, PPT17, Por17, Pre20]. science [Rab19, RSM20, RRH23a, Roq22, SR20, Sha22, Str20, Sys20, Tam15, Ton19, Vig21, Zil14, Aga13, Dau25, Far95b, Kus16, Mor09c, NG22, SJ21, Urb23, Not17, Pal93, Stu77]. Sciences [Car63, Car88, Eyl64, Fig76, Gol90, Har63c, Hes63, Ili98, Jar00b, Jar04, Kni70, Koh07, Lau93, Lav70, Men64, Mor80a, Nav06, Pye82, Slo94, Tur75, Wil62, Woo85, You66, Chi15a, Wil67, Out83, Rab13]. Scientia [Cuo97]. Scientific [AA85, Alb96, Ano77a, Bal12, Bar03, Bra88, Bro98b, Bry77, Buc62a, Car79, Chr74, Chu03, Coc63, Coh99, Cor14, Cro75, Cro95, Csi10, Cur85, Dal97g, Dob77, Fox88, Fyf22, Gas85, Gau79, Gay08, Gei81, Gin08, Gol90, Gun78, Haj15, Hei73, Hoc90a, Hol81, HSS10, Høy90, Ili98, Jac73, Jub85, Kat13, Kjæ10, Kno99, Koh91, Lac09, Lau68, Lev84, Lig15, LC98, Mar82, Mas02, McG66, Mil08b, Mye92, Ols90, Out76, Pag81, Per87, Qui62, Ras89, Ree09, Rei07, Rud80, San79, Sch95, Sch78, Sch63a, Sch63b, ST74, Sha91a, SP90, Spa04, Stu77, SC98, SL86, Tur70, Tur78, Was64, Whi63a, Whi71, Wig63, Wil80b, Woo13, dB13, Car16, Chi23, Die22b, Fen15, Gao15, Goo23, Kel66a, Oli24, SR20, Sib20, Uys24]. Scientifique [Kel66a]. Scientism [Hak95, Ols78]. Scientist [Hun95, Whi75, Mer20]. Scientists [Dal97g, Dav95, Has79, Maa10, Mor93, Nev86, Rei81, Whi84, GT15, HPtH24]. Scottish [Chr74, Eme88, Fri03, Mor74, Woo90]. Scriptural [O’C07]. Scrupulosity [Por01]. Sea [Con95]. Seamen [Hel20]. Search [Cro80, Csi10, Sch77, Sio88, Sur19, Whi75, Ano20]. Seasons [Ano11a]. Second [Coo82, Pag75b]. Secondary [Old84, Sim12b]. Secret [Keb05]. Secretive [BZ09]. Secrets [Eam84, Kas11, Leo24]. Section [SV18]. Sects [Byn73]. Seeing [RC13, Ste96, Ber14]. Seeking [CG21]. Seen [Sha92]. seine [Mau64]. Seismicity [Con95]. Selected [Wil80b]. Selection [Bru07, Par88, Rus74, Sta63]. Self [Fyf22, Gol02, Wei88, Win97, Cap23, Maa20, Por20, Sys20]. Self-help [Fyf22]. Self-Made [Wei88]. self-sufficiency [Cap23]. self-tracking [Sys20]. Seminar [Cla89, Mye92]. Senescence [Aga13]. sensation [Ber16]. sense [BP19]. Senses [Kuu12, Ber14]. Sensitive [Noa04]. Sensory [Sta21]. Separate [CP94]. Seriality [Csi10, HSS10, Hün10, Ram10, Som10]. Serialized [Kaa14]. Series [HM10, Sch10a, Sch10b, Tre10]. Seriously [Cun91]. serotonin [Hew16]. Service [tH04]. services [Eve23]. Set [Col81, Jam83]. Settler [Dea03]. Seurat [Ste91]. Seventeenth [Bro13, LP23, Car16, Coo18, JS24, ND22, Bos05, Clu08, Dob77, For84a, Geo80, Gol87, Hal63, Hal66, JJ76, Liv88, Mil08c, Rob66, Rup90, Sha91b, Sou92, Ste02, Suz95, Wes67, Yeo07]. Seventeenth-century [LP23, Car16, ND22, Bos05, Clu08, Ste02, Yeo07]. Seventieth [Nor66]. Sewage [Ver90]. Sex [Cro01, GC10]. Sexological [Cro08]. Sexologist [Cro01]. Sextants [Bar17]. Sexual [Cro01, GC10, Pre20]. Sexuality [Sen92]. Sfida [Kel76]. shadow [Wal22]. Shall [AT12]. Shape [Ili93, Not17]. Shaped [Gra11, Tam16]. Shaping [Sen00a, Bau19]. Shared [HV19]. shastri [Sin22]. Shatir [Has79]. Shattering [MH23]. Sheep [Mil81]. Shell [Feu93]. Shield [Rou79]. Shifting [Bal12]. Shining [Smi15]. ship [Eve23, Sam23]. Shipbuilding [BD23]. shipwrights [Eve23]. Shock [Feu93]. Shooting [Cuo97]. Short [Duk63, Shr66]. Shouldering [Rig17]. Showing [Bar20b]. showmanship [Uys24]. Shut [Tur86]. Siberian [Tan98]. Siblings [Dal98a]. sich [Row98]. Sickness [Wor94]. Siècle [Car63, Hoy01, Rap69, Slo94, Buk20, Wil67, Sen92, Kel66a]. Sierra [Eve23]. Sigerist [Poy63]. Significance [Büt79, Ung77, Whi74]. Signs [Kel76, Kno99]. Silver [GP24]. Simple [Ste95, Are19]. simulation [Haj15]. since [Her62b, Rei94, Stu77]. Single [Can01, Tei75]. Singularity [Bar79b, GP24]. Sinology [Fan00]. Sir [Coc63, Mil81, Rus94, SP90, Wal02]. sites [Sav17]. Situating [SS15]. Sixteenth [Lew62, Sto18, Bri09, Eam84, Geo80, Liv88, Rob66, Tar14, Wes80]. sixteenth-century [Sto18, Tar14]. Sixth [Ewe63, Fal21]. skilling [Lee19]. Skills [Boa13, Gro13]. Sky [Whi71]. Slash [Rab19]. Slash-and-burn [Rab19]. Slater [vH03]. Slaves [Bri20]. Sleeping [Wor94]. Small [Gro13]. Smeaton [Mor18]. Smith [Cle89]. snakes [Hal15]. Snapshot [Cla09]. Snapshot-Idealization [Cla09]. Snobs [Bro92c]. Snow [Bra88, Edg05]. Social [Bia89, Bro77a, Car79, Col81, Dal98b, Elw03, Fen15, Fox88, Gay08, Gou74, Har63c, Hen83, Hes62, Hes63, Jac73, Pic84, Pic88, Por92, Pum91, Ree06, Rud80, Rup90, Sou63, Tei90, Tre10, Wei86, Woo85, Zui07, Bia19, CG21, Chi15a, Gil20, She15]. Socialism [Tre10]. Sociality [Ree06]. Socialized [Hoc90a]. Socially [Lee19]. Societies [Sch63a, FSB^{+}23, Fyf22]. Society [Bal89, Dal94, Dal95, Gou74, Gow73, Hal63, Hal73b, Hoc90a, Hom03, Mor77, Sho86, Web74, Yat73, tH04, Som21, Hal66, HW86, Hun88, Ili92b, Jac75, Jac80, Mil89, Mor09a, Ram08, RH99, Sar04, Ste65, Web67]. Sociobiological [McG90]. Sociological [Sha82]. Sociology [AA87, Sha92]. Sociomaterial [Par20]. soft [Sha22]. software [She19]. soil [Are19]. Solar [Wal99]. Solitary [Dal94]. Solomon [HW86]. Solution [Pep76]. Solutions [Pel01, Mau64]. Solves [Bel09]. Some [Ano76f, Bow09, Bro71, Cro75, Fre71, GPP^{+}08, Gue79, McG73, Men64, Moo92, Pep76, Rag07, Rya66, San79, Sho86]. Son [Lew62, Sor01]. Song [dP18]. Sorbière [Sar04]. Soul [Sen00a]. Souls [Siv01, Str20]. Source [Smi81, Gia15]. Sources [Ano76f, Bra77, Fra73b, Fra73a, Fre71, Geo80, Hal66, Lev67, Mad63, Men64, Pye95, Rap64, Smi82, Tei80, Tur69, Wil62, GGN16]. South [Hyu21, Lee23, PC18]. Southern [Fre95]. southwest [Gur21]. Soviet [Bro71]. Space [Dik91, Whi96, Sim17]. Spaces [Sch07, Ste07]. Spain [Flo22, GC10]. Spanish [Can17, Goo83, LN17, Mel96]. spatial [BB21]. spatialization [Fle21]. Speaking [Moo92]. Special [Rup90, SV18, Rob20a, RRH23b, Mil75]. Specialties [Gei81]. Spectacle [Sch83]. Spectro [Jam86]. Spectro-Chemistry [Jam86]. Spectroscopy [Jam85, Sut86]. Speech [Ogb13]. Speed [Goo04]. Spencer [Cal98, Elw03]. Sphere [Bro98a, Bud13, Cos02, Wit99, Rap21]. Spheres [Ait81, CP94]. Spirit [Aic09, Gra11, Urb23]. Spiritual [Has79, Wil85]. Spiritualism [Noa04, Wil85]. spirituality [PC18]. Spiritualized [Siv01]. Sprachgebiet [Fig73b]. Sprat [Mor09a]. spread [Sch24]. Spreading [CG82]. Spring [Pal96, Bel22]. Square [GCM05, GCM06]. Squibs [Bro92c]. Stability [Van06]. Stabilization [SH18]. Stabilizing [Kim92b]. Stadt [Bee75]. Stage [Kin95, Lac09]. Stages [Hop05]. Standardization [Hün10]. Standards [Hop05, Pel01]. Stars [Whi83, Glu24]. State [Ash98, JJ76, Mae07, Rey08, Tur91, Wad01, tH04, Cha25, Fen15, PC18, Tur73]. Statement [Sab87]. States [Pal94, Por20, Rei94, She19, SM24, Tuc20]. Statesman [Bon14]. Station [Pum91, Ton19]. Statistics [Mar00, Mar11, SP07]. Statistik [Ham93]. Status [Bia89, Keb05, Mar98, Zui07]. Steam [Hom03, Kel66b, Mil12, BD23]. Steamship [SHW03]. Steel [Luc95]. Stellar [Whi83, Wil83]. Steps [Sch65]. sterilization [Wal22]. Stewart [Por96, Por96]. Stiintifica [Chi23]. Stimulated [Dan83]. Stoic [DV01]. Stone [Sch10b, Bru17]. stop [Smi14]. Storage [Kat13]. Stories [Jar03, PC18, Phi15]. Story [Tur78]. straight [WB14]. Stranger [Nic88, Roq22]. stranger-science [Roq22]. Strata [Por87b]. Strategies [HW86, Löw92]. strategy [Vig21]. Streams [Rob05]. Strength [Löw92]. Structure [Fox88, Hes62, Hol81, Kim92a, Pic84, Ksi15]. Structures [Sho87]. Struggle [NG11, Sch68, LJG^{+}24, Zel19]. Struggles [Rus09]. Struggling [Cha25, Flo22]. Student [Lew62]. Students [Cur12]. Studien [Pag74a]. Studies [Bow09, Bro81b, Can82, Col77, Dav78, Gol90, Gue79, Kno79, Loe70, Loe65, Mad63, Nor66, Pag74b, Rud80, Sch82a, Slo94, Sme64b, Tei80, Tur69, Tur75, Whi83, Wil80b, Wil74, Wil02, Por17, Sta21, WGvA22, Yil23]. Study [Ano11a, Bro81b, Car68, Dav75, Har96b, Hei81, Hom95, McE97, McE13, Nic87, Par88, Pep67, Pum91, PD04, Rup83, Smi88, SL86, Tan67, Tur75, Ung77, Wes80, Fen15, JS24, Ste20]. Studying [Fag21, Tom91a]. Stukeley [RH99]. Sturgeon [Mor92b]. Style [Mar11]. Styles [Cro95, Ili98, LS17]. Sub [Høy90]. Sub-Scientific [Høy90]. Subjective [Sto22, Str20]. Subjectivities [GC10]. Sublime [Has79]. Subsequent [Sab87]. Subtilissimi [Pag75b]. Success [Mar98]. Such [Wil83, Fle18]. Sudhoffs [Pag74a]. sufficiency [Cap23]. Suggestion [LeB01]. Suggestions [Eyl66]. suivi [For84b]. Sun [Tur89]. Sun-Dials [Tur89]. Sunspot [May12]. sunspots [Rap14]. Superlative [Lon81]. Supernaturalism [Hut83]. Support [Cro07]. surface [Fle18]. Surfaces [Cor14]. Surgeons [Sch95, Woo13]. Surgery [Dal95, Sch07, Sen00a]. Surgical [Ano76f]. Survey [Bra77, Cro07, Mad63, Sch84b, Sec86, Tur69, Hyu21]. survival [Flo22]. Surviving [Tur86]. Suzhou [Sch18]. Swedish [Bec04, Bec22, Cur12, Frä74]. Sweeps [Hos05]. syllabus [ABC^{+}23]. Symmetries [Spa04]. Symposium [Bee75, Hal73a, Pal96, Ano74a]. synecdoche [Dav19]. Syntheses [Hak91]. Synthesis [Smi94, Van06]. synthetic [Ber16]. System [Bia90, Ham11, Nut74, Tem63, Bau19]. Systematics [Joh05]. Systems [Feu93].

T. [Hen83, McG90]. Tables [Hop05, Coo18, Maa20]. Tadini [Gia15]. Takes [Pal97]. Taking [Cun91]. Talk [Cro08]. talked [HPtH24]. Talking [Ogb13]. Tambopata [GR22]. Tangled [Bil22]. Taqi [Has79]. Tartaglia [Kel76, Cuo97, Kel63, Kel76]. Taste [Bro77a]. Teacher [Bar79a, McE13]. Teaching [Ber14, Rap14, Whi17]. Teams [RTL00]. Technical [Ash01, Car79, RTL00]. Technicians [RTL00]. Techniques [MB23, cW18]. Technocratic [Edg05]. Technocrats [May05]. Technological [Mat71]. technologies [Mar24, BB19]. Technologizing [Mor99]. Technology [Ano09a, Bru07, Car62, Car68, Car79, Car91, For85, GPP^{+}08, Gue69, Hal73a, HM10, HP07, Hoc90a, Hol25, Jen81, Kat13, Lev70, Luc95, May02, Mer00, Pic94, Por80a, Rav62b, Sim12a, Sou63, Tha70, BB19, Fle21, Par20, VT24, Mor93]. Tedesca [Spe62]. Telenaturalist [Gou11b]. Telescope [AT12, Van75, Sec85]. tell [LS17]. Telling [Tur78]. Temple [Mor10]. Temporal [Dea16]. Ten [Dal99, Sme63]. Tension [Wil80b]. tensions [Gil20]. Tenth [Eva78]. Teodoro [Sil12]. Terminology [Pag74a, Sur19, Ano20, Tam15]. Terms [Gra87]. Terpsichore [Hua16]. territorial [Coo18]. Terror [Out83]. Test [Sch09b]. Testing [Mil12, Cha25]. text [SV18]. Textbook [Duc05]. Textbooks [Cla97a, Cla97b, Sim12b]. Texte [Whi64b]. Texts [Ste05]. Textual [Fan00, Ara19, Die24]. textual-visual [Die24]. Th. [Kon19]. Theatre [Lac09, Whi66a]. Theatres [Rup90]. Theatrical [Hua16]. Theft [Gou25]. Their [Bro71, Lau93, Por87b, Rud81, HPtH24, Lig15]. Thelesmos [Kel10]. Them [Ste02]. Themes [AA85, Ste91]. Theology [Esh01, Gra11, Hen09, Osl98, Tur75, Oak18]. Theorie [Tur73]. Theories [Car77, Fig75, Har98, Hes62, Hom85, Lau68, Sch82a, Sio88, Wes67, Wil74]. Theory [Buc70, D’A90, Dal97c, Ebr00, Fox88, Han70, Har96b, Hen86, Hes62, Hom77, Hos63a, Mil75, Nic88, Old84, Old97, Pal94, Rey08, Rud69, Sta63, Swe93, Ung77, Mau64, Tur73]. Therapy [Far95a]. There [Coh99, Rus09, Wil83]. Theresa [Fag21]. Thermochemistry [Dol84]. Thermodynamics [Dol84]. Theses [Ano62b, Ano63b, Ano64b, Ano65b, Ano66b, Ano67b, Ano68a, Ano69a, Ano70a, Ano73i, Rad83]. Thesis [Noa05, Wes88]. Thing [End09, Wil83]. Things [Del11, For80, Hes62, Jor03, Bun19, Whi17]. Thinking [Can85, Cor14, Cro95, Ili98, MA09, Rot87]. Third [Dal97g]. Thirty [Mil08a]. Thomas [Dea03, Eng78, Haw11, Hen82, Mor09a, Pep67, Por87a, Sar04, Tan67, Whi96, Whi75, Wil80b]. Thomson [Bel14, Nav06]. Thought [Bos05, Cla64, Geo85, Ras89, Smi88, Was64, Whi71]. Thoughts [Har83]. Thousand [Dav95, Mor09c]. Threat [Ano09a]. Three [Pic11b, Tur78]. Thunderstorms [Esp24]. Tides [Ros24]. Tigris [Col24]. Time [Buc09, Dea16, Edd01, Har62, Rup98, Sio88, Sza90, Gil20]. Timing [Glu24]. Tissot [Don01]. Today [Slo94]. Toil [Por80a, Lee19]. Tokyo [Kik12]. Tollet [Far02]. Tom [Sec85]. tomb [Rig17]. tong [SV18]. Too [End09, LS17, She15]. Tool [ST74]. Tools [Bud13]. Toothache [Kin95]. Tops [Sec85]. Tory [Fri03]. Total [Ack67]. toutes [Rob14b]. town [Sch24]. townsfolks [Tet23]. Traces [Lin06, Rhe11, CP21]. tracking [Sys20]. Tract [Ano76f]. Tractatus [Mol69]. Trade [Joh98, Nev76, Sch06, SM24]. Trademarks [BP19]. Tradescant [Tur86]. Trading [Kin14]. Tradition [Chu03, Dal97d, Dol88, Duc05, Eam84, Gin00, Ili98, Lan09, Wil80b]. traduites [Whi64a]. trail [Oli24, Uys24]. Training [Duc05, VFD21]. Trajectories [Sta13]. tranquilizing [Uys24]. Trans [AA87]. Trans-Disciplinary [AA87]. Transactions [Ste65]. Transatlantic [Pet10, Pos15]. transfer [BD23]. Transformation [SC98, Suz95, BD23, Kra24, MB23]. transforming [SH18]. Transfusion [Pel01]. Transit [Mas03]. Transition [GC10, Pag66]. Transits [Hos62]. Translating [Chi15b]. Translation [Ano11a]. translational [Yil23]. Transmission [Dol77, Ras89]. Transnational [AA06, LJG^{+}24]. Transplant [Sch95]. transport [Tet23]. transportation [MB23]. Travel [Del05]. travellers [Fle18]. treasures [MP14]. Treat [GG90]. Treatise [Mol69, Ste02, VB87]. Treatment [Wil67]. Trends [Høy96, VFD21, FSB^{+}23]. Trial [Geo85, Gua91]. trials [GH20]. Triangle [Swe93]. Triangulating [Sta86]. Trilogy [Cha08]. Trismegistus [Par84]. Triumph [Rus74]. Trochet [Goo76]. Trouble [Dal97i]. True [Mil04]. Trust [Cro08, Mar00, Mil12, ES15]. Trustworthy [SHW03]. Truth [Bat98a, Bat98b, Cro80]. Tubal [Mer00]. Tuberculosis [Dal98b, Wad01]. Turn [Hal68, Hua16, Urb23]. Turns [SZDC13]. Turpentine [Mae07]. Tutankhamun [Rig17]. Twentieth [SS80, Bar20a, BB21, Cha18, Cha25, Fle21, Por20, Tam15, cW18, Dam13, GG04, Har98, Sta13, vH94]. twentieth-century [Cha18, Cha25, Dam13, Har98]. Two [Ano62d, Kel62b, Por05, Rav62b, Woo85, Bru17, Mau64]. Tycho [Bon08, Chr98, Tho73]. types [Pos15].

U.S. [DT10]. Ukrainian [Ano20, Sur19]. Unbound [Mat71]. uncontrollable [Mar24]. Undead [Ree09]. Undergraduate [Bro92c, Sch84a]. Underground [Sho94, Esp24]. Understand [Hol92]. Understanding [Eng78, Hen83, Pye95]. understood [Are19]. Uneasy [Ben88]. UNESCO [Mou21, NG22]. Unfinished [Hos05]. Unger [Drö15]. Unification [Pau94]. Uniformity [Mol69, Rud62]. Unit [Gou11a]. United [Pal94, Por20, Rei94, She19, SM24, Tuc20, Wad01]. Unity [Sio88]. Universal [Mor09a]. Universalis [Wil65a, Kno79]. Universe [Arm62, KS03, Wat74, All17]. Universities [For94b, Fra73b, Fra73a, Gas85, Mor77, Pal97]. University [Mac21, Car16, Edd01, Hop97, Mag99a, Mag99b, Mor77, PO02, Tur70]. Unknown [Kno79]. Unnamed [Dor23]. Unpublished [Pep67, Whi63a]. Unravels [AN13]. unskilled [Dor23]. Unsolved [Sch74]. Uppsala [Jon74]. Urania [Hua16]. urban [Hol25]. Use [BC98, GCM06, Han79, Ste02, Zui07, Ile24, Mar24, Rab19, Tur73]. Useful [Ber07, Bru07, McE13, Fag21, GP24, Gur21, Whi17]. Uses [Gol90, Jar00b, Lau93, Sen00b, Gur21, Whi17]. USSR [Kon19, Tan98]. Utility [Nea99].

v [Ash01, Kim92a, May71, Woo88]. Vacation [Ank07]. Valency [Fis73]. Valley [Mil89]. Vallisneri [Rap91]. Value [BC00]. Values [Car79]. Vanilla [Ber16]. Variants [Por12]. variation [SH18]. Vascular [Nut74]. Vasile [Chi23]. vectors [Mea24]. Vegetality [Coo82]. Vegetation [Nic87]. Venel [Lev84]. Venice [McE13]. Ventilation [Sam23]. Ventriloquized [Jac95c]. Venus [Hos62, Qur04]. Vernacular [Mea24, Gur21]. Verse [Dav62]. Versus [Cla09, Buc62a, Dol84, GP18, Lon78b, Lon78a]. Vesalian [O’M65]. Vesalius [Cla65]. vessels [Cap23]. Veterinary [Woo13]. Vices [Ili92a]. Victorian [Mor94, Win94, And99, Bar03, Can64, Dol98, Hal15, Hig85, Jam85, Leo24, McG90, MJ01, Mor80b, Mor99, Nav06, Nev86, Noa05, Ram10, SS15, Smi75, Smi20, Wil77]. Victorians [Sut86]. victory [Pag16]. Vie [Slo94, Wil67]. Vienna [Buk20]. Viète [Par88]. View [Coh99, Gra78, Gru64, Hal75, Høy96, Hut12, Mil75]. Viewed [Cha08, Win03]. Views [Arm62, Hes64]. Vincenzio [Bos05]. Vindicated [Sut86]. virtue [GH20]. Virtues [Hut91, HPtH24]. Virus [vH94]. vis [Mor18, Ter04]. visibility [Bar20b]. Visible [Sho94]. Vision [DT10, Hom85]. Visions [Cro88, Kno99]. Visual [dMF23, Hop05, Kus16, Rud76, Die24]. Visualizing [Car17]. viva [Mor18, Ter04]. Viviani [Bos05]. Vocation [Out83, Pic88]. Vocations [Hos03c]. voiced [Zil14]. Voices [RH99]. Void [Pal05]. Völlig [Pag73]. Volume [HT73, HH73, Whi73]. Volumes [Tur78]. Voluntarism [Har02, Har09]. Voluntarist [Hen09, Oak18]. vor [Row98]. Voyages [Wer23].

W. [Mar13]. Walker [Edd01]. Wallace [Rus74]. Wallis [Ram08]. Walter [Yat73]. War [Dal95, Mil08a, Sha81, Sto22, Car17, Far23, GR22]. Warehouse [Ili92b]. Warfare [Har96a]. Warner [Leo24]. wartime [Mou21]. Was [Wil83, vW04, Mun21, Sar04, Sta21]. Waste [WS89a, WS89b, WS90].

Water [Del11, Kin77, Por91, Zel19]. Water-Clocks [Kin77]. waterwheel [Mor18]. Watt [Mil00, Mil04]. wax [JS24]. Way [Cha11, Löw11, Wea83, Whi74]. Wayne [Hew16]. Ways [Gra87, Pic11a, Por01]. Weather [And99]. Wednesday [Mye92]. Weighing [Bia19, Biv20]. Weil [Bee75]. Weimar [Har96b, Hen80]. Weininger [Sen92]. Wellcome [Key63]. Were [Old98, Pye77, PPT17]. Werke [Whi62b]. Werner [Eyl64]. Wertheimer [Mil75]. Wesley [Don01]. West [Bro75, Osb16]. western [Tet23, Cro88, Dal97d, Dal97e, Has87, Hoy01, Jac95a, JS24, Mat71, Pye93, Rav62b, Sch65, Ste91]. Wheatstone [Hom85]. Wheelwright [Alp91]. Where [Yox79]. Whewell [Mer00, Mor92a, Mor94]. Which [Hoc91, Noa04, Pag16]. Whig [Ols78]. Whiggish [Alv13]. Whiggism [Hal83, McE97]. Whigs [Jar03]. white [JS24]. Whittaker [Nav21]. Who [KS03, Pug04, Sar04, Pye77]. Whole [Pag63]. Wider [Geo85]. Wilberforce [Kaa14]. Wilhelm [Jac06, Har98]. Will [Tan67]. William [Bro94, SP90, Ack67, Dav75, Hes62, Hos63b, Hos03c, Hos05, Hos11, Ili97, McC13, Mer00, Mor92a, Mor94, Mor92b, Noa04, Pag69, Pag70, Par84, RH99, Ste20, Sti64, Wea83, Wil65a]. Willis [Haw11]. Wilson [Oli24]. Wissenschaft [Phi15]. Wissenschaften [Whi64b]. Wizards [Pal94]. Wolffianism [Jon74]. Wolffianismens [Jon74]. Woltereck [Har96b]. Woman [Far02, Fle00, Tom91b]. Womb [Dal97c]. Women [Ben88, Fle00, Jac95c, Pag81, Row99, Sho86, Pet23, VFD21]. won [Pag16]. Wonder [Lac09]. Wonders [Tur86]. Woodhouse [Phi06]. Woodward [Rou79]. Wool [Mil81]. Woolley [Hew16]. Word [CG82, Ram08, Rob16b]. Work [Alp91, Bos05, Bro84, Dav75, Has79, Joh96, Kel63, Mag99a, Mag99b, Por80a, Pug04, RTL00, WS89a, WS89b, WS90, Alg23, RRH23b, SS15]. Workers [Don14, Bar20b, Far23]. Working [MJ01, Pic11a]. Working-Class [MJ01]. Works [Byn88, HT73, HH73, Por87a, Por96, Por01, SHW03, Whi73]. World [Buc62b, Cro01, Del11, Far23, Gin00, Gra78, Gru64, Has87, Mat63, Rud80, Sti92, Tem63, Tur86, Whi62c, Wil85, Zui07, CV17, Del19, GRW22, Gou25, Lav22, LS17, Shi19, GP24, Poy63]. World-Maps [Gin00]. Worlds [Buc09]. Worldview [Bru87]. worrying [Smi14]. Wörterbuch [Pag75a, Pag73]. Wörterbuchs [Pag73]. Wotton [Ili97]. Wounds [Wer00]. Writing [Buc09, Kni93, Lin06, Rab13, Sch95, dB13]. Writings [Bro71, Mic89a, Mic89b]. Wurrld [SS80]. Wynne [Noa05].

xiv [Kus16]. XIXe [Car63]. XVe [Kel66a]. XVIII [Slo94, Wil67]. XVIIP [Rap69].

Year [Dal97d, Doe03, Mor09c]. Years [Byn75, Dav95, Mil08a, Ost93, Por82, Rus74, Sme63, Tur91, Wil02, Tuc20]. York [SJ21]. Young [O’C07, She15]. Young-Earth [O’C07].

Zabarella [Pal07]. Zaccaria [Not17]. zaolian [She15]. ZEITGEIST [Sto86]. Zen [cW18]. Zeno [Mau64]. zenonischen [Mau64]. Zhengxiang [Fen15].

Zhu [Fra24]. zone [Mou21, Muk16, Hyu21]. zoo [Bar22]. Zoological [Hal15]. Zoologie [Cai65]. Zoology [Des85b, Des85a, Cai65]. Zoophyte [Gib12]. Zoos [Dal97a]. Zorndorf [Sto22]. zur [Pag74a, Whi64b]. Zustand [Tur73].

References

Abir-Am:1985:TGO

[AA85] Pnina Abir-Am. Themes, genres and orders of legitimation in the consolidation of new scientific disciplines: Deconstructing the historiography of molecular biology. History of Science (UK), 23(1):73– 117, March 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/1/
73.full.pdf+html.

Ascher:1986:E

[AA86] Marcia Ascher and Robert Ascher. Ethnomathematics. History of Science (UK), 24(2):125–144, June 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/24/2/125.full.pdf+html.

Abir-Am:1987:BGT

[AA87] Pnina G. Abir-Am. The biotheoretical gathering, trans-disciplinary authority and the incipient legitimation of molecular biology in the 1930s: New perspective on the historical sociology of science. History of Science (UK), 25(1):1–70, March 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/25/1/1.full.pdf+html.

Abir-Am:2006:MBR

[AA06] Pnina G. Abir-Am. Molecular biology and its recent historiography: a transnational quest for the ‘big picture’. History of Science (UK), 44 (1):95–118, March 2006. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
44/1/95.full.pdf+html.

Anthony:2023:MLI

[ABC^{+}23] Patrick Anthony, Juliana Broad, Xan Chacko, Zachary Dorner, Judith Kaplan, and Duygu Yildirim. (Un)making labor invisible: a syllabus. History of Science (UK), 61(4):??, December 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231180954.

Ackerknecht:1967:ERT

[Ack67] Erwin H. Ackerknecht. Essay review: The total Harvey: William Harvey’s biological ideas. History of Science (UK), 6(1):169–171, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/6/1/169.full.
pdf+html.

Agar:2013:SES

[Aga13] Jon Agar. Sacrificial experts? Science, senescence and saving the British Nuclear Project. History of Science (UK), 51(1):63–84, March 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/51/1/63.full.
pdf+html.

Aicardi:2009:ASP

[Aic09] Christine Aicardi. The analytic spirit and the Paris institution for the deaf-mutes, 1760–1830. History of Science (UK), 47(2):175–221, June 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/47/2/175.full.
pdf+html.

Aiton:1971:ERC

[Ait71] E. J. Aiton. Essay review: The concept of force: Force in Newton’s physics. History of Science (UK), 10(1):88–102, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/10/1/88.full.pdf+html.

Aiton:1973:ERN

[Ait73] E. J. Aiton. Essay review: Newton’s Principia: Introduction to Newton’s ‘Principia’, Isaac Newton’s ‘philosophiae naturalis Principia mathematica’. History of Science (UK), 11(3):217–230, September 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/11/3/217.full.
pdf+html.

Aiton:1975:BRM

[Ait75] E. J. Aiton. Book review: The mathematics of Newton’s “Principia”: The mathematical papers of Isaac Newton. History of Science (UK), 13(4):301–303, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/13/4/301.full.pdf+html.

Aiton:1976:JKL

[Ait76] E. J. Aiton. Johannes Kepler in the light of recent research. History of Science (UK), 14(2):77–100, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/14/2/77.full.pdf+html.

Aiton:1981:CSC

[Ait81] E. J. Aiton. Celestial spheres and circles. History of Science (UK), 19(2):75–114, June 1981. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
19/2/75.full.pdf+html.

Alborn:1996:BII

[Alb96] Timothy L. Alborn. The business of induction: Industry and genius in the language of British scientific reform, 1820–1840. History of Science (UK), 34(1):91–121, March 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/34/1/91.full.pdf+html.

Algazi:2023:KLF

[Alg23] Gadi Algazi. Kepler’s labors: Figurations of scholarly work c. 1600. History of Science (UK), 61(4):??, December 1, 2023. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753231180287.

Allchin:1976:BRI

[All76] F. R. Allchin. Book review: Indian science: a concise history of science in India. History of Science (UK), 14(2):141–142, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/14/2/141.full.pdf+html.

Allegro:2017:BUF

[All17] James J. Allegro. The bottom of the universe: Flat earth science in the Age of Encounter. History of Science (UK), 55(1):61–85, March 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://journals.sagepub.com/doi/full/10.1177/
0073275316681799.

Alpern:1991:ERA

[Alp91] Andrew Alpern. Essay review: Astronomy for the people, wheelwright of the heavens: The life & work of James Ferguson, FRS. History of Science (UK), 29(3):329–331, September 1991. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/29/3/329.full.pdf+html.

Alvargonzalez:2013:HSE

[Alv13] David Alvargonzález. Is the history of science essentially Whiggish? History of Science (UK), 51(1):85–99, March 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/51/1/85.full.pdf+html.

Amery:2024:LMC

[Ame24] Fiona Amery. From laboratory to mountaintop: Creating an artificial aurora in the late nineteenth century. History of Science (UK), 62(4): 591–623, December 2024. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753241229147.

Abou-Nemeh:2013:NPM

[AN13] S. Catherine Abou-Nemeh. The natural philosopher and the microscope: Nicolas Hartsoeker unravels Nature’s “Admirable OEconomy”. History of Science (UK), 51(1):1–32, March 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/51/1/1.full.pdf+html.

Anderson:1999:WPS

[And99] Katharine Anderson. The weather prophets: Science and reputation in Victorian meteorology. History of Science (UK), 37(2):179– 216, June 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/2/
179.full.pdf+html.

Andrews:2003:GMB

[And03] Jonathan Andrews. Grand master of bedlam: Roy Porter and the history of psychiatry. History of Science (UK), 41(3):269–286, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/41/3/269.full.
pdf+html.

Angulo:2012:PCA

[Ang12] A. J. Angulo. The Polytechnic comes to America: How French approaches to science instruction influenced mid-Nineteenth Century American higher education. History of Science (UK), 50(3):315–338, September 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/50/3/
315.full.pdf+html.

Anker:2007:SVH

[Ank07] Peder Anker. Science as a vacation: a history of ecology in Norway. History of Science (UK), 45(4):455–479, December 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/45/4/455.full.pdf+html.

Anonymous:1962:BR

[Ano62a] Anonymous. Books received. History of Science (UK), 1(1):132– 133, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/1/1/
132.full.pdf+html.

Anonymous:1962:DTP

[Ano62b] Anonymous. Doctoral theses in preparation. History of Science (UK), 1(1):131, March 1962. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/1/
1/131.full.pdf+html.

Anonymous:1962:EN

[Ano62c] Anonymous. Editorial note. History of Science (UK), 1(1):vii, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/1/1/vii.full.
pdf+html.

Anonymous:1962:ERP

[Ano62d] Anonymous. Essay review: “Popularising dead science”: Two books for schools: The major achievements of science. History of Science (UK), 1(1):103–106, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/103.full.pdf+html.

Anonymous:1962:NB

[Ano62e] Anonymous. Notices of books. History of Science (UK), 1(1): 124–127, March 1962. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/1/
1/124.full.pdf+html.

Anonymous:1963:BR

[Ano63a] Anonymous. Books received. History of Science (UK), 2(1):176– 177, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/2/1/
176.full.pdf+html.

Anonymous:1963:DTP

[Ano63b] Anonymous. Doctoral theses in preparation. History of Science (UK), 2(1):175, March 1963. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/2/
1/175.full.pdf+html.

Anonymous:1963:NB

[Ano63c] Anonymous. Notices of books. History of Science (UK), 2(1):174, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/2/1/
174.full.pdf+html.

Anonymous:1964:BR

[Ano64a] Anonymous. Books received. History of Science (UK), 3(1):155, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/3/1/
155.full.pdf+html.

Anonymous:1964:DTP

[Ano64b] Anonymous. Doctoral theses in preparation. History of Science (UK), 3(1):154, March 1964. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/3/
1/154.full.pdf+html.

Anonymous:1964:NB

[Ano64c] Anonymous. Notices of books. History of Science (UK), 3(1):153, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/3/1/
153.full.pdf+html.

Anonymous:1965:BR

[Ano65a] Anonymous. Books received. History of Science (UK), 4(1):153– 154, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/4/1/
153.full.pdf+html.

Anonymous:1965:DTP

[Ano65b] Anonymous. Doctoral theses in preparation. History of Science (UK), 4(1):155, March 1965. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/4/
1/155.full.pdf+html.

Anonymous:1966:BR

[Ano66a] Anonymous. Books received. History of Science (UK), 5(1):160– 161, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/5/1/
160.full.pdf+html.

Anonymous:1966:DTP

[Ano66b] Anonymous. Doctoral theses in preparation. History of Science (UK), 5(1):161, March 1966. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/5/
1/161.full.pdf+html.

Anonymous:1966:NB

[Ano66c] Anonymous. Notices of books. History of Science (UK), 5(1):159, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/5/1/
159.full.pdf+html.

Anonymous:1967:BR

[Ano67a] Anonymous. Books received. History of Science (UK), 6(1):176– 177, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/6/1/
176.full.pdf+html.

Anonymous:1967:DTP

[Ano67b] Anonymous. Doctoral theses in preparation or recently completed. History of Science (UK), 6(1):177–178, March 1967. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/6/1/177.full.pdf+html.

Anonymous:1967:NB

[Ano67c] Anonymous. Notices of books. History of Science (UK), 6(1): 172–176, March 1967. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/6/
1/172.full.pdf+html.

Anonymous:1968:DTP

[Ano68a] Anonymous. Doctoral theses in preparation or recently completed. History of Science (UK), 7(1):147–148, March 1968. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/7/1/147.full.pdf+html.

Anonymous:1968:NB

[Ano68b] Anonymous. Notices of books. History of Science (UK), 7(1): 144–146, March 1968. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/7/
1/144.full.pdf+html.

Anonymous:1969:DTP

[Ano69a] Anonymous. Doctoral theses in preparation or recently completed. History of Science (UK), 8(1):166–167, March 1969. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/8/1/166.full.pdf+html.

Anonymous:1969:NB

[Ano69b] Anonymous. Notices of books. History of Science (UK), 8(1): 158–165, March 1969. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/8/
1/158.full.pdf+html.

Anonymous:1970:DTP

[Ano70a] Anonymous. Doctoral theses in preparation or recently completed. History of Science (UK), 9(1):146–147, March 1970. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/9/1/146.full.pdf+html.

Anonymous:1970:NB

[Ano70b] Anonymous. Notices of books. History of Science (UK), 9(1): 140–145, March 1970. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/9/
1/140.full.pdf+html.

Anonymous:1971:NB

[Ano71] Anonymous. Notices of books. History of Science (UK), 10(1):133– 137, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/10/1/
133.full.pdf+html.

Anonymous:1973:EF

[Ano73a] Anonymous. Editorial foreword. History of Science (UK), 11(1): i, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/1/
i.full.pdf+html.

Anonymous:1973:NCa

[Ano73b] Anonymous. Notes on contributors. History of Science (UK), 11(1): 74, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/1/
74.full.pdf+html.

Anonymous:1973:NCb

[Ano73c] Anonymous. Notes on contributors. History of Science (UK), 11(2): 158, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/2/
158.full.pdf+html.

Anonymous:1973:NCc

[Ano73d] Anonymous. Notes on contributors. History of Science (UK), 11(3): 238, September 1973. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
11/3/238.full.pdf+html.

Anonymous:1973:NCd

[Ano73e] Anonymous. Notes on contributors. History of Science (UK), 11(4): 317, December 1973. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
11/4/317.full.pdf+html.

Anonymous:1973:NBa

[Ano73f] Anonymous. Notices of books. History of Science (UK), 11(1):71– 73, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/1/
71.full.pdf+html.

Anonymous:1973:NBb

[Ano73g] Anonymous. Notices of books. History of Science (UK), 11(2):152– 155, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/2/
152.full.pdf+html.

Anonymous:1973:NBc

[Ano73h] Anonymous. Notices of books. History of Science (UK), 11(4):313– 316, December 1973. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
11/4/313.full.pdf+html.

Anonymous:1973:RCB

[Ano73i] Anonymous. Recently completed British doctoral theses. History of Science (UK), 11(2):156–157, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/11/2/156.full.pdf+html.

Anonymous:1974:CLC

[Ano74a] Anonymous. The Charles Lyell Centenary Symposium. History of Science (UK), 12(2):154, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/12/2/154.full.pdf+html.

Anonymous:1974:EF

[Ano74b] Anonymous. Editorial foreword. History of Science (UK), 12(1): i, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/1/
i.full.pdf+html.

Anonymous:1974:NCa

[Ano74c] Anonymous. Notes on contributors. History of Science (UK), 12(1): 80, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/1/
80.full.pdf+html.

Anonymous:1974:NCb

[Ano74d] Anonymous. Notes on contributors. History of Science (UK), 12(2): 157, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/2/
157.full.pdf+html.

Anonymous:1974:NCc

[Ano74e] Anonymous. Notes on contributors. History of Science (UK), 12(3): 234, September 1974. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
12/3/234.full.pdf+html.

Anonymous:1974:NCd

[Ano74f] Anonymous. Notes on contributors. History of Science (UK), 12(4): 306, December 1974. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
12/4/306.full.pdf+html.

Anonymous:1974:NBa

[Ano74g] Anonymous. Notices of books. History of Science (UK), 12(1):77– 79, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/1/
77.full.pdf+html.

Anonymous:1974:NBb

[Ano74h] Anonymous. Notices of books. History of Science (UK), 12(2):155– 156, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/2/
155.full.pdf+html.

Anonymous:1974:NBc

[Ano74i] Anonymous. Notices of books. History of Science (UK), 12(3):231– 233, September 1974. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
12/3/231.full.pdf+html.

Anonymous:1975:NCa

[Ano75a] Anonymous. Notes on contributors. History of Science (UK), 13(1): 74, March 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/1/
74.full.pdf+html.

Anonymous:1975:NCb

[Ano75b] Anonymous. Notes on contributors. History of Science (UK), 13(2): 152, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/2/
152.full.pdf+html.

Anonymous:1975:NCc

[Ano75c] Anonymous. Notes on contributors. History of Science (UK), 13(3): 232, September 1975. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
13/3/232.full.pdf+html.

Anonymous:1975:NCd

[Ano75d] Anonymous. Notes on contributors. History of Science (UK), 13(4): 306, December 1975. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
13/4/306.full.pdf+html.

Anonymous:1975:NBa

[Ano75e] Anonymous. Notices of books. History of Science (UK), 13(1):72– 74, March 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/1/
72.full.pdf+html.

Anonymous:1975:NBb

[Ano75f] Anonymous. Notices of books. History of Science (UK), 13(2):146– 152, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/2/
146.full.pdf+html.

Anonymous:1975:NBc

[Ano75g] Anonymous. Notices of books. History of Science (UK), 13(3):231– 232, September 1975. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
13/3/231.full.pdf+html.

Anonymous:1975:NBd

[Ano75h] Anonymous. Notices of books. History of Science (UK), 13(4):305, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/4/
305.full.pdf+html.

Anonymous:1976:NCa

[Ano76a] Anonymous. Notes on contributors. History of Science (UK), 14(2): 148, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/14/2/
148.full.pdf+html.

Anonymous:1976:NCb

[Ano76b] Anonymous. Notes on contributors. History of Science (UK), 14(3): 210, September 1976. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
14/3/210.full.pdf+html.

Anonymous:1976:NCc

[Ano76c] Anonymous. Notes on contributors. History of Science (UK), 14(4): 280, December 1976. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
14/4/280.full.pdf+html.

Anonymous:1976:NBa

[Ano76d] Anonymous. Notices of books. History of Science (UK), 14(2):144– 147, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/14/2/
144.full.pdf+html.

Anonymous:1976:NBb

[Ano76e] Anonymous. Notices of books. History of Science (UK), 14(4):277– 279, December 1976. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
14/4/277.full.pdf+html.

Anonymous:1976:SSP

[Ano76f] Anonymous. Some sources and procedures for editing a medieval Arabic surgical tract. History of Science (UK), 14(4):245–264, December 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/14/4/245.full.
pdf+html.

Anonymous:1977:CSA

[Ano77a] Anonymous. Contemporary scientific archives centre. History of Science (UK), 15(3):231–232, September 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/15/3/231.full.pdf+html.

Anonymous:1977:ERN

[Ano77b] Anonymous. Essay review: Newton as alchemist: The foundations of Newton’s alchemy, or “the hunting of the greene lyon”. History of Science (UK), 15(2):102–137, June 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/15/2/102.full.pdf+html.

Anonymous:1977:NCa

[Ano77c] Anonymous. Notes on contributors. History of Science (UK), 15(1): 76, March 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/15/1/
76.full.pdf+html.

Anonymous:1977:NCb

[Ano77d] Anonymous. Notes on contributors. History of Science (UK), 15(2): 154, June 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/15/2/
154.full.pdf+html.

Anonymous:1977:NCc

[Ano77e] Anonymous. Notes on contributors. History of Science (UK), 15(3): 232, September 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/3/232.full.pdf+html.

Anonymous:1977:NCd

[Ano77f] Anonymous. Notes on contributors. History of Science (UK), 15(4): 307, December 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/4/307.full.pdf+html.

Anonymous:1977:NBa

[Ano77g] Anonymous. Notices of books. History of Science (UK), 15(1):73– 75, March 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/15/1/
73.full.pdf+html.

Anonymous:1977:NBb

[Ano77h] Anonymous. Notices of books. History of Science (UK), 15(3):228– 230, September 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/3/228.full.pdf+html.

Anonymous:1977:NBc

[Ano77i] Anonymous. Notices of books. History of Science (UK), 15(4):298– 306, December 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/4/298.full.pdf+html.

Anonymous:1978:NCa

[Ano78a] Anonymous. Notes on contributors. History of Science (UK), 16(1): 76, March 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/16/1/
76.full.pdf+html.

Anonymous:1978:NCb

[Ano78b] Anonymous. Notes on contributors. History of Science (UK), 16(2): 152, June 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/16/2/
152.full.pdf+html.

Anonymous:1978:NCc

[Ano78c] Anonymous. Notes on contributors. History of Science (UK), 16(3): 230, September 1978. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
16/3/230.full.pdf+html.

Anonymous:1978:NCd

[Ano78d] Anonymous. Notes on contributors. History of Science (UK), 16(4): 304, December 1978. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
16/4/304.full.pdf+html.

Anonymous:1978:NB

[Ano78e] Anonymous. Notices of books. History of Science (UK), 16(3):226– 229, September 1978. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
16/3/226.full.pdf+html.

Anonymous:1979:NCa

[Ano79a] Anonymous. Notes on contributors. History of Science (UK), 17(1): 74, March 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/1/
74.full.pdf+html.

Anonymous:1979:NCb

[Ano79b] Anonymous. Notes on contributors. History of Science (UK), 17(2): 150, June 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/2/
150.full.pdf+html.

Anonymous:1979:NCc

[Ano79c] Anonymous. Notes on contributors. History of Science (UK), 17(3): 226, September 1979. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
17/3/226.full.pdf+html.

Anonymous:1979:NCd

[Ano79d] Anonymous. Notes on contributors. History of Science (UK), 17(4): 300, December 1979. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
17/4/300.full.pdf+html.

Anonymous:1979:NBa

[Ano79e] Anonymous. Notices of books. History of Science (UK), 17(1): 73, March 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/1/
73.full.pdf+html.

Anonymous:1979:NBb

[Ano79f] Anonymous. Notices of books. History of Science (UK), 17(3):220– 225, September 1979. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
17/3/220.full.pdf+html.

Anonymous:1979:NBc

[Ano79g] Anonymous. Notices of books. History of Science (UK), 17(4):297– 299, December 1979. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
17/4/297.full.pdf+html.

Anonymous:1980:NCa

[Ano80a] Anonymous. Notes on contributors. History of Science (UK), 18(1): 78, March 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/18/1/
78.full.pdf+html.

Anonymous:1980:NCb

[Ano80b] Anonymous. Notes on contributors. History of Science (UK), 18(2): 154, June 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/18/2/
154.full.pdf+html.

Anonymous:1980:NCc

[Ano80c] Anonymous. Notes on contributors. History of Science (UK), 18(4): 308, December 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/4/308.full.pdf+html.

Anonymous:1980:NBa

[Ano80d] Anonymous. Notices of books. History of Science (UK), 18(1):75– 77, March 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/18/1/
75.full.pdf+html.

Anonymous:1980:NBb

[Ano80e] Anonymous. Notices of books. History of Science (UK), 18(2):151– 154, June 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/18/2/
151.full.pdf+html.

Anonymous:1980:NBc

[Ano80f] Anonymous. Notices of books. History of Science (UK), 18(3):230– 231, September 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/3/230.full.pdf+html.

Anonymous:1980:NBd

[Ano80g] Anonymous. Notices of books. History of Science (UK), 18(4):304– 307, December 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/4/304.full.pdf+html.

Anonymous:1981:NCa

[Ano81a] Anonymous. Notes on contributors. History of Science (UK), 19(1): 74, March 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/19/1/
74.full.pdf+html.

Anonymous:1981:NCb

[Ano81b] Anonymous. Notes on contributors. History of Science (UK), 19(2): 154, June 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/19/2/
154.full.pdf+html.

Anonymous:1981:NCc

[Ano81c] Anonymous. Notes on contributors. History of Science (UK), 19(3): 234, September 1981. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
19/3/234.full.pdf+html.

Anonymous:1981:NCd

[Ano81d] Anonymous. Notes on contributors. History of Science (UK), 19(4): 309, December 1981. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
19/4/309.full.pdf+html.

Anonymous:1981:NBa

[Ano81e] Anonymous. Notices of books. History of Science (UK), 19(1):70– 74, March 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/19/1/
70.full.pdf+html.

Anonymous:1981:NBb

[Ano81f] Anonymous. Notices of books. History of Science (UK), 19(3):226– 233, September 1981. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
19/3/226.full.pdf+html.

Anonymous:1981:NBc

[Ano81g] Anonymous. Notices of books. History of Science (UK), 19(4):308, December 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/19/4/
308.full.pdf+html.

Anonymous:1982:NCa

[Ano82a] Anonymous. Notes on contributors. History of Science (UK), 20(1): 75, March 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/20/1/
75.full.pdf+html.

Anonymous:1982:NCb

[Ano82b] Anonymous. Notes on contributors. History of Science (UK), 20(2): 156, June 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/20/2/
156.full.pdf+html.

Anonymous:1982:NCc

[Ano82c] Anonymous. Notes on contributors. History of Science (UK), 20(3): 234, September 1982. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
20/3/234.full.pdf+html.

Anonymous:1982:NCd

[Ano82d] Anonymous. Notes on contributors. History of Science (UK), 20(4): 312, December 1982. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
20/4/312.full.pdf+html.

Anonymous:1982:NBa

[Ano82e] Anonymous. Notices of books. History of Science (UK), 20(2):148– 155, June 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/20/2/
148.full.pdf+html.

Anonymous:1982:NBb

[Ano82f] Anonymous. Notices of books. History of Science (UK), 20(3):233– 234, September 1982. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
20/3/233.full.pdf+html.

Anonymous:1982:NBc

[Ano82g] Anonymous. Notices of books. History of Science (UK), 20(4):310– 311, December 1982. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
20/4/310.full.pdf+html.

Anonymous:1983:NCa

[Ano83a] Anonymous. Notes on contributors. History of Science (UK), 21(1): 110, March 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/1/
110.full.pdf+html.

Anonymous:1983:NCb

[Ano83b] Anonymous. Notes on contributors. History of Science (UK), 21(2): 222, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/2/
222.full.pdf+html.

Anonymous:1983:NCc

[Ano83c] Anonymous. Notes on contributors. History of Science (UK), 21(3): 334, September 1983. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
21/3/334.full.pdf+html.

Anonymous:1983:NCd

[Ano83d] Anonymous. Notes on contributors. History of Science (UK), 21(4): 444, December 1983. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
21/4/444.full.pdf+html.

Anonymous:1983:NBa

[Ano83e] Anonymous. Notices of books. History of Science (UK), 21(1): 109, March 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/1/
109.full.pdf+html.

Anonymous:1983:NBb

[Ano83f] Anonymous. Notices of books. History of Science (UK), 21(2):215– 221, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/2/
215.full.pdf+html.

Anonymous:1983:NBc

[Ano83g] Anonymous. Notices of books. History of Science (UK), 21(3):333– 334, September 1983. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
21/3/333.full.pdf+html.

Anonymous:1983:NBd

[Ano83h] Anonymous. Notices of books. History of Science (UK), 21(4):434– 443, December 1983. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
21/4/434.full.pdf+html.

Anonymous:1984:AP

[Ano84a] Anonymous. The Author’s preface. History of Science (UK), 22(1): 47–58, March 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/1/
47.full.pdf+html.

Anonymous:1984:NBd

[Ano84b] Anonymous. Notes of books. History of Science (UK), 22(4):428–431, December 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/4/
428.full.pdf+html.

Anonymous:1984:NCa

[Ano84c] Anonymous. Notes on contributors. History of Science (UK), 22(1): 110, March 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/1/
110.full.pdf+html.

Anonymous:1984:NCb

[Ano84d] Anonymous. Notes on contributors. History of Science (UK), 22(2): 216, June 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/2/
216.full.pdf+html.

Anonymous:1984:NCc

[Ano84e] Anonymous. Notes on contributors. History of Science (UK), 22(3): 324, September 1984. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
22/3/324.full.pdf+html.

Anonymous:1984:NCd

[Ano84f] Anonymous. Notes on contributors. History of Science (UK), 22(4): 432, December 1984. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
22/4/432.full.pdf+html.

Anonymous:1984:NBa

[Ano84g] Anonymous. Notices of books. History of Science (UK), 22(1):98– 110, March 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/1/
98.full.pdf+html.

Anonymous:1984:NBb

[Ano84h] Anonymous. Notices of books. History of Science (UK), 22(2):210– 216, June 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/2/
210.full.pdf+html.

Anonymous:1984:NBc

[Ano84i] Anonymous. Notices of books. History of Science (UK), 22(3):321– 324, September 1984. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
22/3/321.full.pdf+html.

Anonymous:1985:C

[Ano85a] Anonymous. Corrigendum. History of Science (UK), 23(4):444, December 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL http://hos.sagepub.com/content/23/4/444.2.
full.pdf+html.

Anonymous:1985:NCa

[Ano85b] Anonymous. Notes on contributors. History of Science (UK), 23(1): 126, March 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/1/
126.full.pdf+html.

Anonymous:1985:NCb

[Ano85c] Anonymous. Notes on contributors. History of Science (UK), 23(2): 222, June 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/2/
222.full.pdf+html.

Anonymous:1985:NCc

[Ano85d] Anonymous. Notes on contributors. History of Science (UK), 23(3): 334, September 1985. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
23/3/334.full.pdf+html.

Anonymous:1985:NCd

[Ano85e] Anonymous. Notes on contributors. History of Science (UK), 23(4): 444, December 1985. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
23/4/444.1.full.pdf+html.

Anonymous:1985:NBa

[Ano85f] Anonymous. Notices of books. History of Science (UK), 23(2):211– 221, June 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/2/
211.full.pdf+html.

Anonymous:1985:NBb

[Ano85g] Anonymous. Notices of books. History of Science (UK), 23(3):326– 334, September 1985. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
23/3/326.full.pdf+html.

Anonymous:1985:NBc

[Ano85h] Anonymous. Notices of books. History of Science (UK), 23(4):440– 443, December 1985. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
23/4/440.full.pdf+html.

Anonymous:1986:NCa

[Ano86a] Anonymous. Notes on contributors. History of Science (UK), 24(1): 110, March 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/1/
110.full.pdf+html.

Anonymous:1986:NCb

[Ano86b] Anonymous. Notes on contributors. History of Science (UK), 24(2): 222, June 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/2/
222.full.pdf+html.

Anonymous:1986:NCc

[Ano86c] Anonymous. Notes on contributors. History of Science (UK), 24(3): 334, September 1986. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
24/3/334.full.pdf+html.

Anonymous:1986:NCd

[Ano86d] Anonymous. Notes on contributors. History of Science (UK), 24(4): 444, December 1986. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
24/4/444.full.pdf+html.

Anonymous:1986:NBa

[Ano86e] Anonymous. Notices of books. History of Science (UK), 24(1):109– 110, March 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/1/
109.full.pdf+html.

Anonymous:1986:NBb

[Ano86f] Anonymous. Notices of books. History of Science (UK), 24(2):216– 221, June 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/2/
216.full.pdf+html.

Anonymous:1986:NBc

[Ano86g] Anonymous. Notices of books. History of Science (UK), 24(3):329– 334, September 1986. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
24/3/329.full.pdf+html.

Anonymous:1986:NBd

[Ano86h] Anonymous. Notices of books. History of Science (UK), 24(4):438– 443, December 1986. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
24/4/438.full.pdf+html.

Anonymous:1987:NCa

[Ano87a] Anonymous. Notes on contributors. History of Science (UK), 25(1): 110, March 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/25/1/
110.full.pdf+html.

Anonymous:1987:NCb

[Ano87b] Anonymous. Notes on contributors. History of Science (UK), 25(2): 222, June 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/25/2/
222.full.pdf+html.

Anonymous:1987:NCc

[Ano87c] Anonymous. Notes on contributors. History of Science (UK), 25(3): 334, September 1987. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
25/3/334.full.pdf+html.

Anonymous:1987:NCd

[Ano87d] Anonymous. Notes on contributors. History of Science (UK), 25(4): 424, December 1987. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
25/4/424.full.pdf+html.

Anonymous:1987:NBa

[Ano87e] Anonymous. Notices of books. History of Science (UK), 25(2):217– 221, June 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/25/2/
217.full.pdf+html.

Anonymous:1987:NBb

[Ano87f] Anonymous. Notices of books. History of Science (UK), 25(3):327– 333, September 1987. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
25/3/327.full.pdf+html.

Anonymous:1988:NCa

[Ano88a] Anonymous. Notes on contributors. History of Science (UK), 26(1): 110, March 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/26/1/
110.full.pdf+html.

Anonymous:1988:NCb

[Ano88b] Anonymous. Notes on contributors. History of Science (UK), 26(2): 222, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/26/2/
222.full.pdf+html.

Anonymous:1988:NCc

[Ano88c] Anonymous. Notes on contributors. History of Science (UK), 26(3): 332, September 1988. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
26/3/332.full.pdf+html.

Anonymous:1988:NCd

[Ano88d] Anonymous. Notes on contributors. History of Science (UK), 26(4): 442, December 1988. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
26/4/442.full.pdf+html.

Anonymous:1988:NB

[Ano88e] Anonymous. Notices of books. History of Science (UK), 26(2):216– 221, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/26/2/
216.full.pdf+html.

Anonymous:1989:C

[Ano89a] Anonymous. Correction. History of Science (UK), 27(1):102, March 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/27/1/102.full.
pdf+html.

Anonymous:1989:NCa

[Ano89b] Anonymous. Notes on contributors. History of Science (UK), 27(1): 110, March 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/27/1/
110.full.pdf+html.

Anonymous:1989:NCb

[Ano89c] Anonymous. Notes on contributors. History of Science (UK), 27(2): 222, June 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/27/2/
222.full.pdf+html.

Anonymous:1989:NCc

[Ano89d] Anonymous. Notes on contributors. History of Science (UK), 27(3): 318, September 1989. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
27/3/318.full.pdf+html.

Anonymous:1989:NCd

[Ano89e] Anonymous. Notes on contributors. History of Science (UK), 27(4): 449, December 1989. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
27/4/449.full.pdf+html.

Anonymous:1989:NBa

[Ano89f] Anonymous. Notices of books. History of Science (UK), 27(1):106– 109, March 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/27/1/
106.full.pdf+html.

Anonymous:1989:NBb

[Ano89g] Anonymous. Notices of books. History of Science (UK), 27(2):219– 221, June 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/27/2/
219.full.pdf+html.

Anonymous:1990:NCa

[Ano90a] Anonymous. Notes on contributors. History of Science (UK), 28(1): 124, March 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/28/1/
124.full.pdf+html.

Anonymous:1990:NCb

[Ano90b] Anonymous. Notes on contributors. History of Science (UK), 28(2): 220, June 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/28/2/
220.full.pdf+html.

Anonymous:1990:NCc

[Ano90c] Anonymous. Notes on contributors. History of Science (UK), 28(3): 332, September 1990. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
28/3/332.full.pdf+html.

Anonymous:1990:NCd

[Ano90d] Anonymous. Notes on contributors. History of Science (UK), 28(4): 442, December 1990. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
28/4/442.full.pdf+html.

Anonymous:1991:NCa

[Ano91a] Anonymous. Notes on contributors. History of Science (UK), 29(1): 110, March 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/29/1/
110.full.pdf+html.

Anonymous:1991:NCb

[Ano91b] Anonymous. Notes on contributors. History of Science (UK), 29(2): 222, June 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/29/2/
222.full.pdf+html.

Anonymous:1991:NCc

[Ano91c] Anonymous. Notes on contributors. History of Science (UK), 29(3): 334, September 1991. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
29/3/334.full.pdf+html.

Anonymous:1991:NCd

[Ano91d] Anonymous. Notes on contributors. History of Science (UK), 29(4): 444, December 1991. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
29/4/444.full.pdf+html.

Anonymous:1991:NBa

[Ano91e] Anonymous. Notices of books. History of Science (UK), 29(1):108– 109, March 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/29/1/
108.full.pdf+html.

Anonymous:1991:NBb

[Ano91f] Anonymous. Notices of books. History of Science (UK), 29(3):332– 334, September 1991. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
29/3/332.full.pdf+html.

Anonymous:1991:NBc

[Ano91g] Anonymous. Notices of books. History of Science (UK), 29(4):435– 443, December 1991. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
29/4/435.full.pdf+html.

Anonymous:1992:NCa

[Ano92a] Anonymous. Notes on contributors. History of Science (UK), 30(1): 118, March 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/1/
118.full.pdf+html.

Anonymous:1992:NCb

[Ano92b] Anonymous. Notes on contributors. History of Science (UK), 30(2): 219, June 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/2/
219.full.pdf+html.

Anonymous:1992:NCc

[Ano92c] Anonymous. Notes on contributors. History of Science (UK), 30(3): 332, September 1992. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
30/3/332.full.pdf+html.

Anonymous:1992:NCd

[Ano92d] Anonymous. Notes on contributors. History of Science (UK), 30(4): 442, December 1992. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
30/4/442.full.pdf+html.

Anonymous:1992:NBa

[Ano92e] Anonymous. Notices of books. History of Science (UK), 30(1):115– 117, March 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/1/
115.full.pdf+html.

Anonymous:1992:NBb

[Ano92f] Anonymous. Notices of books. History of Science (UK), 30(3):329– 331, September 1992. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
30/3/329.full.pdf+html.

Anonymous:1993:NCa

[Ano93a] Anonymous. Notes on contributors. History of Science (UK), 31(1): 98, March 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/31/1/
98.full.pdf+html.

Anonymous:1993:NCb

[Ano93b] Anonymous. Notes on contributors. History of Science (UK), 31(2): 226, June 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/31/2/
226.full.pdf+html.

Anonymous:1993:NCc

[Ano93c] Anonymous. Notes on contributors. History of Science (UK), 31(3): 334, September 1993. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
31/3/334.full.pdf+html.

Anonymous:1993:NCd

[Ano93d] Anonymous. Notes on contributors. History of Science (UK), 31(4): 441, December 1993. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
31/4/441.full.pdf+html.

Anonymous:1994:NCa

[Ano94a] Anonymous. Notes on contributors. History of Science (UK), 32(1): 110, March 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/32/1/
110.full.pdf+html.

Anonymous:1994:NCb

[Ano94b] Anonymous. Notes on contributors. History of Science (UK), 32(2): 236, June 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/32/2/
236.full.pdf+html.

Anonymous:1994:NCc

[Ano94c] Anonymous. Notes on contributors. History of Science (UK), 32(3): 360, September 1994. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
32/3/360.full.pdf+html.

Anonymous:1994:NCd

[Ano94d] Anonymous. Notes on contributors. History of Science (UK), 32(4): 486–488, December 1994. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
32/4/486.full.pdf+html.

Anonymous:1995:NCa

[Ano95a] Anonymous. Notes on contributors. History of Science (UK), 33(1): 124, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/1/
124.full.pdf+html.

Anonymous:1995:NCb

[Ano95b] Anonymous. Notes on contributors. History of Science (UK), 33(2): 252, June 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/2/
252.full.pdf+html.

Anonymous:1995:NCc

[Ano95c] Anonymous. Notes on contributors. History of Science (UK), 33(3): 374, September 1995. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
33/3/374.full.pdf+html.

Anonymous:1995:NCd

[Ano95d] Anonymous. Notes on contributors. History of Science (UK), 33(4): 497, December 1995. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
33/4/497.full.pdf+html.

Anonymous:1995:NBa

[Ano95e] Anonymous. Notices of books. History of Science (UK), 33(1):119– 123, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/1/
119.full.pdf+html.

Anonymous:1995:NBb

[Ano95f] Anonymous. Notices of books. History of Science (UK), 33(3):372– 373, September 1995. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
33/3/372.full.pdf+html.

Anonymous:1996:NCa

[Ano96a] Anonymous. Notes on contributors. History of Science (UK), 34(1): 122, March 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/34/1/
122.full.pdf+html.

Anonymous:1996:NCb

[Ano96b] Anonymous. Notes on contributors. History of Science (UK), 34(2): 249, June 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/34/2/
249.full.pdf+html.

Anonymous:1996:NCc

[Ano96c] Anonymous. Notes on contributors. History of Science (UK), 34(3): 378, September 1996. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
34/3/378.full.pdf+html.

Anonymous:1996:NCd

[Ano96d] Anonymous. Notes on contributors. History of Science (UK), 34(4): 504, December 1996. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
34/4/504.full.pdf+html.

Anonymous:1997:NCa

[Ano97a] Anonymous. Notes on contributors. History of Science (UK), 35(1): 120, March 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/35/1/
120.full.pdf+html.

Anonymous:1997:NCb

[Ano97b] Anonymous. Notes on contributors. History of Science (UK), 35(2): 244, June 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/35/2/
244.full.pdf+html.

Anonymous:1997:NCc

[Ano97c] Anonymous. Notes on contributors. History of Science (UK), 35(3): 366, September 1997. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
35/3/366.full.pdf+html.

Anonymous:1997:NCd

[Ano97d] Anonymous. Notes on contributors. History of Science (UK), 35(4): 491, December 1997. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
35/4/491.full.pdf+html.

Anonymous:1998:NCa

[Ano98a] Anonymous. Notes on contributors. History of Science (UK), 36(1): 122, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/36/1/
122.full.pdf+html.

Anonymous:1998:NCb

[Ano98b] Anonymous. Notes on contributors. History of Science (UK), 36(2): 244, June 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/36/2/
244.full.pdf+html.

Anonymous:1998:NCc

[Ano98c] Anonymous. Notes on contributors. History of Science (UK), 36(3): 372, September 1998. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
36/3/372.full.pdf+html.

Anonymous:1998:NCd

[Ano98d] Anonymous. Notes on contributors. History of Science (UK), 36(4): 498–499, December 1998. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
36/4/498.full.pdf+html.

Anonymous:1998:NBa

[Ano98e] Anonymous. Notices of books. History of Science (UK), 36(2):233– 243, June 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/36/2/
233.full.pdf+html.

Anonymous:1998:NBb

[Ano98f] Anonymous. Notices of books. History of Science (UK), 36(3):369– 371, September 1998. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
36/3/369.full.pdf+html.

Anonymous:1999:NCa

[Ano99a] Anonymous. Notes on contributors. History of Science (UK), 37(1): 122, March 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/1/
122.full.pdf+html.

Anonymous:1999:NCb

[Ano99b] Anonymous. Notes on contributors. History of Science (UK), 37(2): 248, June 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/2/
248.full.pdf+html.

Anonymous:1999:NCc

[Ano99c] Anonymous. Notes on contributors. History of Science (UK), 37(3): 376, September 1999. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
37/3/376.full.pdf+html.

Anonymous:1999:NCd

[Ano99d] Anonymous. Notes on contributors. History of Science (UK), 37(4): 503, December 1999. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
37/4/503.full.pdf+html.

Anonymous:1999:NBa

[Ano99e] Anonymous. Notices of books. History of Science (UK), 37(1):116– 121, March 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/1/
116.full.pdf+html.

Anonymous:1999:NBb

[Ano99f] Anonymous. Notices of books. History of Science (UK), 37(2):237– 247, June 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/2/
237.full.pdf+html.

Anonymous:1999:NBc

[Ano99g] Anonymous. Notices of books. History of Science (UK), 37(3):365– 376, September 1999. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
37/3/365.full.pdf+html.

Anonymous:1999:NBd

[Ano99h] Anonymous. Notices of books. History of Science (UK), 37(4):490– 502, December 1999. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
37/4/490.full.pdf+html.

Anonymous:2000:NCa

[Ano00a] Anonymous. Notes on contributors. History of Science (UK), 38(1): 126, March 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/1/
126.full.pdf+html.

Anonymous:2000:NCb

[Ano00b] Anonymous. Notes on contributors. History of Science (UK), 38(2): 249–250, June 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/2/
249.full.pdf+html.

Anonymous:2000:NCc

[Ano00c] Anonymous. Notes on contributors. History of Science (UK), 38(3): 362, September 2000. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
38/3/362.full.pdf+html.

Anonymous:2000:NCd

[Ano00d] Anonymous. Notes on contributors. History of Science (UK), 38(4): 489, December 2000. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
38/4/489.full.pdf+html.

Anonymous:2000:NBa

[Ano00e] Anonymous. Notices of books. History of Science (UK), 38(1):115– 125, March 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/1/
115.full.pdf+html.

Anonymous:2000:NBb

[Ano00f] Anonymous. Notices of books. History of Science (UK), 38(3):360– 361, September 2000. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
38/3/360.full.pdf+html.

Anonymous:2001:NCa

[Ano01a] Anonymous. Notes on contributors. History of Science (UK), 39(1): 126, March 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/39/1/
126.full.pdf+html.

Anonymous:2001:NCb

[Ano01b] Anonymous. Notes on contributors. History of Science (UK), 39(2): 254, June 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/39/2/
254.full.pdf+html.

Anonymous:2001:NCc

[Ano01c] Anonymous. Notes on contributors. History of Science (UK), 39(3): 382, September 2001. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
39/3/382.full.pdf+html.

Anonymous:2001:NCd

[Ano01d] Anonymous. Notes on contributors. History of Science (UK), 39(4): 507, December 2001. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
39/4/507.full.pdf+html.

Anonymous:2001:NBa

[Ano01e] Anonymous. Notices of books. History of Science (UK), 39(1):121– 125, March 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/39/1/
121.full.pdf+html.

Anonymous:2001:NBb

[Ano01f] Anonymous. Notices of books. History of Science (UK), 39(2):249– 253, June 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/39/2/
249.full.pdf+html.

Anonymous:2001:NBc

[Ano01g] Anonymous. Notices of books. History of Science (UK), 39(4):501– 506, December 2001. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
39/4/501.full.pdf+html.

Anonymous:2002:NCa

[Ano02a] Anonymous. Notes on contributors. History of Science (UK), 40(1): 124, March 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/40/1/
124.full.pdf+html.

Anonymous:2002:NCb

[Ano02b] Anonymous. Notes on contributors. History of Science (UK), 40(2): 250, June 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/40/2/
250.full.pdf+html.

Anonymous:2002:NCc

[Ano02c] Anonymous. Notes on contributors. History of Science (UK), 40(3): 376, September 2002. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
40/3/376.full.pdf+html.

Anonymous:2002:NCd

[Ano02d] Anonymous. Notes on contributors. History of Science (UK), 40(4): 501, December 2002. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
40/4/501.full.pdf+html.

Anonymous:2002:NBa

[Ano02e] Anonymous. Notices of books. History of Science (UK), 40(1):113– 123, March 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/40/1/
113.full.pdf+html.

Anonymous:2002:NBb

[Ano02f] Anonymous. Notices of books. History of Science (UK), 40(2):245– 249, June 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/40/2/
245.full.pdf+html.

Anonymous:2002:NBc

[Ano02g] Anonymous. Notices of books. History of Science (UK), 40(3):369– 375, September 2002. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
40/3/369.full.pdf+html.

Anonymous:2003:NCa

[Ano03a] Anonymous. Notes on contributors. History of Science (UK), 41(1): 124, March 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/41/1/
124.full.pdf+html.

Anonymous:2003:NCb

[Ano03b] Anonymous. Notes on contributors. History of Science (UK), 41(2): 248, June 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/41/2/
248.full.pdf+html.

Anonymous:2003:NCc

[Ano03c] Anonymous. Notes on contributors. History of Science (UK), 41(3): 369–370, September 2003. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
41/3/369.full.pdf+html.

Anonymous:2003:NCd

[Ano03d] Anonymous. Notes on contributors. History of Science (UK), 41(4): 472, December 2003. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
41/4/472.full.pdf+html.

Anonymous:2003:NBa

[Ano03e] Anonymous. Notices of books. History of Science (UK), 41(1):120– 123, March 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/41/1/
120.full.pdf+html.

Anonymous:2003:NBb

[Ano03f] Anonymous. Notices of books. History of Science (UK), 41(2):245– 247, June 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/41/2/
245.full.pdf+html.

Anonymous:2003:NBc

[Ano03g] Anonymous. Notices of books. History of Science (UK), 41(3):355– 368, September 2003. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
41/3/355.full.pdf+html.

Anonymous:2004:NCa

[Ano04a] Anonymous. Notes on contributors. History of Science (UK), 42(1): 136, March 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/42/1/
136.full.pdf+html.

Anonymous:2004:NCb

[Ano04b] Anonymous. Notes on contributors. History of Science (UK), 42(2): 260, June 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/42/2/
260.full.pdf+html.

Anonymous:2004:NCc

[Ano04c] Anonymous. Notes on contributors. History of Science (UK), 42(3): 370, September 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/3/370.full.pdf+html.

Anonymous:2004:NCd

[Ano04d] Anonymous. Notes on contributors. History of Science (UK), 42(4): 495, December 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/4/495.full.pdf+html.

Anonymous:2004:NBa

[Ano04e] Anonymous. Notices of books. History of Science (UK), 42(1):133– 135, March 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/42/1/
133.full.pdf+html.

Anonymous:2004:NBb

[Ano04f] Anonymous. Notices of books. History of Science (UK), 42(2):258– 259, June 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/42/2/
258.full.pdf+html.

Anonymous:2004:NBc

[Ano04g] Anonymous. Notices of books. History of Science (UK), 42(3):367– 369, September 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/3/367.full.pdf+html.

Anonymous:2005:NCa

[Ano05a] Anonymous. Notes on contributors. History of Science (UK), 43(1): 108, March 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/43/1/
108.full.pdf+html.

Anonymous:2005:NCb

[Ano05b] Anonymous. Notes on contributors. History of Science (UK), 43(2): 216, June 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/43/2/
216.full.pdf+html.

Anonymous:2005:NCc

[Ano05c] Anonymous. Notes on contributors. History of Science (UK), 43(3): 342, September 2005. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
43/3/342.full.pdf+html.

Anonymous:2005:NCd

[Ano05d] Anonymous. Notes on contributors. History of Science (UK), 43(4): 467, December 2005. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
43/4/467.full.pdf+html.

Anonymous:2005:NBa

[Ano05e] Anonymous. Notices of books. History of Science (UK), 43(1): 107, March 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/43/1/
107.full.pdf+html.

Anonymous:2005:NBb

[Ano05f] Anonymous. Notices of books. History of Science (UK), 43(2):209– 215, June 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/43/2/
209.full.pdf+html.

Anonymous:2006:NCa

[Ano06a] Anonymous. Notes on contributors. History of Science (UK), 44(1): 122, March 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/44/1/
122.full.pdf+html.

Anonymous:2006:NCb

[Ano06b] Anonymous. Notes on contributors. History of Science (UK), 44(2): 279, June 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/44/2/
279.full.pdf+html.

Anonymous:2006:NCc

[Ano06c] Anonymous. Notes on contributors. History of Science (UK), 44(3): 372, September 2006. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
44/3/372.full.pdf+html.

Anonymous:2006:NCd

[Ano06d] Anonymous. Notes on contributors. History of Science (UK), 44(4): 497, December 2006. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
44/4/497.full.pdf+html.

Anonymous:2006:NB

[Ano06e] Anonymous. Notices of books. History of Science (UK), 44(1):119– 121, March 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/44/1/
119.full.pdf+html.

Anonymous:2007:NCa

[Ano07a] Anonymous. Notes on contributors. History of Science (UK), 45(1): 122, March 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/1/
122.full.pdf+html.

Anonymous:2007:NCb

[Ano07b] Anonymous. Notes on contributors. History of Science (UK), 45(2): 228–229, June 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/2/
228.full.pdf+html.

Anonymous:2007:NCc

[Ano07c] Anonymous. Notes on contributors. History of Science (UK), 45(3): 356, September 2007. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
45/3/356.full.pdf+html.

Anonymous:2007:NCd

[Ano07d] Anonymous. Notes on contributors. History of Science (UK), 45(4): 481, December 2007. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
45/4/481.full.pdf+html.

Anonymous:2007:NBa

[Ano07e] Anonymous. Notices of books. History of Science (UK), 45(1):119– 121, March 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/1/
119.full.pdf+html.

Anonymous:2007:NBb

[Ano07f] Anonymous. Notices of books. History of Science (UK), 45(2):223– 228, June 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/2/
223.full.pdf+html.

Anonymous:2008:NCa

[Ano08a] Anonymous. Notes on contributors. History of Science (UK), 46(1): 122, March 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/46/1/
122.full.pdf+html.

Anonymous:2008:NCb

[Ano08b] Anonymous. Notes on contributors. History of Science (UK), 46(2): 247–248, June 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/46/2/
247.full.pdf+html.

Anonymous:2008:NCc

[Ano08c] Anonymous. Notes on contributors. History of Science (UK), 46(3): 373, September 2008. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
46/3/373.full.pdf+html.

Anonymous:2008:NCd

[Ano08d] Anonymous. Notes on contributors. History of Science (UK), 46(4): 499, December 2008. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
46/4/499.full.pdf+html.

Anonymous:2008:NBa

[Ano08e] Anonymous. Notices of books. History of Science (UK), 46(1):115– 121, March 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/46/1/
115.full.pdf+html.

Anonymous:2008:NBb

[Ano08f] Anonymous. Notices of books. History of Science (UK), 46(2):245– 246, June 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/46/2/
245.full.pdf+html.

Anonymous:2008:NBc

[Ano08g] Anonymous. Notices of books. History of Science (UK), 46(3):371– 372, September 2008. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
46/3/371.full.pdf+html.

Anonymous:2008:NBd

[Ano08h] Anonymous. Notices of books. History of Science (UK), 46(4):497– 498, December 2008. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
46/4/497.full.pdf+html.

Anonymous:2009:JUT

[Ano09a] Anonymous. Journals under threat: a joint response from history of science, technology and medicine Editors. History of Science (UK), 47(1):v–viii, March 2009. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
47/1/v.full.pdf+html.

Anonymous:2009:NCa

[Ano09b] Anonymous. Notes on contributors. History of Science (UK), 47(1): 116, March 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/1/
116.full.pdf+html.

Anonymous:2009:NCb

[Ano09c] Anonymous. Notes on contributors. History of Science (UK), 47(2): 242, June 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/2/
242.full.pdf+html.

Anonymous:2009:NCc

[Ano09d] Anonymous. Notes on contributors. History of Science (UK), 47(3): 368, September 2009. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
47/3/368.full.pdf+html.

Anonymous:2009:NCd

[Ano09e] Anonymous. Notes on contributors. History of Science (UK), 47(4): 393, December 2009. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
47/4/393.full.pdf+html.

Anonymous:2009:NBa

[Ano09f] Anonymous. Notices of books. History of Science (UK), 47(1):114– 115, March 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/1/
114.full.pdf+html.

Anonymous:2009:NBb

[Ano09g] Anonymous. Notices of books. History of Science (UK), 47(2):233– 241, June 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/2/
233.full.pdf+html.

Anonymous:2009:NBc

[Ano09h] Anonymous. Notices of books. History of Science (UK), 47(3):367, September 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/3/
367.full.pdf+html.

Anonymous:2010:NCa

[Ano10a] Anonymous. Notes on contributors. History of Science (UK), 48(1): 124, March 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/48/1/
124.full.pdf+html.

Anonymous:2010:NCb

[Ano10b] Anonymous. Notes on contributors. History of Science (UK), 48(2): 250, June 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/48/2/
250.full.pdf+html.

Anonymous:2010:NCc

[Ano10c] Anonymous. Notes on contributors. History of Science (UK), 48 (3–4):499–500, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Anonymous:2010:NBa

[Ano10d] Anonymous. Notices of books. History of Science (UK), 48(1): 123, March 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/48/1/
123.full.pdf+html.

Anonymous:2010:NBb

[Ano10e] Anonymous. Notices of books. History of Science (UK), 48(3–4):495– 498, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Anonymous:2011:ERC

[Ano11a] Anonymous. Essay review: The Chinese astronomical reform of 1280, granting the seasons: The Chinese astronomical reform of 1280, with a study of its many dimensions and an annotated translation of its records. History of Science (UK), 49(1):109–114, March 2011. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/49/1/109.full.pdf+html.

Anonymous:2011:NCa

[Ano11b] Anonymous. Notes on contributors. History of Science (UK), 49(1): 124, March 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/49/1/
124.full.pdf+html.

Anonymous:2011:NCb

[Ano11c] Anonymous. Notes on contributors. History of Science (UK), 49(2): 234, June 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/49/2/
234.full.pdf+html.

Anonymous:2011:NCc

[Ano11d] Anonymous. Notes on contributors. History of Science (UK), 49(3): 375–376, September 2011. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
49/3/375.full.pdf+html.

Anonymous:2011:NBa

[Ano11e] Anonymous. Notices of books. History of Science (UK), 49(1):121– 123, March 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/49/1/
121.full.pdf+html.

Anonymous:2011:NBb

[Ano11f] Anonymous. Notices of books. History of Science (UK), 49(2):232– 233, June 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/49/2/
232.full.pdf+html.

Anonymous:2011:NBc

[Ano11g] Anonymous. Notices of books. History of Science (UK), 49(4):479– 484, December 2011. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
49/4/479.full.pdf+html.

Anonymous:2012:NCa

[Ano12a] Anonymous. Notes on contributors. History of Science (UK), 50(1): 123, March 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/50/1/
123.full.pdf+html.

Anonymous:2012:NCb

[Ano12b] Anonymous. Notes on contributors. History of Science (UK), 50(2): 250, June 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/50/2/
250.full.pdf+html.

Anonymous:2012:NCc

[Ano12c] Anonymous. Notes on contributors. History of Science (UK), 50(3): 375, September 2012. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
50/3/375.full.pdf+html.

Anonymous:2012:NCd

[Ano12d] Anonymous. Notes on contributors. History of Science (UK), 50(4): 501, December 2012. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
50/4/501.full.pdf+html.

Anonymous:2013:NCa

[Ano13a] Anonymous. Notes on contributors. History of Science (UK), 51(1): 124, March 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/51/1/
124.full.pdf+html.

Anonymous:2013:NCb

[Ano13b] Anonymous. Notes on contributors. History of Science (UK), 51(2): 250, June 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/51/2/
250.full.pdf+html.

Anonymous:2013:NCc

[Ano13c] Anonymous. Notes on contributors. History of Science (UK), 51(3): 375, September 2013. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
51/3/375.full.pdf+html.

Anonymous:2013:NCd

[Ano13d] Anonymous. Notes on contributors. History of Science (UK), 51(4): 501, December 2013. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
51/4/501.full.pdf+html.

Anonymous:2014:NC

[Ano14] Anonymous. Notes on contributors. History of Science (UK), 52(1): 124, March 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/52/1/
124.full.pdf+html.

Anonymous:2020:CST

[Ano20] Anonymous. Corrigendum to “Science and terminology in-between empires: Ukrainian science in a search for its language in the nineteenth century”. History of Science (UK), 58(1):NP1, March 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275319900876. See [Sur19].

Anonymous:2021:CDM

[Ano21] Anonymous. Corrigendum to “Defending metropolitan identity through colonial politics: The role of Portuguese naturalists (1870– 91)”. History of Science (UK), 59(4):522, December 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
00732753211051476. See [GM18].

Applebaum:1996:KAA

[App96] Wilbur Applebaum. Keplerian astronomy after Kepler: Researches and problems. History of Science (UK), 34(4):451–504, December 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/34/4/451.full.
pdf+html.

Andretta:2020:BPH

[APT20] Elisa Andretta and José Pardo-Tomás. Books, plants, herbaria: Diego Hurtado de Mendoza and his circle in Italy (1539–1554). History of Science (UK), 58(1):3–27, March 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275319838891.

Arapostathis:2019:MLB

[Ara19] Stathis Arapostathis. Marconi’s legal battles: Discursive, textual, and material entanglements. History of Science (UK), 57(1):97–118, March 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275318801394.

Arend:2019:SCE

[Are19] Jan Arend. “Simple, clear, and easily understood by the farmer … ”: On expert-layman communication in American soil science, 1920s–50s. History of Science (UK), 57(3):324–345, September 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318813465.

Armitage:1962:BRC

[Arm62] A. Armitage. Book review: Changing views of the universe. History of Science (UK), 1(1):118–119, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/118.full.pdf+html.

Ashworth:1998:JHG

[Ash98] William J. Ashworth. John Herschel, George Airy, and the roaming eye of the state. History of Science (UK), 36(2):151–178, June 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/36/2/151.full.
pdf+html.

Ash:2001:QVN

[Ash01] Eric H. Ash. Queen v. Northumberland, and the control of technical expertise. History of Science (UK), 39(2):215–240, June 2001. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/39/2/215.full.pdf+html.

Ashworth:2008:GRS

[Ash08] William J. Ashworth. The ghost of Rostow: Science, culture and the British industrial revolution. History of Science (UK), 46(3):249–274, September 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/46/3/
249.full.pdf+html.

Ashworth:2014:BIR

[Ash14] William J. Ashworth. The British industrial revolution and the ideological revolution: Science, neoliberalism and history. History of Science (UK), 52(2):178–199, June 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/52/2/178.full.pdf+html.

Ait-Touati:2012:GMT

[AT12] Frédérique Aït-Touati. “Give me a telescope and i Shall move the Earth”: Hooke’s attempt to prove the motion of the Earth from observations. History of Science (UK), 50(1):75–91, March 2012. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/50/1/75.full.pdf+html.

Baldwin:1989:ASP

[Bal89] N. J. Baldwin. The archive of the society for the protection of science and learning. History of Science (UK), 27(1):103–105, March 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/27/1/103.full.
pdf+html.

Baldwin:2012:SGN

[Bal12] Melinda Baldwin. The shifting ground of Nature: Establishing an organ of scientific communication in Britain, 1869–1900. History of Science (UK), 50(2):125–154, June 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/50/2/125.full.pdf+html.

Barber:1964:BRL

[Bar64] W. H. Barber. Book review: The language of science and the language of literature, 1700–1740. History of Science (UK), 3(1): 149–150, March 1964. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/3/
1/149.full.pdf+html.

Bartholomew:1979:ERD

[Bar79a] Michael Bartholomew. Essay review: Darwin’s teacher: Henslow of Hitcham: Botanist, educationalist and clergyman. History of Science (UK), 17(2):140, June 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/17/2/140.full.pdf+html.

Bartholomew:1979:SL

[Bar79b] Michael Bartholomew. The singularity of Lyell. History of Science (UK), 17(4):276–293, December 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/17/4/276.full.pdf+html.

Barton:2003:MSL

[Bar03] Ruth Barton. ‘Men of science’: Language, identity and professionalization in the mid-Victorian scientific community. History of Science (UK), 41(1):73–119, March 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/41/1/73.full.pdf+html.

Barford:2017:DSN

[Bar17] Megan Barford. D.176: Sextants, numbers, and the Hydrographic Office of the Admiralty. History of Science (UK), 55(4):431–456, December 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://journals.sagepub.com/doi/full/
10.1177/0073275317712817.

Barany:2020:IPM

[Bar20a] Michael J. Barany. Impersonation and personification in midtwentieth century mathematics. History of Science (UK), 58(4):417– 436, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275320924571.

Barnett:2020:SHF

[Bar20b] Lydia Barnett. Showing and hiding: The flickering visibility of earth workers in the archives of earth science. History of Science (UK), 58 (3):245–274, September 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275319874982.

Baruzzi:2022:AMB

[Bar22] Miquel Carandell Baruzzi. Animals for the mayor: Barcelona’s zoo in the making of local policies and national narratives (1957–73). History of Science (UK), 60(3):405–429, September 1, 2022. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753211026135.

Bates:1998:CCHa

[Bat98a] Don Bates. Closing the circle: How Harvey and his contemporaries played the game of truth, Part 1. History of Science (UK), 36(2):213– 232, June 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/36/2/
213.full.pdf+html.

Bates:1998:CCHb

[Bat98b] Don Bates. Closing the circle: How Harvey and his contemporaries played the game of truth, Part 2. History of Science (UK), 36(3):245– 267, September 1998. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
36/3/245.full.pdf+html.

Baudry:2019:EIS

[Bau19] Jérôme Baudry. Examining inventions, shaping property: The savants and the French patent system. History of Science (UK), 57(1):62–80, March 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275318767233.

Biagioli:2019:TLL

[BB19] Mario Biagioli and Marius Buning. “Technologies of the law/law as a technology”. History of Science (UK), 57(1):3–17, March 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318816163.

Bergeron:2021:SIS

[BB21] Andrée Bergeron and Charlotte Bigg. The spatial inscription of science in the twentieth century. History of Science (UK), 59(2):121–132, June 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275320988399.

Ben-Chaim:1998:DUN

[BC98] Michael Ben-Chaim. Doctrine and use: Newton’s “gift of preaching”. History of Science (UK), 36(3):269–298, September 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/36/3/269.full.pdf+html.

Ben-Chaim:2000:VFB

[BC00] Michael Ben-Chaim. The value of facts in Boyle’s experimental philosophy. History of Science (UK), 38(1):57–77, March 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/38/1/57.full.pdf+html.

Brandon:2023:NKM

[BD23] Pepijn Brandon and Marten Dondorp. Nodes of knowledge, managing transfer: Shipbuilding and repair during the transformation from sail to steam. History of Science (UK), 61(1):19–39, March 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
0073275320971100.

Brice:2001:CHL

[BdMF01] William R. Brice and Silvia F. de M. Figueirôa. Charles Hartt, Louis Agassiz, and the controversy over Pleistocene glaciation in Brazil. History of Science (UK), 39(2):161–184, June 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/39/2/161.full.pdf+html.

Beckman:2004:NPC

[Bec04] Jenny Beckman. Nature’s palace: Constructing the Swedish museum of natural history. History of Science (UK), 42(1):85–111, March 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/42/1/85.full.
pdf+html.

Beckman:2022:CCS

[Bec22] Jenny Beckman. Competition and coordination in Swedish botanical publication, 1820–79: Eleven editions of Hartman’s Handbook. History of Science (UK), 60(2):211–231, June 1, 2022. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275320987414.

Beer:1975:BRK

[Bee75] Arthur Beer. Book review: Kepler Symposium: Internationales Kepler-Symposium Weil der Stadt. History of Science (UK), 13(1): 70–71, March 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/1/
70.full.pdf+html.

Bellon:2009:CDS

[Bel09] Richard Bellon. Charles Darwin solves the “riddle of the flower”; or, why don’t historians of biology know about the birds and the bees? History of Science (UK), 47(4):373–406, December 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/47/4/373.full.pdf+html.

Belknap:2014:TLG

[Bel14] Geoffrey Belknap. Through the looking glass: Photography, science and Imperial motivations in John Thomson’s photographic expeditions. History of Science (UK), 52(1):73–97, March 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/52/1/73.full.pdf+html.

Belteki:2022:SOR

[Bel22] Daniel Belteki. The spring of order: Robert Main’s management of astronomical labor at the Royal Observatory, Greenwich. History of Science (UK), 60(4):575–593, December 1, 2022. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753211028434.

Bennett:1986:MPM

[Ben86] J. A. Bennett. The mechanics’ philosophy and the mechanical philosophy. History of Science (UK), 24(1):1–28, March 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/24/1/1.full.pdf+html.

Benjamin:1988:ERW

[Ben88] Marina Benjamin. Essay review: Women in science: Uneasy careers and intimate lives: Women in science, 1789–1979. History of Science (UK), 26(4):439–441, December 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/26/4/439.full.pdf+html.

Bennett:2006:CCE

[Ben06] Jim Bennett. Catadioptrics and commerce in Eighteenth-Century London. History of Science (UK), 44(2):247–278, June 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/44/2/247.full.pdf+html.

Bennett:2022:MH

[Ben22] Jim Bennett. Michael Hoskin (1930–2021). History of Science (UK), 60(2):280–283, June 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/00732753221083469.

Berman:1966:BRB

[Ber66] R. Berman. Book review: Britain and atomic energy 1939–45. History of Science (UK), 5(1):154–156, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/5/1/154.full.pdf+html.

Berendzen:1973:EGS

[Ber73] Richard Berendzen. On the exponential growth of science. History of Science (UK), 11(4):283–285, December 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/11/4/283.full.pdf+html.

Bernal:1994:RRP

[Ber94] Martin Bernal. Response to Robert Palter. History of Science (UK), 32(4):445–468, December 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/32/4/445.full.pdf+html. See [Pal93].

Berg:2007:GUK

[Ber07] Maxine Berg. The Genesis of ‘useful knowledge’. History of Science (UK), 45(2):123–133, June 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/45/2/123.full.pdf+html.

Berkowitz:2014:CBS

[Ber14] Carin Berkowitz. Charles Bell’s seeing hand: Teaching anatomy to the senses in Britain, 1750–1840. History of Science (UK), 52(4):377–400, December 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/52/4/
377.full.pdf+html.

Berenstein:2016:MGS

[Ber16] Nadia Berenstein. Making a global sensation: Vanilla flavor, synthetic chemistry, and the meanings of purity. History of Science (UK), 54(4): 399–424, December 2016. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275316681802.

Biagioli:1989:SSI

[Bia89] Mario Biagioli. The social status of Italian mathematicians, 1450– 1600. History of Science (UK), 27(1):41–95, March 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/27/1/41.full.pdf+html.

Biagioli:1990:GSP

[Bia90] Mario Biagioli. Galileo’s system of patronage. History of Science (UK), 28(1):1–62, March 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/28/1/1.full.pdf+html.

Biagioli:2006:PPL

[Bia06] Mario Biagioli. From print to patents: Living on instruments in early modern Europe. History of Science (UK), 44(2):139–186, June 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/44/2/139.full.
pdf+html.

Biagioli:2019:WIP

[Bia19] Mario Biagioli. Weighing intellectual property: Can we balance the social costs and benefits of patenting? History of Science (UK), 57(1): 140–163, March 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275318797787.

Biggs:1969:MAM

[Big69] Robert Biggs. Medicine in ancient Mesopotamia. History of Science (UK), 8(1):94–105, March 1969. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/8/1/94.full.pdf+html.

Bil:2022:TCB

[Bil22] Geoff Bil. Tangled compositions: Botany, agency, and authorship aboard HMS Endeavour. History of Science (UK), 60(2):183–210, June 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275320971109.

Bivins:1999:EEE

[Biv99] Roberta Bivins. Expectations and expertise: Early British responses to Chinese medicine. History of Science (UK), 37(4):459–489, December 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/37/4/459.full.
pdf+html.

Bivins:2020:WUA

[Biv20] Roberta Bivins. Weighing on us all? Quantification and cultural responses to obesity in NHS Britain. History of Science (UK), 58(2): 216–242, July 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275319842965.

Boantza:2013:RFN

[Boa13] Victor D. Boantza. The rise and fall of nitrous air eudiometry: Enlightenment ideals, embodied skills, and the conflicts of experimental philosophy. History of Science (UK), 51(4):377–412, December 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/51/4/377.full.
pdf+html.

Boner:2008:LLF

[Bon08] Patrick J. Boner. Life in the liquid fields: Kepler, Tycho and Gilbert on the nature of the Heavens and Earth. History of Science (UK), 46(3):275–297, September 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/46/3/275.full.pdf+html.

Boner:2014:SSH

[Bon14] Patrick J. Boner. Statesman and scholar: Herwart von Hohenburg as patron and author in the Republic of Letters. History of Science (UK), 52(1):29–51, March 2014. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
52/1/29.full.pdf+html.

Boschiero:2005:PGT

[Bos05] Luciano Boschiero. Post-Galilean thought and experiment in Seventeenth-Century Italy: The life and work of Vincenzio Viviani. History of Science (UK), 43(1):77–100, March 2005. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/43/1/77.full.pdf+html.

Bowler:1974:EE

[Bow74] Peter J. Bowler. Evolutionism in the enlightenment. History of Science (UK), 12(3):159–183, September 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/12/3/159.full.pdf+html.

Bowler:2009:EPD

[Bow09] Peter J. Bowler. The eclipse of pseudo-Darwinism? Reflections on some recent developments in Darwin studies. History of Science (UK), 47(4):431–443, December 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/4/431.full.pdf+html.

Boyer:1967:ERM

[Boy67] Carl B. Boyer. Essay review: The making of a mathematician: The mathematical papers of Isaac Newton. History of Science (UK), 6 (1):97–106, March 1967. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/6/
1/97.full.pdf+html.

Bellido:2019:LPT

[BP19] Jose Bellido and Alain Pottage. Lexical properties: Trademarks, dictionaries, and the sense of the generic. History of Science (UK), 57(1): 119–139, March 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275318766160.

Bracegirdle:1977:HHB

[Bra77] Brian Bracegirdle. The history of histology: a brief survey of sources. History of Science (UK), 15(2):77–101, June 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/15/2/77.full.pdf+html.

Brand:1988:SPC

[Bra88] J. C. D. Brand. The scientific papers of C. P. Snow. History of Science (UK), 26(2):111–127, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/26/2/111.full.pdf+html.

Brazelton:2023:AIR

[Bra23] Mary Augusta Brazelton. Aviation infrastructures in the Republic of China, 1920–37. History of Science (UK), 61(1):102–120, March 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
0073275321995638.

Brioist:2009:FDG

[Bri09] Pascal Brioist. “Familiar demonstrations in geometry”: French and Italian engineers and Euclid in the Sixteenth Century. History of Science (UK), 47(1):1–26, March 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/1/1.full.pdf+html.

Brixius:2020:EES

[Bri20] Dorit Brixius. From ethnobotany to emancipation: Slaves, plant knowledge, and gardens on eighteenth-century Isle de France. History of Science (UK), 58(1):51–75, March 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275319835431.

Brock:1962:BRM

[Bro62] W. H. Brock. Book review: The making of modern science. History of Science (UK), 1(1):117–118, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/117.full.pdf+html.

Brock:1967:ERA

[Bro67] W. H. Brock. Essay review: an attempt to establish the first principles of the history of chemistry: History of analytical chemistry. History of Science (UK), 6(1):156–169, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/6/1/156.full.pdf+html.

Brozek:1971:PPB

[Bro71] Josef Brozek. The psychology and physiology of behaviour: Some recent Soviet writings on their history. History of Science (UK), 10(1): 56–87, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/10/1/
56.full.pdf+html.

Brozek:1975:CWE

[Bro75] Josef Brozek. Contemporary West European historiography of psychology. History of Science (UK), 13(1):29–60, March 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/13/1/29.full.pdf+html.

Brock:1977:ERT

[Bro77a] W. H. Brock. Essay review: a taste for naturalists: The naturalist in Britain. a social history. History of Science (UK), 15(4):287–294, December 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/15/4/
287.full.pdf+html.

Brozek:1977:CEE

[Bro77b] Joseph Brozek. Contemporary East European historiography of psychology. History of Science (UK), 15(4):233–251, December 1977. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/15/4/233.full.pdf+html.

Brock:1981:LON

[Bro81a] W. H. Brock. Liebigiana: Old and new perspectives. History of Science (UK), 19(3):201–218, September 1981. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/19/3/201.full.pdf+html.

Brooke:1981:AHF

[Bro81b] John Hedley Brooke. Avogadro’s hypothesis and its fate: a case-study in the failure of case-studies. History of Science (UK), 19(4):235–273, December 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/19/4/
235.full.pdf+html.

Browne:1981:BRC

[Bro81c] E. J. Browne. Book review: Collecting natural history: British natural history books 1495–1900: a handlist. History of Science (UK), 19(3): 223–224, September 1981. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
19/3/223.full.pdf+html.

Browne:1984:ERB

[Bro84] Janet Browne. Essay review: Botany and botanists: Landmarks of botanical history, Linnaeus: The man and his work. History of Science (UK), 22(2):207–209, June 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/22/2/207.full.pdf+html.

Brock:1992:ERB

[Bro92a] W. H. Brock. Essay review: a biochemical ferment, a documentary history of biochemistry 1770–1940. History of Science (UK), 30(3): 325–328, September 1992. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
30/3/325.full.pdf+html.

Brock:1992:ERE

[Bro92b] W. H. Brock. Essay review: Essays on chemical ideas, ideas in chemistry: a history of the science. History of Science (UK), 30(4):439–442, December 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/4/
439.full.pdf+html.

Browne:1992:SSS

[Bro92c] Janet Browne. Squibs and snobs: Science in humorous British undergraduate magazines around 1830. History of Science (UK), 30(2):

165–197, June 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/2/
165.full.pdf+html.

Brock:1994:MFD

[Bro94] Helen Brock. The many facets of Dr William Hunter (1718–83). History of Science (UK), 32(4):387–408, December 1994. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/32/4/387.full.pdf+html.

Broman:1998:HPS

[Bro98a] Thomas Broman. The Habermasian public sphere and “science in the Enlightenment”. History of Science (UK), 36(2):123–149, June 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/36/2/123.full.
pdf+html.

Browne:1998:BRO

[Bro98b] Janet Browne. Book review: Ornithologists organized, discovering birds: The emergence of ornithology as a scientific discipline, 1760– 1850. History of Science (UK), 36(3):359–360, September 1998. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/36/3/359.full.pdf+html.

Broman:2013:CCN

[Bro13] Thomas Broman. Criticism and the circulation of news: The scholarly Press in the Late Seventeenth Century. History of Science (UK), 51 (2):125–150, June 2013. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
51/2/125.full.pdf+html.

Brush:1987:NHE

[Bru87] Stephen G. Brush. The nebular hypothesis and the evolutionary worldview. History of Science (UK), 25(3):245–278, September 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/25/3/245.full.
pdf+html.

Bruland:2007:TSU

[Bru07] Kristine Bruland. Technology selection and useful knowledge: a comment. History of Science (UK), 45(2):179–183, June 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/45/2/179.full.pdf+html.

Brusius:2017:HTB

[Bru17] Mirjam Brusius. Hitting two birds with one stone: an afterword on archeology and the history of science. History of Science (UK), 55(3): 383–391, September 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317727975.

Bryden:1977:BRS

[Bry77] David Bryden. Book review: Scholars and objects: Scientific instruments. History of Science (UK), 15(4):297–298, December 1977. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/15/4/297.full.pdf+html.

Buchdahl:1962:ERI

[Buc62a] Gerd Buchdahl. Essay review: Inward nature versus objectivity: The edge of objectivity. an essay in the history of scientific ideas. History of Science (UK), 1(1):90–95, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/90.full.pdf+html.

Buchdahl:1962:ERP

[Buc62b] Gerd Buchdahl. Essay review: On the presuppositions of historians of science: The mechanisation of the world picture. History of Science (UK), 1(1):67–77, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/67.full.pdf+html.

Buchdahl:1962:HPS

[Buc62c] Gerd Buchdahl. History and philosophy of science at Cambridge. History of Science (UK), 1(1):62–66, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/62.full.pdf+html.

Buchdahl:1965:RHS

[Buc65] Gerd Buchdahl. A revolution in historiography of science. History of Science (UK), 4(1):55–69, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/4/1/55.full.pdf+html.

Buchdahl:1970:ERH

[Buc70] Gerd Buchdahl. Essay review: History and methodology: Fact and theory. an aspect of philosophy of science. History of Science (UK), 9 (1):93–101, March 1970. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/9/
1/93.full.pdf+html.

Buckland:2009:ERC

[Buc09] Adelene Buckland. Essay review: Collectors of worlds: Writing the history of geohistory, bursting the limits of time: The reconstruction of geohistory in the age of revolution, worlds before Adam: The reconstruction of geohistory in the age of reform. History of Science (UK), 47(4):485–492, December 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/4/485.full.pdf+html.

Bud:2013:FPS

[Bud13] Robert Bud. Framed in the public sphere: Tools for the conceptual history of “applied science” — a review paper. History of Science (UK), 51(4):413–433, December 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/51/4/413.full.pdf+html.

Buklijas:2020:PPP

[Buk20] Tatjana Buklijas. Publicity, politics, and professoriate in fin-de-siècle Vienna: The misconduct of the embryologist Samuel Leopold Schenk. History of Science (UK), 58(4):458–484, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275320914140.

Bulstrode:2018:RAM

[Bul18] Jenny Bulstrode. Riotous assemblage and the materials of regulation. History of Science (UK), 56(3):278–313, September 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//journals.sagepub.com/doi/full/10.1177/0073275318776187.

Buning:2019:MTN

[Bun19] Marius Buning. Making things new: Invention privileges and the configuration of priority. History of Science (UK), 57(1):81–96, March 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318799947.

Buttner:1979:SRR

[Büt79] M. Büttner. The significance of the reformation for the reorientation of geography in Lutheran Germany. History of Science (UK), 17(3):

151–169, September 1979. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
17/3/151.full.pdf+html.

Bycroft:2022:ISC

[BWM22] Michael Bycroft and Alexander Wragge-Morley. Introduction: Science and connoisseurship in the European Enlightenment. History of Science (UK), 60(4):439–457, December 1, 2022. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753211049039.

Bycroft:2022:HCG

[Byc22] Michael Bycroft. The hand of the connoisseur: Gems and hardness in Enlightenment mineralogy. History of Science (UK), 60(4):500– 523, December 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753221087558.

Bynum:1973:BRM

[Byn73] W. F. Bynum. Book review: Medicine in America: American physicians in the Nineteenth Century: From sects to science. History of Science (UK), 11(4):312, December 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/11/4/312.full.pdf+html.

Bynum:1975:GCB

[Byn75] William F. Bynum. The great chain of being after forty years: an appraisal. History of Science (UK), 13(1):1–28, March 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/13/1/1.full.pdf+html.

Bynum:1988:ERD

[Byn88] W. F. Bynum. Essay review: Darwin’s opera: The works of Charles Darwin. History of Science (UK), 26(2):213–215, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/26/2/213.full.pdf+html.

Ben-Zaken:2009:NGB

[BZ09] Avner Ben-Zaken. From Naples to Goa and back: a secretive Galilean messenger and a radical hermeneutist. History of Science (UK), 47(2): 147–174, June 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/2/
147.full.pdf+html.

Ben-Zaken:2011:RPR

[BZ11] Avner Ben-Zaken. The revolving planets and the revolving clocks: Circulating mechanical objects in the Mediterranean. History of Science (UK), 49(2):125–148, June 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/49/2/125.full.pdf+html.

Cain:1965:BRH

[Cai65] A. J. Cain. Book review: Histoire de la Zoologie des origines à Linné (French) [History of zoology from its origins to Linnaeus]. History of Science (UK), 4(1):143–146, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/4/1/143.full.pdf+html.

Callender:1988:GMO

[Cal88] L. A. Callender. Gregor Mendel: an opponent of descent with modification. History of Science (UK), 26(1):41–75, March 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/26/1/41.full.pdf+html.

Call:1998:ADA

[Cal98] Lewis Call. Anti-Darwin, anti-Spencer: Friedrich Nietzsche’s critique of Darwin and “Darwinism”. History of Science (UK), 36(1): 1–22, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/36/1/1.
full.pdf+html.

Cannon:1964:HDE

[Can64] Walter F. Cannon. History in depth: The Early Victorian period. History of Science (UK), 3(1):20–38, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/3/1/20.full.pdf+html.

Cantor:1976:MHA

[Can76] G. N. Cantor. Method in history: For and against. History of Science (UK), 14(4):265–276, December 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/14/4/265.full.pdf+html.

Cantor:1978:HGO

[Can78] G. N. Cantor. The historiography of ‘Georgian’ optics. History of Science (UK), 16(1):1–21, March 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/16/1/1.full.pdf+html.

Cantor:1982:ERE

[Can82] G. N. Cantor. Essay review: The Eighteenth Century problem: The ferment of knowledge: Studies in the historiography of Eighteenth Century science. History of Science (UK), 20(1):44–63, March 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/20/1/44.full.pdf+html.

Cantor:1985:ERE

[Can85] Geoffrey Cantor. Essay review: Eighteenth Century materialism: Thinking matter: Materialism in Eighteenth-Century Britain. History of Science (UK), 23(2):201–206, June 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/23/2/201.full.pdf+html.

Caneva:1997:PNR

[Can97] Kenneth L. Caneva. Physics and naturphilosophie: a reconnaissance. History of Science (UK), 35(1):35–106, March 1997. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/35/1/35.full.pdf+html.

Canales:2001:SER

[Can01] Jimena Canales. The single eye: Re-evaluating Ancien Régime science. History of Science (UK), 39(1):71–94, March 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/39/1/71.full.pdf+html.

Candiani:2017:RKC

[Can17] Vera S. Candiani. Reframing knowledge in colonization: Plebeians and municipalities in the environmental expertise of the Spanish Atlantic. History of Science (UK), 55(2):234–252, June 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//journals.sagepub.com/doi/full/10.1177/0073275317706041.

Cappelletti:1976:LFR

[Cap76] Vincenzo Cappelletti. The latest Freud research. History of Science (UK), 14(1):1–16, March 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/14/1/1.full.pdf+html.

Caputo:2023:EMF

[Cap23] Sara Caputo. Exploration and mortification: Fragile infrastructures, imperial narratives, and the self-sufficiency of British naval “discovery” vessels, 1760–1815. History of Science (UK), 61(1):40–59, March 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
0073275320970042.

Cardwell:1962:STE

[Car62] D. S. L. Cardwell. Science and technology in the Eighteenth Century. History of Science (UK), 1(1):30–43, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/30.full.pdf+html.

Cardwell:1963:ERS

[Car63] D. S. L. Cardwell. Essay review: Science in the Nineteenth Century: Histoire générale des sciences, la science contemporaine, le XIXe siècle. History of Science (UK), 2(1):140–145, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/2/1/140.full.pdf+html.

Cardwell:1968:ASH

[Car68] D. S. L. Cardwell. The academic study of the history of technology. History of Science (UK), 7(1):112–124, March 1968. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/7/1/112.full.pdf+html.

Cardwell:1977:ERT

[Car77] Donald Cardwell. Essay review: Theories of heat and the rise of physics: Joseph Fourier: The man and the physicist. History of Science (UK), 15(2):138–145, June 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/15/2/138.full.pdf+html.

Cardwell:1979:ERP

[Car79] Donald Cardwell. Essay review: The philosophy of technology: The dynamics of science and technology: Social values, technical norms and scientific criteria in the development of knowledge. History of Science (UK), 17(4):293–295, December 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/17/4/293.full.pdf+html.

Caron:1988:BLS

[Car88] Joseph A. Caron. ‘Biology’ in the life sciences: a historiographical contribution. History of Science (UK), 26(3):223–268, September 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/26/3/223.full.
pdf+html.

Cardwell:1991:ERE

[Car91] Donald Cardwell. Essay review: Essays on the history of technology, an encyclopaedia of the history of technology. History of Science (UK), 29(1):105–107, March 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/29/1/105.full.pdf+html.

Carolino:2016:SPA

[Car16] Luís Miguel Carolino. Science, patronage, and academies in early seventeenth-century Portugal: The scientific academy of the nobleman and university professor André de almada. History of Science (UK), 54(2):107–137, June 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/54/2/107.full.pdf+html.

Carruthers:2017:VMP

[Car17] William Carruthers. Visualizing a monumental past: Archeology, Nasser’s Egypt, and the early Cold War. History of Science (UK), 55 (3):273–301, September 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.sagepub.

com/doi/full/10.1177/0073275316681800.

Casini:1984:NCS

[Cas84] Paolo Casini. Newton: The classical scholia. History of Science (UK), 22(1):1–46, March 1984. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
22/1/1.full.pdf+html.

Christianidis:2002:HKN

[CDG02] Jean Christianidis, Dimitris Dialetis, and Kostas Gavroglu. Having a knack for the non-intuitive: Aristarchus’s heliocentrism through Archimedes’s geocentrism. History of Science (UK), 40(2):147–168, June 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL http://hos.sagepub.com/content/40/2/147.
full.pdf+html.

Christie:1982:SWN

[CG82] J. R. R. Christie and J. V. Golinski. The spreading of the word: New directions in the historiography of chemistry 1600–1800. History of Science (UK), 20(4):235–266, December 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/20/4/235.full.pdf+html.

Charles:2021:SMF

[CG21] Loïc Charles and Yann Giraud. Seeking the “museum of the future”: Public exhibitions of science, industry, and the social, 1910–1940. History of Science (UK), 59(2):133–154, June 1, 2021. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275317751316.

Charette:2006:LIA

[Cha06] François Charette. The locales of Islamic astronomical instrumentation. History of Science (UK), 44(2):123–138, June 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/44/2/123.full.pdf+html.

Chapman:2008:ERH

[Cha08] Allan Chapman. Essay review: a Herschel trilogy: The Herschel Partnership: As Viewed by Caroline, Caroline Herschel’s Autobiographies, The Herschels of Hanover. History of Science (UK), 46(3):365–368, September 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/46/3/
365.full.pdf+html.

Chakrabarti:2010:BBA

[Cha10] Pratik Chakrabarti. Beasts of burden: Animals and laboratory research in colonial India. History of Science (UK), 48(2):125–151, June 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/48/2/125.full.
pdf+html.

Chang:2011:CDW

[Cha11] Hasok Chang. Compositionism as a dominant way of knowing in modern chemistry. History of Science (UK), 49(3):247–268, September 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/49/3/247.full.
pdf+html.

Chacko:2018:WLG

[Cha18] Xan Sarah Chacko. When life gives you lemons: Frank Meyer, authority, and credit in early twentieth-century plant hunting. History of Science (UK), 56(4):432–469, December 1, 2018. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275318784124.

Chang:2025:SEF

[Cha25] Pang-Yen Chang. Struggling with exactitude in a fragmented state: Intelligence testing in early twentieth-century China. History of Science (UK), 63(1):73–100, March 2025. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753241235432.

Chilton:1965:BRH

[Chi65] D. Chilton. Book review: The history of the barometer. History of Science (UK), 4(1):141–143, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/4/1/141.full.pdf+html.

Chiang:2015:OSH

[Chi15a] Howard Chiang. Ordering the social: History of the human sciences in modern China. History of Science (UK), 53(1):4–8, March 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/53/1/4.full.pdf+html.

Chiang:2015:TCP

[Chi15b] Howard Chiang. Translating culture and psychiatry across the Pacific: How koro became culture-bound. History of Science (UK), 53(1):102– 119, March 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/53/1/
102.full.pdf+html.

Chiriac:2023:RSV

[Chi23] Alexandra Chiriac. The Revista Stiintifica “Vasile Adamachi” and its role in forming national and international scientific awareness of Greater Romania, 1910–1933. History of Science (UK), 61(2):266– 284, June 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/00732753211054979.

Christie:1974:ODS

[Chr74] John R. R. Christie. The origins and development of the Scottish scientific community, 1680–1760. History of Science (UK), 12(2):122– 141, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/2/
122.full.pdf+html.

Christie:1987:NRH

[Chr87] J. R. R. Christie. Narrative and rhetoric in Hélène Metzger’s historiography of Eighteenth Century chemistry. History of Science (UK), 25 (1):99–109, March 1987. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
25/1/99.full.pdf+html.

Christianson:1998:TBS

[Chr98] J. R. Christianson. Tycho Brahe in Scandinavian scholarship. History of Science (UK), 36(4):467–484, December 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/4/467.full.pdf+html.

Chu:2003:ROG

[Chu03] Pingyi Chu. Remembering our grand tradition: The historical memory of the scientific exchanges between China and Europe, 1600–1800. History of Science (UK), 41(2):193–215, June 2003. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/41/2/193.full.pdf+html.

Clarke:1964:BRG

[Cla64] Edwin Clarke. Book review: The growth of medical thought. History of Science (UK), 3(1):143–146, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/3/1/143.full.pdf+html.

Clarke:1965:BRA

[Cla65] Edwin Clarke. Book review: Andreas Vesalius of Brussels 1514–1564. History of Science (UK), 4(1):139–140, March 1965. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/4/1/139.full.pdf+html.

Clark:1989:DOR

[Cla89] William Clark. On the dialectical origins of the research seminar. History of Science (UK), 27(2):111–154, June 1989. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/27/2/111.full.pdf+html.

Clarke:1993:DPS

[Cla93] Desmond M. Clarke. Dormitive powers and scholastic qualities: a reply to Hutchison. History of Science (UK), 31(3):317–327, September 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/31/3/317.full.

pdf+html. See [Hut91].

Clark:1997:GPTa

[Cla97a] William Clark. German physics textbooks in the Goethezeit, Part 1. History of Science (UK), 35(2):219–239, June 1997. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/35/2/219.full.pdf+html.

Clark:1997:GPTb

[Cla97b] William Clark. German physics textbooks in the Goethezeit, Part 2. History of Science (UK), 35(3):295–363, September 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/35/3/295.full.pdf+html.

Claessens:2009:CPL

[Cla09] Guy Claessens. Clavius, Proclus, and the limits of interpretation: Snapshot-idealization versus projectionism. History of Science (UK), 47(3):317–336, September 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/3/317.full.pdf+html.

Cleaver:1989:ASA

[Cle89] K. C. Cleaver. Adam Smith on astronomy. History of Science (UK), 27(2):211–218, June 1989. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
27/2/211.full.pdf+html.

Clucas:2008:GBR

[Clu08] Stephen Clucas. Galileo, Bruno and the rhetoric of dialogue in Seventeenth-Century natural philosophy. History of Science (UK), 46(4):405–429, December 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/46/4/405.full.pdf+html.

Cockcroft:1963:BRC

[Coc63] J. D. Cockcroft. Book review: The Collected Papers of Lord Rutherford of Nelson. Under the Scientific Direction of Sir James Chadwick. History of Science (UK), 2(1):164–165, March 1963. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/2/1/164.full.pdf+html.

Cohen:1999:SRT

[Coh99] H. Floris Cohen. The scientific revolution: Has there been a British view? — A personal assessment. History of Science (UK), 37(1):107– 112, March 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/1/
107.full.pdf+html.

Coleman:1977:OS

[Col77] Janet Coleman. Ockham studies. History of Science (UK), 15(1):67– 72, March 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/15/1/
67.full.pdf+html.

Collins:1981:PCS

[Col81] H. M. Collins. The place of the ‘core-set’ in modern science: Social contingency with methodological propriety in science. History of Science (UK), 19(1):6–19, March 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/19/1/6.full.pdf+html.

Colp:1989:CDP

[Col89] Ralph Colp, Jr. Charles Darwin’s past and future biographies. History of Science (UK), 27(2):167–197, June 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/27/2/167.full.pdf+html.

Colp:2000:MDI

[Col00] Ralph Colp, Jr. More on Darwin’s illness. History of Science (UK), 38 (2):219–236, June 2000. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
38/2/219.full.pdf+html.

Cole:2024:NTP

[Col24] Camille Lyans Cole. Nafia for the Tigris: The privy purse and the infrastructure of development in late Ottoman Iraq, 1882–1914. History of Science (UK), 62(4):488–510, December 2024. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/0073275321999265.

Conrad:1995:BRE

[Con95] Lawrence I. Conrad. Book review: Earthquakes in the middle East, the seismicity of Egypt, Arabia and the red sea. History of Science (UK), 33(3):369–370, September 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/33/3/369.full.pdf+html.

Cooter:1976:PPP

[Coo76] R. J. Cooter. Phrenology: The provocation of progress. History of Science (UK), 14(4):211–234, December 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/14/4/211.full.pdf+html.

Cooter:1982:ERA

[Coo82] Roger Cooter. Essay review: The archaeology of plants: Nature’s second kingdom. explorations of vegetality in the Eighteenth Century. History of Science (UK), 20(4):304–309, December 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/20/4/304.full.pdf+html.

Cooper:2018:PPP

[Coo18] Alix Cooper. Placing plants on paper: Lists, herbaria, and tables as experiments with territorial inventory at the mid-seventeenth-century Gotha court. History of Science (UK), 56(3):257–277, September 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://journals.sagepub.com/doi/full/10.1177/
0073275318776515.

Corr:2014:ESN

[Cor14] John Corr. The Enlightenment surfaces in Nineteenth-Century Mexico: Scientific thinking attempts to deliver order and progress. History of Science (UK), 52(1):98–123, March 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/52/1/98.full.pdf+html.

Costa:2002:MME

[Cos02] Shelley Costa. Marketing mathematics in Early Eighteenth-Century England: Henry Beighton, certainty, and the public sphere. History of Science (UK), 40(2):211–232, June 2002. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/40/2/211.full.pdf+html.

Cooter:1994:SSP

[CP94] Roger Cooter and Stephen Pumfrey. Separate spheres and public places: Reflections on the history of science popularization and science in popular culture. History of Science (UK), 32(3):237–267, September 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/32/3/
237.full.pdf+html.

Chiapperino:2021:TBH

[CP21] Luca Chiapperino and Francesco Panese. On the traces of the biosocial: Historicizing “plasticity” in contemporary epigenetics. History of Science (UK), 59(1):3–44, March 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275319876839.

Crombie:1962:HPS

[Cro62] A. C. Crombie. History and philosophy of science at Oxford. History of Science (UK), 1(1):57–61, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/57.full.pdf+html.

Crosland:1964:BRC

[Cro64] Maurice Crosland. Book review: Chymia. History of Science (UK), 3(1):152, March 1964. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/3/
1/152.full.pdf+html.

Crombie:1975:SAS

[Cro75] A. C. Crombie. Some attitudes to scientific progress: Ancient, medieval and early modern. History of Science (UK), 13(3):213–230, September 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/3/
213.full.pdf+html.

Crombie:1980:SAR

[Cro80] A. C. Crombie. Science and the arts in the Renaissance: The search for truth and certainty, old and new. History of Science (UK), 18(4): 233–246, December 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/4/233.full.pdf+html.

Crombie:1988:DMW

[Cro88] Alistair C. Crombie. Designed in the mind: Western visions of science, nature and humankind. History of Science (UK), 26(1):1– 12, March 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/26/1/1.
full.pdf+html.

Crombie:1995:CSE

[Cro95] A. C. Crombie. Commitments and styles of European scientific thinking. History of Science (UK), 33(2):225–238, June 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/33/2/225.full.pdf+html.

Crozier:1998:BRK

[Cro98] Ivan Crozier. Book review: Kinsey exposed: Alfred C. Kinsey: a public/private life. History of Science (UK), 36(1):115–118, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/36/1/115.full.
pdf+html.

Crozier:2001:BSN

[Cro01] Ivan Crozier. Becoming a sexologist: Norman Haire, the 1929 London World League for Sexual Reform Congress, and organizing medical knowledge about sex in Interwar England. History of Science (UK), 39(3):299–329, September 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/39/3/299.full.pdf+html.

Crosland:2006:SEN

[Cro06] Maurice Crosland. A science empire in Napoleonic France. History of Science (UK), 44(1):29–48, March 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/44/1/29.full.pdf+html.

Crosland:2007:FSM

[Cro07] Maurice Crosland. The financial support of men of science in France c. 1660–c. 1800: a survey. History of Science (UK), 45(3):327–355, September 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/3/
327.full.pdf+html.

Crozier:2008:PTC

[Cro08] Ivan Crozier. Pillow talk: Credibility, trust and the sexological case history. History of Science (UK), 46(4):375–404, December 2008. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/46/4/375.full.pdf+html.

Csiszar:2010:SSO

[Csi10] Alex Csiszar. Seriality and the search for order: Scientific print and its problems during the Late Nineteenth Century. History of Science (UK), 48(3–4):399–434, September/December 2010. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Cullen:1993:PHL

[Cul93] Christopher Cullen. Patients and healers in Late Imperial China: Evidence from the Jinpingmei. History of Science (UK), 31(2):99– 150, June 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/31/2/
99.full.pdf+html.

Cunningham:1991:HPG

[Cun91] Andrew Cunningham. How the Principia got its name; or, taking natural philosophy seriously. History of Science (UK), 29(4):377–392, December 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/29/4/
377.full.pdf+html.

Cuomo:1997:SBN

[Cuo97] Serafina Cuomo. Shooting by the book: Notes on Niccolò Tartaglia’s Nova Scientia. History of Science (UK), 35(2):155–188, June 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/35/2/155.full.
pdf+html.

Curry:1985:ERR

[Cur85] Patrick Curry. Essay review: Revisions of science and magic: From Paracelsus to Newton: Magic and the making of modern science, occult and scientific mentalities in the Renaissance. History of Science (UK), 23(3):299–325, September 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/23/3/299.full.pdf+html.

Curtis:2012:SNO

[Cur12] Stephan Curtis. Swedish in name only: The international education of Nineteenth-Century Swedish medical students and practitioners. History of Science (UK), 50(3):257–288, September 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/50/3/257.full.pdf+html.

Carruthers:2017:DAR

[CV17] William Carruthers and Stéphane Van Damme. Disassembling archeology, reassembling the modern world. History of Science (UK), 55(3): 255–272, September 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317719849.

Wu:2018:TND

[cW18] Yu chuan Wu. Techniques for nothingness: Debate over the comparability of hypnosis and Zen in early-twentieth-century Japan. History of Science (UK), 56(4):470–496, December 1, 2018. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275317743120.

DAgostino:1990:BHB

[D’A90] S. D’Agostino. Boltzmann and Hertz on the bild-conception of physical theory. History of Science (UK), 28(4):380–398, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/28/4/380.full.
pdf+html.

Dacome:2001:LCP

[Dac01] Lucia Dacome. Living with the chair: Private excreta, collective health and medical authority in the Eighteenth Century. History of Science (UK), 39(4):467–500, December 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/39/4/467.full.pdf+html.

Dally:1994:ERC

[Dal94] Ann Dally. Essay review: Confronting madness, the most solitary of afflictions: Madness and society in Britain, 1700–1900. History of Science (UK), 32(4):479–483, December 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/32/4/479.full.pdf+html.

Dally:1995:BRC

[Dal95] Ann Dally. Book review: Correcting deformities in children, surgery and society in peace and war: Orthopaedics and the organization of modern medicine, 1880–1948. History of Science (UK), 33(1):114– 116, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/1/
114.full.pdf+html.

Dally:1997:BRB

[Dal97a] Ann Dally. Book review: Biology made easy, Aristotle to zoos: a philosophical dictionary of biology. History of Science (UK), 35(3): 365–366, September 1997. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
35/3/365.full.pdf+html.

Dally:1997:BRD

[Dal97b] Ann Dally. Book review: Duchenne muscular dystrophy, the history of a genetic disease: Duchenne muscular dystrophy. History of Science (UK), 35(4):489–491, December 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/35/4/489.full.pdf+html.

Dally:1997:BRI

[Dal97c] Ann Dally. Book review: Images of childbirth., reproducing the womb: Images of childbirth in science, feminist theory, and literature. History of Science (UK), 35(1):113–114, March 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/35/1/113.full.pdf+html.

Dally:1997:BRM

[Dal97d] Ann Dally. Book review: Millenarianism, the great year: Astrology, millenarianism and history in the Western tradition. History of Science (UK), 35(1):117–119, March 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/35/1/117.full.pdf+html.

Dally:1997:BRP

[Dal97e] Ann Dally. Book review: Politics and science, the politics of Western science, 1640–1990. History of Science (UK), 35(3):364–365, September 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/35/3/364.full.
pdf+html.

Dally:1997:BRQ

[Dal97f] Ann Dally. Book review: Quantifying ability, measuring the mind: Education and psychology in England, c. 1860–1990. History of Science (UK), 35(4):485–487, December 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/35/4/485.full.pdf+html.

Dally:1997:BRS

[Dal97g] Ann Dally. Book review: Scientists and intellectuals, the third culture: Beyond the scientific revolution. History of Science (UK), 35(1):111– 112, March 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/35/1/
111.full.pdf+html.

Dally:1997:BRE

[Dal97h] Ann Dally. Book review: The economics of medical practice, making a medical living: Doctors and patients in the English market for medicine, 1720–1911. History of Science (UK), 35(4):487–489, December 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/35/4/
487.full.pdf+html.

Dally:1997:ERA

[Dal97i] Ann Dally. Essay review: Ambivalence towards science, the trouble with science. History of Science (UK), 35(2):241–243, June 1997. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/35/2/241.full.pdf+html.

Dally:1998:BRR

[Dal98a] Ann Dally. Book review: Revolutionary siblings: Birth order, family dynamics, and creative lives. History of Science (UK), 36(1):118– 120, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/36/1/
118.full.pdf+html.

Dally:1998:BRT

[Dal98b] Ann Dally. Book review: Tuberculosis in France: The making of a social disease: Tuberculosis in Nineteenth-Century France. History of Science (UK), 36(1):120–121, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/1/120.full.pdf+html.

Dally:1999:BRD

[Dal99] Ann Dally. Book review: Discoveries ranked, Medicine’s ten greatest discoveries. History of Science (UK), 37(1):113–114, March 1999. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/37/1/113.full.pdf+html.

Damodaran:2013:GRS

[Dam13] Vinita Damodaran. Gender, race and science in Twentieth-Century India: E. K. Janaki Ammal and the history of science. History of Science (UK), 51(3):283–307, September 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/51/3/283.full.pdf+html.

Danziger:1983:OSS

[Dan83] Kurt Danziger. Origins of the schema of stimulated motion: Towards a pre-history of modern psychology. History of Science (UK), 21(2): 183–210, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/2/
183.full.pdf+html.

Daum:2025:HSH

[Dau25] Andreas W. Daum. Humboldtian Science and Humboldt’s science. History of Science (UK), 63(1):29–51, March 2025. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753241252478.

Davie:1962:ERS

[Dav62] Donald Davie. Essay review: Science and poetry: a book of science verse. History of Science (UK), 1(1):100–102, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/1/1/100.full.pdf+html.

Davies:1975:ERR

[Dav75] Gordon L. Davies. Essay review: Research by debate: The geomorphology of William Morris Davis: The history of the study of landforms, or the development of geomorphology, the life and work of William Morris Davis. History of Science (UK), 13(2):139–145, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/13/2/139.full.
pdf+html.

Davis:1978:HSM

[Dav78] Audrey B. Davis. Historical studies of medical instruments. History of Science (UK), 16(2):107–133, June 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/16/2/107.full.pdf+html.

Davies:1995:TYS

[Dav95] Mansel Davies. A thousand years of science and scientists: 988 to 1988. History of Science (UK), 33(2):239–251, June 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/33/2/239.full.pdf+html.

Davis:2019:PPS

[Dav19] Frederick Rowe Davis. Pesticides and the perils of synecdoche in the history of science and environmental history. History of Science (UK), 57(4):469–492, December 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275319848964.

deBont:2013:WLB

[dB13] Raf de Bont. “Writing in letters of blood”: Manners in scientific dispute in Nineteenth-Century Britain and the German lands. History of Science (UK), 51(3):309–335, September 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/51/3/309.full.pdf+html.

DeVos:2017:MCI

[De 17] Paula De Vos. Methodological challenges involved in compiling the Nahua pharmacopeia. History of Science (UK), 55(2):210–233, June 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://journals.sagepub.com/doi/full/10.1177/
0073275317712139.

Dean:2001:HPS

[Dea01] Michael Emmans Dean. Homeopathy and “the progress of science”. History of Science (UK), 39(3):255–283, September 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/39/3/255.full.pdf+html.

Dean:2003:ISS

[Dea03] Katrina Dean. Inscribing settler science: Ernest Rutherford, Thomas Laby and the making of careers in physics. History of Science (UK), 41(2):217–240, June 2003. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
41/2/217.full.pdf+html.

Dear:2012:HRS

[Dea12] Peter Dear. Historiography of not-so-recent science. History of Science (UK), 50(2):197–211, June 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/50/2/197.full.pdf+html.

Dear:2016:DDT

[Dea16] Peter Dear. Darwin and deep time: Temporal scales and the Naturalist’s imagination. History of Science (UK), 54(1):3–18, March 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/54/1/3.full.pdf+html.

Debus:1963:BRH

[Deb63] Allen G. Debus. Book review: a history of chemistry. History of Science (UK), 2(1):165–166, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/2/1/165.full.pdf+html.

Debus:1967:ERA

[Deb67] A. G. Debus. Essay review: Alchemy and the historian of science: Elias Ashmole. History of Science (UK), 6(1):128–138, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/6/1/128.full.pdf+html.

Debus:1974:CPC

[Deb74] Allen G. Debus. The chemical philosophers: Chemical medicine from Paracelsus to Van Helmont. History of Science (UK), 12(4):235–259, December 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/4/
235.full.pdf+html.

Delbourgo:2005:ERL

[Del05] James Delbourgo. Essay review: Leviathan and the Atlantic: The cultural geography of colonial American literatures: Empire, travel, modernity. History of Science (UK), 43(1):101–107, March 2005. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/43/1/101.full.pdf+html.

Delbourgo:2011:DTC

[Del11] James Delbourgo. Divers things: Collecting the world under water. History of Science (UK), 49(2):149–185, June 2011. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/49/2/149.full.pdf+html.

Delbourgo:2019:KWN

[Del19] James Delbourgo. The knowing world: a new global history of science. History of Science (UK), 57(3):373–399, September 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275319831582.

Desmond:1985:MIZb

[Des85a] Adrian Desmond. The making of institutional zoology in London 1822–1836: Part 2. History of Science (UK), 23(3):223–250, September 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/23/3/223.full.
pdf+html.

Desmond:1985:MIZa

[Des85b] Adrian Desmond. The making of institutional zoology in London 1822–1836: Part I. History of Science (UK), 23(2):153–185, June 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/23/2/153.full.
pdf+html.

Dietz:2022:IBP

[Die22a] Bettina Dietz. Iterative books: Posthumous publishing in eighteenthcentury botany. History of Science (UK), 60(2):166–182, June 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275320970831.

Dietz:2022:THS

[Die22b] Bettina Dietz. Towards a history of scientific publishing. History of Science (UK), 60(2):155–165, June 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/00732753221074609.

Dietz:2024:HMP

[Die24] Bettina Dietz. Herbaria as manuscripts: Philology, ethnobotany, and the textual-visual mesh of early modern botany. History of Science (UK), 62(1):??, March 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/abs/10.1177/00732753231181285.

Dikotter:1991:DRM

[Dik91] Frank Dikötter. The discourse of race and the medicalization of public and private space in modern China (1895–1949). History of Science (UK), 29(4):411–420, December 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/29/4/411.full.pdf+html.

Figueiroa:2023:SEH

[dMF23] Silvia F de M Figueirôa. Scaling down the Earth’s history: Visual materials for popular education by Nérée Boubée (1806–1862). History of Science (UK), 61(3):383–408, September 1, 2023. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753221089812.

Dobbs:1977:ERM

[Dob77] B. J. T. Dobbs. Essay review: Multiple perspectives: The Seventeenth Century scientific revolution Then and now: Reason, experiment, and mysticism in the scientific revolution. History of Science (UK), 15(4): 273–286, December 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/4/273.full.pdf+html.

Doel:2003:OHA

[Doe03] Ronald E. Doel. Oral history of American science: a forty-year review. History of Science (UK), 41(4):349–378, December 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/41/4/349.full.pdf+html.

Dolby:1977:TS

[Dol77] R. G. A. Dolby. The transmission of science. History of Science (UK), 15(1):1–43, March 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/1/1.full.pdf+html.

Dolby:1984:TVT

[Dol84] R. G. A. Dolby. Thermochemistry versus thermodynamics: The Nineteenth Century controversy. History of Science (UK), 22(4):375–400, December 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/4/
375.full.pdf+html.

Dols:1988:ERI

[Dol88] Michael W. Dols. Essay review: Islam and medicine: Health and medicine in the Islamic tradition: Change and identity. History of Science (UK), 26(4):417–425, December 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/26/4/417.full.pdf+html.

Dolan:1998:RNC

[Dol98] Brian P. Dolan. Representing novelty: Charles Babbage, Charles Lyell, and experiments in Early Victorian geology. History of Science (UK), 36(3):299–327, September 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/36/3/299.full.pdf+html.

Donat:2001:RJW

[Don01] James G. Donat. The Rev. John Wesley’s extractions from Dr Tissot: a Methodist imprimatur. History of Science (UK), 39(3):285–298, September 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/39/3/
285.full.pdf+html.

Donnelly:2014:OAM

[Don14] Kevin Donnelly. The other average man: Science workers in Quetelet’s Belgium. History of Science (UK), 52(4):401–428, December 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/52/4/401.full.
pdf+html.

Dooley:1995:CRI

[Doo95] Brendan Dooley. The communications revolution in Italian science. History of Science (UK), 33(4):469–496, December 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/33/4/469.full.pdf+html.

Dorner:2023:UUT

[Dor23] Zachary Dorner. Unnamed, not unskilled: Toward a new labor history of pharmacy. History of Science (UK), 61(4):??, December 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231187477.

Dougherty:1988:ERC

[Dou88] F. W. P. Dougherty. Essay review: Correspondence of Lichtenberg: Briefwechsel. History of Science (UK), 26(1):77–91, March 1988. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/26/1/77.full.pdf+html.

dePee:2018:CFI

[dP18] Christian de Pee. Circulation and flow: Immanent metaphors in the financial debates of Northern Song China (960–1127 CE). History of Science (UK), 56(2):168–195, June 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275317724706.

Droscher:2015:GMF

[Drö15] Ariane Dröscher. Gregor Mendel, Franz Unger, Carl Nägeli and the magic of numbers. History of Science (UK), 53(4):492–508, December 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Dolan:2010:PPP

[DT10] Brian Dolan and Allison Tillack. Pixels, patterns and problems of vision: The adaptation of computer-aided diagnosis for mammography in radiological practice in the U.S. History of Science (UK), 48(2): 227–249, June 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/48/2/
227.full.pdf+html.

Ducheyne:2005:NTA

[Duc05] Steffen Ducheyne. Newton’s training in the Aristotelian textbook tradition: From effects to causes and back. History of Science (UK), 43(3):217–237, September 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/43/3/217.full.pdf+html.

Dukes:1963:BRS

[Duk63] Cuthbert E. Dukes. Book review: a short history of clinical pathology. History of Science (UK), 2(1):172, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/2/1/172.1.full.pdf+html.

DeSmet:2001:NBG

[DV01] Rudolf De Smet and Karin Verelst. Newton’s Scholium Generale: The Platonic and Stoic legacy — Philo, Justus Lipsius and the Cambridge Platonists. History of Science (UK), 39(1):1–30, March 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/39/1/1.full.pdf+html.

Eamon:1984:ADA

[Eam84] William Eamon. Arcana disclosed: The advent of printing, the books of secrets tradition and the development of experimental science in the Sixteenth Century. History of Science (UK), 22(2):111–150, June 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/22/2/111.full.
pdf+html.

Eastwood:1986:ERM

[Eas86] Bruce Eastwood. Essay review: Medieval science illustrated: Album of science: Antiquity and the Middle Ages. History of Science (UK), 24(2):183–208, June 1986. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
24/2/183.full.pdf+html.

Ebrahimnejad:2000:TPN

[Ebr00] Hormoz Ebrahimnejad. Theory and practice in Nineteenth-Century Persian medicine: Intellectual and institutional reforms. History of Science (UK), 38(2):171–178, June 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/38/2/171.full.pdf+html.

Ebrahimnejad:2002:RMI

[Ebr02] Hormoz Ebrahimnejad. Religion and medicine in Iran: From relationship to dissociation. History of Science (UK), 40(1):91–112, March 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/40/1/91.full.
pdf+html.

Eddy:2001:GMT

[Edd01] M. D. Eddy. Geology, mineralogy and time in John Walker’s University of Edinburgh natural history lectures (1779–1803). History of Science (UK), 39(1):95–119, March 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/39/1/95.full.pdf+html.

Edge:1979:QMC

[Edg79] David Edge. Quantitative measures of communication in science: a critical review. History of Science (UK), 17(2):102–134, June 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/17/2/102.full.
pdf+html.

Edgerton:2005:CPS

[Edg05] David Edgerton. C. P. Snow as anti-historian of British science: Revisiting the technocratic moment, 1959–1964. History of Science (UK), 43(2):187–208, June 2005. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
43/2/187.full.pdf+html.

Egerton:1977:BGH

[Ege77] Frank N. Egerton. A bibliographical guide to the history of general ecology and population ecology. History of Science (UK), 15(3):189– 215, September 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/3/189.full.pdf+html.

Elwick:2003:HSD

[Elw03] James Elwick. Herbert Spencer and the disunity of the social organism. History of Science (UK), 41(1):35–72, March 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/41/1/35.full.pdf+html.

Emerson:1988:SOC

[Eme88] Roger L. Emerson. Science and the origins and concerns of the Scottish Enlightenment. History of Science (UK), 26(4):333–366, December 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/26/4/333.full.
pdf+html.

Endersby:2009:ERT

[End09] Jim Endersby. Essay review: Too much of a good thing? History of Science (UK), 47(4):475–484, December 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/47/4/475.full.pdf+html.

Eng:1978:THH

[Eng78] Erling Eng. Thomas Henry Huxley’s understanding of ‘evolution’. History of Science (UK), 16(4):291–303, December 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/16/4/291.full.pdf+html.

Eriksen:2022:MFP

[Eri22] Christoffer Basse Eriksen. Magnifying the first points of life: Harvey and Descartes on generation and scale. History of Science (UK), 60 (4):524–545, December 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/abs/10.1177/00732753211033476.

Easterby-Smith:2015:RBO

[ES15] Sarah Easterby-Smith. Reputation in a box. Objects, communication and trust in late 18th-century botanical networks. History of Science (UK), 53(2):180–208, June 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/53/2/180.full.pdf+html.

Eshet:2001:RPS

[Esh01] Dan Eshet. Rereading Priestley: Science at the intersection of theology and politics. History of Science (UK), 39(2):127–159, June 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/39/2/127.full.
pdf+html.

Esposito:2024:TUG

[Esp24] Salvatore Esposito. Thunderstorms underground: Giuseppe Saverio Poli and the electric earthquake. History of Science (UK), 62(1):??, March 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/00732753231178142.

Evans:1978:IBA

[Eva78] Gillian R. Evans. Introductions to Boethius’s “Arithmetica” of the Tenth to the Fourteenth Century. History of Science (UK), 16(1):22– 41, March 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/16/1/
22.full.pdf+html.

Everill:2023:SSC

[Eve23] Bronwen Everill. “For the services of shipwrights, coopers, and grumettas”: Freetown’s ship repair cluster in nineteenth-century Sierra Leone. History of Science (UK), 61(1):60–76, March 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
0073275320945117.

Ewen:1963:SFC

[Ewe63] A. H. Ewen. A sixth form course in the history and philosophy of science. History of Science (UK), 2(1):84–90, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/2/1/84.full.pdf+html.

Eyles:1964:ERA

[Eyl64] V. A. Eyles. Essay review: Abraham Gottlob Werner (1749–1817) and his position in the history of the mineralogical and geological sciences: On the external characters of minerals. History of Science (UK), 3 (1):102–115, March 1964. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/3/
1/102.full.pdf+html.

Eyles:1966:HGS

[Eyl66] V. A. Eyles. The history of geology: Suggestions for further research. History of Science (UK), 5(1):77–86, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/5/1/77.full.pdf+html.

Fagnani:2021:SUP

[Fag21] Martino Lorenzo Fagnani. Studying “useful plants” from Maria Theresa to Napoleon: Continuity and invisibility in agricultural science, northern Italy, the late eighteenth to early nineteenth century. History of Science (UK), 59(4):373–406, December 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275321992914.

Falconer:2021:PPJ

[Fal21] Isobel Falconer. Phases of physics in J. D. Forbes’ dissertation sixth for the Encyclopaedia Britannica (1856). History of Science (UK), 59 (1):47–72, March 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275318811443.

Fan:2000:HDT

[Fan00] Fa-Ti Fan. Hybrid discourse and textual practice: Sinology and natural history in the Nineteenth Century. History of Science (UK), 38(1):

25–56, March 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/1/
25.full.pdf+html.

Fara:1995:ATA

[Far95a] Patricia Fara. An atttractive therapy: Animal magnetism in Eighteenth-Century England. History of Science (UK), 33(2):127– 177, June 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/2/
127.full.pdf+html.

Fara:1995:FKG

[Far95b] Patricia Fara. Fit for a king? The George III Gallery at the Science Museum. History of Science (UK), 33(3):359–367, September 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/33/3/359.full.
pdf+html.

Fara:2002:ETN

[Far02] Patricia Fara. Elizabeth Tollet: a new Newtonian woman. History of Science (UK), 40(2):169–187, June 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/40/2/169.full.pdf+html.

Fara:2023:CCM

[Far23] Patricia Fara. Chemical ‘canaries’: Munitions workers in the First World War. History of Science (UK), 61(4):??, December 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231173063.

Felten:2018:HSH

[Fel18] Sebastian Felten. The history of science and the history of bureaucratic knowledge: Saxon mining, circa 1770. History of Science (UK), 56(4):403–431, December 1, 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275318792451.

Feng:2015:DCS

[Fen15] John H. Feng. Disciplining China with the scientific study of the state: Lu zhengxiang and the Chinese social and political science association, 1915–1920. History of Science (UK), 53(1):9–20, March 2015.

CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/53/1/9.full.pdf+html.

Feudtner:1993:MDR

[Feu93] Chris Feudtner. “Minds the dead have ravished”: 1 shell shock, history, and the ecology of disease-systems. History of Science (UK), 31(4):377–420, December 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/31/4/377.full.pdf+html.

Figala:1973:BRP

[Fig73a] Karin Figala. Book review: Pharmaceutical chemistry: Geschichte der pharmazeutischen Chemie (German) [History of pharmaceutical chemistry]. History of Science (UK), 11(2):150–151, June 1973. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/11/2/150.full.pdf+html.

Figala:1973:ERP

[Fig73b] Karin Figala. Essay review: Projekt der Katalogisierung von alchemischen Handschriften im deutschen Sprachgebiet. (German) [Project of cataloging of alchemical manuscripts in the Germanlanguage area]. History of Science (UK), 11(1):70, March 1973. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/11/1/70.full.pdf+html.

Figlio:1975:TPP

[Fig75] Karl M. Figlio. Theories of perception and the physiology of mind in the Late Eighteenth Century. History of Science (UK), 13(3):177–212, September 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/3/
177.full.pdf+html.

Figlio:1976:MOH

[Fig76] Karl M. Figlio. The metaphor of organization: an historiographical perspective on the bio-medical sciences of the Early Nineteenth Century. History of Science (UK), 14(1):17–53, March 1976. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/14/1/17.full.pdf+html.

Figala:1979:ERN

[Fig79] Karin Figala. Essay review: Newton and the scholars: Newton and Newtoniana 1672–1975. a bibliography. History of Science (UK), 17 (2):147–149, June 1979. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
17/2/147.full.pdf+html.

Findlen:1993:CER

[Fin93] Paula Findlen. Controlling the experiment: Rhetoric, court patronage and the experimental method of Francesco Redi. History of Science (UK), 31(1):35–64, March 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/31/1/35.full.pdf+html.

Finkelstein:2000:CKS

[Fin00] Gabriel Finkelstein. “Conquerors of the Künlün”? The Schlagintweit Mission to High Asia, 1854–57. History of Science (UK), 38(2):179– 218, June 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/2/
179.full.pdf+html.

Fisher:1973:ERN

[Fis73] N. W. Fisher. Essay review: The nature of the chemical atom: The history of valency. History of Science (UK), 11(1):53–61, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/11/1/53.full.pdf+html.

Fisher:1982:ACHa

[Fis82a] Nicholas Fisher. Avogadro, the chemists, and historians of chemistry: Part 1. History of Science (UK), 20(2):77–102, June 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/20/2/77.full.pdf+html.

Fisher:1982:ACHb

[Fis82b] Nicholas Fisher. Avogadro, the chemists, and historians of chemistry: Part 2. History of Science (UK), 20(3):212–231, September 1982. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/20/3/212.full.pdf+html.

Flemming:2000:BRW

[Fle00] Rebecca Flemming. Book review: Women in Greek antiquity, Hippocrates’ woman: Reading the female body in Ancient Greece. History of Science (UK), 38(2):243–245, June 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/38/2/243.full.pdf+html.

Fleetwood:2018:NFT

[Fle18] Lachlan Fleetwood. “No former travellers having attained such a height on the Earth’s surface”: Instruments, inscriptions, and bodies in the Himalaya, 1800–1830. History of Science (UK), 56(1):3– 34, March 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://journals.sagepub.com/doi/full/
10.1177/0073275317732254.

Fleming:2021:EEM

[Fle21] Martha Fleming. Embodied ephemeralities: Methodologies and historiographies for investigating the display and spatialization of science and technology in the twentieth century. History of Science (UK), 59 (2):197–219, June 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275319858528.

Florensa:2022:SSP

[Flo22] Clara Florensa. Struggling for survival: The popularization of Darwinism and the elite’s fight for power in Franco’s Spain (1939-1967). History of Science (UK), 60(3):348–382, September 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753221091032.

Froggatt:1971:GLA

[FN71] P. Froggatt and N. C. Nevin. Galton’s ‘law of ancestral heredity’: Its influence on the early development of human genetics. History of Science (UK), 10(1):1–27, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/10/1/1.full.pdf+html.

Florensa:2022:ISP

[FNG22] Clara Florensa and Agustí Nieto-Galan. Introduction: Science popularization, dictatorships, and democracies. History of Science (UK), 60(3):329–347, September 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753221091029.

Forbes:1974:HSF

[For74] Eric G. Forbes. History of science in the Federal Republic of Germany. History of Science (UK), 12(2):147–151, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/12/2/147.full.pdf+html.

Forrester:1980:ERM

[For80] John Forrester. Essay review: Michel Foucault and the history of psychoanalysis: The order of things. History of Science (UK), 18(4): 286–303, December 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/4/286.full.pdf+html.

Fores:1983:SNP

[For83] Michael Fores. Science and the ‘neolithic paradox’. History of Science (UK), 21(2):141–163, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/21/2/141.full.pdf+html.

Fores:1984:CSS

[For84a] Michael Fores. Constructed science and the Seventeenth Century ‘revolution’. History of Science (UK), 22(3):217–244, September 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/22/3/217.full.
pdf+html.

Forrester:1984:ERP

[For84b] John Forrester. Essay review: The pasteurization of France: Les microbes. guerre et paix, suivi de irréductions. History of Science (UK), 22(4):425–427, December 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/22/4/425.full.pdf+html.

Fores:1985:NHC

[For85] Michael Fores. Newton on a horse: a critique of the historiographies of ‘technology’ and ‘modernity’. History of Science (UK), 23(4):351–378, December 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/4/
351.full.pdf+html.

Fores:1994:HMD

[For94a] Michael Fores. Homo et machina: Of divided labour, revolutions and a history of marvels. History of Science (UK), 32(1):63–87, March 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/32/1/63.full.pdf+html.

Forgan:1994:ADM

[For94b] Sophie Forgan. The architecture of display: Museums, universities and objects in Nineteenth-Century Britain. History of Science (UK), 32(2):139–162, June 1994. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
32/2/139.full.pdf+html.

Fox:1983:ERR

[Fox83] Robert Fox. Essay review: Referencing the history of science: Dictionary of the history of science. History of Science (UK), 21(2):213– 214, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/2/
213.full.pdf+html.

Fox:1988:STC

[Fox88] Nicholas J. Fox. Scientific theory choice and social structure: The case of Joseph Lister’s antisepsis, humoral theory and asepsis. History of Science (UK), 26(4):367–397, December 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/26/4/367.full.pdf+html.

Frank:1973:SMUb

[Fra73a] Robert G. Frank, Jr. Science, medicine and the universities of Early Modern England: Background and sources, Part 2. History of Science (UK), 11(4):239–269, December 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/11/4/239.full.pdf+html.

Frank:1973:SMUa

[Fra73b] Robert G. Frank, Jr. Science, medicine and the universities of Early Modern England: Background and sources, Part I. History of Science (UK), 11(3):194–216, September 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/11/3/194.full.pdf+html.

Frangsmyr:1974:SSE

[Frä74] Tore Frängsmyr. Swedish science in the Eighteenth Century. History of Science (UK), 12(1):29–42, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/12/1/29.full.pdf+html.

Frank:2024:NCZ

[Fra24] Mark E. Frank. National climate: Zhu Kezhen and the framing of the atmosphere in modern China. History of Science (UK), 62(4): 562–590, December 2024. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753231157453.

French:1971:SPS

[Fre71] Richard D. French. Some problems and sources in the foundations of modern physiology in Great Britain. History of Science (UK), 10(1): 28–55, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/10/1/
28.full.pdf+html.

Freudenthal:1995:SMJ

[Fre95] Gad Freudenthal. Science in the Medieval Jewish culture of southern France. History of Science (UK), 33(1):23–58, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/33/1/23.full.pdf+html.

Friesen:2003:APD

[Fri03] John Friesen. Archibald Pitcairne, David Gregory and the Scottish origins of English Tory Newtonianism, 1688–1715. History of Science (UK), 41(2):163–191, June 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/41/2/163.full.pdf+html.

Fancy:2023:CDE

[FSB^{+}23] Nahyan Fancy, Justin Stearns, Sonja Brentjes, A. Tunç Şen, Scott Trigg, Noah Gardiner, Nükhet Varlık Rutgers, Matthew Melvin-Koushki, and S. Nomanul Haq. Current debates and emerging trends in the history of science in premodern Islamicate societies. History of Science (UK), 61(2):123–178, June 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753231154690.

Felismino:2016:PID

[FTC16] David Felismino, Conceição Tavares, and Ana Carneiro. The power of islands and of discipleship: Francisco de Arruda Furtado (1854–1887) and the making of a disciple of Darwin. History of Science (UK), 54 (2):138–168, June 2016. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
54/2/138.full.pdf+html.

Fyfe:2022:SHL

[Fyf22] Aileen Fyfe. Self-help for learned journals: Scientific societies and the commerce of publishing in the 1950s. History of Science (UK), 60(2):

255–279, June 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275321999901.

Gomes:2025:FMC

[GÁ25] Inês Gomes and Frederico Ágoas. Fire management and community restraint: The rise of forestry science and the governance of commons. History of Science (UK), 63(1):52–72, March 2025. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753241304144.

Gao:2015:PCP

[Gao15] Zhipeng Gao. Pavlovianism in China: Politics and differentiation across scientific disciplines in the Maoist era. History of Science (UK), 53(1):57–85, March 2015. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
53/1/57.full.pdf+html.

Gascoigne:1985:USR

[Gas85] John Gascoigne. The universities and the scientific revolution: The case of Newton and restoration Cambridge. History of Science (UK), 23(4):391–434, December 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/23/4/391.full.pdf+html.

Gaukroger:1979:ERS

[Gau79] S. W. Gaukroger. Essay review: Science and matter: Scientific materialism in Nineteenth Century Germany. History of Science (UK), 17(3):214–216, September 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/17/3/214.full.pdf+html.

Gauvin:2006:AMD

[Gau06] Jean-François Gauvin. Artisans, machines, and Descartes’s organon. History of Science (UK), 44(2):187–216, June 2006. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/44/2/187.full.pdf+html.

Gay:2003:SOL

[Gay03] Hannah Gay. Science and opportunity in London, 1871–85: The diary of Herbert McLeod. History of Science (UK), 41(4):427–458, December 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/41/4/427.full.
pdf+html.

Gay:2008:SSC

[Gay08] Hannah Gay. Science, scientific careers and social exchange in London: The diary of Herbert McLeod, 1885–1900. History of Science (UK), 46(4):457–496, December 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/46/4/457.full.pdf+html.

Garcia:2010:STG

[GC10] Francisco Vázquez García and Richard Cleminson. Subjectivities in transition: Gender and sexual identities in cases of ‘sex change’ and ‘hermaphroditism’ in Spain, c. 1500–1800. History of Science (UK), 48(1):1–38, March 2010. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
48/1/1.full.pdf+html.

Gal:2005:AIS

[GCM05] Ofer Gal and Raz Chen-Morris. The archaeology of the inverse square law: (1) metaphysical images and mathematical practices. History of Science (UK), 43(4):391–414, December 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/43/4/391.full.pdf+html.

Gal:2006:AIS

[GCM06] Ofer Gal and Raz Chen-Morris. The archaeology of the inverse square law: (2) the use and non-use of mathematics. History of Science (UK), 44(1):49–67, March 2006. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
44/1/49.full.pdf+html.

Geison:1981:SCE

[Gei81] Gerald L. Geison. Scientific change, emerging specialties, and research schools. History of Science (UK), 19(1):20–40, March 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/19/1/20.full.pdf+html.

George:1980:SBD

[Geo80] Wilma George. Sources and background to discoveries of new animals in the Sixteenth and Seventeenth Centuries. History of Science (UK), 18(2):79–104, June 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/2/79.full.pdf+html.

George:1985:ERD

[Geo85] Wilma George. Essay review: Darwin on trial: Dimensions of Darwinism, the wider domain of evolutionary thought. History of Science (UK), 23(1):122–125, March 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/23/1/122.full.pdf+html.

Grattan-Guinness:1990:DHS

[GG90] I. Grattan-Guinness. Does history of science treat of the history of science? The case of mathematics. History of Science (UK), 28(2): 149–173, June 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/28/2/
149.full.pdf+html.

Gay:1997:BSS

[GG97] Hannah Gay and John W. Gay. Brothers in science: Science and fraternal culture in Nineteenth-Century Britain. History of Science (UK), 35(4):425–453, December 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/35/4/425.full.pdf+html.

Grattan-Guinness:2004:DTR

[GG04] I. Grattan-Guinness. Decline, Then recovery: an overview of activity in the history of mathematics during the Twentieth century. History of Science (UK), 42(3):279–312, September 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/42/3/279.full.pdf+html.

Gorman:2016:LIG

[GGN16] Hugh S. Gorman, Valoree S. Gagnon, and Emma S. Norman. Local impacts, global sources: The governance of boundary-crossing chemicals. History of Science (UK), 54(4):443–459, December 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://journals.sagepub.com/doi/full/10.1177/
0073275316681804.

Gabriel:2020:CTO

[GH20] Joseph M. Gabriel and Bennett Holman. Clinical trials and the origins of pharmaceutical fraud: Parke, Davis & Company, virtue epistemology, and the history of the fundamental antagonism. History of Science (UK), 58(4):533–558, December 1, 2020. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275320942435.

Giannini:2015:GTF

[Gia15] Giulia Giannini. Gianantonio Tadini and falling bodies: a new documentary source for the reconstruction of the history of experimental proofs on the Earth’s rotation. History of Science (UK), 53(3):320– 337, September 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Gibson:2012:BAE

[Gib12] Susannah Gibson. On being an animal, or, the Eighteenth-Century zoophyte controversy in Britain. History of Science (UK), 50(4):453– 476, December 2012. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
50/4/453.full.pdf+html.

Gilman:1983:ERP

[Gil83] Sander L. Gilman. Essay review: Photography of the insane: Invention de l’hysterie: Charcot et l’iconographie photographique de la Salpêtrière. History of Science (UK), 21(4):432–434, December 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/21/4/432.full.
pdf+html.

Gillin:2020:MMG

[Gil20] Edward J. Gillin. Mechanics and mathematicians: George Biddell Airy and the social tensions in constructing time at Parliament, 1845–1860. History of Science (UK), 58(3):301–325, September 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275319879279.

Gillin:2024:MKN

[Gil24] Edward J. Gillin. Mining knowledge: Nineteenth-century Cornish electrical science and the controversies of clay. History of Science (UK), 62(2):202–226, June 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753231189965.

Gingerich:2000:BRD

[Gin00] Owen Gingerich. Book review: Direction and distance to Mecca, world-maps for finding the direction and distance to Mecca: Innovation and tradition in Islamic science. History of Science (UK), 38(2): 245–247, June 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/2/
245.full.pdf+html.

Gingras:2001:WDM

[Gin01] Yves Gingras. What did mathematics do to physics? History of Science (UK), 39(4):383–416, December 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/39/4/383.full.pdf+html.

Gingras:2008:CCS

[Gin08] Yves Gingras. The collective construction of scientific memory: The Einstein–Poincaré connection and its discontents, 1905–2005. History of Science (UK), 46(1):75–114, March 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/46/1/75.full.pdf+html.

Gliboff:1999:GML

[Gli99] Sander Gliboff. Gregor Mendel and the laws of evolution. History of Science (UK), 37(2):217–235, June 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/37/2/217.full.pdf+html.

Gluch:2024:TSC

[Glu24] Sibylle Gluch. Timing the stars: Clocks and complexities of precision in eighteenth-century observatories. History of Science (UK), 62(3): 329–365, September 2024. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753231193819.

Graber:1989:DDF

[GM89] Robert Bates Graber and Lynate Pettengill Miles. In defence of Darwin’s father. History of Science (UK), 27(1):97–102, March 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/27/1/97.full.pdf+html.

Gamito-Marques:2018:DMI

[GM18] Daniel Gamito-Marques. Defending metropolitan identity through colonial politics: The role of Portuguese naturalists (1870–91). History of Science (UK), 56(2):224–253, June 2018. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//journals.sagepub.com/doi/full/10.1177/0073275317722240.

See corrigendum [Ano21].

Golinski:1986:ERS

[Gol86] J. V. Golinski. Essay review: Science in the Enlightenment: Science and the Enlightenment. History of Science (UK), 24(4):411–424, December 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/4/
411.full.pdf+html.

Golinski:1987:HMI

[Gol87] Jan Golinski. Hélène Metzger and the interpretation of Seventeenth Century chemistry. History of Science (UK), 25(1):85–97, March 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/25/1/85.full.
pdf+html.

Golinski:1990:ERE

[Gol90] Jan Golinski. Essay review: Experiment in scientific practice, the uses of experiment: Studies in the natural sciences. History of Science (UK), 28(2):203–209, June 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/28/2/203.full.pdf+html.

Golinski:2002:CSM

[Gol02] Jan Golinski. The care of the self and the masculine birth of science. History of Science (UK), 40(2):125–145, June 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/40/2/125.full.pdf+html.

Golinski:2003:LE

[Gol03] Jan Golinski. A legacy of enlightenment. History of Science (UK), 41(3):345–350, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/41/3/345.full.pdf+html.

Golinski:2011:SER

[Gol11] Jan Golinski. Science in the Enlightenment, revisited. History of Science (UK), 49(2):217–231, June 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/49/2/217.full.pdf+html.

Golinski:2016:RKR

[Gol16] Jan Golinski. Richard Kirwan and the Royal Irish Academy: Provincial metropolitanism and the crisis of the 1790s. History of Science (UK), 54(3):257–275, September 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/54/3/257.full.pdf+html.

Gold:2019:AEG

[Gol19] Meira Gold. Ancient Egypt and the geological antiquity of man, 1847–1863. History of Science (UK), 57(2):194–230, June 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318795944.

Goodman:1976:BRD

[Goo76] D. C. Goodman. Book review: Dutrochet and Physiology: Henri Dutrochet (Henri du Trochet 1776–1847). Le Matérialisme Mécaniste et la Physiologie Générale. History of Science (UK), 14(2):143– 144, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/14/2/
143.full.pdf+html.

Goodman:1983:SCS

[Goo83] David Goodman. Science and the clergy in the Spanish Enlightenment. History of Science (UK), 21(2):111–140, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/21/2/111.full.pdf+html.

Goodey:2004:IAS

[Goo04] C. F. Goodey. Intellectual ability and speed of performance: Galen to Galton. History of Science (UK), 42(4):465–495, December 2004. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/42/4/465.full.pdf+html.

Gooday:2008:LEA

[Goo08] Graeme Gooday. Liars, experts and authorities. History of Science (UK), 46(4):431–456, December 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/46/4/431.full.pdf+html.

Goodrum:2009:HHO

[Goo09] Matthew R. Goodrum. The history of human origins research and its place in the history of science: Research problems and historiography.

History of Science (UK), 47(3):337–357, September 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/47/3/337.full.pdf+html.

Goodman:2023:AGF

[Goo23] Dena Goodman. Affective geographies: Family and friendship in the production of scientific knowledge. History of Science (UK), 61(2): 236–265, June 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753211059957.

Gordin:2015:HOC

[Gor15] Michael D. Gordin. Hydrogen oxygenovich: Crafting Russian as a language of science in the late nineteenth century. History of Science (UK), 53(4):417–437, December 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Gould:1974:ERB

[Gou74] Stephen Jay Gould. Essay review: On biological and social determinism: Genetics and American society. History of Science (UK), 12(3): 212–220, September 1974. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
12/3/212.full.pdf+html.

Gouyon:2011:BNH

[Gou11a] Jean-Baptiste Gouyon. The BBC natural history unit: Instituting natural history film-making in Britain. History of Science (UK), 49(4):425–451, December 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/49/4/425.full.pdf+html.

Gouyon:2011:KAF

[Gou11b] Jean-Baptiste Gouyon. From Kearton to Attenborough: Fashioning the telenaturalist’s identity. History of Science (UK), 49(1):25– 60, March 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/49/1/
25.full.pdf+html.

Goucher:2025:GMI

[Gou25] Candice Goucher. Gendering the memory of iron: Theft, lineage, and African metallurgists in the Atlantic world. History of Science (UK), 63(1):3–28, March 2025. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753251323307.

Gowing:1973:BRS

[Gow73] Margaret Gowing. Book review: Science and society: Science and society 1600–1900. History of Science (UK), 11(2):143–145, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/11/2/143.full.
pdf+html.

Geroulanos:2018:EVI

[GP18] Stefanos Geroulanos and Jamie Phillips. Eurasianism versus IndoGermanism: Linguistics and mythology in the 1930s’ controversies over European prehistory. History of Science (UK), 56(3):343–378, September 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://journals.sagepub.com/doi/full/
10.1177/0073275318776422.

Guerrero:2024:SRN

[GP24] Saul Guerrero and David Pretel. Silver refining in the New World: a singularity in the history of useful knowledge. History of Science (UK), 62(2):175–201, June 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753231185027.

Gavroglu:2008:STE

[GPP^{+}08] Kostas Gavroglu, Manolis Patiniotis, Faidra Papanelopoulou, Ana Simões, Ana Carneiro, Maria Paula Diogo, José Ramón Bertomeu Sánchez, Antonio García Belmar, and Agustí Nieto-Galan. Science and technology in the European periphery: Some historiographical reflections. History of Science (UK), 46(2):153–175, June 2008. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/46/2/153.full.pdf+html.

Gil-Riano:2022:RMR

[GR22] Sebastián Gil-Riaño. Risky migrations: Race, Latin eugenics, and Cold War development in the International Labor Organization’s Puno-Tambopata project in Peru, 1930–60. History of Science (UK), 60(1):41–68, March 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/00732753211035283.

Grant:1978:ALM

[Gra78] Edward Grant. Aristotelianism and the longevity of the medieval world view. History of Science (UK), 16(2):93–106, June 1978. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/16/2/93.full.pdf+html.

Grant:1987:WIT

[Gra87] Edward Grant. Ways to interpret the terms ‘Aristotelian’ and ‘Aristotelianism’ in medieval and Renaissance natural philosophy. History of Science (UK), 25(4):335–358, December 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/25/4/335.full.pdf+html.

Grant:2011:HTI

[Gra11] Edward Grant. How theology, imagination, and the spirit of inquiry shaped natural philosophy in the Late Middle Ages. History of Science (UK), 49(1):89–108, March 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/49/1/89.full.pdf+html.

Grosslight:2013:SSB

[Gro13] Justin Grosslight. Small skills, big networks: Marin Mersenne as mathematical intelligencer. History of Science (UK), 51(3):337–374, September 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/51/3/
337.full.pdf+html.

Gruber:1964:ERD

[Gru64] Jacob W. Gruber. Essay review: Darwinism and its critics: Darwin and the modern world view. History of Science (UK), 3(1): 115–123, March 1964. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/3/
1/115.full.pdf+html.

Gruber:1981:RBA

[Gru81] Howard E. Gruber. On the relation between AHA experiences and the construction of ideas. History of Science (UK), 19(1):41–59, March 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/19/1/41.full.
pdf+html.

Gil-Riano:2022:IRS

[GRW22] Sebastián Gil-Riaño and Sarah Walsh. Introduction: Race science in the Latin world. History of Science (UK), 60(1):4–17, March 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
00732753211053517.

Godin:2016:CIR

[GS16] Benoît Godin and Désirée Schauz. The changing identity of research: a cultural and conceptual history. History of Science (UK), 54(3):276– 306, September 2016. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
54/3/276.full.pdf+html.

Gordin:2015:ILS

[GT15] Michael D. Gordin and Kostas Tampakis. Introduction: The languages of scientists. History of Science (UK), 53(4):365–377, December 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Guarnieri:1991:ATC

[Gua91] Patrizia Guarnieri. Alienists on trial: Conflict and convergence between psychiatry and law (1876–1913). History of Science (UK), 29(4):393–410, December 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/29/4/393.full.pdf+html.

Guerra:1969:AST

[Gue69] Francisco Guerra. Aztec science and technology. History of Science (UK), 8(1):32–52, March 1969. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/8/1/32.full.pdf+html.

Guerlac:1979:SAF

[Gue79] Henry Guerlac. Some areas for further Newtonian studies. History of Science (UK), 17(2):75–101, June 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/17/2/75.full.pdf+html.

Guerrini:2016:GK

[Gue16] Anita Guerrini. The ghastly kitchen. History of Science (UK), 54(1): 71–97, March 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/54/1/
71.full.pdf+html.

Guicciardini:2017:PNO

[Gui17] Niccolò Guicciardini. The publication of Newton’s Opera Omnia in Geneva and Lausanne (1739–1761): a chapter in the reception of Newtonianism. History of Science (UK), 55(4):457–489, December 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://journals.sagepub.com/doi/full/10.1177/
0073275317716369.

Guntau:1978:EGS

[Gun78] Martin Guntau. The emergence of geology as a scientific discipline. History of Science (UK), 16(4):280–290, December 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/16/4/280.full.pdf+html.

Gurevitch:2021:UUK

[Gur21] Eric Moses Gurevitch. The uses of useful knowledge and the languages of vernacular science: Perspectives from southwest India. History of Science (UK), 59(3):256–286, September 1, 2021. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275320931976.

Hajek:2015:FSS

[Haj15] Kim M. Hajek. The fear of simulation: Scientific authority in late 19th-century French disputes over hypnotism. History of Science (UK), 53(3):237–263, September 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Hakfoort:1991:MSH

[Hak91] C. Hakfoort. The missing syntheses in the historiography of science. History of Science (UK), 29(2):207–216, June 1991. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/29/2/207.full.pdf+html.

Hakfoort:1995:HSC

[Hak95] Casper Hakfoort. The historiography of scientism: a critical review. History of Science (UK), 33(4):375–395, December 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/33/4/375.full.pdf+html.

Hall:1963:MRS

[Hal63] A. Rupert Hall. Merton revisited or science and society in the Seventeenth Century. History of Science (UK), 2(1):1–16, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/2/1/1.full.pdf+html.

Hall:1966:SHR

[Hal66] Marie Boas Hall. Sources for the history of the Royal Society in the Seventeenth Century. History of Science (UK), 5(1):62–76, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/5/1/62.
full.pdf+html.

Hall:1968:ERN

[Hal68] A. Rupert Hall. Essay review: Newton at the turn of the century: The correspondence of Isaac Newton. History of Science (UK), 7(1): 134–143, March 1968. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/7/
1/134.full.pdf+html.

Hall:1970:ERR

[Hal70] Marie Boas Hall. Essay review: Robert Boyle: The life of the Honourable Robert Boyle F.R.S. History of Science (UK), 9(1):139, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/9/1/
139.full.pdf+html.

Hall:1973:ERO

[Hal73a] A. Rupert Hall. Essay review: The 1969 Oklahoma symposium: Perspectives in the history of science and technology. History of Science (UK), 11(2):124–129, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/11/2/124.full.pdf+html.

Hall:1973:ERS

[Hal73b] Marie Boas Hall. Essay review: Science, humanism and society: Science in a Renaissance society. History of Science (UK), 11(1):49– 52, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/1/
49.full.pdf+html.

Hall:1975:NFN

[Hal75] A. Rupert Hall. Newton in France: a new view. History of Science (UK), 13(4):233–250, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/13/4/233.full.pdf+html.

Hall:1983:W

[Hal83] A. Rupert Hall. On Whiggism. History of Science (UK), 21(1):45– 59, March 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/1/
45.full.pdf+html.

Hall:1991:CBAa

[Hal91a] J. J. Hall. The classification of birds, in Aristotle and early modern naturalists (i). History of Science (UK), 29(2):111–151, June 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/29/2/111.full.
pdf+html.

Hall:1991:CBAb

[Hal91b] J. J. Hall. The classification of birds, in Aristotle and early modern naturalists (II). History of Science (UK), 29(3):223–243, September 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/29/3/223.full.
pdf+html.

Hall:2002:PEN

[Hal02] A. Rupert Hall. Pitfalls in the editing of Newton’s papers. History of Science (UK), 40(4):407–424, December 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/40/4/407.full.pdf+html.

Hall:2015:ESE

[Hal15] James R. Hall. Encountering snakes in early Victorian London: The first reptile house at the Zoological Gardens. History of Science (UK), 53(3):338–361, September 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Hamm:1993:BSA

[Ham93] Ernst P. Hamm. Bureaucratic statistik or actualism? K. E. A. Von Hoff’s history and the history of geology. History of Science (UK), 31 (2):151–176, June 1993. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
31/2/151.full.pdf+html.

Hamlin:2011:BCS

[Ham11] Christopher Hamlin. Bacteriology as a cultural system: Analysis and its discontents. History of Science (UK), 49(3):269–298, September 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/49/3/269.full.
pdf+html.

Hankins:1970:ERC

[Han70] Thomas L. Hankins. Essay review: The concept of hard bodies in the history of physics: The conflict between atomism and conservation theory 1644 to 1860. History of Science (UK), 9(1):119–128, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/9/1/119.full.
pdf+html.

Hankins:1979:DBU

[Han79] Thomas L. Hankins. In defence of biography: The use of biography in the history of science. History of Science (UK), 17(1):1– 16, March 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/1/1.
full.pdf+html.

Harre:1962:BRN

[Har62] R. Harré. Book review: The natural philosophy of time. History of Science (UK), 1(1):121–123, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/121.full.pdf+html.

Harre:1963:BRPa

[Har63a] R. Harré. Book review: Philosophical impact of contemporary physics. History of Science (UK), 2(1):168–169, March 1963. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/2/1/168.full.pdf+html.

Harre:1963:BRPb

[Har63b] R. Harré. Book review: Physical science: Men and concepts. History of Science (UK), 2(1):173, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/2/1/173.full.pdf+html.

Harre:1963:BRC

[Har63c] R. Harré. Book review: The concept and the Rôle of the model in mathematics and natural and social sciences. History of Science (UK), 2(1):172–173, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/2/1/172.3.full.pdf+html.

Harman:1983:ERB

[Har83] P. M. Harman. Essay review: Biography: Fact or fiction?: Night thoughts of a classical physicist. History of Science (UK), 21(4):430– 432, December 1983. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
21/4/430.full.pdf+html.

Harrison:1996:MMM

[Har96a] Mark Harrison. Medicine and the management of modern warfare. History of Science (UK), 34(4):379–410, December 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/34/4/379.full.pdf+html.

Harwood:1996:WCB

[Har96b] Jonathan Harwood. Weimar culture and biological theory: a study of Richard Woltereck (1877–1944). History of Science (UK), 34(3):347– 377, September 1996. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
34/3/347.full.pdf+html.

Harwood:1998:ERH

[Har98] Jonathan Harwood. Essay review: Holistic theories of mind in Early Twentieth-century Germany, reenchanted science: Holism in German culture from Wilhelm II to Hitler, gestalt psychology in German culture, 1890–1967: Holism and the quest for objectivity. History of Science (UK), 36(4):485–498, December 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/4/485.full.pdf+html.

Harrison:2002:VEM

[Har02] Peter Harrison. Voluntarism and early modern science. History of Science (UK), 40(1):63–89, March 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/40/1/63.full.pdf+html. See response [Hen09].

Harrison:2009:VOM

[Har09] Peter Harrison. Voluntarism and the origins of modern science: a reply to John Henry. History of Science (UK), 47(2):223–231, June 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/47/2/223.full.
pdf+html.

Hassani:1979:ERA

[Has79] A. M. Hassani. Essay review: Arab scientists revisited: Ibn Ash-Shatir and Taqi Ed-Din: The life and work of Ibn Al-Shatir, an Arab astronomer of the 14th Century, Taqi Al-Din and Arabic mechanical engineering, with the sublime methods of spiritual machines. History of Science (UK), 17(2):135–140, June 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/17/2/135.full.pdf+html.

Hassani:1987:ERD

[Has87] A. M. Hassani. Essay review: Darwin among the Arabs: Western science in the Arab world: The impact of Darwinism, 1860–1930. History of Science (UK), 25(3):321–323, September 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/25/3/321.full.pdf+html.

Hawkins:2011:PPT

[Haw11] Michael Hawkins. Piss profits: Thomas Willis, his Diatribae Duae and the formation of his professional identity. History of Science (UK), 49(1):1–24, March 2011. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
49/1/1.full.pdf+html.

Heilbron:1968:QHA

[Hei68] J. L. Heilbron. Quantum historiography and the archive for history of quantum physics. History of Science (UK), 7(1):90–111, March 1968. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/7/1/90.full.pdf+html.

Heimann:1973:NNP

[Hei73] P. M. Heimann. Newtonian natural philosophy and the scientific revolution. History of Science (UK), 11(1):1–7, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/11/1/1.full.pdf+html.

Heimann:1978:ERS

[Hei78] P. M. Heimann. Essay review: Science and the English Enlightenment: The Newtonians and the English revolution 1689–1720, reason, ridicule and religion. the Age of Enlightenment in England 1660–1750. History of Science (UK), 16(2):143–151, June 1978. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/16/2/143.full.pdf+html.

Heimann:1979:ERN

[Hei79] P. M. Heimann. Essay review: Newton and continental controversy: The correspondence of Isaac Newton. History of Science (UK), 17(2): 144–147, June 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/2/
144.full.pdf+html.

Heimann:1981:ERE

[Hei81] P. M. Heimann. Essay review: Electricity and natural philosophy: Electricity in the 17th and 18th centuries. a study of early modern physics. History of Science (UK), 19(3):219–222, September 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/19/3/219.full.
pdf+html.

Hellman:1964:BRR

[Hel64] C. Doris Hellman. Book review: La révolution astronomique: Copernic, Kepler, Borelli. History of Science (UK), 3(1):134–139, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/3/1/134.full.
pdf+html.

Hellawell:2020:BMP

[Hel20] Philippa Hellawell. “The best and most practical philosophers”: Seamen and the authority of experience in early modern science. History of Science (UK), 58(1):28–50, March 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275319842425.

Hendry:1980:WCQ

[Hen80] John Hendry. Weimar culture and quantum causality. History of Science (UK), 18(3):155–180, September 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/18/3/155.full.pdf+html.

Henry:1982:THA

[Hen82] John Henry. Thomas Harriot and atomism: a reappraisal. History of Science (UK), 20(4):267–296, December 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/20/4/267.full.pdf+html.

Hendry:1983:ERU

[Hen83] John Hendry. Essay review: Understanding science: T. S. Kuhn and social science. History of Science (UK), 21(4):415–424, December 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/21/4/415.full.
pdf+html.

Henry:1986:OQE

[Hen86] John Henry. Occult qualities and the experimental philosophy: Active principles in pre-Newtonian matter theory. History of Science (UK), 24(4):335–381, December 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/24/4/335.full.pdf+html.

Henry:1993:ERH

[Hen93] John Henry. Essay review: Henry more and Newton’s gravity, Henry more: Magic, religion and experiment. History of Science (UK), 31(1): 83–97, March 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/31/1/
83.full.pdf+html.

Henry:2008:FRO

[Hen08] John Henry. The fragmentation of Renaissance occultism and the decline of magic. History of Science (UK), 46(1):1–48, March 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/46/1/1.full.pdf+html.

Henry:2009:VTO

[Hen09] John Henry. Voluntarist theology at the origins of modern science: a response to Peter Harrison. History of Science (UK), 47(1):79– 113, March 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/1/
79.full.pdf+html. See [Har02].

Herivel:1962:BRN

[Her62a] J. W. Herivel. Book review: Newtonian science. History of Science (UK), 1(1):112–113, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/112.full.pdf+html.

Herivel:1962:BRS

[Her62b] J. W. Herivel. Book review: Science since Babylon. History of Science (UK), 1(1):114–115, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/114.full.pdf+html. See comment [Stu77].

Hesse:1962:BRW

[Hes62] Mary B. Hesse. Book review: William Herschel, the development of physical theory in the Middle Ages, pioneers of prehistory in England, the image of Newton and Locke in the age of reason, social anthropology, the structure of chemistry, theories and things. History of Science (UK), 1(1):115–117, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/115.full.pdf+html.

Hesse:1963:ERM

[Hes63] Mary Hesse. Essay review: Measurement in science: Quantification: a history of the meaning of measurement in the natural and social sciences. History of Science (UK), 2(1):152–155, March 1963. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/2/1/152.full.pdf+html.

Hesse:1964:ERC

[Hes64] Mary Hesse. Essay review: Changing views of matter: The concept of matter. History of Science (UK), 3(1):79–84, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/3/1/79.full.pdf+html.

Hewitt:2016:RLH

[Hew16] Kim Hewitt. Rehabilitating LSD history in postwar America: Dilworth Wayne Woolley and the serotonin hypothesis of mental illness. History of Science (UK), 54(3):307–330, September 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/54/3/307.full.pdf+html.

Hall:1973:ERMb

[HH73] A. Rupert Hall and Marie Boas Hall. Essay review: Multi-volume works in progress (3): The correspondence of Henry Oldenburg (?1618–77). History of Science (UK), 11(3):236–237, September 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/11/3/236.full.
pdf+html.

Higgs:1985:CHJ

[Hig85] Edward Higgs. Counting heads and jobs: Science as an occupation in the Victorian census. History of Science (UK), 23(4):335–349, December 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/23/4/335.full.
pdf+html.

Hess:2010:CSM

[HM10] Volker Hess and J. Andrew Mendelsohn. Case and series: Medical knowledge and paper technology, 1600–1900. History of Science (UK), 48(3–4):287–314, September/December 2010. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Hoch:1990:ERS

[Hoc90a] Paul K. Hoch. Essay review: a socialized history of science: Science as power: Discourse and ideology in modern society, science, technology and the military, scientific knowledge socialized. History of Science (UK), 28(2):193–202, June 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/28/2/193.full.pdf+html.

Hoch:1990:ERH

[Hoc90b] Paul K. Hoch. Essay review: An historical philosophy of science? History of Science (UK), 28(2):211–219, June 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/28/2/211.full.pdf+html.

Hoch:1991:ERS

[Hoc91] Paul K. Hoch. Essay review: Science in context — but which context?, literature and science as modes of expression. History of Science (UK), 29(2):217–221, June 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/29/2/217.full.pdf+html.

Holmes:1975:ERB

[Hol75] Frederic L. Holmes. Essay review: Biochemistry and the historian: Molecules and life. History of Science (UK), 13(2):114–121, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/13/2/114.full.
pdf+html.

Holmes:1981:FSS

[Hol81] F. L. Holmes. The fine structure of scientific creativity. History of Science (UK), 19(1):60–70, March 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/19/1/60.full.pdf+html.

Holmes:1992:DWU

[Hol92] Frederic L. Holmes. Do we understand historically how experimental knowledge is acquired? History of Science (UK), 30(2):119–136, June 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/30/2/119.full.
pdf+html.

Holmes:2025:AAT

[Hol25] Matthew Holmes. Avian architects: Technology, domestication, and animal minds in urban America. History of Science (UK), 63(1): 101–119, March 2025. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753241235433.

Home:1977:NTM

[Hom77] R. W. Home. ‘Newtonianism’ and the theory of the magnet. History of Science (UK), 15(4):252–266, December 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/15/4/252.full.pdf+html.

Home:1985:ERP

[Hom85] R. W. Home. Essay review: Post-Newtonian optics: Optics after Newton: Theories of light in Britain and Ireland, 1704–1840, Brewster and Wheatstone on vision. History of Science (UK), 23(2):207–211, June 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/23/2/207.full.
pdf+html.

Home:1995:HSR

[Hom95] R. W. Home. Humboldtian science revisited: an Australian case study. History of Science (UK), 33(1):1–22, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/33/1/1.full.pdf+html.

Home:2003:BRC

[Hom03] R. W. Home. Book review: Coffee house science: Discussing chemistry and steam: The minutes of a coffee house philosophical society 1780– 1787. History of Science (UK), 41(2):241–243, June 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/41/2/241.full.pdf+html.

Hopwood:1997:BBU

[Hop97] Nick Hopwood. Biology between University and proletariat: The making of a red Professor. History of Science (UK), 35(4):367–424, December 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/35/4/
367.full.pdf+html.

Hopwood:2005:VSD

[Hop05] Nick Hopwood. Visual standards and disciplinary change: Normal plates, tables and stages in embryology. History of Science (UK), 43(3):239–303, September 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/43/3/239.full.pdf+html.

Hoskin:1962:BRT

[Hos62] M. A. Hoskin. Book review: The transits of Venus. History of Science (UK), 1(1):120–121, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/120.full.pdf+html.

Hoskin:1963:BRP

[Hos63a] M. A. Hoskin. Book review: The planet Saturn: a history of observation, theory and discovery. History of Science (UK), 2(1): 170–171, March 1963. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/2/
1/170.full.pdf+html.

Hoskin:1963:ERE

[Hos63b] M. A. Hoskin. Essay review: The elder Herschel: William Herschel. History of Science (UK), 2(1):145–148, March 1963. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/2/1/145.full.pdf+html.

Hoskin:1966:BRB

[Hos66] M. A. Hoskin. Book review: Bibliographie Générale de l’astronomie. History of Science (UK), 5(1):152, March 1966. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/5/1/152.full.pdf+html.

Hoskin:1973:BRC

[Hos73] M. A. Hoskin. Book review: Copernicus facsimile: The manuscript of Nicholas Copernicus’ “On the Revolutions”: Facsimile. History of Science (UK), 11(4):311, December 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/11/4/311.full.pdf+html.

Hoskin:2002:CHA

[Hos02] Michael Hoskin. Caroline Herschel: Assistant astronomer or astronomical assistant? History of Science (UK), 40(4):425–444, December 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/40/4/425.full.
pdf+html.

Hoskin:2003:BRC

[Hos03a] Michael Hoskin. Book review: a Copernicus census: an annotated census of Copernicus’, Nuremberg, 1543 and Basel, 1566. History of Science (UK), 41(2):243–244, June 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/41/2/243.full.pdf+html.

Hoskin:2003:RPD

[Hos03b] Michael Hoskin. Roy Porter (31 December 1946–4 March 2002). History of Science (UK), 41(3):249, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/41/3/249.full.pdf+html.

Hoskin:2003:VCW

[Hos03c] Michael Hoskin. Vocations in conflict: William Herschel in Bath, 1766–1782. History of Science (UK), 41(3):315–333, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/41/3/315.full.
pdf+html.

Hoskin:2005:UBW

[Hos05] Michael Hoskin. Unfinished business: William Herschel’s sweeps for nebulae. History of Science (UK), 43(3):305–320, September 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/43/3/305.full.
pdf+html.

Hoskin:2011:WHH

[Hos11] Michael Hoskin. William Herschel and Herschelian reflectors. History of Science (UK), 49(1):115–120, March 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/49/1/115.full.pdf+html.

Hoyrup:1990:SSM

[Høy90] Jens Høyrup. Sub-scientific mathematics: Observations on a premodern phenomenon. History of Science (UK), 28(1):63–87, March 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/28/1/63.full.
pdf+html.

Hoyrup:1996:CTH

[Høy96] Jens Høyrup. Changing trends in the historiography of Mesopotamian mathematics: an insider’s view. History of Science (UK), 34(1):1– 32, March 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/34/1/1.
full.pdf+html.

Hoyt:2001:RPF

[Hoy01] David L. Hoyt. The reanimation of the primitive: Fin-de-siècle ethnographic discourse in Western Europe. History of Science (UK), 39(3): 331–354, September 2001. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
39/3/331.full.pdf+html.

Hilaire-Perez:2007:TPC

[HP07] Liliane Hilaire-Pérez. Technology as a public culture in the Eighteenth Century: The artisans’ legacy. History of Science (UK), 45(2):135– 153, June 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/2/
135.full.pdf+html.

Hajek:2024:OHI

[HPtH24] Kim M. Hajek, Herman Paul, and Sjang ten Hagen. Objectivity, honesty, and integrity: How American scientists talked about their virtues, 1945–2000. History of Science (UK), 62(3):442–469, September 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231206773.

Hopwood:2010:SSO

[HSS10] Nick Hopwood, Simon Schaffer, and Jim Secord. Seriality and scientific objects in the Nineteenth Century. History of Science (UK), 48 (3–4):251–285, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Hall:1973:ERMa

[HT73] A. Rupert Hall and Laura Tilling. Essay review: Multi-volume works in progress (1): The correspondence of Isaac Newton. History of Science (UK), 11(1):68–70, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/11/1/68.full.pdf+html.

Huang:2016:WUM

[Hua16] Hsiang-Fu Huang. When Urania meets Terpsichore: a theatrical turn for astronomy lectures in early nineteenth-century Britain. History of Science (UK), 54(1):45–70, March 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/54/1/45.full.pdf+html.

Hunter:1988:PNS

[Hun88] Michael Hunter. Promoting the new science: Henry Oldenburg and the early Royal Society. History of Science (UK), 26(2):165–181, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/26/2/165.full.
pdf+html.

Hunter:1995:HBB

[Hun95] Michael Hunter. How Boyle became a scientist. History of Science (UK), 33(1):59–103, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/33/1/59.full.pdf+html.

Huntelmann:2010:SSP

[Hün10] Axel C. Hüntelmann. Seriality and standardization in the production of “606”. History of Science (UK), 48(3–4):435–460, September/ December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Hutchison:1983:SMP

[Hut83] Keith Hutchison. Supernaturalism and the mechanical philosophy. History of Science (UK), 21(3):297–333, September 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/21/3/297.full.pdf+html.

Hutchison:1991:DVS

[Hut91] Keith Hutchison. Dormitive virtues, scholastic qualities, and the new philosophies. History of Science (UK), 29(3):245–278, September 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/29/3/245.full.

pdf+html. See reply [Cla93].

Hutchison:2012:AVH

[Hut12] Keith Hutchison. An Angel’s view of heaven: The mystical heliocentricity of medieval geocentric cosmology. History of Science (UK), 50 (1):33–74, March 2012. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
50/1/33.full.pdf+html.

Hersey:2019:SGB

[HV19] Mark D. Hersey and Jeremy Vetter. Shared ground: Between environmental history and the history of science. History of Science (UK), 57 (4):403–440, December 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275319851013.

Hunter:1986:TSH

[HW86] Michael Hunter and Paul B. Wood. Towards Solomon’s house: Rival strategies for reforming the Early Royal Society. History of Science (UK), 24(1):49–108, March 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/24/1/49.full.pdf+html.

Hyun:2021:BSB

[Hyu21] Jaehwan Hyun. Brokering science, blaming culture: The US–South Korea ecological survey in the Demilitarized Zone, 1963–8. History of Science (UK), 59(3):315–343, September 1, 2021. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275320974209.

Ileri:2024:PEU

[Ile24] Nurcin Ileri. The politics of electricity use and non-use in late Ottoman Istanbul. History of Science (UK), 62(4):511–538, December 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231211175.

Iliffe:1992:ERR

[Ili92a] Rob Iliffe. Essay review: Rhetorical vices: Outlines of a Feyerabendian history of science, farewell to reason. History of Science (UK), 30 (2):199–219, June 1992. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
30/2/199.full.pdf+html.

Iliffe:1992:WPP

[Ili92b] Rob Iliffe. “In the warehouse”: Privacy, property and priority in the Early Royal Society. History of Science (UK), 30(1):29–68, March 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/30/1/29.full.
pdf+html.

Iliffe:1993:ADM

[Ili93] Rob Iliffe. “Aplatisseur du monde et de Cassini”: Maupertuis, precision measurement, and the shape of the Earth in the 1730s. History of Science (UK), 31(4):335–375, December 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/31/4/335.full.pdf+html.

Iliffe:1997:ERB

[Ili97] Rob Iliffe. Essay review: Boyle’s industry, Robert Boyle reconsidered, Robert Boyle reconsidered, Robert Boyle: By himself and his friends; with a fragment of William Wotton’s lost, the diffident naturalist: Robert Boyle and the philosophy of experiment. History of Science (UK), 35(4):455–484, December 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/35/4/455.full.pdf+html.

Iliffe:1998:ERR

[Ili98] Rob Iliffe. Essay review: Rational artistry, styles of scientific thinking in the European tradition: The history of argument and explanation especially in the mathematical and biomedical sciences and arts. History of Science (UK), 36(3):329–357, September 1998. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/36/3/329.full.pdf+html.

Ishizuka:2006:EAF

[Ish06] Hisao Ishizuka. The elasticity of the animal fibre: Movement and life in Enlightenment medicine. History of Science (UK), 44(4):435–468, December 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/44/4/
435.full.pdf+html.

Jackson:1973:ERE

[Jac73] Anthony Jackson. Essay review: The end of science and the ends of history of science: Scientific knowledge and its social problems. History of Science (UK), 11(4):292–305, December 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/11/4/292.full.pdf+html.

Jacob:1974:EN

[Jac74] M. C. Jacob. Early Newtonianism. History of Science (UK), 12(2): 142–146, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/2/
142.full.pdf+html.

Jacob:1975:RRO

[Jac75] J. R. Jacob. Restoration, reformation and the origins of the Royal Society. History of Science (UK), 13(3):155–176, September 1975. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/13/3/155.full.pdf+html.

Jacob:1978:NFP

[Jac78] Margaret C. Jacob. Newton and the French prophets: New evidence. History of Science (UK), 16(2):134–142, June 1978. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/16/2/134.full.pdf+html.

Jacob:1980:RIR

[Jac80] J. R. Jacob. Restoration ideologies and the Royal Society. History of Science (UK), 18(1):25–38, March 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/18/1/25.full.pdf+html.

Jacyna:1983:IJH

[Jac83] L. S. Jacyna. Images of John Hunter in the Nineteenth Century. History of Science (UK), 21(1):85–108, March 1983. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/21/1/85.full.pdf+html.

Jacyna:1987:MSM

[Jac87] L. S. Jacyna. Medical science and moral science: The cultural relations of physiology in Restoration France. History of Science (UK), 25 (2):111–146, June 1987. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
25/2/111.full.pdf+html.

Jacob:1995:RIM

[Jac95a] Margaret C. Jacob. Reflections on the ideological meanings of Western science from Boyle and Newton to the postmodernists. History of Science (UK), 33(3):333–357, September 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/33/3/333.full.pdf+html.

Jacyna:1995:BRC

[Jac95b] L. S. Jacyna. Book review: Citizen and patient, the citizen-patient in revolutionary and Imperial Paris. History of Science (UK), 33(1):112– 113, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/1/
112.full.pdf+html.

Jacyna:1995:BRH

[Jac95c] L. S. Jacyna. Book review: Hysteria in women, ventriloquized bodies: Narratives of hysteria in Nineteenth-Century France. History of Science (UK), 33(3):371–372, September 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/33/3/371.full.pdf+html.

Jacob:2003:BCE

[Jac03] Margaret C. Jacob. Being cheerfully enlightened. History of Science (UK), 41(3):287–292, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/41/3/287.full.pdf+html.

Jackson:2004:PMM

[Jac04] Myles W. Jackson. Physics, machines and musical pedagogy in Nineteenth-Century Germany. History of Science (UK), 42(4):371– 418, December 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/4/371.full.pdf+html.

Jackson:2006:RER

[Jac06] Catherine M. Jackson. Re-examining the research school: August Wilhelm Hofmann and the re-creation of a Liebigian research school in London. History of Science (UK), 44(3):281–319, September 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/44/3/281.full.
pdf+html.

Jacob:2007:MSF

[Jac07] Margaret C. Jacob. Mechanical science on the factory floor: The early industrial revolution in Leeds. History of Science (UK), 45(2):197– 221, June 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/2/
197.full.pdf+html.

Jackson:2017:EFA

[Jac17] Catherine M. Jackson. Emil Fischer and the “art of chemical experimentation”. History of Science (UK), 55(1):86–120, March 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://journals.sagepub.com/doi/full/10.1177/
0073275316685714.

James:1983:DNA

[Jam83] Frank A. J. L. James. The debate on the nature of the absorption of light, 1830–1835: a core-set analysis. History of Science (UK), 21(4): 335–368, December 1983. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
21/4/335.full.pdf+html.

James:1985:CVM

[Jam85] Frank A. J. L. James. The creation of a Victorian myth: The historiography of spectroscopy. History of Science (UK), 23(1):1–24, March 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/1/
1.full.pdf+html.

James:1986:ERS

[Jam86] Frank A. J. L. James. Essay review: Spectro-chemistry and myth: a rejoinder. History of Science (UK), 24(4):433–437, December 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/24/4/433.full.
pdf+html.

Jankovic:2000:PNN

[Jan00] Vladimir Jankovic. The place of nature and the nature of place: The chorographic challenge to the history of British provincial science. History of Science (UK), 38(1):79–113, March 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/38/1/79.full.pdf+html.

Jardine:2000:KKK

[Jar00a] Nick Jardine. Koyré’s Kepler/Kepler’s Koyré. History of Science (UK), 38(4):363–376, December 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/38/4/363.full.pdf+html.

Jardine:2000:UAA

[Jar00b] Nick Jardine. Uses and abuses of anachronism in the history of the sciences. History of Science (UK), 38(3):251–270, September 2000. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/38/3/251.full.pdf+html.

Jardine:2003:WSH

[Jar03] Nick Jardine. Whigs and stories: Herbert Butterfield and the historiography of science. History of Science (UK), 41(2):125–140, June 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/41/2/125.full.
pdf+html.

Jardine:2004:EEM

[Jar04] Nick Jardine. Etics and emics (not to mention anemics and emetics) in the history of the sciences. History of Science (UK), 42(3):261–278, September 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/42/3/
261.full.pdf+html.

Jennings:1981:BRD

[Jen81] Richard C. Jennings. Book review: Of dictionaries and men: Asimov’s biographical encyclopedia of science and technology. History of Science (UK), 19(3):222, September 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/19/3/222.full.pdf+html.

Jacob:1976:SCS

[JJ76] J. R. Jacob and M. C. Jacob. Seventeenth Century science and religion: The state of the argument. History of Science (UK), 14(3):196– 207, September 1976. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
14/3/196.full.pdf+html.

Johnston:1996:MLW

[Joh96] Sean F. Johnston. Making light work: Practices and practitioners of photometry. History of Science (UK), 34(3):273–302, September 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/34/3/273.full.
pdf+html.

Johns:1998:PPE

[Joh98] Adrian Johns. Prudence and pedantry in early modern cosmology: The trade of Al Ross. History of Science (UK), 36(1):23–59, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/36/1/23.full.
pdf+html.

Johnson:2005:EMK

[Joh05] Kristin Johnson. Ernst Mayr, Karl Jordan, and the history of systematics. History of Science (UK), 43(1):1–35, March 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/43/1/1.full.pdf+html.

Jonsson:1974:BRW

[Jon74] Inge Jonsson. Book review: Wolffianism and natural philosophy: Wolffianismens genombrott i Uppsala. History of Science (UK), 12(2): 152–154, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/2/
152.full.pdf+html.

Jones:2016:MCS

[Jon16] Peter M. Jones. Making chemistry the ‘science’ of agriculture, c. 1760– 1840. History of Science (UK), 54(2):169–194, June 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/54/2/169.full.pdf+html.

Jordanova:1978:ERH

[Jor78] L. J. Jordanova. Essay review: The historiography of the Claude Bernard industry: Claude Bernard and animal chemistry, Durkheim, Bernard and epistemology, Claude Bernard’s revised edition of his Introduction à l’étude de la médecine experimentale. History of Science (UK), 16(3):214–221, September 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/16/3/214.full.pdf+html.

Jordanova:2003:PPT

[Jor03] Ludmilla Jordanova. Portraits, people and things: Richard Mead and medical identity. History of Science (UK), 41(3):293–313, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/41/3/293.full.
pdf+html.

Jiang:2024:BSW

[JS24] Xue Jiang and Tao Shi. The borderline of science: Western exploration and study of Chinese insect white wax from the seventeenth to the nineteenth century. History of Science (UK), 62(1):??, March 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231180264.

Jubb:1985:SRB

[Jub85] Michael Jubb. Scientific records in the British public record Office. History of Science (UK), 23(4):379–389, December 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/23/4/379.full.pdf+html.

Kaalund:2014:OSR

[Kaa14] Nanna Katrine Lüders Kaalund. Oxford serialized: Revisiting the Huxley–Wilberforce debate through the periodical press. History of Science (UK), 52(4):429–453, December 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/52/4/429.full.pdf+html.

Kalmus:1983:SOM

[Kal83] H. Kalmus. The scholastic origins of Mendel’s concepts. History of Science (UK), 21(1):61–83, March 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/21/1/61.full.pdf+html.

Kang:2019:GIP

[Kan19] Hyo Yoon Kang. Ghosts of inventions: Patent law’s digital mediations. History of Science (UK), 57(1):38–61, March 1, 2019.

CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318770773.

Kassler:1982:MME

[Kas82] Jamie Croy Kassler. Music as a model in early science. History of Science (UK), 20(2):103–139, June 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/20/2/103.full.pdf+html.

Kassler:1984:MMI

[Kas84] Jamie C. Kassler. Man — a musical instrument: Models of the brain and mental functioning before the computer. History of Science (UK), 22(1):59–92, March 1984. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
22/1/59.full.pdf+html.

Kassell:2011:SRA

[Kas11] Lauren Kassell. Secrets revealed: Alchemical books in Early-Modern England. History of Science (UK), 49(1):61–A38, March 2011. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/49/1/61.full.pdf+html.

Katzir:2013:SPT

[Kat13] Shaul Katzir. Scientific practice for technology: Hermann Aron’s development of the storage battery. History of Science (UK), 51(4): 481–500, December 2013. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
51/4/481.full.pdf+html.

Keblusek:2005:KIS

[Keb05] Marika Keblusek. Keeping it secret: The identity and status of an early-modern inventor. History of Science (UK), 43(1):37–56, March 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/43/1/37.full.
pdf+html.

Keene:2014:FSN

[Kee14] Melanie Keene. Familiar science in Nineteenth-Century Britain. History of Science (UK), 52(1):53–71, March 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/52/1/53.full.pdf+html.

Keightley:1977:MAC

[Kei77] David N. Keightley. On the misuse of ancient Chinese inscriptions: an astronomical fantasy. History of Science (UK), 15(4):267–272, December 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/15/4/
267.full.pdf+html.

Keller:1962:BRM

[Kel62a] A. G. Keller. Book review: Man the maker. History of Science (UK), 1 (1):110–111, March 1962. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/1/
1/110.full.pdf+html.

Keller:1962:BRT

[Kel62b] A. G. Keller. Book review: The two Netherlanders. History of Science (UK), 1(1):119–120, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/119.full.pdf+html.

Keller:1963:ERM

[Kel63] A. G. Keller. Essay review: Mathematics at work: Niccoló Tartaglia, Quesiti et Inventioni Diverse. History of Science (UK), 2(1):155– 159, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/2/1/
155.full.pdf+html.

Keller:1966:BRB

[Kel66a] A. G. Keller. Book review: La Bibliothèque Scientifique d’un Imprimeur Humaniste au XVe Siècle. (French) [The scientific library of a humanist printer in the 15th Century]. History of Science (UK), 5 (1):146–147, March 1966. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/5/
1/146.full.pdf+html.

Keller:1966:BRS

[Kel66b] A. G. Keller. Book review: Steam power in the Eighteenth Century. History of Science (UK), 5(1):153, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/5/1/153.full.pdf+html.

Keller:1976:BRR

[Kel76] Alex Keller. Book review: Renaissance mathematical duels: Lodovico Ferrari e Niccoló Tartaglia. Cartelli di Sfida Matematica. (Italian)

[Lodovico Ferrari and Niccoló Tartaglia: Signs of mathematical challenges]. History of Science (UK), 14(3):208–209, September 1976. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/14/3/208.full.pdf+html.

Keller:2010:DLI

[Kel10] Vera Keller. Drebbel’s living instruments, Hartmann’s microcosm, and Libavius’s Thelesmos: Epistemic machines before Descartes. History of Science (UK), 48(1):39–74, March 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/48/1/39.full.pdf+html.

Keynes:1963:BRC

[Key63] Geoffrey Keynes. Book review: a catalogue of printed books in the Wellcome Historical Medical Library: Books printed before 1641. History of Science (UK), 2(1):162–164, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/2/1/162.full.pdf+html.

Kikuchi:2012:CNO

[Kik12] Yoshiyuki Kikuchi. Cross-national odyssey of a chemist: Edward Divers at London, Galway and Tokyo. History of Science (UK), 50(3): 289–314, September 2012. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
50/3/289.full.pdf+html.

Kim:1992:LCLa

[Kim92a] Mi Gyung Kim. The layers of chemical language, i: Constitution of bodies v. structure of matter. History of Science (UK), 30(1):69– 96, March 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/1/
69.full.pdf+html.

Kim:1992:LCLb

[Kim92b] Mi Gyung Kim. The layers of chemical language, II: Stabilizing atoms and molecules in the practice of organic chemistry. History of Science (UK), 30(4):397–437, December 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/30/4/397.full.pdf+html.

King:1975:ERM

[Kin75] David A. King. Essay review: Medieval mechanical devices: The Book of Knowledge of Ingenious Mechanical Devices: Kitab fi maarifat Al-H .iyal Al-Handasiya by Ibn Al-Razzaz Al-Jazari. History of Science (UK), 13(4):284–289, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/13/4/284.full.pdf+html.

King:1977:BRA

[Kin77] David A. King. Book review: On Arabic water-clocks: On the construction of water-clocks. History of Science (UK), 15(4):295–297, December 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/15/4/
295.full.pdf+html.

King:1979:ERI

[Kin79] David A. King. Essay review: Islamic mathematics: The Muslim contribution to mathematics. History of Science (UK), 17(4):295–296, December 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/4/
295.full.pdf+html.

King:1995:CTS

[Kin95] Roger King. Curing toothache on the stage? The importance of reading pictures in context. History of Science (UK), 33(4):396–416, December 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/4/
396.full.pdf+html.

Kingston:2014:TPA

[Kin14] Ralph Kingston. Trading places: Accumulation as mediation in French ministry map depots, 1798–1810. History of Science (UK), 52(3):247–276, September 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/52/3/247.full.pdf+html.

Kjaergaard:2010:DES

[Kjæ10] Peter C. Kjærgaard. The Darwin enterprise: From scientific Icon to global product. History of Science (UK), 48(1):105–122, March 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/48/1/105.full.
pdf+html.

Knight:1966:ERR

[Kni66] D. M. Knight. Essay review: Reduction in physiology: Heat and life. History of Science (UK), 5(1):134–140, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/5/1/134.full.pdf+html.

Knight:1970:PSR

[Kni70] D. M. Knight. The physical sciences and the romantic movement. History of Science (UK), 9(1):54–75, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/9/1/54.full.pdf+html.

Knight:1993:ERW

[Kni93] David Knight. Essay review: Writing the history of chemistry, the Fontana history of chemistry. History of Science (UK), 31(3):329–334, September 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/31/3/
329.full.pdf+html.

Knight:1995:BRE

[Kni95] David Knight. Book review: Exhibiting chemistry, chemistry imagined: Reflections on science. History of Science (UK), 33(1):111– 112, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/33/1/
111.full.pdf+html.

Knobloch:1979:MUU

[Kno79] Eberhard Knobloch. Musurgia Universalis: Unknown combinatorial studies in the age of Baroque absolutism. History of Science (UK), 17(4):258–275, December 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/17/4/258.full.pdf+html.

Knox:1996:DBN

[Kno96] Kevin C. Knox. Dephlogisticating the Bible: Natural philosophy and religious controversy in Late Georgian Cambridge. History of Science (UK), 34(2):167–200, June 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/34/2/167.full.pdf+html.

Knox:1999:LVP

[Kno99] Kevin C. Knox. Lunatick visions: Prophecy, signs and scientific knowledge in 1790s London. History of Science (UK), 37(4):427–458, December 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/4/
427.full.pdf+html.

Kohler:1991:DEG

[Koh91] Robert E. Kohler. Drosophila and evolutionary genetics: The moral economy of scientific practice. History of Science (UK), 29(4):335– 375, December 1991. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
29/4/335.full.pdf+html.

Kohler:2002:LNL

[Koh02a] Robert E. Kohler. Labscapes: Naturalizing the lab. History of Science (UK), 40(4):473–501, December 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/40/4/473.full.pdf+html.

Kohler:2002:PPF

[Koh02b] Robert E. Kohler. Place and practice in field biology. History of Science (UK), 40(2):189–210, June 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/40/2/189.full.pdf+html.

Kohler:2007:FKC

[Koh07] Robert E. Kohler. Finders, keepers: Collecting sciences and collecting practice. History of Science (UK), 45(4):428–454, December 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/45/4/428.full.
pdf+html.

Kollerstrom:2006:HHBa

[Kol06a] Nicholas Kollerstrom. An hiatus in history: The British claim for Neptune’s co-prediction, 1845–1846: Part 1. History of Science (UK), 44 (1):1–28, March 2006. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
44/1/1.full.pdf+html.

Kollerstrom:2006:HHBb

[Kol06b] Nicholas Kollerstrom. An hiatus in history: The British claim for Neptune’s co-prediction, 1845–1846: Part 2. History of Science (UK), 44(3):349–371, September 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/44/3/349.full.pdf+html.

Konashev:2019:TDD

[Kon19] Mikhail B. Konashev. Th. Dobzhansky and the development of evolutionary biology in the USSR. History of Science (UK), 57(3):346– 371, September 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275319827213.

Krasnodebski:2024:BGC

[Kra24] Marcin Krasnodebski. Beyond green chemistry: Radical environmental transformation through Sanfte Chemie (1985–1995). History of Science (UK), 62(2):280–304, June 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753231194801.

Kragh:2003:WDE

[KS03] Helge Kragh and Robert W. Smith. Who discovered the expanding universe? History of Science (UK), 41(2):141–162, June 2003. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/41/2/141.full.pdf+html.

Ksiazkiewicz:2015:PPN

[Ksi15] Allison Ksiazkiewicz. A philosophical pursuit: Natural models and the practical arts in establishing the structure of the earth. History of Science (UK), 53(2):125–154, June 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/53/2/125.full.pdf+html.

Kumar:2023:IBH

[Kum23] S. Prashant Kumar. The instrumental Brahmin and the “halfcaste” computer: Astronomy and colonial rule in Madras, 1791– 1835. History of Science (UK), 61(3):308–337, September 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753221090435.

Kusukawa:2016:CJM

[Kus16] Sachiko Kusukawa. Classics from this journal: Martin Rudwick’s ‘The Emergence of a Visual Language for Geological Science 1760-1840’, History of Science, xiv: 3, 1976, pp. 149–195. History of Science (UK), 54(1):98–104, March 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/54/1/98.full.pdf+html.

Kuukkanen:2012:SL

[Kuu12] Jouni-Matti Kuukkanen. Senses of localism. History of Science (UK), 50(4):477–500, December 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/50/4/477.full.pdf+html.

Labouff:2021:PSP

[Lab21] Nicole Labouff. Public science in the private garden: Noblewomen horticulturalists and the making of British botany c. 1785– 1810. History of Science (UK), 59(3):223–255, September 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275320961908.

Lachapelle:2009:SSA

[Lac09] Sofie Lachapelle. Science on stage: Amusing physics and scientific wonder at the Nineteenth-Century French theatre. History of Science (UK), 47(3):297–315, September 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/3/297.full.pdf+html.

Langermann:2009:IAG

[Lan09] Y. Tzvi Langermann. Islamic atomism and the Galenic tradition. History of Science (UK), 47(3):277–295, September 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/47/3/277.full.pdf+html.

Lash:1973:BRS

[Las73] Nicholas Lash. Book review: Science and the reformation: Religion and the rise of modern science. History of Science (UK), 11(2):145– 148, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/2/
145.full.pdf+html.

Laudan:1966:ERM

[Lau66] Laurens Laudan. Essay review: Method and the mechanical philosophy: Matter and method. History of Science (UK), 5(1):117– 124, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/5/1/
117.full.pdf+html.

Laudan:1968:TSM

[Lau68] Laurens Laudan. Theories of scientific method from Plato to Mach: a bibliographical review. History of Science (UK), 7(1):1–63, March 1968. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/7/1/1.full.pdf+
html.

Laudan:1993:HST

[Lau93] Rachel Laudan. Histories of the sciences and their uses: a review to 1913. History of Science (UK), 31(1):1–34, March 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/31/1/1.full.pdf+html.

Lavers:1970:ERC

[Lav70] Annette Lavers. Essay review: For a “committed” history of science: Études d’histoire et de philosophie des sciences. History of Science (UK), 9(1):101–105, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/9/1/101.full.pdf+html.

Laveaga:2022:RSL

[Lav22] Gabriela Soto Laveaga. Race science in the Latin world: an afterword. History of Science (UK), 60(1):96–102, March 1, 2022. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/00732753211070072.

Lawrence:1990:ERH

[Law90] Christopher Lawrence. Essay review: History of modern science considered: Companion to the history of modern science. History of Science (UK), 28(3):327–331, September 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/28/3/327.full.pdf+html.

Lawrence:1995:ERG

[Law95] Christopher Lawrence. Essay review: Good companions, companion encyclopedia of the history of medicine. History of Science (UK), 33 (1):105–109, March 1995. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
33/1/105.full.pdf+html.

Lux:1998:CCO

[LC98] David S. Lux and Harold J. Cook. Closed circles or open networks?: Communicating at a distance during the scientific revolution. History of Science (UK), 36(2):179–211, June 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/2/179.full.pdf+html.

LeRoux:2016:CIE

[Le 16] Thomas Le Roux. Chemistry and industrial and environmental governance in France, 1770–1830. History of Science (UK), 54(2):195– 222, June 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/54/2/
195.full.pdf+html.

LeBlanc:2001:OCD

[LeB01] André LeBlanc. The origins of the concept of dissociation: Paul Janet, his nephew Pierre, and the problem of post-hypnotic suggestion. History of Science (UK), 39(1):57–69, March 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/39/1/57.full.pdf+html.

Lee:2019:SST

[Lee19] Jung Lee. Socially skilling toil: New artisanship in papermaking in late Choson Korea. History of Science (UK), 57(2):167–193, June 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318804692.

Lee:2023:PLC

[Lee23] Juyoung Lee. Preparatory labor for chemical fertilizer: Rural modernity and the practices of South Korean farmers in the 1960s. History of Science (UK), 61(4):??, December 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753231188253.

Leonard:2024:BMC

[Leo24] Zak Leonard. A benefactor to mankind? Captain Warner’s secrets and the politics of invention in early Victorian Britain. History of Science (UK), 62(1):??, March 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753231157953.

Levene:1967:SCD

[Lev67] John R. Levene. Sources of confusion in Descartes’s illustrations, with reference to the history of contact lenses. History of Science (UK), 6 (1):90–96, March 1967. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/6/
1/90.full.pdf+html.

Levere:1970:RRI

[Lev70] Trevor H. Levere. Relations and rivalry: Interactions between Britain and the Netherlands in Eighteenth-Century science and technology. History of Science (UK), 9(1):42–53, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/9/1/42.full.pdf+html.

Levin:1984:VLF

[Lev84] A. Levin. Venel, Lavoisier, Fourcroy, Cabanis and the idea of scientific revolution: The French political context and the general patterns of conceptualization of scientific change. History of Science (UK), 22(3): 303–320, September 1984. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
22/3/303.full.pdf+html.

Lewis:1962:BRB

[Lew62] R. G. Lewis. Book review: Beloved Son Felix. The Journal of Felix Platter, a Medical Student in Montpellier in the Sixteenth Century. History of Science (UK), 1(1):113–114, March 1962. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/1/1/113.full.pdf+html.

Lightman:2015:SNT

[Lig15] Bernard Lightman. Scientific naturalists and their language games. History of Science (UK), 53(4):395–416, December 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Link:2006:TMA

[Lin06] David Link. Traces of the mouth: Andrei Andreyevich Markov’s mathematization of writing. History of Science (UK), 44(3):321–348, September 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/44/3/
321.full.pdf+html.

Livingstone:1984:HSH

[Liv84] David N. Livingstone. The history of science and the history of geography: Interactions and implications. History of Science (UK), 22(3): 271–302, September 1984. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
22/3/271.full.pdf+html.

Livingstone:1987:HAP

[Liv87] David N. Livingstone. Human acclimatization: Perspectives on a contested field of inquiry in science, medicine and geography. History of Science (UK), 25(4):359–394, December 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/25/4/359.full.pdf+html.

Livingstone:1988:SMR

[Liv88] David N. Livingstone. Science, magic and religion: a contextual reassessment of geography in the Sixteenth and Seventeenth Centuries. History of Science (UK), 26(3):269–294, September 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/26/3/269.full.pdf+html.

Livingstone:1991:DDC

[Liv91] David N. Livingstone. Of design and dining clubs: Geography in America and Britain, 1770–1860. History of Science (UK), 29(2): 153–183, June 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/29/2/
153.full.pdf+html.

Livesey:2005:BPE

[Liv05] James Livesey. Botany and provincial enlightenment in Montpellier: Antoine Banal Père and fils 1750–1800. History of Science (UK), 43 (1):57–76, March 2005. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
43/1/57.full.pdf+html.

Liskova:2024:SND

[LJG^{+}24] Kateřina Lišková, Natalia Jarska, Annina Gagyiova, José Luis Aguilar López-Barajas, and Šárka Caitlín Rábová. Saving newborns, defining livebirth: The struggle to reduce infant mortality in East-Central Europe in comparative and transnational perspectives, 1945–1965. History of Science (UK), 62(2):252–279, June 1, 2024. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753231187486.

Lloyd:1968:ERA

[Llo68] G. E. R. Lloyd. Essay review: Ancient medicine and modern controversies: Ancient medicine. History of Science (UK), 7(1):125– 129, March 1968. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/7/1/
125.full.pdf+html.

Lanuza-Navarro:2017:ACS

[LN17] Tayra M. C. Lanuza-Navarro. Astrology in court: The Spanish Inquisition, authority, and expertise. History of Science (UK), 55(2): 187–209, June 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317710537.

Loewenberg:1965:DDS

[Loe65] Bert James Loewenberg. Darwin and Darwin studies, 1959–63. History of Science (UK), 4(1):15–54, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/4/1/15.full.pdf+html.

Loewe:1970:ERC

[Loe70] Michael Loewe. Essay review: Chinese alchemy: Chinese alchemy: Preliminary studies. History of Science (UK), 9(1):90–93, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/9/1/90.full.
pdf+html.

Lohne:1967:ICN

[Loh67] J. A. Lohne. The increasing corruption of Newton’s diagrams. History of Science (UK), 6(1):69–89, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/6/1/69.full.pdf+html.

Lonie:1978:CVCb

[Lon78a] I. M. Lonie. Cos versus Cnidus and the historians: Part 2. History of Science (UK), 16(2):77–92, June 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/16/2/77.full.pdf+html.

Lonie:1978:CVCa

[Lon78b] I. M. Lonie. Cos versus Cnidus and the historians: Part I. History of Science (UK), 16(1):42–75, March 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/16/1/42.full.pdf+html.

Longrigg:1980:GPA

[Lon80] James Longrigg. The great plague of Athens. History of Science (UK), 18(3):209–225, September 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/18/3/209.full.pdf+html.

Longrigg:1981:SAC

[Lon81] James Longrigg. Superlative achievement and comparative neglect: Alexandrian medical science and modern historical research. History of Science (UK), 19(3):155–200, September 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/19/3/155.full.pdf+html.

Longrigg:1989:PPH

[Lon89] James Longrigg. Presocratic philosophy and Hippocratic medicine. History of Science (UK), 27(1):1–39, March 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/27/1/1.full.pdf+html.

Lowy:1992:SLC

[Löw92] Ilana Löwy. The strength of loose concepts — boundary concepts, federative experimental strategies and disciplinary growth: The case of immunology. History of Science (UK), 30(4):371–396, December 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/30/4/371.full.
pdf+html.

Lowy:2011:LBC

[Löw11] Ilana Löwy. Labelled bodies: Classification of diseases and the medical way of knowing. History of Science (UK), 49(3):299–315, September 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/49/3/299.full.
pdf+html.

Leonard:2023:PME

[LP23] Alice Leonard and Sarah E. Parker. “Put a mark on the errors”: Seventeenth-century medicine and science. History of Science (UK), 61(3):287–307, September 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753221135046.

Leitao:2017:TMT

[LS17] Henrique Leitão and Antonio Sánchez. Too much to tell: Narrative styles of the first descriptions of the natural world of the Indies. History of Science (UK), 55(2):167–186, June 2017. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

journals.sagepub.com/doi/full/10.1177/0073275317712864.

Lucier:1995:BRS

[Luc95] Paul Lucier. Book review: Steel industries compared, enterprise and technology: The German and British steel industries 1865–1895. History of Science (UK), 33(1):116–118, March 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/33/1/116.full.pdf+html.

Lucier:1999:PAG

[Luc99] Paul Lucier. A plea for applied geology. History of Science (UK), 37(3):283–318, September 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/37/3/283.full.pdf+html.

Luscombe:1968:MME

[Lus68] D. E. Luscombe. Medicine in medieval England. History of Science (UK), 7(1):129–133, March 1968. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/7/1/129.full.pdf+html.

Moro-Abadia:2009:TAP

[MA09] Oscar Moro-Abadía. Thinking about ‘presentism’ from a Historian’s perspective: Herbert Butterfield and Hélène Metzger. History of Science (UK), 47(1):55–77, March 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/1/55.full.pdf+html.

Maas:2010:CSD

[Maa10] Ad Maas. Civil scientists: Dutch scientists between 1750 and 1875. History of Science (UK), 48(1):75–103, March 2010. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/48/1/75.full.pdf+html.

Maas:2020:MSF

[Maa20] Harro Maas. Monitoring the self: François-Marc-Louis Naville and his moral tables. History of Science (UK), 58(2):117–141, July 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275319876851.

Macdonald:2021:UPO

[Mac21] Lee T. Macdonald. University physicists and the origins of the National Physical Laboratory, 1830–1900. History of Science (UK), 59 (1):73–92, March 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275318811445.

Maddison:1963:EAM

[Mad63] Francis Maddison. Early astronomical and mathematical instruments: a brief survey of sources and modern studies. History of Science (UK), 2(1):17–50, March 1963. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/2/
1/17.full.pdf+html.

Maerker:2007:THE

[Mae07] Anna Maerker. “Turpentine hides everything”: Autonomy and organization in anatomical model production for the state in Late Eighteenth-Century Florence. History of Science (UK), 45(3):257– 286, September 2007. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
45/3/257.full.pdf+html.

Magnello:1999:NCBa

[Mag99a] M. Eileen Magnello. The non-correlation of biometrics and Eugenics: Rival forms of laboratory work in Karl Pearson’s career at University College London, Part 1. History of Science (UK), 37(1):79– 106, March 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/1/
79.full.pdf+html.

Magnello:1999:NCBb

[Mag99b] M. Eileen Magnello. The non-correlation of biometrics and eugenics: Rival forms of laboratory work in Karl Pearson’s career at University College London, Part 2. History of Science (UK), 37(2):123–150, June 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/37/2/123.full.
pdf+html.

MaltaRomeiras:2018:GCJ

[Mal18] Francisco Malta Romeiras. For the greater credibility: Jesuit science and education in modern Portugal (1858–1910). History of Science (UK), 56(1):97–119, March 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.sagepub.

com/doi/full/10.1177/0073275317738599.

Martin:1982:ERD

[Mar82] R. Niall D. Martin. Essay review: Darwin and Duhem: The edge of contingency: French Catholic reaction to scientific changes from Darwin to Duhem. History of Science (UK), 20(1):64–74, March 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/20/1/64.full.pdf+html.

Martin:1987:SDG

[Mar87] R. Niall D. Martin. Saving Duhem and Galileo: Duhemian methodology and the saving of the phenomena. History of Science (UK), 25(3): 301–319, September 1987. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
25/3/301.full.pdf+html.

Marsden:1998:BHC

[Mar98] Ben Marsden. Blowing hot and cold: Reports and retorts on the status of the air-engine as success or failure, 1830–1855. History of Science (UK), 36(4):373–420, December 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/4/373.full.pdf+html.

Marks:2000:TMM

[Mar00] Harry M. Marks. Trust and mistrust in the marketplace: Statistics and clinical research, 1945–1960. History of Science (UK), 38(3):343– 355, September 2000. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
38/3/343.full.pdf+html.

Martin:2011:CFS

[Mar11] Shirley A. Martin. Not as the crow flies: ‘styles’ of educational measurement in the reception of inferential statistics at Iowa and Minnesota. History of Science (UK), 49(2):187–215, June 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/49/2/187.full.pdf+html.

Marsden:2013:RRW

[Mar13] Ben Marsden. Ranking Rankine: W. J. M. Rankine (1820–72) and the making of ‘engineering science’ revisited. History of Science (UK), 51(4):434–456, December 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/51/4/434.full.pdf+html.

Martykanova:2024:MUO

[Mar24] Darina Martykánová. Mastering the uncontrollable: The Ottomans and the use of modern technologies. History of Science (UK), 62(4): 473–487, December 2024. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753241298134.

Mason:2002:GSD

[Mas02] Stephen Mason. Galileo’s scientific discoveries, cosmological confrontations, and the aftermath. History of Science (UK), 40(4):377– 406, December 2002. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
40/4/377.full.pdf+html.

Mason:2003:RRP

[Mas03] Stephen Mason. Religious reform and the pulmonary transit of the blood. History of Science (UK), 41(4):459–471, December 2003. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/41/4/459.full.pdf+html.

Matson:1963:BRP

[Mat63] W. I. Matson. Book review: The physical world of late antiquity. History of Science (UK), 2(1):171–172, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/2/1/171.full.pdf+html.

Mathias:1971:ERT

[Mat71] Peter Mathias. Essay review: Technological change on the grand scale: The unbound Prometheus: Technological change and industrial development in Western Europe from 1750 to the present. History of Science (UK), 10(1):113–117, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/10/1/113.full.pdf+html.

Mau:1964:BRB

[Mau64] Jürgen Mau. Book review: Die Bedeutung der Bewegungslehre des Aristoteles für seine beiden Lösungen der zenonischen Paradoxie (German) [The importance of the theory of motion of Aristotle for his two solutions of Zeno’s paradox]. History of Science (UK), 3(1): 127–131, March 1964. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/3/
1/127.full.pdf+html.

May:1971:ERM

[May71] Kenneth O. May. Essay review: Mathematics in Russia: Istoriya Matematiki v Rossii Do 1917 Goda (history of mathematics in Russia to 1917). History of Science (UK), 10(1):122–127, March 1971. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/10/1/122.full.pdf+html.

Mayer:2002:FME

[May02] A. K. Mayer. Fatal mutilations: Educationism and the British background to the 1931 international Congress for the history of science and technology. History of Science (UK), 40(4):445–472, December 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/40/4/445.full.
pdf+html.

Mayer:2005:RTS

[May05] Anna-K. Mayer. Reluctant technocrats: Science promotion in the neglect-of-science debate of 1916–1918. History of Science (UK), 43 (2):139–159, June 2005. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
43/2/139.full.pdf+html.

Mayer:2012:IRC

[May12] Thomas F. Mayer. An interim report on a census of Galileo’s sunspot letters. History of Science (UK), 50(2):155–196, June 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/50/2/155.full.pdf+html.

Margocsy:2023:TRC

[MB23] Dániel Margócsy and Mary Augusta Brazelton. Techniques of repair, the circulation of knowledge, and environmental transformation: Towards a new history of transportation. History of Science (UK), 61 (1):3–18, March 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753211046450.

McCrea:1985:ERR

[McC85] W. H. McCrea. Essay review: Recent history of astronomy: The general history of astronomy. History of Science (UK), 23(1):119– 122, March 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/1/
119.full.pdf+html.

McCormick:2013:GMP

[McC13] Ted McCormick. Governing model populations: Queries, quantification, and William Petty’s “scale of salubrity”. History of Science (UK), 51(2):179–197, June 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/51/2/179.full.pdf+html.

McEvoy:1997:PWC

[McE97] John G. McEvoy. Positivism, Whiggism, and the chemical revolution: a study in the historiography of chemistry. History of Science (UK), 35(1):1–33, March 1997. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
35/1/1.full.pdf+html.

McEvoy:2013:FTN

[McE13] Timothy McEvoy. Finding a teacher of navigation abroad in Eighteenth-Century Venice: a study of the circulation of useful knowledge. History of Science (UK), 51(1):100–123, March 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/51/1/100.full.pdf+html.

McGuire:1966:ERI

[McG66] J. E. McGuire. Essay review: Intellectual history or scientific biography?: Michael Faraday. a biography. History of Science (UK), 5(1): 140–144, March 1966. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/5/
1/140.full.pdf+html.

McGuire:1973:NDF

[McG73] J. E. McGuire. Newton and the demonic furies: Some current problems and approaches in history of science. History of Science (UK), 11(1):21–48, March 1973. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
11/1/21.full.pdf+html.

McGeachie:1985:ERD

[McG85] James McGeachie. Essay review: Darwin and George Eliot: Plotting and organicism: Darwin’s plots: Evolutionary narrative in Darwin, George Eliot and Nineteenth-Century fiction, George Eliot and Nineteenth-Century science: The make-believe of a beginning. History of Science (UK), 23(2):187–200, June 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/23/2/187.full.pdf+html.

McGeachie:1990:ERP

[McG90] James McGeachie. Essay review: From parson-Hunter to eco-prophet: Evolution and ethics: T. H. Huxley’s “evolution and ethics” with new essays on its Victorian and sociobiological context. History of Science (UK), 28(4):429–442, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/28/4/429.full.pdf+html.

Meade:2024:SAM

[Mea24] Ruselle Meade. Science across the Meiji divide: Vernacular literary genres as vectors of science in modern Japan. History of Science (UK), 62(2):227–251, June 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/abs/10.1177/00732753231170413.

Meli:1996:NPP

[Mel96] Domenico Bertoloni Meli. The neoterics and political power in Spanish Italy: Giovanni Alfonso Borelli and his circle. History of Science (UK), 34(1):57–89, March 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/34/1/57.full.pdf+html.

Mendelsohn:1964:BSN

[Men64] Everett Mendelsohn. The biological sciences in the Nineteenth Century: Some problems and sources. History of Science (UK), 3(1): 39–59, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/3/1/39.
full.pdf+html.

Mertens:2000:TCF

[Mer00] Joost Mertens. From Tubal Cain to Faraday: William Whewell as a philosopher of technology. History of Science (UK), 38(3):321–342, September 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/3/
321.full.pdf+html.

Mercelis:2020:SAR

[Mer20] Joris Mercelis. The scientist and the advertisement: Reklamegutachten in imperial Germany. History of Science (UK), 58(4):507–532, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275320916971.

MacLeod:1973:QAB

[MF73] Roy M. MacLeod and James R. Friday. The quest for archives of British men of science. History of Science (UK), 11(1):8–20, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/1/
8.full.pdf+html.

Marcus:2023:SCC

[MH23] Hannah Marcus and Crystal Hall. Shattering crystal with crystal: Galileo’s rhetoric, lenses, and the epistemology of metaphor. History of Science (UK), 61(2):179–213, June 1, 2023. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753211041858.

Micale:1989:HHRa

[Mic89a] Mark S. Micale. Hysteria and its historiography: a review of past and present writings (i). History of Science (UK), 27(3):223–261, September 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/27/3/
223.full.pdf+html.

Micale:1989:HHRb

[Mic89b] Mark S. Micale. Hysteria and its historiography: a review of past and present writings (II). History of Science (UK), 27(4):319–351, December 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/27/4/
319.full.pdf+html.

Miller:1975:AEM

[Mil75] Arthur I. Miller. Albert Einstein and Max Wertheimer: a gestalt psychologist’s view of the genesis of Special Relativity theory. History of Science (UK), 13(2):75–103, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/13/2/75.full.pdf+html.

Miller:1981:ERS

[Mil81] David Philip Miller. Essay review: Sir Joseph Banks: an historiographical perspective: The sheep and wool correspondence of Sir Joseph Banks 1781–1820, Sir Joseph Banks: 18th century explorer, botanist and entrepreneur. History of Science (UK), 19(4):284–292, December 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/19/4/
284.full.pdf+html.

Miller:1989:VDR

[Mil89] David P. Miller. “Into the valley of darkness”: Reflections on the Royal Society in the Eighteenth Century. History of Science (UK), 27(2):155–166, June 1989. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
27/2/155.full.pdf+html.

Miller:2000:PJC

[Mil00] David Philip Miller. “Puffing Jamie”: The commercial and ideological importance of being a ‘philosopher’ in the case of the reputation of James Watt (1736–1819). History of Science (UK), 38(1): 1–24, March 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/1/1.
full.pdf+html.

Miller:2004:TMJ

[Mil04] David Philip Miller. True myths: James Watt’s kettle, his condenser, and his chemistry. History of Science (UK), 42(3):333–360, September 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/42/3/333.full.
pdf+html.

Miller:2008:TYW

[Mil08a] David Marshall Miller. The thirty years war and the Galileo affair. History of Science (UK), 46(1):49–74, March 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/46/1/49.full.pdf+html.

Miller:2008:IPN

[Mil08b] David Philip Miller. Intellectual property and narratives of discovery/ invention: The League of Nations’ draft convention on ‘scientific property’ and its fate. History of Science (UK), 46(3):299–342, September 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/46/3/299.full.
pdf+html.

Miller:2008:BNJ

[Mil08c] Gordon L. Miller. Beasts of the New Jerusalem: John Jonston’s natural history and the launching of millenarian pedagogy in the Seventeenth Century. History of Science (UK), 46(2):203–243, June 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/46/2/203.full.
pdf+html.

Miller:2012:TPT

[Mil12] David Philip Miller. Testing power and trust: The steam indicator, the ‘Reynolds Controversy’, and the relations of engineering science and practice in Late Nineteenth-Century Britain. History of Science (UK), 50(2):212–250, June 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/50/2/212.full.pdf+html.

McLaughlin-Jenkins:2001:CKS

[MJ01] Erin McLaughlin-Jenkins. Common knowledge: Science and the Late Victorian working-class Press. History of Science (UK), 39(4):445– 465, December 2001. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
39/4/445.full.pdf+html.

Mokyr:2007:KEI

[Mok07] Joel Mokyr. Knowledge, Enlightenment, and the Industrial Revolution: Reflections on the gifts of Athena. History of Science (UK), 45 (2):185–196, June 2007. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
45/2/185.full.pdf+html.

Molland:1969:ERO

[Mol69] A. G. Molland. Essay review: Oresme redivivus: Nicole Oresme and the medieval geometry of qualities and motions. a treatise on the uniformity and difformity of intensities known as Tractatus de configurationibus qualitatum et motuum. History of Science (UK), 8(1): 106–119, March 1969. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/8/
1/106.full.pdf+html.

Molland:1976:AP

[Mol76] A. G. Molland. Ancestors of physics. History of Science (UK), 14(1): 54–75, March 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/14/1/
54.full.pdf+html.

Molland:1981:ERA

[Mol81] George Molland. Essay review: Archimedean fortunes: Archimedes in the Middle Ages. History of Science (UK), 19(2):143–147, June 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/19/2/143.full.
pdf+html.

Moore:1992:ERS

[Moo92] James Moore. Essay review: Speaking of “science and religion” — Then and now, science and religion: Some historical perspectives. History of Science (UK), 30(3):311–323, September 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/30/3/311.full.pdf+html.

Morrell:1974:RHS

[Mor74] J. B. Morrell. Reflections on the history of Scottish science. History of Science (UK), 12(2):81–94, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/12/2/81.full.pdf+html.

Morrell:1977:ERS

[Mor77] J. B. Morrell. Essay review: Science and the universities: The University in society. History of Science (UK), 15(2):145–152, June 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/15/2/145.full.
pdf+html.

Moravia:1980:ESM

[Mor80a] Sergio Moravia. The Enlightenment and the sciences of man. History of Science (UK), 18(4):247–268, December 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/18/4/247.full.pdf+html.

Morrell:1980:ERS

[Mor80b] J. B. Morrell. Essay review: Savants and clergymen: Science in culture: The Early Victorian period. History of Science (UK), 18(1): 39–45, March 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/18/1/
39.full.pdf+html.

Morrell:1992:ERJ

[Mor92a] Jack Morrell. Essay review: The judge and purifier of all, William Whewell: Philosopher of science, William Whewell: a composite portrait. History of Science (UK), 30(1):97–114, March 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/30/1/97.full.pdf+html.

Morus:1992:DEL

[Mor92b] Iwan Rhys Morus. Different experimental lives: Michael Faraday and William Sturgeon. History of Science (UK), 30(1):1–28, March 1992.

CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/30/1/1.full.pdf+html.

Morrell:1993:ERH

[Mor93] Jack Morrell. Essay review: Hustlers and patrons of science, Millikan’s school: a history of the California Institute of Technology, partners in science: Foundations and natural scientists 1900–1945. History of Science (UK), 31(1):65–82, March 1993. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/31/1/65.full.pdf+html.

Morrell:1994:ERW

[Mor94] Jack Morrell. Essay review: William Whewell: Rough diamond, defining science: William Whewell, natural knowledge, and public debate in Early Victorian Britain. History of Science (UK), 32(3):345–359, September 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/32/3/
345.full.pdf+html.

Morus:1999:MMT

[Mor99] Iwan Rhys Morus. The measure of man: Technologizing the Victorian body. History of Science (UK), 37(3):249–282, September 1999. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/37/3/249.full.pdf+html.

Morgan:2009:SEI

[Mor09a] John Morgan. Science, England’s ‘interest’ and universal monarchy: The making of Thomas Sprat’s history of the Royal Society. History of Science (UK), 47(1):27–54, March 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/47/1/27.full.pdf+html.

Morus:2009:E

[Mor09b] Iwan Rhys Morus. Editorial. History of Science (UK), 47(4):369–371, December 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/47/4/
369.full.pdf+html.

Morus:2009:ERW

[Mor09c] Iwan Rhys Morus. Essay review: What is the history of science really like? Science: a four thousand year history. History of Science (UK), 47(3):359–366, September 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/3/359.full.pdf+html.

Moreira:2010:WES

[Mor10] Raquel Delgado Moreira. “What Ezekiel says”: Newton as a Temple scholar. History of Science (UK), 48(2):153–180, June 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/48/2/153.full.pdf+html.

Morus:2014:E

[Mor14] Iwan Rhys Morus. Editorial. History of Science (UK), 52(2):129, June 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/52/2/129.full.
pdf+html.

Mori:2016:PSE

[Mor16] Jennifer C. Mori. Popular science in eighteenth century almanacs: The editorial career of Henry Andrews of Royston, 1780–1820. History of Science (UK), 54(1):19–44, March 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/54/1/19.full.pdf+html.

Morris:2018:JSV

[Mor18] Andrew M. A. Morris. John Smeaton and the vis viva controversy: Measuring waterwheel efficiency and the influence of industry on practical mechanics in Britain 1759–1808. History of Science (UK), 56 (2):196–223, June 2018. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317745455.

Mougey:2021:BUS

[Mou21] Thomas Mougey. Building UNESCO science from the “dark zone”: Joseph Needham, Empire, and the wartime reorganization of international science from China, 1942–6. History of Science (UK), 59 (4):461–491, December 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275320987428.

Marcaida:2014:GTP

[MP14] José Ramón Marcaida and Juan Pimentel. Green treasures and paper floras: the business of Mutis in New Granada (1783–1808). History of Science (UK), 52(3):277–296, September 2014. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/52/3/277.full.pdf+html.

Mody:2020:IRI

[MSR20] Cyrus C. M. Mody, H. Otto Sibum, and Lissa L. Roberts. Integrating research integrity into the history of science. History of Science (UK), 58(4):369–385, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275320952257.

Mukharji:2016:PCC

[Muk16] Projit Bihari Mukharji. Parachemistries: Colonial chemopolitics in a zone of contest. History of Science (UK), 54(4):362–382, December 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://journals.sagepub.com/doi/full/10.1177/
0073275316681803.

Munns:2021:ABW

[Mun21] David P. D. Munns. The age of biology: When plant physiology was in the center of American life science. History of Science (UK), 59 (4):492–521, December 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275320954123.

Musselman:2000:LCJ

[Mus00] Elizabeth Green Musselman. Local colour: John Dalton and the politics of colour blindness. History of Science (UK), 38(4):401–424, December 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/38/4/
401.full.pdf+html.

Myers:1992:HPS

[Mye92] Greg Myers. History and philosophy of science seminar 4:00 Wednesday, seminar room 2 “fictions for facts: The form and authority of the scientific dialogue”. History of Science (UK), 30(3):221–247, September 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/30/3/221.full.
pdf+html.

Navarro:2006:IIL

[Nav06] Jaume Navarro. Imperial incursions in Late-Victorian Cambridge: J. J. Thomson and the domains of the physical sciences. History of Science (UK), 44(4):469–495, December 2006. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/44/4/469.full.pdf+html.

Navarro:2021:WEH

[Nav21] Jaume Navarro. Whittaker, Einstein, and the history of the aether: Alternative interpretation, blunder, or bigotry? History of Science (UK), 59(3):287–314, September 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275320968408.

Niermeier-Dohoney:2022:RCA

[ND22] Justin Niermeier-Dohoney. “Rusticall chymistry”: Alchemy, saltpeter projects, and experimental fertilizers in seventeenth-century English agriculture. History of Science (UK), 60(4):546–574, December 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753211033159.

Neal:1999:RUA

[Nea99] Katherine Neal. The rhetoric of utility: Avoiding occult associations for mathematics through profitability and pleasure. History of Science (UK), 37(2):151–178, June 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/37/2/151.full.pdf+html.

Needham:1973:BRB

[Nee73] Dorothy Needham. Book review: The beginnings of biochemistry: a history of biochemistry. History of Science (UK), 11(2):148–150, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/11/2/148.full.
pdf+html.

Netz:2002:CCT

[Net02] Reviel Netz. Counter culture: Towards a history of Greek numeracy. History of Science (UK), 40(3):321–352, September 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/40/3/321.full.pdf+html.

Neve:1976:BRT

[Nev76] Michael Neve. Book review: The trade in lunacy: Madness and morals. ideas on insanity in the Nineteenth Century. History of Science (UK), 14(2):138–139, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/14/2/138.full.pdf+html.

Neve:1986:ERL

[Nev86] Michael Neve. Essay review: Letters of Victorian scientists, gentlemen of science: Early correspondence of the British association for the advancement of science. History of Science (UK), 24(3):327–328, September 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/3/
327.full.pdf+html.

Nieto-Galan:2011:AGR

[NG11] Agustí Nieto-Galan. Antonio Gramsci revisited: Historians of science, intellectuals, and the struggle for hegemony. History of Science (UK), 49(4):453–478, December 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/49/4/453.full.pdf+html.

Nieto-Galan:2022:PMU

[NG22] Agustí Nieto-Galan. A puzzling marriage? UNESCO and the Madrid Festival of Science (1955). History of Science (UK), 60(3):383–404, September 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/0073275321991288.

Nicolson:1987:AHH

[Nic87] Malcolm Nicolson. Alexander von Humboldt, Humboldtian science and the origins of the study of vegetation. History of Science (UK), 25(2):167–194, June 1987. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
25/2/167.full.pdf+html.

Nicolson:1988:ERN

[Nic88] Malcolm Nicolson. Essay review: No longer a stranger? A decade in the history of ecology: Modeling nature: Episodes in the history of population ecology, the background of ecology: Concept and theory, saving the prairies: The life cycle of the founding school of American plant ecology 1895–1955, Nature’s economy: a history of ecological ideas, Nature’s economy: The roots of ecology. History of Science (UK), 26(2):183–200, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/26/2/183.full.pdf+html.

Noakes:2004:BWB

[Noa04] Richard Noakes. The “bridge which is between physical and psychical research”: William Fletcher Barrett, sensitive flames, and spiritualism. History of Science (UK), 42(4):419–464, December 2004. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/42/4/419.full.pdf+html.

Noakes:2005:ERP

[Noa05] Richard Noakes. Ethers, religion and politics in late-Victorian physics: Beyond the Wynne thesis. History of Science (UK), 43(4):415–455, December 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/43/4/
415.full.pdf+html.

North:1966:BRL

[Nor66] J. D. North. Book review: Logic and language studies dedicated to Professor Rudolf Carnap on the occasion of his seventieth birthday, Boston studies in the philosophy of science. History of Science (UK), 5 (1):156–158, March 1966. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/5/
1/156.full.pdf+html.

North:1996:ACC

[Nor96] J. D. North. Alistair Cameron Crombie (1915–1996). History of Science (UK), 34(2):245–248, June 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/34/2/245.full.pdf+html.

Nothaft:2017:ZLS

[Not17] C. Philipp E. Nothaft. Zaccaria Lilio and the shape of the earth: a brief response to Allegro’s “Flat earth science”. History of Science (UK), 55(4):490–498, December 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275317714933.

Nutton:1974:ERG

[Nut74] Vivian Nutton. Essay review: Greek medicine dissected: Greek medicine, the heart and the vascular system in Ancient Greek medicine: From Alcmaeon to Galen. History of Science (UK), 12(1): 59–69, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/1/
59.full.pdf+html.

Oakley:2018:VTE

[Oak18] Francis Oakley. Voluntarist theology and early-modern science: The matter of the divine power, absolute and ordained. History of Science (UK), 56(1):72–96, March 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.sagepub.

com/doi/full/10.1177/0073275317722241.

OConnor:2007:YEC

[O’C07] Ralph O’Connor. Young-Earth creationists in Early Nineteenth-Century Britain? Towards a reassessment of ‘scriptural geology’. History of Science (UK), 45(4):357–403, December 2007. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/45/4/357.full.pdf+html.

Orozco-Echeverri:2019:JCM

[OEMB19] Sergio H. Orozco-Echeverri and Sebastián Molina-Betancur. José Celestino Mutis’ appropriation of Newton’s experimental physics in New Granada (1761–1808). History of Science (UK), 57(3):291–323, September 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL https://journals.sagepub.com/doi/full/
10.1177/0073275319827203.

Ogborn:2013:TPB

[Ogb13] Miles Ogborn. Talking plants: Botany and speech in Eighteenth-Century Jamaica. History of Science (UK), 51(3):251–282, September 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/51/3/251.full.
pdf+html.

Olby:1979:MNM

[Olb79] Robert Olby. Mendel no Mendelian? History of Science (UK), 17(1): 53–72, March 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/1/
53.full.pdf+html.

Oldroyd:1979:HRHa

[Old79a] D. R. Oldroyd. Historicism and the rise of historical geology, Part 1. History of Science (UK), 17(3):191–213, September 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/17/3/191.full.pdf+html.

Oldroyd:1979:HRHb

[Old79b] D. R. Oldroyd. Historicism and the rise of historical geology, Part 2. History of Science (UK), 17(4):227–257, December 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/17/4/227.full.pdf+html.

Oldroyd:1984:HDD

[Old84] David R. Oldroyd. How did Darwin arrive at his theory? The secondary literature to 1982. History of Science (UK), 22(4):325–374, December 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/4/
325.full.pdf+html.

Oldroyd:1986:GGA

[Old86] David R. Oldroyd. Grid/group analysis for historians of science? History of Science (UK), 24(2):145–171, June 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/24/2/145.full.pdf+html.

Oldroyd:1997:ERP

[Old97] David Oldroyd. Essay review: Peripheral Darwinism, Darwin’s laboratory: Evolutionary theory and the natural history of the Pacific. History of Science (UK), 35(1):107–110, March 1997. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/35/1/107.full.pdf+html.

Oldroyd:1998:BRH

[Old98] David Oldroyd. Book review: History of mankind and history of the Earth: When geologists were historians: 1665–1750. History of Science (UK), 36(3):362–364, September 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/3/362.full.pdf+html.

Oliver:2024:MMW

[Oli24] Kendrick Oliver. The mule on the Mount Wilson trail: George Ellery Hale, American scientific cosmology, and cosmologies of American science. History of Science (UK), 62(1):??, March 1, 2024. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753231179330.

Olson:1978:SSA

[Ols78] Richard Olson. Science, scientism and anti-science in Hellenic Athens: a new Whig interpretation. History of Science (UK), 16(3):179–199, September 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/16/3/
179.full.pdf+html.

Olson:1990:HRF

[Ols90] Richard Olson. Historical reflections on feminist critiques of science: The scientific background to modern feminism. History of Science (UK), 28(2):125–147, June 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/28/2/125.full.pdf+html.

OMalley:1965:RVL

[O’M65] C. D. O’Malley. A review of Vesalian literature. History of Science (UK), 4(1):1–14, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/4/1/1.full.pdf+html.

Ortolano:2005:FRL

[Ort05] Guy Ortolano. F. R. Leavis, science, and the abiding crisis of modern civilization. History of Science (UK), 43(2):161–185, June 2005. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/43/2/161.full.pdf+html.

Osborn:2016:BCP

[Osb16] Emily Lynn Osborn. From bauxite to cooking pots: Aluminum, chemistry, and West African artisanal production. History of Science (UK), 54(4):425–442, December 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275316681806.

Osler:1998:MMS

[Osl98] Margaret J. Osler. Mixing metaphors: Science and religion or natural philosophy and theology in early modern Europe. History of Science (UK), 36(1):91–113, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/36/1/91.full.pdf+html.

Oster:1993:BCS

[Ost93] Malcolm Oster. Biography, culture, and science: The formative years of Robert Boyle. History of Science (UK), 31(2):177–226, June 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/31/2/177.full.
pdf+html.

Outram:1976:SBC

[Out76] Dorinda Outram. Scientific biography and the case of Georges Cuvier: With a critical bibliography. History of Science (UK), 14(2):101– 137, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/14/2/
101.full.pdf+html.

Outram:1978:LNP

[Out78] Dorinda Outram. The language of natural power: The “Éloges” of Georges Cuvier and the public language of Nineteenth Century science. History of Science (UK), 16(3):153–178, September 1978. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/16/3/153.full.pdf+html.

Outram:1983:OVP

[Out83] Dorinda Outram. The ordeal of vocation: The Paris Academy of Sciences and the terror, 1793–95. History of Science (UK), 21(3):251– 273, September 1983. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
21/3/251.full.pdf+html.

Pagel:1963:ERH

[Pag63] Walter Pagel. Essay review: an Harveyan prelude to Harvey: Lectures on the whole of anatomy. History of Science (UK), 2(1): 114–125, March 1963. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/2/
1/114.full.pdf+html.

Pagel:1966:ERT

[Pag66] Walter Pagel. Essay review: Transition in chemistry: The English Paracelsians. History of Science (UK), 5(1):100–104, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/5/1/100.full.pdf+html.

Pagel:1969:WHR

[Pag69] Walter Pagel. William Harvey revisited. History of Science (UK), 8 (1):1–31, March 1969. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/8/
1/1.full.pdf+html.

Pagel:1970:WHR

[Pag70] Walter Pagel. William Harvey revisited. History of Science (UK), 9 (1):1–41, March 1970. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/9/
1/1.full.pdf+html.

Page:1971:ERG

[Pag71] Leroy E. Page. Essay review: Geomorphology in Britain: The Earth in decay. History of Science (UK), 10(1):118–122, March 1971. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/10/1/118.full.pdf+html.

Pagel:1973:ERH

[Pag73] Walter Pagel. Essay review: The historical dimension to philosophy: Historisches Wörterbuch der Philosophie. Völlig neu bearbeitete Ausgabe des “Wörterbuchs der philosophischen Begriffe” von Rudolph Eisler. History of Science (UK), 11(3):231–235, September 1973. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/11/3/231.full.pdf+html.

Pagel:1974:ERB

[Pag74a] Walter Pagel. Essay review: The History of Mineral Terminology: Studien zur Geschichte der Mineralnamen in Pharmazie, Chemie und Medizin von den Anfängen bis auf Paracelsus, Sudhoffs Archiv. History of Science (UK), 12(1):70–76, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/12/1/70.full.pdf+html.

Pagel:1974:RPS

[Pag74b] Walter Pagel. Recent Paracelsian studies. History of Science (UK), 12(3):200–211, September 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/12/3/200.full.pdf+html.

Pagel:1975:BRE

[Pag75a] Walter Pagel. Book review: Encyclopaedias of philosophy: Historisches Wörterbuch der Philosophie (German) [Historical dictionary of philosophy]. History of Science (UK), 13(1):71–72, March 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/13/1/71.full.pdf+html.

Pagel:1975:ERC

[Pag75b] Walter Pagel. Essay review: The “claim” of Cesalpino and the first and second editions of his “Peripatetic Questions”: Andreae Caesalpini Aretini Medici Clarissimi atque Philosophi Subtilissimi Peritissimi Perpateticarum Questionum. Libri Quinque. History of Science (UK), 13(2):130–138, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/13/2/130.full.pdf+html.

Pagel:1981:ERM

[Pag81] Walter Pagel. Essay review: Mother nature and the scientific revolution: The death of nature: Women, ecology and the scientific revolution. History of Science (UK), 19(2):148–153, June 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/19/2/148.full.pdf+html.

Page:2016:GVW

[Pag16] Arnaud Page. “The greatest victory which the chemist has won in the fight (…) against nature”: Nitrogenous fertilizers in Great Britain and the British Empire, 1910s–1950s. History of Science (UK), 54(4): 383–398, December 2016. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275316681801.

Palter:1993:BAA

[Pal93] Robert Palter. Black Athena, Afro-Centrism, and the History of Science. History of Science (UK), 31(3):227–287, September 1993. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/31/3/227.full.pdf+html. See

response [Ber94].

Palladino:1994:WDM

[Pal94] Paolo Palladino. Wizards and devotees: On the Mendelian theory of inheritance and the professionalization of agricultural science in Great Britain and the United States, 1880–1930. History of Science (UK), 32(4):409–444, December 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/32/4/409.full.pdf+html.

Palladino:1996:PII

[Pal96] Paolo Palladino. People, institutions, and ideas: American and British geneticists at the Cold Spring Harbor symposium on quantitative biology, June 1955. History of Science (UK), 34(4):411–450, December 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/34/4/
411.full.pdf+html.

Palti:1997:BRE

[Pal97] Elías José Palti. Book review: The emergence of morphology, biology takes form: Animal morphology and the German universities, 1800– 1900. History of Science (UK), 35(1):114–116, March 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/35/1/114.full.pdf+html.

Palmieri:2005:GCI

[Pal05] Paolo Palmieri. Galileo’s construction of idealized fall in the void. History of Science (UK), 43(4):343–389, December 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/43/4/343.full.pdf+html.

Palmieri:2007:SAG

[Pal07] Paolo Palmieri. Science and authority in Giacomo Zabarella. History of Science (UK), 45(4):404–427, December 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/45/4/404.full.pdf+html.

Parry:1984:PSC

[Par84] G. J. R. Parry. Puritanism, science and capitalism: William Harrison and the rejection of Hermes Trismegistus. History of Science (UK), 22(3):245–270, September 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/22/3/245.full.pdf+html.

Parshall:1988:AAA

[Par88] Karen Hunger Parshall. The art of algebra from Al-Khwarizmi to Viète: a study in the natural selection of ideas. History of Science (UK), 26(2):129–164, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/26/2/129.full.pdf+html.

Park:2020:MMF

[Par20] Buhm Soon Park. Making matters of fraud: Sociomaterial technology in the case of Hwang and Schatten. History of Science (UK), 58 (4):393–416, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275320921687.

Paul:1994:ERG

[Pau94] Diane B. Paul. Essay review: German Rassenhygiene, health, race and German politics between national unification and Nazism, 1870–1945. History of Science (UK), 32(4):484–486, December 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/32/4/484.full.pdf+html.

Park:2018:SSS

[PC18] Hyung Wook Park and Kyuhoon Cho. Science, state, and spirituality: Stories of four creationists in South Korea. History of Science (UK), 56(1):35–71, March 2018. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317740268.

Pumfrey:2004:SPE

[PD04] Stephen Pumfrey and Frances Dawbarn. Science and patronage in England, 1570–1625: a preliminary study. History of Science (UK), 42(2):137–188, June 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/2/137.full.pdf+html.

Peierls:1965:ERC

[Pei65] R. E. Peierls. Essay review: The concept of the positron. History of Science (UK), 4(1):124–129, March 1965. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/4/1/124.full.pdf+html;   https://
ui.adsabs.harvard.edu/#abs/1965HisSc...4..124P.

Pelis:2001:BSF

[Pel01] Kim Pelis. Blood standards and failed fluids: Clinic, lab, and transfusion solutions in London, 1868–1916. History of Science (UK), 39(2): 185–213, June 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/39/2/
185.full.pdf+html.

Pepper:1967:STH

[Pep67] Jon V. Pepper. The study of Thomas Harriot’s manuscripts: II. Harriot’s unpublished papers. History of Science (UK), 6(1):17– 40, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/6/1/17.
full.pdf+html.

Pepper:1976:SCH

[Pep76] Jon V. Pepper. Some clarifications of Harriot’s solution of Mercator’s problem. History of Science (UK), 14(4):235–244, December 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/14/4/235.full.
pdf+html.

Perrin:1987:RRC

[Per87] C. E. Perrin. Revolution or reform: The chemical revolution and Eighteenth Century concepts of scientific change. History of Science (UK), 25(4):395–423, December 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/25/4/395.full.pdf+html.

Pethers:2010:BMP

[Pet10] Matthew Pethers. “Balloon madness”: Politics, public entertainment, the transatlantic science of flight, and Late Eighteenth-Century America. History of Science (UK), 48(2):181–226, June 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/48/2/181.full.pdf+html.

Petrocelli:2023:MMW

[Pet23] Carla Petrocelli. Maszyny Matematyczne, women, and computing: The birth of computers in the Polish communist era. History of Science (UK), 61(3):409–435, September 1, 2023. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753221121502.

Phillips:2006:RWE

[Phi06] Christopher Phillips. Robert Woodhouse and the evolution of Cambridge mathematics. History of Science (UK), 44(1):69–93, March 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/44/1/69.full.
pdf+html.

Phillips:2015:FBG

[Phi15] Denise Phillips. Francis Bacon and the Germans: Stories from when ‘science’ meant ‘ Wissenschaft’. History of Science (UK), 53(4):378– 394, December 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Pickstone:1981:BLB

[Pic81] John V. Pickstone. Bureaucracy, liberalism and the body in postrevolutionary France: Bichat’s physiology and the Paris school of medicine. History of Science (UK), 19(2):115–142, June 1981. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/19/2/115.full.pdf+html.

Pickstone:1984:FFK

[Pic84] John V. Pickstone. Ferriar’s fever to Kay’s cholera: Disease and social structure in Cottonopolis. History of Science (UK), 22(4):401–419, December 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/22/4/
401.full.pdf+html.

Pickstone:1988:ERS

[Pic88] J. V. Pickstone. Essay review: Science in France: Lamarck, science and medicine in France: The emergence of experimental physiology 1790–1855, death is a social disease: Public health and political economy in early industrial France, Georges Cuvier: Vocation, science and authority in Post-Revolutionary France, Georges Cuvier: Vocation, science and authority in Post-Revolutionary France. History of Science (UK), 26(2):201–211, June 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/26/2/201.full.pdf+html.

Pickstone:1994:MSP

[Pic94] John V. Pickstone. Museological science? The place of the analytical/comparative in Nineteenth-Century science, technology and medicine. History of Science (UK), 32(2):111–138, June 1994. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/32/2/111.full.pdf+html.

Pickstone:1995:PPK

[Pic95] John V. Pickstone. Past and present knowledges in the practice of the history of science. History of Science (UK), 33(2):203–224, June 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/33/2/203.full.
pdf+html.

Pickstone:1999:HMW

[Pic99] John V. Pickstone. How might we map the cultural fields of science? Politics and organisms in Restoration France. History of Science (UK), 37(3):347–364, September 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/37/3/347.full.pdf+html.

Pickstone:2011:BIW

[Pic11a] John V. Pickstone. A brief introduction to ways of knowing and ways of working. History of Science (UK), 49(3):235–245, September 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/49/3/235.full.
pdf+html.

Pickstone:2011:NHA

[Pic11b] John V. Pickstone. Natural histories, analyses and experimentation: Three afterwords. History of Science (UK), 49(3):349–374, September 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/49/3/349.full.
pdf+html.

Piegorsch:1986:GMC

[Pie86] Walter W. Piegorsch. The Gregor Mendel controversy: Early issues of goodness-of-fit and recent issues of genetic linkage. History of Science (UK), 24(2):173–182, June 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/24/2/173.full.pdf+html.

Pozzo:2002:PSK

[PO02] Riccardo Pozzo and Michael Oberhausen. The place of science in Kant’s University. History of Science (UK), 40(3):353–368, September 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/40/3/353.full.
pdf+html.

Porter:1973:ERH

[Por73] Roy Porter. Essay review: The history of palaeontology: The meaning of fossils. History of Science (UK), 11(2):130–138, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/11/2/130.full.pdf+html.

Porter:1976:BRB

[Por76] Roy Porter. Book review: Books and plants: British botanical and horticultural literature before 1800. History of Science (UK), 14(2):

140–141, June 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/14/2/
140.full.pdf+html.

Porter:1979:ERP

[Por79] Roy Porter. Essay review: Physics & geology: Lord Kelvin and the age of the Earth. History of Science (UK), 17(3):216–220, September 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/17/3/216.full.
pdf+html.

Porter:1980:ERDb

[Por80a] Roy Porter. Essay review: Danger: Science at work: Technology and toil in Nineteenth Century Britain. History of Science (UK), 18(4): 303–304, December 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/4/303.full.pdf+html.

Porter:1980:ERDa

[Por80b] Roy Porter. Essay review: Dictionaries and the history of science. History of Science (UK), 18(2):148–150, June 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/18/2/148.full.pdf+html.

Porter:1981:BRB

[Por81] Roy Porter. Book review: The bibliography of geology: Geologists and the history of geology. an international bibliography from the origins to 1978. History of Science (UK), 19(3):224–226, September 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/19/3/224.full.
pdf+html.

Porter:1982:ERO

[Por82] Roy Porter. Essay review: The omniologists: Gentlemen of science: Early years of the British association for the advancement of science. History of Science (UK), 20(3):232–233, September 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/20/3/232.full.pdf+html.

Porter:1987:ERM

[Por87a] Roy Porter. Essay review: Malthus and Darwin: The works of Thomas Robert Malthus. History of Science (UK), 25(2):215–216, June 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL http://hos.sagepub.com/content/25/2/215.
full.pdf+html.

Porter:1987:ERS

[Por87b] Roy Porter. Essay review: Strata and their data. History of Science (UK), 25(3):324–326, September 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/25/3/324.full.pdf+html.

Porter:1991:ERD

[Por91] Dorothy Porter. Essay review: Drinking water, a science of impurity: Water analysis in Nineteenth Century Britain. History of Science (UK), 29(4):429–432, December 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/29/4/429.full.pdf+html.

Porter:1992:CDJ

[Por92] Dorothy Porter. Changing disciplines: John Ryle and the making of social medicine in Britain in the 1940s. History of Science (UK), 30(2): 137–164, June 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/2/
137.full.pdf+html.

Porter:1996:ERD

[Por96] Roy Porter. Essay review: Dugald Stewart reprinted, the Collected Works of Dugald Stewart. History of Science (UK), 34(2):241–244, June 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL http://hos.sagepub.com/content/34/2/241.
full.pdf+html.

Porter:2001:BRJ

[Por01] Roy Porter. Book review: To justify the ways of Boyle to man: The works of Robert Boyle, Robert Boyle (1627–91): Scrupulosity and science. History of Science (UK), 39(2):241–248, June 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/39/2/241.full.pdf+html.

Porter:2005:IHT

[Por05] Theodore M. Porter. Introduction: Historicizing the two cultures. History of Science (UK), 43(2):109–114, June 2005. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/43/2/109.full.pdf+html.

Porter:2012:DDA

[Por12] Dorothy Porter. Darwinian disease archaeology: Genomic variants and the eugenic debate. History of Science (UK), 50(4):432–452, December 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/50/4/
432.full.pdf+html.

Portuondo:2017:ISR

[Por17] María M. Portuondo. Iberian science: Reflections and studies. History of Science (UK), 55(2):123–132, June 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275317701146.

Porter:2020:GCM

[Por20] Jim Wynter Porter. Guidance counseling in the mid-twentieth century United States: Measurement, grouping, and the making of the intelligent self. History of Science (UK), 58(2):191–215, July 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275319874977.

Poskett:2015:NTT

[Pos15] James Poskett. National types: The transatlantic publication and reception of Crania Americana (1839). History of Science (UK), 53 (3):264–295, September 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Poynter:1962:HM

[Poy62] F. N. L. Poynter. The history of medicine in 1960–61. History of Science (UK), 1(1):44–56, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/44.full.pdf+html.

Poynter:1963:ERH

[Poy63] F. N. L. Poynter. Essay review: Henry Sigerist and medicine in the Ancient World. History of Science (UK), 2(1):148–152, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/2/1/148.full.pdf+html.

Poynter:1964:BRN

[Poy64] F. N. L. Poynter. Book review: Naissance de la clinique: Une archéologie du regard médical. History of Science (UK), 3(1):140– 143, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/3/1/
140.full.pdf+html.

Pimentel:2017:YWW

[PPT17] Juan Pimentel and José Pardo-Tomás. And yet, we were modern. The paradoxes of Iberian science after the Grand Narratives. History of Science (UK), 55(2):133–147, June 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275316684353.

Pretsell:2020:EQG

[Pre20] Douglas Pretsell. The evolution of the questionnaire in German sexual science: a methodological narrative. History of Science (UK), 58(3): 326–349, September 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275319881014.

Present:2022:PVM

[Pre22] Pieter Present. Petrus van Musschenbroek (1692–1761) and the early Leiden jar: a discussion of the neglected manuscripts. History of Science (UK), 60(1):103–129, March 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/00732753211000186.

Pugliese:2004:ERL

[Pug04] Patri J. Pugliese. Essay review: Light on Hooke: London’s Leonardo: The life and work of Robert Hooke, the curious life of Robert Hooke: The man who measured London. History of Science (UK), 42(3):361– 366, September 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/3/361.full.pdf+html.

Pumfrey:1991:IAH

[Pum91] Stephen Pumfrey. Ideas above his station: a social study of Hooke’s curatorship of experiments. History of Science (UK), 29(1):1–44, March 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/29/1/
1.full.pdf+html.

Pyenson:2011:EMC

[PV11] Lewis Pyenson and Christophe Verbruggen. Elements of the modernist creed in Henri Pirenne and George Sarton. History of Science (UK), 49(4):377–394, December 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/49/4/377.full.pdf+html.

Yen:2015:PCC

[pY15] Hsiao pei Yen. From palaeoanthropology in China to Chinese palaeoanthropology: Science, imperialism and nationalism in North China, 1920–1939. History of Science (UK), 53(1):21–56, March 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/53/1/21.full.pdf+html.

Pyenson:1977:WGW

[Pye77] Lewis Pyenson. “Who the guys were”: Prosopography in the history of science. History of Science (UK), 15(3):155–188, September 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/15/3/155.full.
pdf+html.

Pyenson:1982:CIE

[Pye82] Lewis Pyenson. Cultural imperialism and exact sciences: German expansion overseas 1900–1930. History of Science (UK), 20(1):1– 43, March 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/20/1/1.
full.pdf+html.

Pyenson:1989:WGH

[Pye89] Lewis Pyenson. What is the good of history of science? History of Science (UK), 27(4):353–389, December 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/27/4/353.full.pdf+html.

Pyenson:1990:END

[Pye90] Lewis Pyenson. Einstein’s natural daughter. History of Science (UK), 28(4):365–379, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/28/4/365.full.pdf+html.

Pyenson:1993:PEI

[Pye93] Lewis Pyenson. Prerogatives of European intellect: Historians of science and the promotion of Western civilization. History of Science (UK), 31(3):289–315, September 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/31/3/289.full.pdf+html.

Pyenson:1995:IRU

[Pye95] Lewis Pyenson. Inventory as a route to understanding: Sarton, Neugebauer, and sources. History of Science (UK), 33(3):253–282, September 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/33/3/253.full.
pdf+html.

Pyenson:2002:CHS

[Pye02a] Lewis Pyenson. Comparative history of science. History of Science (UK), 40(1):1–33, March 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/40/1/1.full.pdf+html.

Pyenson:2002:ENS

[Pye02b] Lewis Pyenson. An end to national science: The meaning and the extension of local knowledge. History of Science (UK), 40(3):251–290, September 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/40/3/
251.full.pdf+html.

Quirk:1962:SM

[Qui62] R. N. Quirk. Scientific manuscripts. History of Science (UK), 1(1): 128–130, March 1962. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/1/
1/128.full.pdf+html.

Qureshi:2004:DSB

[Qur04] Sadiah Qureshi. Displaying Sara Baartman, the ‘Hottentot Venus’. History of Science (UK), 42(2):233–257, June 2004. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/42/2/233.full.pdf+html.

Rabinovitch:2013:CBS

[Rab13] Oded Rabinovitch. Chameleons between science and literature: Observation, writing, and the Early Parisian Academy of Sciences in the literary field. History of Science (UK), 51(1):33–62, March 2013. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/51/1/33.full.pdf+html.

Raby:2019:SBE

[Rab19] Megan Raby. “Slash-and-burn ecology”: Field science as land use. History of Science (UK), 57(4):441–468, December 1, 2019.

CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318819656.

Radder:1983:KFT

[Rad83] Hans Radder. Kramers and the Forman theses. History of Science (UK), 21(2):165–182, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/21/2/165.full.pdf+html.

Ragep:2007:CHI

[Rag07] F. Jamil Ragep. Copernicus and his Islamic predecessors: Some historical remarks. History of Science (UK), 45(1):65–81, March 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/45/1/65.full.pdf+html.

Rampelt:2008:LWJ

[Ram08] Jason M. Rampelt. The last word: John Wallis on the origin of the Royal Society. History of Science (UK), 46(2):177–201, June 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/46/2/177.full.
pdf+html.

Ramalingam:2010:NHD

[Ram10] Chitra Ramalingam. Natural history in the dark: Seriality and the electric discharge in Victorian physics. History of Science (UK), 48 (3–4):371–398, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Rappaport:1964:PSH

[Rap64] Rhoda Rappaport. Problems and sources in the history of geology, 1749–1810. History of Science (UK), 3(1):60–78, March 1964. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/3/1/60.full.pdf+html.

Rappaport:1969:ERG

[Rap69] Rhoda Rappaport. Essay review: Government patronage of science in Eighteenth-Century France: Agronomie et agronomes en France au XVIIP siècle. History of Science (UK), 8(1):119–136, March 1969. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/8/1/119.full.pdf+html.

Rappaport:1991:IEC

[Rap91] Rhoda Rappaport. Italy and Europe: The case of Antonio Vallisneri (1661–1730). History of Science (UK), 29(1):73–98, March 1991. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/29/1/73.full.pdf+html.

Raphael:2014:TSD

[Rap14] Renee Raphael. Teaching sunspots: Disciplinary identity and scholarly practice in the Collegio Romano. History of Science (UK), 52(2): 130–152, June 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/52/2/
130.full.pdf+html.

Raposo:2021:SDC

[Rap21] Pedro M. P. Raposo. The sphere and the dome: The Calouste Gulbenkian Planetarium in Lisbon and the imperial myth of the Estado Novo. History of Science (UK), 59(2):179–196, June 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318797809.

Rashed:1989:PTG

[Ras89] Roshdi Rashed. Problems of the transmission of Greek scientific thought into Arabic: Examples from mathematics and optics. History of Science (UK), 27(2):199–209, June 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/27/2/199.full.pdf+html.

Rasmussen:1997:MCB

[Ras97] Nicolas Rasmussen. The mid-century biophysics bubble: Hiroshima and the biological revolution in America, revisited. History of Science (UK), 35(3):245–293, September 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/35/3/245.full.pdf+html.

Ravetz:1962:ERA

[Rav62a] J. R. Ravetz. Essay review: Action at a distance: Forces and fields. History of Science (UK), 1(1):78–82, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/78.full.pdf+html.

Ravetz:1962:ERP

[Rav62b] J. R. Ravetz. Essay review: “Popularising dead science”: Two books for schools: a history of Western technology. History of Science (UK), 1(1):106–107, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/106.full.pdf+html.

Ravetz:1963:BRH

[Rav63a] J. R. Ravetz. Book review: a history of astronomy. History of Science (UK), 2(1):166–167, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/2/1/166.full.pdf+html.

Ravetz:1963:BRR

[Rav63b] J. R. Ravetz. Book review: Roger Joseph Boscovich. History of Science (UK), 2(1):167–168, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/2/1/167.full.pdf+html.

Ravetz:1966:BRN

[Rav66a] J. R. Ravetz. Book review: The natural philosopher. History of Science (UK), 5(1):158, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/5/1/158.full.pdf+html.

Ravetz:1966:BRP

[Rav66b] J. R. Ravetz. Book review: The philosophy of Francis Bacon. History of Science (UK), 5(1):147–150, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/5/1/147.full.pdf+html.

Root-Bernstein:1983:MM

[RB83] Robert Scott Root-Bernstein. Mendel and methodology. History of Science (UK), 21(3):275–295, September 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/21/3/275.full.pdf+html.

Ruiz-Castell:2013:SII

[RC13] Pedro Ruiz-Castell. Seeing the invisible: The introduction and development of electron microscopy in Britain, 1935–1945. History of Science (UK), 51(2):221–249, June 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/51/2/221.full.pdf+html.

Rees:2006:EBS

[Ree06] Amanda Rees. Ecology, biology and social life: Explaining the origins of primate sociality. History of Science (UK), 44(4):409–434, December 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/44/4/
409.full.pdf+html.

Rees:2009:UDI

[Ree09] Amanda Rees. The undead Darwin: Iconic narrative, scientific controversy and the history of science. History of Science (UK), 47(4): 445–457, December 2009. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
47/4/445.full.pdf+html.

Reingold:1981:SSH

[Rei81] Nathan Reingold. Science, scientists, and historians of science. History of Science (UK), 19(4):274–283, December 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/19/4/274.full.pdf+html.

Reingold:1994:SGU

[Rei94] Nathan Reingold. Science and government in the United States since 1945. History of Science (UK), 32(4):361–386, December 1994. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/32/4/361.full.pdf+html.

Reis:2007:SDP

[Rei07] Fernando Egídio Reis. Scientific dissemination in Portuguese encyclopaedic periodicals, 1779–1820. History of Science (UK), 45(1):83– 118, March 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/1/
83.full.pdf+html.

Ren:2023:GCL

[Ren23] Chao Ren. Global circulation of low-end expertise: Knowledge, hierarchy, and labor migration in a Burmese oilfield. History of Science (UK), 61(4):??, December 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753231189442.

Reynolds:2008:EHT

[Rey08] Andrew Reynolds. Ernst Haeckel and the theory of the cell state: Remarks on the history of a bio-political metaphor. History of Science (UK), 46(2):123–152, June 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/46/2/123.full.pdf+html.

Roll-Hansen:1988:PEG

[RH88] Nils Roll-Hansen. The progress of Eugenics: Growth of knowledge and change in ideology. History of Science (UK), 26(3):295–331, September 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/26/3/295.full.
pdf+html.

Rousseau:1999:VCR

[RH99] G. S. Rousseau and David Haycock. Voices calling for reform: The Royal Society in the mid-Eighteenth Century — Martin Folkes, John Hill, and William Stukeley. History of Science (UK), 37(4):377–406, December 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/4/
377.full.pdf+html.

Rousseau:2000:JCC

[RH00] G. S. Rousseau and David Haycock. The Jew of Crane Court: Emanuel Mendes Da Costa (1717–91), natural history and natural excess. History of Science (UK), 38(2):127–170, June 2000. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/38/2/127.full.pdf+html.

Rheinberger:2011:IET

[Rhe11] Hans-Jörg Rheinberger. Infra-experimentality: From traces to data, from data to patterning facts. History of Science (UK), 49(3):337–348, September 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/49/3/
337.full.pdf+html.

Rice:1996:AMH

[Ric96] Adrian Rice. Augustus De Morgan: Historian of science. History of Science (UK), 34(2):201–240, June 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/34/2/201.full.pdf+html.

Rieppel:2015:PCP

[Rie15] Lukas Rieppel. Plaster cast publishing in nineteenth-century paleontology. History of Science (UK), 53(4):456–491, December 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Riggs:2017:SPP

[Rig17] Christina Riggs. Shouldering the past: Photography, archaeology, and collective effort at the tomb of Tutankhamun. History of Science (UK), 55(3):336–363, September 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275316676282.

Roberts:1966:ERR

[Rob66] R. S. Roberts. Essay review: The Royal College of Physicians of London in the Sixteenth and Seventeenth Centuries: a history of the Royal College of Physicians of London. History of Science (UK), 5 (1):87–100, March 1966. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/5/
1/87.full.pdf+html.

Robinson:2005:CSG

[Rob05] Kate Robinson. The celestial streams of Giulio Camillo. History of Science (UK), 43(3):321–341, September 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/43/3/321.full.pdf+html.

Roberts:2014:AMG

[Rob14a] Lissa Roberts. Accumulation and management in global historical perspective: an introduction. History of Science (UK), 52(3):227–246, September 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/52/3/
227.full.pdf+html.

Roberts:2014:CTC

[Rob14b] Lissa Roberts. ‘Le centre de toutes choses’: Constructing and managing centralization on the Isle de France. History of Science (UK), 52(3):319–342, September 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/52/3/319.full.pdf+html.

Roberts:2015:E

[Rob15] Lissa L. Roberts. Editorial. History of Science (UK), 53(2):123– 124, June 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/53/2/
123.full.pdf+html.

Roberts:2016:EGH

[Rob16a] Lissa Roberts. Exploring global history through the lens of history of chemistry: Materials, identities and governance. History of Science (UK), 54(4):335–361, December 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275316681805.

Roberts:2016:IWE

[Rob16b] Lissa Roberts. An introductory word from the Editor. History of Science (UK), 54(4):333–334, December 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275316682249.

Roberts:2019:I

[Rob19] Lissa L. Roberts. Introduction. History of Science (UK), 57(3):372, September 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL https://journals.sagepub.com/doi/full/
10.1177/0073275319870358.

Roberts:2020:CSI

[Rob20a] Lissa L. Roberts. Contributions to this special issue. History of Science (UK), 58(4):386–392, December 1, 2020. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275320952261.

Roberts:2020:HRI

[Rob20b] Lissa L. Roberts. Historicizing research integrity and fraud. History of Science (UK), 58(4):353, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275320952617.

Roberts:2021:PPB

[Rob21] Lissa L. Roberts. Phases of physics: Building the discipline during the long nineteenth century. History of Science (UK), 59(1):45–46, March 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275321992612.

Roberts:2022:NE

[Rob22] Lissa L. Roberts. A note from the Editor. History of Science (UK), 60(1):3, March 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/00732753211070642.

Roos:2008:LGM

[Roo08] Anna Marie Roos. Lodestones and gallstones: The magnetic iatrochemistry of Martin Lister (1639–1712). History of Science (UK), 46(3):343–364, September 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/46/3/343.full.pdf+html.

Roque:2022:LSS

[Roq22] Ricardo Roque. The Latin stranger-science, or l’anthropologie among the Lusitanians. History of Science (UK), 60(1):69–95, March 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318755291.

Rose:2024:GHD

[Ros24] Edwin D. Rose. George Howard Darwin and the “public” interpretation of The Tides. History of Science (UK), 62(1):??, March 1, 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231181548.

Rothenberg:1987:EBC

[Rot87] Albert Rothenberg. Einstein, Bohr, and creative thinking in science. History of Science (UK), 25(2):147–166, June 1987. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/25/2/147.full.pdf+html.

Rousseau:1979:ERS

[Rou79] G. S. Rousseau. Essay review: Science and antiquarianism: Dr Woodward’s shield: History, science, and satire in Augustan England. History of Science (UK), 17(2):142–144, June 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/17/2/142.full.pdf+html.

Rowold:1998:BRE

[Row98] Katharina Rowold. Book review: an Eighteenth-Century female Professor, “Und sie fürchtet sich vor niemandem”: Die Physikern Laura Bassi (1711–1778). History of Science (UK), 36(3):361–362, September 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/36/3/361.full.
pdf+html.

Rowold:1999:BRW

[Row99] Katharina Rowold. Book review: Women and science, Osiris. History of Science (UK), 37(1):114–115, March 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/37/1/114.full.pdf+html.

Rashed:2012:ONH

[RP12] Roshdi Rashed and Lewis Pyenson. Otto Neugebauer, historian. History of Science (UK), 50(4):402–431, December 2012. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/50/4/402.full.pdf+html.

Roberts:2023:HSL

[RRH23a] Lissa Roberts, Seth Rockman, and Alexandra Hui. Historiographies of science and labor: From past perspectives to future possibilities. History of Science (UK), 61(4):??, December 1, 2023. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753231209023.

Roberts:2023:SWI

[RRH23b] Lissa Roberts, Seth Rockman, and Alexandra Hui. Science and/ as work: an introduction to this special issue. History of Science (UK), 61(4):??, December 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/abs/10.1177/00732753231206580.

Robb-Smith:1962:BRE

[RS62] A. H. T. Robb-Smith. Book review: The evolution of medical practice in Britain. History of Science (UK), 1(1):111–112, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/1/1/111.full.pdf+html.

Roberts:2020:IHS

[RSM20] Lissa L. Roberts, H. Otto Sibum, and Cyrus C. M. Mody. Integrating the history of science into broader discussions of research integrity and fraud. History of Science (UK), 58(4):354–368, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275320952268.

Russell:2000:MLH

[RTL00] N. C. Russell, E. M. Tansey, and P. V. Lear. Missing links in the history and practice of science: Teams, technicians and technical work.

History of Science (UK), 38(2):237–241, June 2000. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/38/2/237.full.pdf+html.

Rudwick:1962:ERP

[Rud62] M. J. S. Rudwick. Essay review: The principle of uniformity: Natural law and divine miracle. History of Science (UK), 1(1):82–86, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/1/1/82.
full.pdf+html.

Rudwick:1963:BRC

[Rud63] M. J. S. Rudwick. Book review: Charles Lyell. History of Science (UK), 2(1):172, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/2/1/172.2.full.pdf+html.

Rudwick:1969:ERG

[Rud69] M. J. S. Rudwick. Essay review: The glacial theory. History of Science (UK), 8(1):136–157, March 1969. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/8/1/136.full.pdf+html.

Rudwick:1976:EVL

[Rud76] Martin J. S. Rudwick. The emergence of a visual language for geological science 1760–1840. History of Science (UK), 14(3):149–195, September 1976. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/14/3/
149.full.pdf+html.

Rudwick:1980:ERS

[Rud80] Martin Rudwick. Essay review: Social order and the natural world: Natural order: Historical studies of scientific culture. History of Science (UK), 18(4):269–285, December 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/18/4/269.full.pdf+html.

Rudwick:1981:ITI

[Rud81] Martin Rudwick. Individuals and their interactions in science past and present: Introduction. History of Science (UK), 19(1):1–5, March 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/19/1/
1.full.pdf+html.

Rudwick:2003:RPH

[Rud03] Martin Rudwick. Roy Porter, historian of geology. History of Science (UK), 41(3):251–256, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/41/3/251.full.pdf+html.

Rupke:1983:SFR

[Rup83] N. A. Rupke. The study of fossils in the romantic philosophy of history and nature. History of Science (UK), 21(4):389–413, December 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/21/4/389.full.
pdf+html.

Rupp:1990:MLD

[Rup90] Jan C. C. Rupp. Matters of life and death: The social and cultural conditions of the rise of anatomical theatres, with special reference to Seventeenth Century Holland. History of Science (UK), 28(3):263– 287, September 1990. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
28/3/263.full.pdf+html.

Rupke:1996:EIC

[Rup96] Nicolaas A. Rupke. Eurocentric ideology of continental drift. History of Science (UK), 34(3):251–272, September 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/34/3/251.full.pdf+html.

Rupke:1998:EHE

[Rup98] Nicolaas A. Rupke. “The end of history” in the early picturing of geological time. History of Science (UK), 36(1):61–90, March 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/36/1/61.full.pdf+html.

Ruse:1974:ERD

[Rus74] Michael Ruse. Essay review: The Darwin industry — a critical evaluation: The triumph of the Darwinian method, Charles Darwin: The years of controversy, Wallace and natural selection. History of Science (UK), 12(1):43–58, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/12/1/43.full.pdf+html.

Rusnock:1994:ERJ

[Rus94] Andrea Rusnock. Essay review: Jonas Moore, Sir Jonas Moore: Practical mathematics and restoration science. History of Science (UK), 32 (1):108–109, March 1994. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
32/1/108.full.pdf+html.

Ruse:2009:DST

[Rus09] Michael Ruse. Darwinian struggles: But is there progress? History of Science (UK), 47(4):407–430, December 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/47/4/407.full.pdf+html.

Ryan:1966:SMR

[Rya66] W. F. Ryan. Science in medieval Russia: Some reflections on a recent book. History of Science (UK), 5(1):52–61, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/5/1/52.full.pdf+html.

Sabra:1987:ASN

[Sab87] A. I. Sabra. The appropriation and subsequent naturalization of Greek science in medieval Islam: a preliminary statement. History of Science (UK), 25(3):223–243, September 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/25/3/223.full.pdf+html.

Sampson:2023:LSV

[Sam23] Paul E Sampson. “The lungs of a ship”: Ventilation, acclimatization, and labor in the maritime environment, 1740–1800. History of Science (UK), 61(2):214–235, June 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753211046449.

Sandler:1979:SRP

[San79] Iris Sandler. Some reflections on the protean nature of the scientific precursor. History of Science (UK), 17(3):170–190, September 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/17/3/170.full.
pdf+html.

Santesmases:2017:CBI

[San17] María Jesús Santesmases. Circulating biomedical images: Bodies and chromosomes in the post-eugenic era. History of Science (UK), 55(4):

395–430, December 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317701145.

Sarasohn:2004:WWT

[Sar04] Lisa T. Sarasohn. Who was then the gentleman?: Samuel Sorbière, Thomas Hobbes, and the Royal Society. History of Science (UK), 42 (2):211–232, June 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/2/211.full.pdf+html.

Savino:2017:CSI

[Sav17] Melania Savino. Connecting sites and images: Archeology as controversial knowledge in modern Izmir. History of Science (UK), 55(3): 364–382, September 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317726419.

Sturdy:1998:SSM

[SC98] Steve Sturdy and Roger Cooter. Science, scientific management, and the transformation of medicine in Britain c. 1870–1950. History of Science (UK), 36(4):421–466, December 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/4/421.full.pdf+html.

Schofield:1963:HSS

[Sch63a] Robert E. Schofield. Histories of scientific societies: Needs and opportunities for research. History of Science (UK), 2(1):70–83, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/2/1/70.
full.pdf+html.

Schramm:1963:ERA

[Sch63b] Matthias Schramm. Essay review: Aristotelianism: Basis and obstacle to scientific progress in the Middle Ages: Augustine to Galileo. History of Science (UK), 2(1):91–113, March 1963. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/2/1/91.full.pdf+html.

Schramm:1965:ERS

[Sch65] Matthias Schramm. Essay review: Steps towards the idea of function: a comparison between Eastern and Western science of the Middle Ages: Augustine to Galileo. History of Science (UK), 4(1):70– 103, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/4/1/70.
full.pdf+html.

Schiller:1968:PSI

[Sch68] J. Schiller. Physiology’s struggle for independence in the first half of the Nineteenth Century. History of Science (UK), 7(1):64–89, March 1968. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/7/1/64.
full.pdf+html.

Schmitt:1973:TRR

[Sch73] Charles B. Schmitt. Towards a reassessment of Renaissance Aristotelianism. History of Science (UK), 11(3):159–193, September 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/11/3/159.full.
pdf+html.

Schiller:1974:QAU

[Sch74] Joseph Schiller. Queries, answers and unsolved problems in Eighteenth Century biology. History of Science (UK), 12(3):184–199, September 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/3/
184.full.pdf+html.

Schaffer:1977:BRC

[Sch77] Simon Schaffer. Book review: Clouds of light: The search for the nebulae. History of Science (UK), 15(2):152–153, June 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/15/2/152.full.pdf+html.

Schmitt:1978:ERR

[Sch78] Charles B. Schmitt. Essay review: Reappraisals in Renaissance science: Hermeticism and the scientific revolution. History of Science (UK), 16(3):200–214, September 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/16/3/200.full.pdf+html.

Schaffer:1982:ERM

[Sch82a] Simon Schaffer. Essay review: The matter of ether: Conceptions of ether: Studies in the history of ether theories 1740–1900. History of Science (UK), 20(4):297–303, December 1982. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/20/4/297.full.pdf+html.

Schaffer:1982:ERN

[Sch82b] Simon Schaffer. Essay review: The Newtonian revolution revisited: The Newtonian revolution. History of Science (UK), 20(2):140–144, June 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL http://hos.sagepub.com/content/20/2/140.
full.pdf+html.

Schaffer:1983:NPP

[Sch83] Simon Schaffer. Natural philosophy and public spectacle in the Eighteenth Century. History of Science (UK), 21(1):1–43, March 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/21/1/1.full.pdf+html.

Schaffer:1984:ERN

[Sch84a] Simon Schaffer. Essay review: Newton’s undergraduate notebook. History of Science (UK), 22(1):93–97, March 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/22/1/93.full.pdf+html.

Schaffer:1984:PQH

[Sch84b] Simon Schaffer. Priestley’s questions: an historiography survey. History of Science (UK), 22(2):151–183, June 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/22/2/151.full.pdf+html.

Schlich:1995:HGS

[Sch95] Thomas Schlich. How gods and saints became transplant surgeons: The scientific article as a model for the writing of history. History of Science (UK), 33(3):311–331, September 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/33/3/311.full.pdf+html.

Schaffer:2003:EBE

[Sch03] Simon Schaffer. Enlightenment brought down to Earth. History of Science (UK), 41(3):257–268, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/41/3/257.full.pdf+html.

Schaffer:2006:ICC

[Sch06] Simon Schaffer. Instruments as cargo in the China trade. History of Science (UK), 44(2):217–246, June 2006. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/44/2/217.full.pdf+html.

Schlich:2007:SSM

[Sch07] Thomas Schlich. Surgery, science and modernity: Operating rooms and laboratories as spaces of control. History of Science (UK), 45(3): 231–256, September 2007. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
45/3/231.full.pdf+html.

Schaffer:2009:NBI

[Sch09a] Simon Schaffer. Newton on the beach: The information order of Principia Mathematica. History of Science (UK), 47(3):243–276, September 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/47/3/243.full.
pdf+html.

Schickore:2009:TOM

[Sch09b] Jutta Schickore. Test objects for microscopes. History of Science (UK), 47(2):117–145, June 2009. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/47/2/117.full.pdf+html.

Schaffer:2010:LS

[Sch10a] Simon Schaffer. Lovejoy’s series. History of Science (UK), 48(3–4): 483–494, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Schlanger:2010:SPA

[Sch10b] Nathan Schlanger. Series in progress: Antiquities of nature, numismatics and stone implements in the emergence of prehistoric archaeology. History of Science (UK), 48(3–4):343–369, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Scheid:2018:PFF

[Sch18] Volker Scheid. Promoting free flow in the networks: Reimagining the body in early modern Suzhou. History of Science (UK), 56(2): 131–167, June 2018. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317709406.

Schweig:2024:PDR

[Sch24] Alexander Schweig. Progressing into disaster: The railroad and the spread of cholera in a provincial Ottoman town. History of Science (UK), 62(4):539–561, December 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753221113151.

Strasser:2011:CER

[SdC11] Bruno J. Strasser and Soraya de Chadarevian. The comparative and the exemplary: Revisiting the early history of molecular biology. History of Science (UK), 49(3):317–336, September 2011. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/49/3/317.full.pdf+html.

Secord:1985:NNT

[Sec85] James A. Secord. Newton in the nursery: Tom Telescope and the philosophy of tops and balls, 1761–1838. History of Science (UK), 23(2): 127–151, June 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/2/
127.full.pdf+html.

Secord:1986:GSG

[Sec86] James A. Secord. The geological survey of Great Britain as a research school, 1839–1855. History of Science (UK), 24(3):223–275, September 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/3/
223.full.pdf+html.

Secord:1994:SPA

[Sec94] Anne Secord. Science in the pub: Artisan botanists in Early Nineteenth-Century Lancashire. History of Science (UK), 32(3):269– 315, September 1994. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
32/3/269.full.pdf+html.

Sengoopta:1992:SSG

[Sen92] Chandak Sengoopta. Science, sexuality, and gender in the Fin de Siècle: Otto Weininger as Baedeker. History of Science (UK), 30(3): 249–279, September 1992. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
30/3/249.full.pdf+html.

Sengoopta:2000:BRB

[Sen00a] Chandak Sengoopta. Book review: Beauty in One’s own eye: Making the body beautiful: a cultural history of aesthetic surgery, creating beauty to cure the soul: Race and psychology in the shaping of aesthetic surgery. History of Science (UK), 38(3):357–359, September 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/38/3/357.full.
pdf+html.

Sengoopta:2000:MOC

[Sen00b] Chandak Sengoopta. The modern ovary: Constructions, meanings, uses. History of Science (UK), 38(4):425–488, December 2000. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/38/4/425.full.pdf+html.

Silva:2018:REE

[SH18] Cibelle Celestino Silva and Peter Heering. Re-examining the early history of the Leiden jar: Stabilization and variation in transforming a phenomenon into a fact. History of Science (UK), 56(3):314–342, September 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://journals.sagepub.com/doi/full/
10.1177/0073275318768418.

Shaffer:1974:CNP

[Sha74a] E. S. Shaffer. Coleridge and natural philosophy: a review of recent literary and historical research. History of Science (UK), 12(4):284–298, December 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/12/4/
284.full.pdf+html.

Shapin:1974:ASE

[Sha74b] Steven Shapin. The audience for science in Eighteenth Century Edinburgh. History of Science (UK), 12(2):95–121, June 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/12/2/95.full.pdf+html.

Shapin:1981:ERL

[Sha81] Steven Shapin. Essay review: Licking Leibniz: Philosophers at war: The quarrel between Newton and Leibniz. History of Science (UK), 19(4):293–305, December 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/19/4/293.full.pdf+html.

Shapin:1982:HSS

[Sha82] Steven Shapin. History of science and its sociological reconstructions. History of Science (UK), 20(3):157–211, September 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/20/3/157.full.pdf+html.

Shapin:1991:SGP

[Sha91a] Steven Shapin. “A scholar and a gentleman”: The problematic identity of the scientific practitioner in Early Modern England. History of Science (UK), 29(3):279–327, September 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/29/3/279.full.pdf+html.

Shapiro:1991:EMI

[Sha91b] Barbara Shapiro. Early modern intellectual life: Humanism, religion and science in Seventeenth Century England. History of Science (UK), 29(1):45–71, March 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/29/1/45.full.pdf+html.

Shapin:1992:DBH

[Sha92] Steven Shapin. Discipline and bounding: The history and sociology of science as seen through the externalism–internalism debate. History of Science (UK), 30(4):333–369, December 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/30/4/333.full.pdf+html.

Shapin:2022:HSS

[Sha22] Steven Shapin. Hard science, soft science: a political history of a disciplinary array. History of Science (UK), 60(3):287–328, September 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753221094739.

Shen:2015:TYD

[She15] Yubin Shen. Too young to date! The origins of zaolian (early love) as a social problem in 20th-century China. History of Science (UK), 53 (1):86–101, March 2015. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
53/1/86.full.pdf+html.

Sherman:2019:IMP

[She19] Brad Sherman. Intangible machines: Patent protection for software in the United States. History of Science (UK), 57(1):18–37, March 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318770781.

Shih:2019:MPN

[Shi19] Ashanti Shih. The most perfect natural laboratory in the world: Making and knowing Hawaii National Park. History of Science (UK), 57 (4):493–517, December 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275319848966.

Shortland:1983:ERL

[Sho83] Michael Shortland. Essay review: Limits to growth: The mismeasure of man. History of Science (UK), 21(4):425–429, December 1983. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/21/4/425.full.pdf+html.

Shortland:1986:ERB

[Sho86] Michael Shortland. Essay review: Bodies of history: Some problems and perspectives: a history of women’s bodies, the body and society, Michel Foucault. History of Science (UK), 24(3):303–326, September 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/24/3/303.full.
pdf+html.

Shortland:1987:ERD

[Sho87] Michael Shortland. Essay review: Darwinian structures: Darwinism and divinity: Essays on evolution and religious belief. History of Science (UK), 25(2):195–213, June 1987. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/25/2/195.full.pdf+html.

Shortland:1994:DVU

[Sho94] Michael Shortland. Darkness visible: Underground culture in the Golden age of geology. History of Science (UK), 32(1):1–61, March 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/32/1/1.full.
pdf+html.

Shryock:1966:BRS

[Shr66] Richard H. Shryock. Book review: a short history of medicine. History of Science (UK), 5(1):150–152, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/5/1/150.full.pdf+html.

Smith:2003:IGO

[SHW03] Crosbie Smith, Ian Higginson, and Phillip Wolstenholme. “Imitations of God’s own works”: Making trustworthy the ocean steamship. History of Science (UK), 41(4):379–426, December 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/41/4/379.full.pdf+html.

Sibum:2020:WEE

[Sib20] H. Otto Sibum. When is enough enough? Accurate measurement and the integrity of scientific research. History of Science (UK), 58 (4):437–457, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275320939696.

Silverstein:1999:ENP

[Sil99] Arthur M. Silverstein. “The end is near!”: The phenomenon of the declaration of closure in a discipline. History of Science (UK), 37(4): 407–425, December 1999. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
37/4/407.full.pdf+html.

Silva:2012:PPS

[Sil12] José Alberto Silva. The Portuguese popularizer of science Teodoro de Almeida: Agendas, publics, and bilingualism. History of Science (UK), 50(1):93–122, March 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/50/1/93.full.pdf+html.

Simon:2012:CNE

[Sim12a] Josep Simon. Cross-national education and the making of science, technology and medicine. History of Science (UK), 50(3):251–256, September 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/50/3/
251.full.pdf+html.

Simon:2012:SMT

[Sim12b] Josep Simon. Secondary matters: Textbooks and the making of physics in Nineteenth-Century France and England. History of Science (UK), 50(3):339–374, September 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/50/3/339.full.pdf+html.

Simpson:2017:CMK

[Sim17] Thomas Simpson. “Clean out of the map”: Knowing and doubting space at India’s high imperial frontiers. History of Science (UK), 55 (1):3–36, March 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275316686580.

Singh:2022:SAP

[Sin22] Charu Singh. The shastri and the air-pump: Experimental fictions and fictions of experiment for Hindi readers in colonial north India. History of Science (UK), 60(2):232–254, June 1, 2022. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275320987421.

Siorvanes:1988:ERS

[Sio88] Lucas Siorvanes. Essay review: Science, philosophy and religion: Philoponus and the search for unity in late antiquity: Time, creation and the continuum: Theories in antiquity and the Early Middle Ages, Philoponus, and the rejection of Aristotelian science. History of Science (UK), 26(1):93–102, March 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/26/1/93.full.pdf+html.

Sivasundaram:2001:NHS

[Siv01] Sujit Sivasundaram. Natural history spiritualized: Civilizing islanders, cultivating breadfruit, and collecting souls. History of Science (UK), 39(4):417–443, December 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/39/4/417.full.pdf+html.

Sastre-Juan:2021:SAP

[SJ21] Jaume Sastre-Juan. “Science in action”: The politics of hands-on display at the new York museum of science and industry. History of Science (UK), 59(2):155–178, June 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275317725239.

Suarez:1986:IES

[SL86] María Matilde Suárez and Walewska Lemoine. From internalism to externalism: a study of academic resistance to new scientific findings. History of Science (UK), 24(4):383–410, December 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/24/4/383.full.pdf+html.

Slaughter:2024:PMP

[Sla24] Aimee Slaughter. Performing the Manhattan Project in Los Alamos. History of Science (UK), 62(2):305–325, June 1, 2024. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753231187011.

Sloan:1990:ERD

[Slo90] Phillip R. Sloan. Essay review: Deconstructing evolution: The politics of evolution: Morphology, medicine, and reform in radical London. History of Science (UK), 28(4):419–428, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/28/4/419.full.pdf+html.

Sloan:1994:ERB

[Slo94] Phillip R. Sloan. Essay review: Buffon studies today, Buffon: Un philosophe au jardin du roi, Buffon 88: Actes du colloque international pour le bicentenaire de la mort de Buffon, les sciences de la vie dans la pensée française du XVIII^{e} siècle. History of Science (UK), 32(4):469–477, December 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/32/4/469.full.pdf+html.

Slater:2017:BBB

[SLT17] John Slater and Maríaluz López-Terrada. Being beyond: The Black Legend and how we got over it. History of Science (UK), 55(2): 148–166, June 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275317694897.

Suay-Matallana:2024:CTE

[SM24] Ignacio Suay-Matallana. Chemistry, trade, and the economy: Exploring the history of customs laboratories in the United States (1870s–1930s). History of Science (UK), 62(3):391–415, September 2024. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753231197292.

Smeaton:1963:NLL

[Sme63] W. A. Smeaton. New light on Lavoisier: The research of the last ten years. History of Science (UK), 2(1):51–69, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/2/1/51.full.pdf+html.

Smeaton:1964:BRHa

[Sme64a] W. A. Smeaton. Book review: a history of chemistry. History of Science (UK), 3(1):148–149, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/3/1/148.full.pdf+html.

Smeaton:1964:BRHb

[Sme64b] W. A. Smeaton. Book review: Historical studies in the language of chemistry. History of Science (UK), 3(1):150–151, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/3/1/150.full.pdf+html.

Smith:1973:BPP

[Smi73] Roger Smith. The background of physiological psychology in natural philosophy. History of Science (UK), 11(2):75–123, June 1973. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/11/2/75.full.pdf+html.

Smith:1975:ERN

[Smi75] Norman A. F. Smith. Essay review: Nineteenth Century civil engineering: The Victorian city. History of Science (UK), 13(2):104– 113, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/2/
104.full.pdf+html.

Smith:1978:NCB

[Smi78] Crosbie Smith. A new chart for British natural philosophy: The development of energy physics in the Nineteenth Century. History of Science (UK), 16(4):231–279, December 1978. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/16/4/231.full.pdf+html.

Smith:1981:ERM

[Smi81] Robert W. Smith. Essay review: Modern astronomy: a source book in astronomy and astrophysics, 1900–1975. History of Science (UK), 19(4):306–307, December 1981. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/19/4/306.full.pdf+html.

Smith:1982:ERH

[Smi82] Roger Smith. Essay review: History and psychology: Historiography of modern psychology: Aims, sources, approaches. History of Science (UK), 20(2):144–147, June 1982. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/20/2/144.full.pdf+html.

Smith:1988:ERO

[Smi88] Roger Smith. Essay review: Origins of neuroscience: Nineteenth-Century origins of neuroscientific concepts, medicine, mind and the double brain: a study in Nineteenth-Century thought. History of Science (UK), 26(4):427–437, December 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/26/4/427.full.pdf+html.

Smith:1994:ERH

[Smi94] Roger Smith. Essay review: History of psychology in synthesis, mental machinery: The origins and consequences of psychological ideas. History of Science (UK), 32(1):103–106, March 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/32/1/103.full.pdf+html.

Smith:2012:LNH

[Smi12] Justin E. H. Smith. Leibniz on natural history and national history. History of Science (UK), 50(4):377–401, December 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/50/4/377.full.pdf+html.

Smith:2014:HLS

[Smi14] Christopher Smith. How I learned to stop worrying and love the bombe: Machine research and development and Bletchley Park. History of Science (UK), 52(2):200–222, June 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/52/2/200.full.pdf+html.

Smith:2015:SLH

[Smi15] Russell Smith. Shining a light on Harriot and Galileo: On the mechanics of reflection and projectile motion. History of Science (UK), 53(3):296–319, September 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Smith:2020:WDW

[Smi20] Elise Smith. “Why do we measure mankind?” marketing anthropometry in late-Victorian Britain. History of Science (UK), 58(2): 142–165, July 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275319842977.

Sommer:2010:SMO

[Som10] Marianne Sommer. Seriality in the making: The Osborn–Knight restorations of evolutionary history. History of Science (UK), 48(3–4): 461–482, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Somsen:2021:PCS

[Som21] Geert Somsen. The princess at the conference: Science, pacifism, and Habsburg society. History of Science (UK), 59(4):434–460, December 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564

(electronic). URL https://journals.sagepub.com/doi/full/10.
1177/0073275320977750.

Somsen:2022:ASP

[Som22] Geert Somsen. Afterword: Science popularization, dictatorships, and democracies. History of Science (UK), 60(3):430–435, September 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753211073422.

Sorrenson:2001:DSP

[Sor01] Richard Sorrenson. Dollond & Son’s pursuit of achromaticity, 1758– 1789. History of Science (UK), 39(1):31–55, March 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/39/1/31.full.pdf+html.

Southern:1963:ERT

[Sou63] R. W. Southern. Essay review: Technology and history: Medieval technology and social change. History of Science (UK), 2(1):130– 135, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/2/1/
130.full.pdf+html.

Southgate:1992:PIP

[Sou92] Beverley C. Southgate. “The power of imagination”: Psychological explanations in mid-Seventeenth-Century England. History of Science (UK), 30(3):281–294, September 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/30/3/281.full.pdf+html.

Sheets-Pyenson:1987:CSD

[SP87] Susan Sheets-Pyenson. Cathedrals of science: The development of colonial natural history museums during the Late Nineteenth Century. History of Science (UK), 25(3):279–300, September 1987. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/25/3/279.full.pdf+html.

Sheets-Pyenson:1990:NDS

[SP90] Susan Sheets-Pyenson. New directions for scientific biography: The case of Sir William Dawson. History of Science (UK), 28(4):399–410, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/28/4/
399.full.pdf+html.

Smith-Peter:2007:DRP

[SP07] Susan Smith-Peter. Defining the Russian people: Konstantin Arsen’ev and Russian statistics before 1861. History of Science (UK), 45(1): 47–64, March 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/45/1/
47.full.pdf+html.

Spary:2004:SS

[Spa04] E. C. Spary. Scientific symmetries. History of Science (UK), 42(1): 1–46, March 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/42/1/1.
full.pdf+html.

Spary:2008:ERH

[Spa08] Emma Spary. Essay review: The history of the indigenous: Inventing the indigenous: Local knowledge and natural history in early modern Europe. History of Science (UK), 46(3):369–370, September 2008. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/46/3/369.full.
pdf+html.

Speiser:1962:BRL

[Spe62] D. Speiser. Book review: Lettere a una Principessa Tedesca. (Italian) [Letters to a German princess]. History of Science (UK), 1(1): 108–110, March 1962. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/1/
1/108.full.pdf+html.

Spelda:2015:CCI

[Spe15] Daniel Spelda. From closed cycles to infinite progress: Early modern historiography of astronomy. History of Science (UK), 53(2):209– 233, June 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/53/2/
209.full.pdf+html.

Shahare:2020:HCS

[SR20] Mahendra Shahare and Lissa L. Roberts. Historicizing the crisis of scientific misconduct in Indian science. History of Science (UK), 58 (4):485–506, December 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275320930908.

Sutherland:1980:FOP

[SS80] Gillian Sutherland and Stephen Sharp. “The fust official psychologist in the wurrld”: Aspects of the professionalization of psychology in Early Twentieth Century Britain. History of Science (UK), 18(3): 181–208, September 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/3/181.full.pdf+html.

Sera-Shriar:2015:APV

[SS15] Efram Sera-Shriar. Anthropometric portraiture and Victorian anthropology: Situating Francis Galton’s photographic work in the late 1870s. History of Science (UK), 53(2):155–179, June 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/53/2/155.full.pdf+html.

Shapin:1974:PRT

[ST74] Steven Shapin and Arnold Thackray. Prosopography as a research tool in history of science: The British scientific community 1700–1900. History of Science (UK), 12(1):1–28, March 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/12/1/1.full.pdf+html.

Stauffer:1963:BRC

[Sta63] R. C. Stauffer. Book review: Charles Darwin: The founder of the theory of evolution and natural selection. History of Science (UK), 2 (1):169–170, March 1963. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/2/
1/169.full.pdf+html.

Star:1986:TCB

[Sta86] Susan Leigh Star. Triangulating clinical and basic research: British localizationists, 1870–1906. History of Science (UK), 24(1):29–48, March 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/1/
29.full.pdf+html.

Staley:2013:THH

[Sta13] Richard Staley. Trajectories in the history and historiography of physics in the Twentieth century. History of Science (UK), 51(2): 151–177, June 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/51/2/
151.full.pdf+html.

Staley:2021:SSW

[Sta21] Richard Staley. Sensory studies, or when physics was psychophysics: Ernst Mach and physics between physiology and psychology, 1860–71. History of Science (UK), 59(1):93–118, March 1, 2021. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275318784104.

Stern:1965:BRM

[Ste65] S. M. Stern. Book review: Mediaeval Arabic bookmaking and its relation to early chemistry and pharmacology (Transactions of the American Philosophical Society). History of Science (UK), 4(1):151, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/4/1/151.full.
pdf+html.

Stevenson:1991:ERA

[Ste91] Christine Stevenson. Essay review: Art and optics, the science of art: Optical themes in Western art from Brunelleschi to Seurat. History of Science (UK), 29(1):99–101, March 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/29/1/99.full.pdf+html.

Steinle:1995:LSC

[Ste95] Friedrich Steinle. Looking for a “simple case”: Faraday and electromagnetic rotation. History of Science (UK), 33(2):179–202, June 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/33/2/179.full.
pdf+html.

Stewart:1996:STS

[Ste96] Larry Stewart. Seeing through the scholium: Religion and reading Newton in the Eighteenth Century. History of Science (UK), 34(2): 123–165, June 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/34/2/
123.full.pdf+html.

Stewart:2002:ERB

[Ste02] Ian G. Stewart. Essay review: “Books & How to Use Them”: Generall Learning: a Seventeenth-Century treatise on the formation of the general scholar by Meric Casaubon. History of Science (UK), 40(2): 233–244, June 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/40/2/
233.full.pdf+html.

Stewart:2005:ERN

[Ste05] Ian G. Stewart. Essay review: a new Novum Organum: a new Bacon?: The Instauratio Magna, Part II: Book review: Novum Organum and associated texts. History of Science (UK), 43(4):457–466, December 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/43/4/457.full.
pdf+html.

Stewart:2007:ESK

[Ste07] Larry Stewart. Experimental spaces and the knowledge economy. History of Science (UK), 45(2):155–177, June 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/45/2/155.full.pdf+html.

Stewart:2020:WFE

[Ste20] Ian B. Stewart. William Frédéric Edwards and the study of human races in France, from the Restoration to the July Monarchy. History of Science (UK), 58(3):275–300, September 1, 2020. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/0073275319889832.

Sticker:1964:ERH

[Sti64] B. Sticker. Essay review: Herschel’s cosmology: William Herschel and the construction of the heavens. History of Science (UK), 3(1): 91–101, March 1964. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/3/
1/91.full.pdf+html.

Stiefel:1992:PCI

[Sti92] Tina Stiefel. The problem of causal inference in the empirical world of Nicholas of Autrecourt. History of Science (UK), 30(3):295–309, September 1992. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/30/3/
295.full.pdf+html.

Stokes:1986:RZ

[Sto86] T. D. Stokes. Reason in the ZEITGEIST. History of Science (UK), 24 (2):111–123, June 1986. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
24/2/111.full.pdf+html.

Stolberg:2018:LAL

[Sto18] Michael Stolberg. Learning anatomy in late sixteenth-century Padua. History of Science (UK), 56(4):381–402, December 1, 2018. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275318794581.

Storring:2022:SPW

[Sto22] Adam L. Storring. Subjective practices of war: The Prussian army and the Zorndorf campaign, 1758. History of Science (UK), 60(4): 458–480, December 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/abs/10.1177/0073275320958950.

Strickland:1995:GDB

[Str95] Stuart Strickland. Galvanic disciplines: The boundaries, objects, and identities of experimental science in the Era of Romanticism. History of Science (UK), 33(4):449–468, December 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/33/4/449.full.pdf+html.

Strang:2020:MSP

[Str20] Cameron B. Strang. Measuring souls: Psychometry, female instruments, and subjective science, 1840–1910. History of Science (UK), 58(1):76–100, March 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.sagepub.

com/doi/full/10.1177/0073275319847065.

Studer:1977:ISG

[Stu77] Kenneth E. Studer. Interpreting scientific growth: a comment on Derek Price’s “Science since Babylon”. History of Science (UK), 15 (1):44–51, March 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/1/44.full.pdf+html. See [Her62b].

Surman:2019:STB

[Sur19] Jan Surman. Science and terminology in-between empires: Ukrainian science in a search for its language in the nineteenth century. History of Science (UK), 57(2):260–287, June 1, 2019. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:

//journals.sagepub.com/doi/full/10.1177/0073275318800637.

See corrigendum [Ano20].

Sutherland:1975:BRH

[Sut75] Gillian Sutherland. Book review: The history of science education: Science for the people. History of Science (UK), 13(4):303–304, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/4/
303.full.pdf+html.

Sutton:1986:ERS

[Sut86] M. A. Sutton. Essay review: Spectroscopy, historiography and myth: The Victorians vindicated. History of Science (UK), 24(4):425–432, December 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/4/
425.full.pdf+html.

Suzuki:1995:DTP

[Suz95] Akihito Suzuki. Dualism and the transformation of psychiatric language in the Seventeenth and Eighteenth Centuries. History of Science (UK), 33(4):417–447, December 1995. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/33/4/417.full.pdf+html.

Scheid:2018:IHS

[SV18] Volker Scheid and Curie Virág. Introduction to history of science special section on tong [Chinese text]. History of Science (UK), 56 (2):123–130, June 2018. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275318770832.

Swetz:1993:RTC

[Swe93] Frank Swetz. Right triangle concepts in ancient China: From application to theory. History of Science (UK), 31(4):421–439, December 1993. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/31/4/421.full.
pdf+html.

Sysling:2020:MST

[Sys20] Fenneke Sysling. Measurement, self-tracking and the history of science: an introduction. History of Science (UK), 58(2):103–116, July 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275319865830.

Szamosi:1990:PMC

[Sza90] Geza Szamosi. Polyphonic music and classical physics: The origin of Newtonian time. History of Science (UK), 28(2):175–191, June 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/28/2/175.full.
pdf+html.

Simoes:2013:HTR

[SZDC13] Ana Simões, Isabel Zilhão, Maria Paula Diogo, and Ana Carneiro. Halley turns republican: How the Portuguese Press presented the 1910 return of Halley’s Comet. History of Science (UK), 51(2):199– 219, June 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/51/2/
199.full.pdf+html.

Talbot:1965:BRS

[Tal65] C. H. Talbot. Book review: The Salernitan questions: an introduction to the history of medieval and Renaissance problem literature. History of Science (UK), 4(1):148–150, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/4/1/148.full.pdf+html.

Tampakis:2015:OFL

[Tam15] Kostas Tampakis. The once and future language: Communication, terminology and the practice of science in nineteenth and early twentieth century Greece. History of Science (UK), 53(4):438–455, December 2015. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Tamborini:2016:IAC

[Tam16] Marco Tamborini. “If the Americans can do it, so can we”: How dinosaur bones shaped German paleontology. History of Science (UK), 54(3):225–256, September 2016. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/54/3/225.full.pdf+html.

Tanner:1967:STH

[Tan67] Rosalind C. H. Tanner. The study of Thomas Harriot’s manuscripts: I. Harriot’s will. History of Science (UK), 6(1):1–16, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/6/1/1.full.pdf+html.

Tansey:1998:BRS

[Tan98] E. M. Tansey. Book review: a science city in the USSR, new Atlantis revisited: Akademgorodok, the Siberian city of science. History of Science (UK), 36(3):365–368, September 1998. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/36/3/365.full.pdf+html.

Tarrant:2014:CSS

[Tar14] Neil Tarrant. Censoring science in Sixteenth-Century Italy: Recent (and not-so-recent) research. History of Science (UK), 52(1): 1–27, March 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/52/1/1.
full.pdf+html.

Taylor:1994:EMD

[Tay94] Alan B. H. Taylor. An episode with May-dew. History of Science (UK), 32(2):163–184, June 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/32/2/163.full.pdf+html.

Teich:1975:ERS

[Tei75] Mikulás Teich. Essay review: a single path to the double helix?: The path to the double helix. History of Science (UK), 13(4):264–283, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/4/
264.full.pdf+html.

Teich:1980:ERH

[Tei80] Mikulas Teich. Essay review: a history of biochemistry: History of the identification of the sources of free energy in organisms, early studies on biosynthesis. History of Science (UK), 18(1):46–67, March 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/18/1/46.full.pdf+html.

Teich:1990:RGJ

[Tei90] Mikulás Teich. Reflecting on the Golden jubilee of Bernal’s the social function of science. History of Science (UK), 28(4):411–418, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/28/4/411.full.
pdf+html.

Teich:2003:HIA

[Tei03] Mikulás Teich. How it all began: From the Enlightenment in national context to revolution in history. History of Science (UK), 41(3):335– 343, September 2003. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
41/3/335.full.pdf+html.

Temkin:1963:ERP

[Tem63] Owsei Temkin. Essay review: Physics in Aristotle: Aristotle’s system of the physical world. History of Science (UK), 2(1):135–139, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/2/1/135.full.
pdf+html.

Terrall:1990:CSF

[Ter90] Mary Terrall. The culture of science in Frederick the Great’s Berlin. History of Science (UK), 28(4):333–364, December 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/28/4/333.full.pdf+html.

Terrall:1995:EDC

[Ter95] Mary Terrall. Émilie Du Châtelet and the gendering of science. History of Science (UK), 33(3):283–310, September 1995. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/33/3/283.full.pdf+html.

Terrall:2004:VVR

[Ter04] Mary Terrall. Vis Viva revisited. History of Science (UK), 42(2):189– 209, June 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/42/2/
189.full.pdf+html.

Tetzlaff:2023:CAP

[Tet23] Stefan Tetzlaff. Contested “automobility”: Peasants, townsfolks, and infrastructures of road transport in interwar central and western India (c. 1919–39). History of Science (UK), 61(1):77–101, March 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
00732753221125055.

teHeesen:2004:NHI

[tH04] Anke te Heesen. From natural historical investment to state service: Collectors and collections of the Berlin society of friends of nature research, c. 1800. History of Science (UK), 42(1):113–131, March 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/42/1/113.full.
pdf+html.

Thackray:1966:ERC

[Tha66] A. W. Thackray. Essay review: The chemistry of history: The development of modern chemistry: a history of chemistry. History of Science (UK), 5(1):124–134, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/5/1/124.full.pdf+html.

Thackray:1970:STI

[Tha70] Arnold Thackray. Science and technology in the industrial revolution. History of Science (UK), 9(1):76–89, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/9/1/76.full.pdf+html.

Thoren:1973:TBP

[Tho73] Victor E. Thoren. Tycho Brahe: Past and future research. History of Science (UK), 11(4):270–282, December 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/11/4/270.full.pdf+html.

Tomaselli:1984:FPD

[Tom84] Sylvana Tomaselli. The first person: Descartes, Locke and mind-body dualism. History of Science (UK), 22(2):185–205, June 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/22/2/185.full.pdf+html.

Tomaselli:1991:ERS

[Tom91a] Sylvana Tomaselli. Essay review: Studying Eighteenth Century psychology, psychology and literature in the Eighteenth Century. History of Science (UK), 29(1):102–104, March 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/29/1/102.full.pdf+html.

Tomaselli:1991:RHS

[Tom91b] Sylvana Tomaselli. Reflections on the history of the science of woman. History of Science (UK), 29(2):185–205, June 1991. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/29/2/185.full.pdf+html.

Tomory:2011:GDI

[Tom11] Leslie Tomory. Gaslight, distillation, and the industrial revolution. History of Science (UK), 49(4):395–424, December 2011. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/49/4/395.full.pdf+html.

Tonn:2019:LDG

[Ton19] Jenna Tonn. Laboratory of domesticity: Gender, race, and science at the Bermuda Biological Station for Research, 1903–30. History of Science (UK), 57(2):231–259, June 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275318797789.

Tresch:2010:OPS

[Tre10] John Tresch. The order of the prophets: Series in Early French social science and socialism. History of Science (UK), 48(3–4):315– 342, September/December 2010. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

Tuchman:2020:BCM

[Tuc20] Arleen Marcia Tuchman. Biometrics and citizenship: Measuring diabetes in the United States in the interwar years. History of Science (UK), 58(2):166–190, July 1, 2020. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/full/10.1177/0073275319869762.

Turner:1969:HOI

[Tur69] G. L’E. Turner. The history of optical instruments: a brief survey of sources and modern studies. History of Science (UK), 8(1):53– 93, March 1969. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/8/1/53.
full.pdf+html.

Turner:1970:ERA

[Tur70] G. L’E. Turner. Essay review: The apparatus of science: The apparatus of science at Harvard 1766–1800. collection of historical scientific instruments, Harvard University. History of Science (UK), 9(1): 129–138, March 1970. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/9/
1/129.full.pdf+html.

Turner:1973:ERH

[Tur73] G. L’E. Turner. Essay review: Book review: The history of the microscope: Das Mikroskop. Theorie, Gebrauch, Geschichte und gegenwärtiger Zustand desselben. (German) [The microscope: Its theory, use, history, and current state]. History of Science (UK), 11(1): 62–67, March 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/1/
62.full.pdf+html.

Turner:1975:BRS

[Tur75] A. J. Turner. Book review: Science and the occult: The occult sciences in the Renaissance: a study in intellectual patterns, the ancient theology: Studies in Christian Platonism from the Fifteenth to the Eighteenth Century. History of Science (UK), 13(4):300–301, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/13/4/300.full.
pdf+html.

Turner:1978:ERH

[Tur78] A. J. Turner. Essay review: The history of Greenwich Observatory: Greenwich Observatory, three volumes by different authors telling the story of Britain’s oldest scientific institution, the Royal Observatory at Greenwich and Herstmonceux 1675–1975, Francis Place and the early history of the Greenwich Observatory. History of Science (UK), 16(3): 222–226, September 1978. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
16/3/222.full.pdf+html.

Turner:1986:ERW

[Tur86] A. J. Turner. Essay review: “A world of wonders in one closet shut”: Elias Ashmole 1617–1692: The founder of the Ashmolean Museum and his world, Tradescant’s rarities: Essays on the foundation of the Ashmolean Museum 1683 with a catalogue of the surviving early collections, the Ashmolean Museum and Oxford science 1683–1983. History of Science (UK), 24(2):209–215, June 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/24/2/209.full.pdf+html.

Turner:1989:SDH

[Tur89] A. J. Turner. Sun-dials: History and classification. History of Science (UK), 27(3):303–318, September 1989. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/27/3/303.full.pdf+html.

Turner:1991:ERA

[Tur91] A. J. Turner. Essay review: Astronomy at Armagh, church, state and astronomy in Ireland: 200 years of Armagh Observatory. History of Science (UK), 29(4):433–434, December 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/29/4/433.full.pdf+html.

Turner:1994:ERM

[Tur94] A. J. Turner. Essay review: Mechanics in the Roberts collection, bibliotheca mechanica. History of Science (UK), 32(1):106–108, March 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/32/1/106.full.
pdf+html.

Unguru:1977:ERI

[Ung77] Sabetai Unguru. Essay review: Incommensurability and irrationality: a new historical interpretation: The evolution of the Euclidean elements. a study of the theory of incommensurable magnitudes and its significance for Early Greek geometry. History of Science (UK), 15(3): 216–227, September 1977. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
15/3/216.full.pdf+html.

Urbanowicz:2023:HSM

[Urb23] Piotr Urbanowicz. Heavenly spirit or material being? Science on electricity at the turn of the 19th century in Poland. History of Science (UK), 61(3):360–382, September 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753221074607.

Uys:2024:TMB

[Uys24] Mia Uys. “On the trail of the mercy bullet”: Pain, scientific showmanship, and the early history of animal tranquilizing, c. 1912–1932. History of Science (UK), 62(3):416–441, September 2024. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/abs/10.1177/00732753231197875.

Valleriani:2012:GAP

[Val12] Matteo Valleriani. Galileo’s abandoned project on acoustic instruments at the Medici Court. History of Science (UK), 50(1):1–31, March 2012. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/50/1/
1.full.pdf+html.

VanHelden:1975:HPI

[Van75] Albert Van Helden. The historical problem of the invention of the telescope. History of Science (UK), 13(4):251–263, December 1975. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/13/4/251.full.pdf+html.

VanDyck:2006:GTS

[Van06] Maarten Van Dyck. Gravitating towards stability: Guidobaldo’s Aristotelian–Archimedean synthesis. History of Science (UK), 44(4): 373–407, December 2006. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
44/4/373.full.pdf+html.

Damme:2017:PMG

[Van17] Stéphane Van Damme. The pillar of metropolitan greatness: The long making of archeological objects in Paris (1711–2001). History of Science (UK), 55(3):302–335, September 2017. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://journals.

sagepub.com/doi/full/10.1177/0073275317698711.

Vincent-Bensaude:1987:HMC

[VB87] Bernadette Vincent-Bensaude. Hélène Metzger’s La Chimie: a popular treatise. History of Science (UK), 25(1):71–84, March 1987. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/25/1/71.full.pdf+html.

vandeRoemer:2004:NNR

[vdR04] Bert van de Roemer. Neat nature: The relation between nature and art in a Dutch cabinet of curiosities from the Early Eighteenth Century. History of Science (UK), 42(1):47–84, March 2004. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/42/1/47.full.pdf+html.

Vernon:1990:PSB

[Ver90] Keith Vernon. Pus, sewage, beer and milk: Microbiology in Britain, 1870–1940. History of Science (UK), 28(3):289–325, September 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/28/3/289.full.
pdf+html.

Verlet:1996:FMN

[Ver96] Loup Verlet. ‘F = MA’ and the Newtonian revolution: an exit from religion through religion. History of Science (UK), 34(3):303–346, September 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/34/3/
303.full.pdf+html.

Vermij:2014:MAA

[Ver14] Rienk Vermij. The marginalization of astrology among Dutch astronomers in the first half of the 17th century. History of Science (UK), 52(2):153–177, June 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/52/2/153.full.pdf+html.

Vigil-Fowler:2021:CBW

[VFD21] Margaret Vigil-Fowler and Sukumar Desai. The community of Black women physicians, 1864–1941: Trends in background, education, and training. History of Science (UK), 59(4):407–433, December 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275320987417.

vanHelvoort:1994:HVR

[vH94] Ton van Helvoort. History of virus research in the Twentieth century: The problem of conceptual continuity. History of Science (UK), 32(2): 185–235, June 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/32/2/
185.full.pdf+html.

vanHelvoort:2003:PSC

[vH03] Ton van Helvoort. ‘Purifying’ science: E. C. Slater and postwar biochemistry in the Netherlands. History of Science (UK), 41(1): 1–34, March 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/41/1/1.
full.pdf+html.

Vigni:2021:FIC

[Vig21] Fabrizio Li Vigni. The failed institutionalization of “complexity science”: a focus on the Santa Fe Institute’s legitimization strategy. History of Science (UK), 59(3):344–369, September 1, 2021. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/full/10.1177/
0073275320938295.

Vlahakis:1999:GES

[Vla99] George N. Vlahakis. The Greek Enlightenment in science: Hermes the scholar and its contribution to science in Early Nineteenth-Century Greece. History of Science (UK), 37(3):319–345, September 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/37/3/319.full.
pdf+html.

Vollgraff:2024:BPG

[VT24] Matthew Vollgraff and Marco Tamborini. Biotechnics and politics: a genealogy of nonhuman technology. History of Science (UK), 62(3): 366–390, September 2024. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
abs/10.1177/00732753231187676.

vanWyhe:2004:WPR

[vW04] John van Wyhe. Was phrenology a reform science? Towards a new generalization for phrenology. History of Science (UK), 42(3):313– 331, September 2004. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
42/3/313.full.pdf+html.

vanWyhe:2009:DOE

[vW09] John van Wyhe. Darwin online and the evolution of the Darwin industry. History of Science (UK), 47(4):459–473, December 2009. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/47/4/459.full.pdf+html.

Waddington:2001:SCT

[Wad01] Keir Waddington. The science of cows: Tuberculosis, research and the state in the United Kingdom, 1890–1914. History of Science (UK), 39(3):355–381, September 2001. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/39/3/355.full.pdf+html.

Walker:1973:ERE

[Wal73] D. P. Walker. Essay review: The elusive Rosicrucians: The Rosicrucian Enlightenment. History of Science (UK), 11(4):306–310, December 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/11/4/306.full.
pdf+html.

Walters:1997:CPS

[Wal97] Alice N. Walters. Conversation pieces: Science and politeness in Eighteenth-Century England. History of Science (UK), 35(2):121– 154, June 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/35/2/
121.full.pdf+html.

Walters:1999:EEE

[Wal99] Alice N. Walters. Ephemeral events: English broadsides of Early Eighteenth-Century solar eclipses. History of Science (UK), 37(1): 1–43, March 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/37/1/1.
full.pdf+html.

Waller:2002:PMF

[Wal02] John C. Waller. Putting method first: Re-appraising the extreme determinism and hard hereditarianism of Sir Francis Galton. History of Science (UK), 40(1):35–62, March 2002. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/40/1/35.full.pdf+html.

Walsh:2022:ESC

[Wal22] Sarah Walsh. The executioner’s shadow: Coerced sterilization and the creation of “Latin” eugenics in Chile. History of Science (UK), 60(1): 18–40, March 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL https://journals.sagepub.com/doi/
full/10.1177/0073275318755533.

Warnock:1965:BRM

[War65] G. J. Warnock. Book review: The myth of metaphor. History of Science (UK), 4(1):146–148, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/4/1/146.full.pdf+html.

Wasserstein:1964:ERS

[Was64] A. Wasserstein. Essay review: Science among the Greeks: The origins of scientific thought. History of Science (UK), 3(1):85–91, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/3/1/85.
full.pdf+html.

Wasserstein:1965:ERG

[Was65] A. Wasserstein. Essay review: Greek science, the Romans and the Middle Ages: Roman science: Origins, development and influence to the Later Middle Ages. History of Science (UK), 4(1):129–138, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/4/1/129.full.
pdf+html.

Waters:1974:ERR

[Wat74] D. W. Waters. Essay review: Renaissance cosmography: a Navigator’s universe. the “Libro de Cosmographia” of 1538 by Pedro de Medina. History of Science (UK), 12(3):227–230, September 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/12/3/227.full.pdf+html.

Walton:2014:WSC

[WB14] Steven A. Walton and Thomas Boothby. What is straight cannot fall: Gothic architecture, scholasticism, and dynamics. History of Science (UK), 52(4):347–376, December 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/52/4/347.full.pdf+html.

Wear:1975:ERH

[Wea75] Andrew Wear. Essay review: The history of Galenism: Galenism. rise and decline of a medical philosophy. History of Science (UK), 13(2): 122–129, June 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/2/
122.full.pdf+html.

Wear:1983:WHW

[Wea83] Andrew Wear. William Harvey and the ‘way of the anatomists’. History of Science (UK), 21(3):223–249, September 1983. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/21/3/223.full.pdf+html.

Webster:1967:ERO

[Web67] C. Webster. Essay review: The origins of the Royal Society: The Royal Society: Concept and creation. History of Science (UK), 6(1): 106–128, March 1967. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/6/
1/106.full.pdf+html.

Weber:1974:SSN

[Web74] Gay Weber. Science and society in Nineteenth Century anthropology. History of Science (UK), 12(4):260–283, December 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/12/4/260.full.pdf+html.

Weber:2014:BFA

[Web14] Andreas Weber. Bitter fruits of accumulation: The case of Caspar Georg Carl Reinwardt (1773–1854). History of Science (UK), 52(3): 297–318, September 2014. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
52/3/297.full.pdf+html.

Weindling:1986:MMS

[Wei86] Paul Weindling. Medicine and modernization: The social history of German health and medicine. History of Science (UK), 24(3):277–301, September 1986. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/24/3/
277.full.pdf+html.

Weisz:1988:SMM

[Wei88] George Weisz. The self-made Mandarin: The Éloges of the French Academy of Medicine, 1824–47. History of Science (UK), 26(1):13– 40, March 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/26/1/
13.full.pdf+html.

Werrett:2000:HNW

[Wer00] Simon Werrett. Healing the nation’s wounds: Royal ritual and experimental philosophy in Restoration England. History of Science (UK), 38(4):377–399, December 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/38/4/377.full.pdf+html.

Werrett:2023:VME

[Wer23] Simon Werrett. Voyages of maintenance: Exploration, infrastructure, and modernity on the Krusenstern–Lisianskii circumnavigation between Russia and Japan from 1803 to 1806. History of Science (UK), 61(3):338–359, September 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https://journals.

sagepub.com/doi/abs/10.1177/00732753211060186.

Westfall:1967:ERS

[Wes67] Richard S. Westfall. Essay review: The science of optics in the Seventeenth Century: Theories of light. History of Science (UK), 6(1): 150–156, March 1967. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/6/
1/150.full.pdf+html.

Westman:1980:ARS

[Wes80] Robert S. Westman. The Astronomer’s role in the Sixteenth Century: a preliminary study. History of Science (UK), 18(2):105–147, June 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/18/2/105.full.
pdf+html.

Westfall:1988:ERG

[Wes88] Richard S. Westfall. Essay review: Galileo heretic: Problems, as they appear to me, with Redondi’s thesis: Galileo heretic. History of Science (UK), 26(4):399–415, December 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/26/4/399.full.pdf+html.

Walter:2022:EHG

[WGvA22] Tilmann Walter, Abdolbaset Ghorbani, and Tinde van Andel. The emperor’s herbarium: The German physician Leonhard Rauwolf (1535-?-96) and his botanical field studies in the Middle East. History of Science (UK), 60(1):130–151, March 1, 2022. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL https:/

/journals.sagepub.com/doi/full/10.1177/00732753211019848.

Whiteside:1962:ERA

[Whi62a] D. T. Whiteside. Essay review: After the Principia: The correspondence of Isaac Newton. History of Science (UK), 1(1):96–100, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/1/1/96.
full.pdf+html.

Whiteside:1962:ERK

[Whi62b] D. T. Whiteside. Essay review: Kepler the mathematician: Gesammelte Werke Band IX: Mathematische Schriften. History of Science (UK), 1(1):86–90, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/1/1/86.full.pdf+html.

Whiteside:1962:EWN

[Whi62c] D. T. Whiteside. The expanding world of Newtonian research. History of Science (UK), 1(1):16–29, March 1962. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/1/1/16.full.pdf+html.

Whiteside:1963:ERS

[Whi63a] D. T. Whiteside. Essay review: Scientific papers of Newton: Unpublished scientific papers of Isaac Newton. History of Science (UK), 2 (1):125–130, March 1963. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/2/
1/125.full.pdf+html.

Whittet:1963:BRH

[Whi63b] T. D. Whittet. Book review: History of pharmacy in Britain. History of Science (UK), 2(1):161–162, March 1963. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/2/1/161.full.pdf+html.

Whiteside:1964:BRB

[Whi64a] D. T. Whiteside. Book review: Les Coniques d’Apollonius de Perge: oeuvres traduites pour la première fois du grec en français, avec une introduction et des notes. History of Science (UK), 3(1):124–127, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/3/1/
124.full.pdf+html.

Whiteside:1964:BRF

[Whi64b] D. T. Whiteside. Book review: Frans van Schooten der Jüngere (Boethius. Texte und Abhandlungen zur Geschichte der exakten Wissenschaften. Band II). History of Science (UK), 3(1):146–148, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/3/1/146.full.
pdf+html.

White:1966:BRT

[Whi66a] Lynn White, Jr. Book review: a theatre of machines. History of Science (UK), 5(1):145–146, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/5/1/145.full.pdf+html.

Whiteside:1966:ERN

[Whi66b] D. T. Whiteside. Essay review: Newtonian dynamics: The background to Newton’s Principia. History of Science (UK), 5(1):104– 117, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/5/1/
104.full.pdf+html.

Whiteside:1967:FLN

[Whi67] D. T. Whiteside. A face-lift for Newton: Current facsimile reprints. History of Science (UK), 6(1):59–68, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/6/1/59.full.pdf+html.

Whitrow:1971:ERW

[Whi71] G. J. Whitrow. Essay review: Why is the sky dark at night?: The paradox of Olbers’ Paradox: a case history of scientific thought. History of Science (UK), 10(1):128–132, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/10/1/128.full.pdf+html.

Whitrow:1973:ERM

[Whi73] Magda Whitrow. Essay review: Multi-volume works in progress (2): The Isis cumulative bibliography. History of Science (UK), 11(2):139– 142, June 1973. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/11/2/
139.full.pdf+html.

Whitrow:1974:ERC

[Whi74] G. J. Whitrow. Essay review: The cosmological significance of the milky way: The milky way: an elusive road for science, the discovery of our galaxy. History of Science (UK), 12(4):299–306, December 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/12/4/299.full.
pdf+html.

Whiteside:1975:ERS

[Whi75] D. T. Whiteside. Essay review: In search of Thomas Harriot: Thomas Harriot: Renaissance scientist. History of Science (UK), 13(1):61– 70, March 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/13/1/
61.full.pdf+html.

Whitteridge:1980:PH

[Whi80] Gweneth Whitteridge. The Piltdown Hippocrates. History of Science (UK), 18(3):226–230, September 1980. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/18/3/226.full.pdf+html.

Whitrow:1983:ERA

[Whi83] G. J. Whitrow. Essay review: Astronomy of the stars: Stellar astronomy: Historical studies. History of Science (UK), 21(2):211–213, June 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/21/2/211.full.
pdf+html.

Whitaker:1984:SSH

[Whi84] M. A. B. Whitaker. Science, scientists and history of science. History of Science (UK), 22(4):421–424, December 1984. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/22/4/421.full.pdf+html.

White:1996:SHS

[Whi96] Paul White. Science at home: The space between Henrietta Heathorn and Thomas Huxley. History of Science (UK), 34(1):33–56, March 1996. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/34/1/33.full.
pdf+html.

White:2005:MCA

[Whi05] Paul White. Ministers of culture: Arnold, Huxley and liberal Anglican reform of learning. History of Science (UK), 43(2):115–138, June 2005. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/43/2/115.full.
pdf+html.

Whitmer:2017:IUT

[Whi17] Kelly J. Whitmer. Imagining uses for things: Teaching “useful knowledge” in the early eighteenth century. History of Science (UK), 55 (1):37–60, March 2017. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://journals.sagepub.com/doi/
full/10.1177/0073275316678872.

Wightman:1963:BRS

[Wig63] William P. D. Wightman. Book review: The scientific Renaissance. History of Science (UK), 2(1):160–161, March 1963. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/2/1/160.full.pdf+html.

Wightman:1964:SR

[Wig64] W. P. D. Wightman. Science and the Renaissance. History of Science (UK), 3(1):1–19, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/3/1/1.full.pdf+html.

Williams:1962:PSF

[Wil62] L. Pearce Williams. The physical sciences in the first half of the Nineteenth Century: Problems and sources. History of Science (UK), 1(1):1–15, March 1962. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/1/
1/1.full.pdf+html.

Wilkie:1965:ERH

[Wil65a] J. S. Wilkie. Essay review: Harvey’s immediate debt to Aristotle and to Galen: The anatomical lectures of William Harvey. prelectiones anatomie universalis: De musculis. History of Science (UK), 4(1): 103–124, March 1965. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/4/
1/103.full.pdf+html.

Williams:1965:BRA

[Wil65b] L. Pearce Williams. Book review: Alexander von Humboldt. History of Science (UK), 4(1):150–151, March 1965. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/4/1/150.full.pdf+html.

Wilkie:1967:ERP

[Wil67] J. S. Wilkie. Essay review: Preformation and epigenesis: a new historical treatment: Les sciences de la vie dans la pensée française du XVIII e siècle. History of Science (UK), 6(1):138–150, March 1967. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic).

URL http://hos.sagepub.com/content/6/1/138.full.pdf+html.

Wilson:1974:ERA

[Wil74] David B. Wilson. Essay review: Aether studies: Nineteenth Century aether theories, the ethereal aether: a history of the Michelson– Morley-Miller aether drift experiments, 1880–1930. History of Science (UK), 12(3):220–227, September 1974. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/12/3/220.full.pdf+html.

Wilson:1977:VSR

[Wil77] David B. Wilson. Victorian science and religion. History of Science (UK), 15(1):52–67, March 1977. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/15/1/52.full.pdf+html.

Wilde:1980:HNP

[Wil80a] C. B. Wilde. Hutchinsonianism, natural philosophy and religious controversy in Eighteenth Century Britain. History of Science (UK), 18 (1):1–24, March 1980. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
18/1/1.full.pdf+html.

Williams:1980:ERE

[Wil80b] L. Pearce Williams. Essay review: The essential Thomas Kuhn: The essential tension: Selected studies in scientific tradition and change. History of Science (UK), 18(1):68–74, March 1980. CODEN HISCAR.

ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/18/1/68.full.pdf+html.

Williams:1983:WTS

[Wil83] M. E. W. Williams. Was there such a thing as stellar astronomy in the Eighteenth Century? History of Science (UK), 21(4):369–388, December 1983. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/21/4/
369.full.pdf+html.

Williams:1985:ERS

[Wil85] Perry Williams. Essay review: Science, religion, and the spiritual world: The other world: Spiritualism and psychical research in England, 1850–1914. History of Science (UK), 23(4):435–440, December 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/23/4/435.full.
pdf+html.

Wilson:2000:HDC

[Wil00] Adrian Wilson. On the history of disease-concepts: The case of pleurisy. History of Science (UK), 38(3):271–319, September 2000. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/38/3/271.full.
pdf+html.

Wilson:2002:EAB

[Wil02] David A. H. Wilson. Experimental animal behaviour studies: The loss of initiative in Britain 100 years ago. History of Science (UK), 40(3): 291–320, September 2002. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/
40/3/291.full.pdf+html.

Winter:1994:MPC

[Win94] Alison Winter. Mesmerism and popular culture in Early Victorian England. History of Science (UK), 32(3):317–343, September 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/32/3/317.full.
pdf+html.

Wintroub:1997:LGF

[Win97] Michael Wintroub. The looking glass of facts: Collecting, rhetoric and citing the self in the experimental natural philosophy of Robert Boyle. History of Science (UK), 35(2):189–217, June 1997. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/35/2/189.full.pdf+html.

Winterburn:2003:BRC

[Win03] Emily Winterburn. Book review: Caroline Herschel’s perspective: Caroline Herschel’s autobiographies, the Herschel partnership: As viewed by Caroline. History of Science (UK), 41(3):351–354, September 2003. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/41/3/351.full.
pdf+html.

Withers:1999:THG

[Wit99] Charles W. J. Withers. Towards a history of geography in the public sphere. History of Science (UK), 37(1):45–78, March 1999. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/37/1/45.full.pdf+html.

Wragge-Morley:2022:MCA

[WM22] Alexander Wragge-Morley. Medicine, connoisseurship, and the animal body. History of Science (UK), 60(4):481–499, December 1, 2022. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL https://journals.sagepub.com/doi/abs/10.1177/
0073275320949001.

Woodward:1985:CHP

[Woo85] William R. Woodward. Committed history and philosophy of the social sciences in the two Germanies. History of Science (UK), 23(1): 25–72, March 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/23/1/
25.full.pdf+html.

Wood:1988:ERB

[Woo88] P. B. Wood. Essay review: Behemoth v. the sceptical chymist: Leviathan and the air-pump: Hobbes, Boyle, and the experimental life. History of Science (UK), 26(1):103–109, March 1988. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/26/1/103.full.pdf+html.

Wood:1990:NHM

[Woo90] P. B. Wood. The natural history of man in the Scottish Enlightenment. History of Science (UK), 28(1):89–123, March 1990. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/28/1/89.full.pdf+html.

Wood:1991:ERE

[Woo91] P. B. Wood. Essay review: The enlightenment of the dons, Cambridge in the Age of the Enlightenment: Science, religion and politics from the Restoration to the French Revolution. History of Science (UK), 29(4):421–428, December 1991. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/29/4/421.full.pdf+html.

Woods:2013:PMS

[Woo13] Abigail Woods. From practical men to scientific experts: British veterinary surgeons and the development of government scientific expertise, c. 1878–1919. History of Science (UK), 51(4):457–480, December 2013. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/51/4/457.full.
pdf+html.

Worboys:1994:CHS

[Wor94] Michael Worboys. The comparative history of sleeping sickness in East and central Africa, 1900–1914. History of Science (UK), 32(1):89– 102, March 1994. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/32/1/
89.full.pdf+html.

Wise:1989:WWPa

[WS89a] M. Norton Wise and Crosbie Smith. Work and waste: Political economy and natural philosophy in Nineteenth Century Britain (i). History of Science (UK), 27(3):263–301, September 1989. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/27/3/263.full.pdf+html.

Wise:1989:WWPb

[WS89b] M. Norton Wise and Crosbie Smith. Work and waste: Political economy and natural philosophy in Nineteenth Century Britain (II). History of Science (UK), 27(4):391–449, December 1989. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/27/4/391.full.pdf+html.

Wise:1990:WWP

[WS90] M. Norton Wise and Crosbie Smith. Work and waste: Political economy and natural philosophy in Nineteenth Century Britain (III). History of Science (UK), 28(3):221–261, September 1990. CODEN HIS-CAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://

hos.sagepub.com/content/28/3/221.full.pdf+html.

Yates:1973:ERS

[Yat73] Frances A. Yates. Essay review: Science in its context: Science, medicine and society in the Renaissance: Essays to honor Walter Pagel. History of Science (UK), 11(4):286–291, December 1973. CO-DEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL

http://hos.sagepub.com/content/11/4/286.full.pdf+html.

Yeo:1985:IMP

[Yeo85] Richard Yeo. An idol of the market-place: Baconianism in Nineteenth Century Britain. History of Science (UK), 23(3):251–298, September 1985. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/23/3/251.full.
pdf+html.

Yeo:2007:BMP

[Yeo07] Richard Yeo. Between memory and paperbooks: Baconianism and natural history in Seventeenth-Century England. History of Science (UK), 45(1):1–46, March 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.sagepub.

com/content/45/1/1.full.pdf+html.

Yildirim:2023:OPN

[Yil23] Duygu Yildirim. Ottoman plants, nature studies, and the attentiveness of translational labor. History of Science (UK), 61(4):??, December 1, 2023. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL https://journals.sagepub.com/doi/abs/
10.1177/00732753231191340.

Yolton:1971:ERP

[Yol71] John W. Yolton. Essay review: Philosophy of science from Descartes to Kant: Metaphysics and the philosophy of science: The classical origins, Descartes to Kant. History of Science (UK), 10(1):102– 113, March 1971. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/10/1/
102.full.pdf+html.

Young:1966:SHB

[You66] Robert M. Young. Scholarship and the history of the behavioural sciences. History of Science (UK), 5(1):1–51, March 1966. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http:

//hos.sagepub.com/content/5/1/1.full.pdf+html.

Youschkevitch:1967:RLM

[You67] A. P. Youschkevitch. Recherches sur l’histoire des mathématiques au moyen age dans les pays d’orient: bilans et perspectives. (French) [Research on the history of mathematics in the Middle Ages in the Eastern countries: assessments and prospects]. History of Science (UK), 6(1):41–58, March 1967. CODEN HISCAR. ISSN 0073-2753 (print),

1753-8564 (electronic). URL http://hos.sagepub.com/content/6/
1/41.full.pdf+html.

Youschkevitch:1970:ERL

[You70] A. P. Youschkevitch. Essay review: L’oeuvre mathématique de Newton de 1667 à 1673: The Mathematical Papers of Isaac Newton. History of Science (UK), 9(1):105–119, March 1970. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/9/1/105.full.pdf+html.

Youschkevitch:1975:ERN

[You75] A. P. Youschkevitch. Essay review: Newton’s mathematical development 1674–1684: The mathematical papers of Isaac Newton. History of Science (UK), 13(4):290–299, December 1975. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/13/4/290.full.pdf+html.

Yoxen:1979:WDS

[Yox79] Edward J. Yoxen. Where does Schrödinger’s What Is Life? belong in the history of molecular biology? History of Science (UK), 17(1): 17–52, March 1979. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL http://hos.sagepub.com/content/17/1/
17.full.pdf+html.

Zelko:2019:ONI

[Zel19] Frank Zelko. Optimizing nature: Invoking the “natural” in the struggle over water fluoridation. History of Science (UK), 57(4):518–539, December 1, 2019. CODEN HISCAR. ISSN 0073-2753 (print), 1753-

8564 (electronic). URL https://journals.sagepub.com/doi/full/
10.1177/0073275318809764.

Zilhao:2014:RFS

[Zil14] Isabel Zilhão. The rise and fall of science for all: Science for children voiced by a Portuguese daily newspaper (1924–1933). History of Science (UK), 52(4):454–488, December 2014. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/52/4/454.full.pdf+html.

Zubov:1964:BRS

[Zub64] V. P. Zubov. Book review: Science and the Renaissance. History of Science (UK), 3(1):131–134, March 1964. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (electronic). URL http://hos.

sagepub.com/content/3/1/131.full.pdf+html.

Zuidervaart:2007:PLW

[Zui07] Huib J. Zuidervaart. “A plague to the learned world”: Pieter Gabry, F.R.S. (1715–1770) and his use of natural philosophy to gain prestige and social status. History of Science (UK), 45(3):287–326, September 2007. CODEN HISCAR. ISSN 0073-2753 (print), 1753-8564 (elec-

tronic). URL http://hos.sagepub.com/content/45/3/287.full.
pdf+html.